### v6

# Various Properties of Various Ultrafilters, Various Graph Width Parameters, and Various Connectivity Systems

Abstract
Filters are collections of sets that are closed under supersets and finite intersections, serving as fundamental tools in topology and set theory. An ultrafilter, a maximal filter on a set, plays a crucial role in these fields by rigorously handling limits, convergence, and compactness. A connectivity system is defined as a pair $(X,f)$, where $X$ is a finite set and $f$ is a symmetric submodular function. Understanding the duality in these parameters elucidates the relationship between different decompositions and measures of a graph’s complexity. In this paper, we delve into ultrafilters on connectivity systems, applying Tukey’s Lemma to these systems. Additionally, we explore prefilters, ultra-prefilters, and subbases within the context of connectivity systems. Furthermore, we introduce and investigate new parameters related to width, length, and depth, enhancing our understanding of these mathematical structures. We also investigate and compare various graph width parameters and their related parameters. (It also includes information similar to that found in surveys, aiming to promote future research on graph width parameters.)
keywords:
Ultrafilter, Filter, Connectivity System, Prefilter, Branch-width
category:
[…]
1 Introduction
1.1 Filter and Graph width parameters
Filters are essential collections of sets in topology and set theory, characterized by their closure under supersets and finite intersections. They can be viewed as a way to "focus" on certain subsets of a space, much like a lens that sharpens specific details while filtering out others. An ultrafilter, being a maximal filter on a set, is particularly significant for addressing fundamental concepts such as limits, convergence, and compactness. Its unique properties render it indispensable in various fields, including non-standard analysis, model theory, social choice theory (social judgments), group theory, boolean algebra, geometry, probability theory, vector theory, semigroup theory, abstract algebra, topology, set theory, infinite combinatorics, fuzzy theory, graph theory, matroid theory, lattice theory, and first-order logic, where it provides powerful tools for both mathematical and logical applications[304, 302, 719, 389, 3, 167, 340, 475, 642, 78, 144, 544, 295, 142, 297, 200].
Graph theory, a key branch of mathematics, delves into the study of networks consisting of nodes and edges, with a focus on their paths, structures, and properties [696, 187, 99]. Among the critical metrics in this field is the "graph width parameter," which measures the maximum width across all cuts or layers within a hierarchical decomposition of the graph. This metric is crucial for analyzing a graph’s complexity and structure, serving as a primary factor in transforming computationally hard graph problems into more tractable ones when the graph class is restricted to having bounded width. Various width parameters have been rigorously explored and are widely recognized in the literature [330, 577, 232, 242, 304, 19, 613, 486, 370, 423, 279, 485, 405, 296, 667, 298]. For further information, see Appendix A, “Various Width Parameters,” and Appendix C, “Comparing Various Graph Parameters (Over 70 Parameters).”
Branch-width is an important graph width parameter, defined through branch decomposition where the leaves correspond to the graph’s edges [330, 577, 84]. It is closely related to tree-width, another significant graph width parameter. Tree-width is determined by a tree decomposition, which represents the graph as a tree structure, grouping vertices into "bags," with the tree-width being the size of the largest bag minus one [613, 608]. Conversely, branch-width involves a branch decomposition where the graph is segmented into a tree-like structure, with the branch-width representing the maximum size of the minimum cut between two parts of the graph. The relationship between branch-width (bw(G)) and tree-width (tw(G)) is given by the inequalities: bw(G) $\leq$ tw(G)+1 $\leq$ $\frac{3}{2}$cdot bw(G) for a graph $G$ with bw(G) $\geq$ 2 [84]. This relationship shows that both parameters are linked, with tree-width generally being larger, but not excessively so. Extending tree-width to the connectivity system framework leads to the concept of branch-width, highlighting their conceptual connection.
Graph width parameters offer several significant advantages in both theoretical and practical contexts:
  1. 1.
    Foundational Role in Theoretical Graph Theory: Graph width parameters play a critical role in the Graph Minors project serving as a fundamental combinatorial tool [330, 577]. This project, which is pivotal in the structure theory of graphs, uses graph width parameters to explore graph properties and their relationships through a series of influential theorems and algorithms.
  2. 2.
    Algorithmic Efficiency: Graph width parameters are conducive to algorithmic applications, particularly in the field of fixed-parameter tractable (FPT) algorithms [267, 305, 257, 638]. Algorithms that are parameterized by Graph width parameters often demonstrate superior efficiency and are widely utilized in computational graph theory. This aspect is particularly valuable in optimizing complex computations and in the development of algorithms that can efficiently solve problems considered intractable by other means.
  3. 3.
    Practical Applicability: In real-world applications, graphs derived from various domains such as programming language [140, 665], road networks [722, 694], and organizational structures [563] often exhibit small width. This characteristic simplifies complex problems, making them more manageable and allowing for the application of advanced graph algorithms. As a result, studying graph width parameters can directly impact the effectiveness and efficiency of practical solutions in engineering, software development, and logistics [688, 152, 708, 634, 140, 186, 439, 189].
A pair $(X,f)$, consisting of a finite set $X$ and a symmetric submodular function $f$, is recognized as a connectivity system [330]. This concept is widely utilized in the analysis of graph structures, particularly in relation to graph width parameters such as branch-width and tree-width [330, 577, 304]. Exploring the duality (the min-max theorem) within these parameters enhances our understanding of the relationships among various graph decompositions and measures of graph complexity. In this context, "duality" refers to a theorem or relationship where the presence (or absence) of one entity implies the absence (or presence) of its corresponding dual entity, often referred to as a minmax theorem [625]. Ultrafilters on connectivity systems are known to exhibit such a dual relationship with branch-width [304]. Concepts like ultrafilters, often termed as obstructions, play a crucial role in determining the values of graph width parameters.
1.2 Our Contribution
This paper outlines our contributions as follows:
  • Section 2: We primarily explain the basic concepts of ultrafilters on connectivity systems and graph-width parameters, along with previously known concepts. We discuss the relationship between ultrafilters on connectivity systems and well-known concepts like Tangle from Graph Minor theory and Matroid commonly used in optimization theory.
  • Section 3: We delve into ultrafilters on connectivity systems, considering Tukey’s Lemma for these systems. Tukey’s Lemma asserts that every non-empty collection of sets, closed under supersets, contains a maximal element. This fundamental result in set theory is often used to prove the existence of ultrafilters and related concepts [719, 389, 3]. Additionally, we explore chains and antichains on connectivity systems.
  • Section 4: We discuss prefilters, ultra-prefilters, and subbases on connectivity systems. Prefilters, ultra-prefilters, and subbases are known concepts used to generate filters and are studied in various fields. This exploration enhances our understanding of these concepts in Set Theory and their significance in different mathematical and logical contexts.
  • Section 5: We examine the connectivity system discussed for finite sets from the perspective of infinite connectivity systems and countable connectivity systems. We analyze the properties of ultrafilters on these connectivity systems.
  • Section 6: We investigate new parameters such as width, length, and depth, as well as the Ultrafilter game on connectivity systems, ultraproduct on connectivity systems, and the Axiom of Choice on connectivity systems. We also consider a Small Set Expansion [702, 39, 714, 137, 251, 70].
  • Appendix: This section presents an investigation and comparison of various graph width parameters and their related parameters. It also includes information similar to that found in surveys, aiming to promote future research on graph width parameters.
Note 1.1.
Please note that preprints and early-stage research may not have undergone peer review. Additionally, as I am an independent researcher, please understand. Sorry.
2 Definitions and Notations in this paper
This section provides mathematical definitions for each concept. Before delving into specific definitions, let’s outline the basic mathematical concepts used in this text.
2.1 Notation in this paper
We explain the notation used in this paper.
2.1.1 Notation for Set theory
In set theory, a set is a collection of distinct elements or objects, considered as an entity and often denoted with curly braces. A subset is a set where all elements are also contained within another set. Boolean algebra $(X,\cup,\cap)$ is a mathematical structure with a set $X$, union ()($\cup$)( ∪ ), and intersection ()($\cap$)( ∩ ), satisfying specific axioms for operations. In this paper, we consider finite sets (except in Section 5).
Note 2.1.
In this paper, we use expressions like $A\subseteq X$ to indicate that $A$ is a subset of $X$, $A\cup B$ to represent the union of two subsets $A$ and $B$ (both of which are subsets of $X$), and $A=\emptyset$ to signify an empty set. Specifically, $A\cap B$ denotes the intersection of subsets $A$ and $B$. Similarly, $A\setminus B$ represents the difference between subsets $A$ and $B$.
Definition 2.2.
The powerset of a set $A$, denoted as $2^{A}$, is the set of all possible subsets of $A$, including the empty set and $A$ itself.
Definition 2.3.
In set theory, a partition of a set is a way of dividing the set into non-overlapping, non-empty subsets, such that every element of the original set is included in exactly one subset. These subsets are called the "blocks" or "parts" of the partition (cf.[559]).
For more detailed information on set theory, please refer to surveys and lecture notes (ex.[497, 679, 272, 483]).
2.1.2 Notation for undirected graph
Except appendix and section 6, we consider a simple undirected graph $G$, where the vertex set is denoted by $V(G)$ and the edge set by $E(G)$. For simplicity, we will often write $G=(V,E)$, with $V$ representing the vertices and $E$ the edges. If $X$ is a subset of the vertices $V(G)$ (or the edges $E(G)$), then $X^{c}$ represents the complement set $V(G)\setminus X$ (or $E(G)\setminus X$, respectively), which includes all elements not in $X$.
Definition 2.4.
The degree of a vertex refers to the number of edges connected to it. For example, if a vertex has three edges connected to it, its degree is three.
Definition 2.5.
A path graph is a type of tree that consists of a sequence of vertices where each vertex is connected to exactly two others, except for the two endpoints, which are connected to only one other vertex. A linear ordering of a graph’s vertices is an arrangement of the vertices in a linear sequence in such a way that each vertex comes before all vertices to which it is connected by edges in the sequence, except possibly the previous vertex.
Example 2.6.
Consider the path graph $P_{4}$ consisting of four vertices $v_{1},v_{2},v_{3},v_{4}$ and three edges $\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\}$. The vertices can be arranged linearly as $v_{1},v_{2},v_{3},v_{4}$, which respects the edge connections, making it a valid path graph. The endpoints $v_{1}$ and $v_{4}$ are connected to only one other vertex, while the intermediate vertices $v_{2}$ and $v_{3}$ are each connected to two vertices.
Definition 2.7.
A tree is a connected, acyclic graph where any two vertices are connected by exactly one path, symbolizing hierarchical relationships. The root of a tree is the topmost node from which all other nodes descend. In a tree, vertices with a degree of 1 are called leaves, while all other vertices are referred to as inner vertices. A ternary tree is a specific type of tree where all inner vertices have a degree of 3.
Example 2.8.
Consider a tree with 7 vertices, labeled $v_{1}$ to $v_{7}$, where $v_{1}$ is the root. The edges of the tree are $\{v_{1},v_{2}\}$, $\{v_{1},v_{3}\}$, $\{v_{1},v_{4}\}$, $\{v_{2},v_{5}\}$, $\{v_{2},v_{6}\}$, and $\{v_{3},v_{7}\}$. This forms a ternary tree, as the root $v_{1}$ has three children $v_{2}$, $v_{3}$, and $v_{4}$, making it an inner vertex with degree 3. The vertices $v_{5}$, $v_{6}$, $v_{7}$, and $v_{4}$ are leaves with a degree of 1.
Definition 2.9.
A subgraph is a graph formed from a subset of a graph’s vertices and edges. A subpath is a continuous segment of a path, consisting of consecutive edges and vertices from the original path, and is useful for analyzing specific portions of larger paths.
Example 2.10.
Consider the original graph $G$ with vertices $V(G)=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ and edges $E(G)=\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\},\{v_{4},v_{5}\}\}$. A subgraph $H$ can be formed by selecting the subset of vertices $V(H)=\{v_{2},v_{3},v_{4}\}$ and the corresponding edges $E(H)=\{\{v_{2},v_{3}\},\{v_{3},v_{4}\}\}$. This subgraph $H$ represents a subpath of the original path in $G$, highlighting a specific segment of the larger path.
Definition 2.11.
The distance between two vertices $u$ and $v$ in a graph $G=(V,E)$ is defined as the length of the shortest path connecting $u$ and $v$ in $G$. Formally, the distance $d_{G}(u,v)$ is given by:
$d_{G}(u,v)=\min\{\ell(P)\mid P\text{ is a path in }G\text{ connecting }u\text{and }v\}$
where $\ell(P)$ denotes the length of the path $P$, which is the number of edges in $P$.
Example 2.12.
Consider the graph $G$ with vertices $V(G)=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ and edges $E(G)=\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\},\{v_{4},v_{5}\},\{v_{1},v_{3}\}\}$.
To find the distance $d_{G}(v_{1},v_{4})$, observe that there are two paths connecting $v_{1}$ and $v_{4}$:
  • Path 1: $v_{1}\to v_{2}\to v_{3}\to v_{4}$, with a length of 3 edges.
  • Path 2: $v_{1}\to v_{3}\to v_{4}$, with a length of 2 edges.
The shortest path is Path 2, so the distance $d_{G}(v_{1},v_{4})$ is 2.
For more detailed information on graphs, please refer to surveys and lecture notes (ex.[601, 98, 188, 682, 97]).
2.1.3 Notation for directed graph
We define a directed graph $D$, where the vertex set is denoted by $V(D)$ and the edge set by $E(D)$. For clarity, we will use the notation $D=(V,E)$, where $V$ represents the vertices and $E$ represents the directed edges. If $X$ is a subset of either the vertex set $V(D)$ or the edge set $E(D)$, then $X^{c}$ denotes the complement, $V(D)\setminus X$ or $E(D)\setminus X$, respectively, consisting of all elements not contained in $X$.
Definition 2.13.
The degree of a vertex in a directed graph is defined by the number of directed edges incident to it, separated into in-degree and out-degree. For example, a vertex with three incoming edges has an in-degree of three.
Definition 2.14.
A path in a directed graph is a sequence of vertices where each vertex is connected to the next by a directed edge. A linear ordering of the vertices of a directed graph arranges them in a sequence such that, for each directed edge, the source vertex precedes the target vertex in the sequence.
Definition 2.15.
A tree in the context of directed graphs is a directed acyclic graph (DAG) that is connected and contains no directed cycles. A directed tree with $n$ vertices always has $n-1$ directed edges. A leaf in a directed tree is a vertex with exactly one incoming or outgoing edge, making it a terminal vertex. In this context, a node is any vertex, which can be either a leaf or an internal vertex depending on its degree. A subcubic directed tree is one in which every vertex has a total degree (in-degree plus out-degree) of at most three, meaning no vertex is connected by more than three directed edges.
For more detailed information on directed graphs, please refer to surveys and lecture notes (ex.[43, 42]).
2.2 Symmetric Submodular Function and Connectivity System
The definition of a symmetric submodular function is presented below. This concept is extensively used and discussed in numerous scholarly articles [478, 409, 41]. While symmetric submodular functions are generally defined over real numbers, this paper specifically considers those restricted to natural numbers. This submodular function is also sometimes referred to as the connectivity function [330]. Additionally, a variant of submodular function known as the submodular partition function is also well-known [22, 734]. Also, $k$-Submodular [401, 692, 691], two-dimensional submodular [533], monotone-submodular [250], and maximum-submodular [262] are also known.
Definition 2.16.
Let $X$ be a finite set. A function $f:2^{X}\rightarrow\mathbb{N}$ is called symmetric submodular if it satisfies the following conditions:
  • $\forall A\subseteq X,\ f(A)=f(X\setminus A)$.
  • $\forall A,B\subseteq X,\ f(A)+f(B)\geq f(A\cap B)+f(A\cup B)$.
In this paper, a pair $(X,f)$ of a finite set $X$ and a symmetric submodular function $f$ is called a connectivity system. This concept is frequently used in discussions of graph width parameters, such as branch-width and tree-width, to analyze graph structures (e.g., [194, 292, 290, 330]).
The following is an example illustrating the concept of a symmetric submodular function.
Example 2.17.
Consider a simple undirected graph $G=(V,E)$ where $V$ is the set of vertices and $E$ is the set of edges. Let $V=\{1,2,3,4\}$ and the edges $E=\{(1,2),(2,3),(3,4),(4,1)\}$, forming a cycle. Define a function $f:2^{V}\rightarrow\mathbb{N}$ as follows:
$f(A)=|E(A,V\setminus A)|$
where $E(A,V\setminus A)$ is the set of edges with one endpoint in $A$ and the other endpoint in $V\setminus A$. The symmetric submodularity condition is satisfied.
Example 2.18.
For a set of random variables $X_{1},X_{2},\dots,X_{n}$, the entropy function $H(S)$ for a subset $S\subseteq\{X_{1},X_{2},\dots,X_{n}\}$ is defined as the joint entropy of the variables in $S$. This function is known to be submodular and symmetric. Generally, the entropy function $H(S)$ itself is not symmetric, as it depends on the specific set of variables $S$. However, the mutual information between sets of variables, which is related to entropy, is symmetric. And note that entropy, in information theory, measures the uncertainty or randomness of a random variable’s outcomes. It quantifies the average amount of information produced by a stochastic source of data (cf.[635, 348]).
Example 2.19.
In the context of network flow, the cut-value function $f(S)$ for a subset $S\subseteq V$ (where $V$ is the set of vertices) is the total capacity of edges crossing from $S$ to $V\setminus S$. This is an example of a symmetric submodular function.
Example 2.20.
Let $\Omega=\{v_{1},v_{2},\dots,v_{n}\}$ be the set of vertices of a directed graph. For any set of vertices $S\subseteq\Omega$, define the function $f(S)$ to denote the number of edges $e=(u,v)$ such that $u\in S$ and $v\in\Omega\setminus S$. This function counts the number of edges leaving the set $S$.
The function $f(S)$ is submodular, meaning that for any two sets $A$ and $B$ with $A\subseteq B\subseteq\Omega$ and any vertex $x\in\Omega\setminus B$, it holds that:
$f(A\cup\{x\})-f(A)\geq f(B\cup\{x\})-f(B).$
However, $f(S)$ is not symmetric. Specifically, $f(S)$ depends on the direction of the edges, so in general, $f(S)\neq f(\Omega\setminus S)$. This asymmetry arises because the number of edges leaving $S$ is typically different from the number of edges entering $S$.
Moreover, this function can be generalized by assigning non-negative weights to the directed edges. In this case, $f(S)$ would represent the total weight of the edges leaving the set $S$, which still maintains the submodular property without being symmetric.
It is known that a symmetric submodular function $f$ satisfies the following useful properties:
Lemma 2.21.
[330] Let $X$ be a finite set. A symmetric submodular function $f$ satisfies:
  1. 1.
    $\forall A\subseteq X,\ f(A)\geq f(\emptyset)=f(X)=0$.
  2. 2.
    $\forall A,B\subseteq X,\ f(A)+f(B)\geq f(A\setminus B)+f(B\setminus A)$.
Proof 2.22.
The following results can be obtained:
  1. 1.
    $f(A)+f(A)=f(A)+f(A)\geq f(A\cup A)+f(A\cap A)=f(X)+f(\emptyset)=f(\emptyset)+f(\emptyset)$.
  2. 2.
    $f(A)+f(B)=f(A)+f(B)\geq f(A\cup B)+f(A\cap B)=f(B\setminus A)+f(A\setminus B)$.
Thus, the proof is completed.
2.3 Ultrafilter on Connectivity System
In this subsection. we consider about ultrafilter on Connectivity System. First, we consider about ultrafilter for set theory.
2.3.1 Ultrafilter for set theory
First, let’s introduce the general concept of an Ultrafilter in set theory. The definition of an Ultrafilter in Set Theory is described as follows.
Definition 2.23.
Let $X$ be a set. A collection $\mathcal{F}\subseteq 2^{X}$ is called a filter on $X$ if it satisfies the following conditions111*1:
  1. (B1)
    If $A$ and $B$ are both in $\mathcal{F}$, then their intersection $A\cap B$ is also in $\mathcal{F}$.
  2. (B2)
    If $A$ is in $\mathcal{F}$ and $A\subseteq B\subseteq X$, then $B$ is also in $\mathcal{F}$.
  3. (B3)
    The empty set $\emptyset$ is not in $\mathcal{F}$.
Definition 2.24.
A maximal filter, which cannot be extended any further while still being a filter, is called an ultrafilter. An ultrafilter satisfies an additional condition:
  1. (BT1)
    For any subset $A\subseteq X$, either $A$ is in $\mathcal{F}$ or its complement $X\setminus A$ is in $\mathcal{F}$, but not both.
Example 2.25.
Consider the set $X=\{1,2,3\}$. The collection =$\mathcal{F}=\{\{1,2,3\},\{1,2\},\{2,3\},\{2\}\}$ is a filter on $X$. It satisfies all the conditions of a filter:
  • The intersection of any two sets in $\mathcal{F}$ is also in $\mathcal{F}$.
  • If $A$ is in $\mathcal{F}$ and $A\subseteq B\subseteq X$, then $B$ is also in $\mathcal{F}$.
  • The empty set is not in $\mathcal{F}$.
Example 2.26.
Another example of a filter is the Fréchet filter on an infinite set $X$. The Fréchet filter consists of all cofinite subsets of $X$ (i.e., subsets whose complements are finite). This collection forms a filter because:
  • The intersection of two cofinite sets is also cofinite.
  • Any superset of a cofinite set is cofinite.
  • The empty set is not cofinite, so it is not in the filter.
Several related concepts to Filters and Ultrafilters have been proposed. Here, I will introduce just one: the Weak Filter. This concept, discussed in the field of logic, is used to interpret defaults through a generalized "most" quantifier in first-order logic [476, 621, 34, 477, 141]. The definition and an example are provided below.
Definition 2.27.
[476] Let $A,B$ be a set. A weak filter over a non-empty set $X$ is a collection $\mathcal{F}_{w}\subseteq 2^{X}$ that satisfies the following conditions:
  1. (WFB0)w
    If $A\in\mathcal{F}_{w}$ and $A\subseteq B\subseteq X$, then $B\in\mathcal{F}_{w}$.
  2. (WBF1)w
    If $A,B\in\mathcal{F}_{w}$, then $A\cap B\neq\emptyset$.
  3. (WFB2)w
    $X\in\mathcal{F}_{w}$.
Definition 2.28.
[476] Let $A,B$ be a set. A weak ultrafilter over a non-empty set $X$ is a complete weak filter, meaning it additionally satisfies the following condition:
  1. (WFB3)w
    For every $A\subseteq X$, $A\notin\mathcal{F}_{w}$ if and only if $X\setminus A\in\mathcal{F}_{w}$.
Example 2.29.
Let $I=\{1,2,3\}$. Consider the collection $\mathcal{F}_{w}=\{\{1,2,3\},\{1,2\},\{2,3\},\{1,3\}\}$. This collection $\mathcal{F}_{w}$ is a weak filter over $X$ because it satisfies the following:
  • For any $A\in\mathcal{F}_{w}$ and any $A\subseteq B\subseteq X$, we have $B\in\mathcal{F}_{w}$.
  • The intersection of any two sets in $\mathcal{F}_{w}$ is non-empty.
  • The entire set $X=\{1,2,3\}$ is in $\mathcal{F}_{w}$.
For more detailed information for ultrafilter, please refer to surveys and lecture notes (ex.[340, 642]).
2.3.2 Ultrafilter on Connectivity system
We introduce some properties of the ultrafilter on the connectivity system $(X,f)$ as an extension of the Ultrafilter on Boolean Algebras. First, we introduce definitions of a filter and an ultrafilter on a connectivity system [304]. These concepts extend the traditional notions of filter and ultrafilter from set theory by incorporating the condition of symmetric submodularity. Also, an ultrafilter on the connectivity system $(X,f)$ is co-Maximal ideal [713] on the connectivity system $(X,f)$.
Definition 2.30 ([304]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a connectivity system, the set family $F\subseteq 2^{X}$ is called a filter of order $k+1$ if the following axioms hold true:
  • (Q0)
    $\forall A\in F,\ f(A)\leq k$
  • (Q1)
    $A\in F,\ B\in F,\ f(A\cap B)\leq k\Rightarrow A\cap B\in F$
  • (Q2)
    $A\in F,\ A\subseteq B\subseteq X,\ f(B)\leq k\Rightarrow B\in F$
  • (Q3)
    $\emptyset\notin F$
An Ultrafilter on a connectivity system is a filter on a connectivity system that satisfies the following condition $(Q4)$. The Ultrafilter on a Connectivity System has a dual relationship with branch-width, a graph width parameter [304]. Note that an ultrafilter on a connectivity system is a non-empty and proper (i.e., $\emptyset\notin F$) family.
  • (Q4)
    $\forall A\subseteq X,\ f(A)\leq k\Rightarrow\text{either }A\in F\text{ or }X\setminus A\in F$
Definition 2.31.
[304] A filter is principal if the filter satisfies the following axiom $(QP5)$ [304]:
  • (QP5)
    $A\in F$ for all $A\subseteq X$ with $|A|=1$ and $f(A)\leq k$.
A filter is non-principal if the filter satisfies the following axiom $(Q5)$ [304]:
  • (Q5)
    $A\notin F$ for all $A\subseteq X$ with $|A|=1$.
Non-principal refers to a filter or ideal that does not contain any singletons (i.e., sets with exactly one element). It is not generated by any finite set.
If a filter is weak [288, 475, 35], the following axiom $(QW1^{\prime})$ holds instead of axiom $(Q1)$. Note that weak filter aims at interpreting defaults via a generalized ‘most’ quantifier in first-order logic [475, 35, 622, 276]. A weak filter on a connectivity system is co-Weak Ideal on a connectivity system:
  • (QW1’)
    $A\in F,\ B\in F,\ f(A\cap B)\leq k\Rightarrow A\cap B\neq\emptyset$
If a filter is quasi [293, 143], the following axiom $(QQ1^{\prime})$ holds instead of axiom $(Q1)$. Note that a quasi-Ultrafilter is used to do an axiomatic analysis of incomplete social judgments [143]:
  • (QQ1’)
    $A\subseteq X,\ B\subseteq X,\ A\notin F,\ B\notin F\Rightarrow A\cup B\notin F$
Note 2.32.
[…]
  • (QS1)
    For any $A\in F$, $e\in X$, if $f(\{e\})\leq k$ and $f(A\cap(X\setminus\{e\}))\leq k$, then $A\cap(X\setminus\{e\})\in F$.
[…]
  • (QSD1)
    For $A\in F$, $e\in X$, if $f(A\setminus\{e\})\leq k$, then $A\setminus\{e\}\in F$.
Note that an equivalent form of a given ultrafilter $U\subseteq 2^{X}$ on a connectivity system is manifested as a two-valued morphism. This relationship is defined through a function $m$ where $m(A)=1$ if $A$ is an element of $U$, and $m(A)=0$ otherwise. Additionally, the function $m$ holds the value 1 for subsets where the value of the submodular function is at most $k$. This suggests that ultrafilters are well-suited for use in two-player games [33, 619], such as Cops and Robbers. Indeed, in literature [304], an ultrafilter on a connectivity system is employed as a winning strategy.
2.3.3 Relation between Tangle and Ultrafilter
A concept closely related to ultrafilters on a connectivity system is the tangle. Tangles are known to be dual to the concepts of tree-decomposition and branch-decomposition, and they are extensively used in the study of graph width parameters [330, 577, 613]. Tangles have played a significant role in algorithms across various fields, including graph minors, width parameters, and graph isomorphism problems. The definition of a tangle on a connectivity system is provided below.
Definition 2.33 ([330]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. A family $T\subseteq 2^{X}$ is a tangle of order $k+1$ on a connectivity system $(X,f)$ if $T$ satisfies the following axioms:
  • (T1)
    $\forall A\in T,\ f(A)\leq k$,
  • (T2)
    $A\subseteq X,\ f(A)\leq k\Rightarrow\text{either }(A\in T)\text{ or }(X\setminus A\in T)$,
  • (T3)
    $A,B,C\in T\Rightarrow A\cup B\cup C\neq X$,
  • (T4)
    $\forall e\in X,\ X\setminus\{e\}\notin T$.
A tangle of order $k+1$ is also a tangle of order $k$. Moreover, if $k$ is too large, a tangle cannot exist. The maximum $k$ for which a tangle can exist is called the tangle number, which coincides with the branch-width. The same principle applies to ultrafilters. (cf.[294, 330]) Actually, in an ultrafilter on a connectivity system, the following properties hold.
Theorem 2.34.
Let $(X,f)$ be a connectivity system. An ultrafilter of order $k+1$ on a connectivity system $(X,f)$ is also an ultrafilter of order $k$ on a connectivity system $(X,f)$.
Proof 2.35.
An ultrafilter of order $k+1$ on a connectivity system $(X,f)$ obviously holds axioms of an ultrafilter of order $k$ on a connectivity system $(X,f)$.
Given a set $Y\subseteq X$, the collection $\{A\subseteq X\mid A\subseteq Y,f(A)\leq k\}$ forms a tangle. This tangle is referred to as a principal tangle generated by the set $Y$ [352]. The set $Y$ is called the generator of this tangle [191].
A family $F$ on a connectivity system $(X,f)$ is a co-tangle if there is a tangle $T$ such that $A\in F$ if and only if $X\setminus A\in T$. Filter on a connectivity system is a co-tangle on the connectivity system. A filter on a connectivity system also satisfies the following conditions [304]:
  • (FT1)
    $A,B,C\in F\Rightarrow A\cap B\cap C\neq\emptyset$.
A Single-Filter on a connectivity system is a co-linear tangle on that system. It is important to note that a linear tangle is a concept dual to linear-width, which will be discussed later. The main difference between a linear tangle and a co-linear tangle lies in one of the sets in axiom (T2), which is replaced by a singleton set $\{e\}$. A linear tangle of order $k+1$ is also a linear tangle of order $k$. Furthermore, if $k$ is too large, a linear tangle cannot exist. In this paper, the largest possible $k$ for which a linear tangle can exist is referred to as the linear tangle number.
Single-Filter on a connectivity system also satisfies the following conditions [290]:
  • (FLT1)
    $A,B\in F,\ e\in X,\ f(\{e\})\leq k\Rightarrow A\cap B\cap(X\setminus\{e\})\neq\emptyset$.
[…]
  • (FLTR1)
    $A,B\in F,\ |A|\neq|B|,\ e\in X,\ f(\{e\})\leq k\Rightarrow A\cap B\cap(X\setminus\{e\})\neq\emptyset$.
Additionally, concepts such as Directed Ultrafilter [299], defined on directed graphs, Edge-Ultrafilter [291], and Ultrafilter Tangle [485, 484, 487, 233], on the Abstract separation system, which abstracts the Connectivity system, have also been studied. It is fascinating to see the connection between seemingly unrelated concepts of Ultrafilter and Tangle when the condition of symmetric submodularity is applied.
It is known that a Bramble can be constructed using a concept essentially equivalent to an Ultrafilter on separation [194]. A Bramble (cf. Strict Bramble[492]) is recognized in graph game theory, particularly in the Cops and Robbers game, as an obstruction to well-known graph width parameters such as escape routes and tree-width [384].
2.4 Branch width on Connectivity System
In this section, we consider about branch width on connectivity system.
2.4.1 Branch width
Branch width is a significant graph width parameter that involves a branch decomposition where the leaves correspond to the graph’s edges. Each edge is assigned a value from a symmetric submodular function, measuring connectivity. Branch width generalizes the width of symmetric submodular functions on graphs, making it crucial in graph theory (e.g., [330, 304]).
Definition 2.36 ([577]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. Let $(X,f)$ be a connectivity system. The pair $(T,\mu)$ is a branch decomposition tree of $(X,f)$ if $T$ is a ternary tree such that $|L(T)|=|X|$ and $\mu$ is a bijection from $L(T)$ to $X$, where $L(T)$ denotes the leaves in $T$. For each $e\in E(T)$, we define $\text{bw}(T,\mu,e)$ as $f\left(\bigcup_{v\in L(T_{1})}\mu(v)\right)$, where $T_{1}$ is a tree obtained by removing $e$ from $T$ (taking into account the symmetry property of $f$). The width of $(T,\mu)$ is defined as the maximum value among $\text{bw}(T,\mu,e)$ for all $e\in E(T)$. The branch-width of $X$, denoted by $\text{bw}(X)$, is defined as the minimum width among all possible branch decomposition trees of $X$.
Example 2.37.
Consider the connectivity system $(X,f)$ where $X=\{a,b,c,d\}$ and $f$ is defined as follows: for any subset $S\subseteq X$, $f(S)$ equals the number of elements in $S$. We construct a branch decomposition tree $T$ as follows:
  • $T$ is a ternary tree with leaves $L(T)=\{v_{a},v_{b},v_{c},v_{d}\}$.
  • The bijection $\mu$ maps $v_{a}$ to $a$, $v_{b}$ to $b$, $v_{c}$ to $c$, and $v_{d}$ to $d$.
To define the width of this branch decomposition tree, we consider each edge $e\in E(T)$ and the corresponding tree $T_{1}$ obtained by removing $e$ from $T$. For instance, remove an edge $e$ that splits $T$ into subtrees $T_{1}$ and $T_{2}$ where $T_{1}$ contains leaves $\{v_{a},v_{b}\}$ and $T_{2}$ contains leaves $\{v_{c},v_{d}\}$. In this case, $\mu(\{v_{a},v_{b}\})=\{a,b\}$ and $\mu(\{v_{c},v_{d}\})=\{c,d\}$. The width $\text{bw}(T,\mu,e)$ is then calculated as:
$\text{bw}(T,\mu,e)=f(\{a,b\})=2$
Repeating this process for all edges $e$ in $T$, we determine the maximum width among all these values. In this simple example, the width $\text{bw}(T,\mu)$ is the maximum value of $f(S)$ for any partition $S$ of $X$, which is 2. Finally, the branch-width $\text{bw}(X)$ of the connectivity system $(X,f)$ is the minimum width among all possible branch decomposition trees of $X$. Here, the minimum width is achieved by our constructed tree $T$, and thus $\text{bw}(X)=2$.
In graph theory, the duality theorem for width parameters, such as tree-width and branch-width, is discussed, highlighting their dual concepts like tangles and branch decompositions[330, 577, 116, 290, 304, 19]. Additionally, obstructions are minimal structures or subgraphs that prevent a graph from having a width parameter below a certain threshold, providing insight into the graph’s complexity. The following duality theorem is known for the branch-width of the connectivity system $(X,f)$ and the ultrafilter of the connectivity system $(X,f)$.
Theorem 2.38 ([304]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. The branch-width of the connectivity system $(X,f)$ is at most $k$ if and only if no (non-principal) ultrafilter of order $k+1$ exists.
Also, the following duality theorem is known for the branch-width of the connectivity system $(X,f)$ and the tangle of the connectivity system $(X,f)$.
Theorem 2.39 ([330]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. The branch-width of the connectivity system $(X,f)$ is at most $k$ if and only if no tangle of order $k+1$ exists.
2.4.2 Linear Branch width
Here, we introduce the concept of linear decomposition, which is a linear version of branch decomposition. Like branch decomposition, linear branch decomposition has been the subject of extensive research [369, 303, 300]. Focusing on linear structures often facilitates deriving results for both general width parameters and linear width parameters.
Definition 2.40.
(cf.[301]) Let $X$ be a finite set and $f$ be a symmetric submodular function. Let $C$ be a caterpillar, defined as a tree where interior vertices have a degree of 3 and leaves have a degree of 1. Consider $C$ as the path $(l_{1},b_{2},b_{3},\ldots,b_{n-1},l_{n})$. For $12\leq i\leq n-$ 1, the subgraph of $C$ induced by $\{b_{i-1},b_{i},b_{i+1}\}$ forms a connectivity system $(X,f)$. The Linear Decomposition of $C$ is a process that partitions the elements of $X$ into the sets $\{e_{1}\},\{e_{2}\},\ldots,\{e_{n-1}\},\{e_{n}\}$. For each $11\leq i\leq n-$ 1, define $w_{i}:=f(\{e_{1},\ldots,e_{i}\})$. The width of the Linear Decomposition is given by $\max\{w_{1},\ldots,w_{n-1},f(e_{1}),\ldots,f(e_{n-1}),f(e_{n})\}$. The linear width of $(X,f)$ is the smallest width among all Linear Decompositions of $(X,f)$.
Example 2.41.
Consider a finite set $X=\{a,b,c,d\}$ and a symmetric submodular function $f$ defined on $2^{X}$. Suppose $C$ is a caterpillar tree, where the path is represented by vertices $l_{1},b_{2},b_{3},l_{n}$ with edges between them. Let’s assume the caterpillar has 4 vertices with the structure $l_{1}-b_{2}-b_{3}-l_{n}$, where $l_{1}$ and $l_{n}$ are leaves, and $b_{2}$ and $b_{3}$ are interior vertices. In this scenario, we define a Linear Decomposition of $X$ by partitioning it into individual elements $\{a\},\{b\},\{c\},\{d\}$. For each partition step $i$, where $1\leq i\leq 3$, the width $w_{i}$ is calculated using the submodular function $f$.
  • For example, $w_{1}=f(\{a\})$, $w_{2}=f(\{a,b\})$, $w_{3}=f(\{a,b,c\})$.
  • If the subgraph of $C$ induced by $\{b_{2},b_{3}\}$ forms the connectivity system with the function $f(\{b_{2}\})=2$, $f(\{b_{3}\})=3$, and $f(\{b_{2},b_{3}\})=2$, then the width of this Linear Decomposition would be the maximum value among these: $\max\{w_{1},w_{2},w_{3},f(\{b_{2}\}),f(\{b_{3}\})\}$.
Finally, the linear branch width of $(X,f)$ is the smallest width obtained by considering all possible Linear Decompositions of $(X,f)$. In this case, the linear width would be determined by minimizing the maximum width $w_{i}$ across different decompositions of the caterpillar tree.
Theorem 2.42 ([290]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. The linear-branch-width of the connectivity system $(X,f)$ is at most $k$ if and only if no (non-principal) Single-Ultrafilter of order $k+1$ exists.
Based on the above, the relationships shown in the following diagram become clear.
Figure 1: Relation about Width Parameter and Obstruction on Connectivity system
3 Tukey’s Lemma and Chain for Connectivity Systems
In this section, we explain some properties of ultrafilters on connectivity systems.
3.1 Chain for Connectivity Systems
First, let’s consider a chain on a connectivity system. Note that a chain in mathematics is crucial because it represents a totally ordered sequence of elements, which is fundamental in studying hierarchical structures, lattice theory, and optimization problems [50, 507, 211, 508, 382, 338, 273].
Definition 3.1.
Let $X$ be a finite set and $f$ be a symmetric submodular function. A chain of order $k+1$ on a connectivity system is a sequence of subsets $\{A_{1},A_{2},\dots,A_{m}\}$ of $X$ such that:
  1. 1.
    $A_{1}\subseteq A_{2}\subseteq\dots\subseteq A_{m}$.
  2. 2.
    For each $i$ $(1\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$.
3.2 Tukey’s Lemma for Connectivity System
We demonstrate the theorem (Tukey’s Lemma for Connectivity Systems) by using a chain of order $k+1$ on a connectivity system. Tukey’s Lemma states that in any non-empty collection of sets closed under taking supersets, there exists a maximal set within that collection. Intuitively, this means you can always find the largest possible set that cannot be expanded further while staying within the collection. Additionally, in set theory, Tukey’s Lemma is known to be closely related to Zorn’s Lemma (e.g., [498, 135, 52]). These lemmas are also widely recognized for their broad range of applications.
Theorem 3.2.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Any filter on a connectivity system $(X,f)$, where $f$ is a symmetric submodular function and $X$ is a non-empty finite set, can be extended to an ultrafilter. Specifically, for any $k$-efficient subset $A\subseteq X$ (i.e., $f(A)\leq k$), there exists an ultrafilter containing the filter.
Proof 3.3.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Here, we consider chains composed of $k$-efficient subsets.
First, we define the initial filter. Let $Q_{0}$ be a filter on the connectivity system $(X,f)$ where $f$ is a symmetric submodular function and $X$ is a non-empty finite set that satisfies the axioms $Q_{0}$ through $Q_{3}$.
Next, define the collection $\mathcal{F}$ as the set of all filters on $(X,f)$ that contain $Q_{0}$. This means $\mathcal{F}$ is the set of all sets $Q$ such that:
  • $Q_{0}\subseteq Q$
  • $Q$ satisfies axioms $Q_{0}$ through $Q_{3}$
$\mathcal{F}$ is partially ordered by inclusion. We need to show that every chain on $(X,f)$ (totally ordered subset on $(X,f)$) in $\mathcal{F}$ has an upper bound in $\mathcal{F}$.
Let $C\subseteq\mathcal{F}$ be a chain in $\mathcal{F}$, where each element of $C$ is a filter that respects the $k$-efficiency condition, meaning $f(A)\leq k$ for every $A$ in the filter.
Define $Q_{C}=\bigcup C$. We need to show that $Q_{C}$ is a filter and an element of $\mathcal{F}$.
Axiom (Q0): This axiom obviously holds.
Axiom (Q2): For any $A,B\in Q_{C}$, there exist $Q_{A},Q_{B}\in C$ such that $A\in Q_{A}$ and $B\in Q_{B}$. Since $C$ is a chain on the connectivity system, one of these filters contains the other; assume without loss of generality that $Q_{A}\subseteq Q_{B}$. Thus, $A,B\in Q_{B}$, and because $Q_{B}$ is a filter, $A\cap B\in Q_{B}\subseteq Q_{C}$. Additionally, since both $A$ and $B$ are in $Q_{C}$, $f(A)\leq k$ and $f(B)\leq k$ by Axiom (Q0). Therefore, $f(A\cap B)\leq k$ must hold because if it did not, $A\cap B$ would not be in $Q_{B}$, contradicting the closure under intersection in Axiom (Q1).
Axiom (Q1): For any $A\in Q_{C}$ and $A\subseteq B\subseteq X$, there exists $Q_{A}\in C$ such that $A\in Q_{A}$. Since $Q_{A}$ is a filter and $f(B)\leq k$, $B\in Q_{A}\subseteq Q_{C}$.
Axiom (Q3): Since $\emptyset\notin Q$ for any $Q\in C$, we have $\emptyset\notin Q_{C}$.
Thus, $Q_{C}$ is a filter on the connectivity system and $Q_{C}\in\mathcal{F}$.
$\mathcal{F}$ contains a maximal element $Q$. This maximal filter $Q$ satisfies axioms (Q0) through (Q3).
To show $Q$ is an ultrafilter, we need to verify Axiom (Q4): For any $A\subseteq X$ such that $f(A)\leq k$, suppose $A\notin Q$. We must show $X\setminus A\in Q$.
Assume $A\notin Q$. If $X\setminus A\notin Q$, then we will show that this leads to a contradiction, meaning $Q$ must contain either $A$ or $X\setminus A$.
  1. 1.
    Constructing $Q^{\prime}=Q\cup\{A\}$: If $A\notin Q$, consider adding $A$ to $Q$. This new set $Q^{\prime}=Q\cup\{A\}$ must be checked against the filter axioms.
  2. 2.
    Closed under Intersection: For $Q^{\prime}$, consider any $B\in Q$:
    • If $A$ were added to $Q$, $B\cap A$ might violate the $k$-efficiency condition. Specifically, since $A\notin Q$ and $Q$ is maximal, there could be an element $B\in Q$ such that $B\cap A$ does not satisfy $f(B\cap A)\leq k$. Since $B\cap A\notin Q$, this would violate Axiom (Q1), meaning $Q^{\prime}$ cannot be a filter.
  3. 3.
    Constructing $Q^{\prime\prime}=Q\cup\{X\setminus A\}$: If $X\setminus A\notin Q$, consider adding $X\setminus A$ to $Q$. This new set $Q^{\prime\prime}=Q\cup\{X\setminus A\}$ must also be checked against the filter axioms.
  4. 4.
    Closed under Intersection: For $Q^{\prime\prime}$, consider any $B\in Q$:
    • If $X\setminus A\notin Q$, then there must exist some $B\in Q$ such that $B\cap(X\setminus A)$ does not satisfy $f(B\cap(X\setminus A))\leq k$. Since $Q$ is maximal, adding $X\setminus A$ would create a set that fails to be closed under intersection, as $B\cap(X\setminus A)\notin Q$. This violates Axiom (Q1), meaning $Q^{\prime\prime}$ cannot be a filter.
Therefore, the assumption that both $A\notin Q$ and $X\setminus A\notin Q$ leads to contradictions. Hence, $Q$ must contain either $A$ or $X\setminus A$. This proof is completed.
We gain the following property of an ultrafilter on a connectivity system.
Theorem 3.4.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In any finite non-empty connectivity system $(X,f)$ where $f$ is a symmetric submodular function, there exists at least one ultrafilter.
Proof 3.5.
Consider the collection $\mathcal{F}$ of all filters on $(X,f)$. This collection is non-empty because the trivial filter $\{X\}$ exists. By Tukey’s Lemma for Connectivity Systems, every filter can be extended to an ultrafilter. Thus, there exists at least one maximal element in $\mathcal{F}$ under inclusion that satisfies all the filter axioms and the ultrafilter condition $(Q4)$. This proof is completed.
3.3 Antichain for Connectivity Systems: Relationship among Ultrafilter, Chain, and Antichain
We consider an antichain on a connectivity system. An antichain is a collection of elements in a partially ordered set where no element is comparable to another, meaning no element in the set precedes or follows any other element in the ordering (cf. [50, 507, 211, 508, 382, 338, 273]).
Definition 3.6.
Let $X$ be a finite set and $f$ a symmetric submodular function. An antichain of order $k+1$ on a connectivity system is defined as a collection of subsets $\{A_{1},A_{2},\dots,A_{m}\}$ of $X$ such that:
  1. 1.
    Antichain Condition: For any two distinct subsets $A_{i}$ and $A_{j}$ in the collection, neither is a subset of the other, i.e., $A_{i}\not\subseteq A_{j}$ and $A_{j}\not\subseteq A_{i}$ for all $1\leq i,j\leq m$ with $i\neq j$.
  2. 2.
    Submodular Condition: For each $i$ $(1\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$.
The relationship between chains and antichains is established by the following theorem. In this paper, the following theorem is called Dilworth’s Theorem [581, 55, 311, 539] on a connectivity system.
Theorem 3.7 (Dilworth’s Theorem for Connectivity Systems).
Let $(X,f)$ be a connectivity system with a symmetric submodular function $f$ and let $\mathcal{A}\subseteq 2^{X}$ be a family of subsets of $X$ such that $f(A)\leq k$ for all $A\in\mathcal{A}$. Then, the size of the largest antichain in $\mathcal{A}$ is equal to the minimum number of chains of order $k+1$ needed to cover $\mathcal{A}$.
Proof 3.8.
Let $X$ be a finite set and $f$ a symmetric submodular function. We proceed by induction on the number of elements in $X$.
For $|X|=1$, the statement trivially holds because any non-empty subset $\mathcal{A}$ of $X$ is both a chain and an antichain.
Assume the statement holds for all connectivity systems with less than $|X|$ elements. Consider a connectivity system $(X,f)$ with $|X|$ elements. Let $\mathcal{A}$ be a family of subsets of $X$ such that $f(A)\leq k$ for all $A\in\mathcal{A}$. We need to prove that the size of the largest antichain in $\mathcal{A}$ is equal to the minimum number of chains required to cover $\mathcal{A}$.
  • Let $\mathcal{A}$ be an antichain such that no element in $\mathcal{A}$ is a subset of any other. Due to the submodularity condition, the maximum value of $f(A\cap B)$ for any $A,B\in\mathcal{A}$ is less than or equal to $k$, ensuring the sets in the antichain are maximally disconnected.
  • Given any partition of $X$ into chains, by the pigeonhole principle, there is at least one element $A$ in $\mathcal{A}$ that intersects every chain. Each of these chains has an associated subset $A$ that satisfies $f(A)\leq k$, ensuring that the number of chains required to cover $\mathcal{A}$ is equal to the largest antichain.
By induction, the theorem holds for all finite connectivity systems. This proof is completed.
The relationship between ultrafilters and antichains is established by the following theorem.
Theorem 3.9.
Let $X$ be a finite set and $f$ be a symmetric submodular function. A maximal antichain on a connectivity system $(X,f)$ intersects with every ultrafilter of order $k+1$.
Proof 3.10.
Let $X$ be a finite set and $f$ a symmetric submodular function. Suppose $\{B_{1},B_{2},\dots,B_{n}\}$ is a maximal antichain of order $k+1$ on the connectivity system $(X,f)$. Let $U$ be an ultrafilter of order $k+1$ on $(X,f)$. Note that by definition, an ultrafilter $U$ satisfies: For any $A\subseteq X$ such that $f(A)\leq k$, either $A\in U$ or $X\setminus A\in U$ (this is the defining property of an ultrafilter on a connectivity system).
Now, consider each set $B_{i}$ in the antichain. Since $B_{i}$ are mutually incomparable and $f(B_{i})\leq k$ for all $i$, at least one of the sets $B_{i}$ must belong to the ultrafilter $U$, otherwise, the union $X\setminus\bigcup_{i=1}^{n}B_{i}$ would belong to $U$, which contradicts the maximality of the antichain (since $U$ would then not intersect with any $B_{i}$).
Hence, every ultrafilter of order $k+1$ intersects with the maximal antichain $\{B_{1},B_{2},\dots,B_{n}\}$. This proof is completed.
The relationship between ultrafilters and chains is established by the following theorem.
Theorem 3.11.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a chain of order $k+1$ on a connectivity system $(X,f)$, every ultrafilter of order $k+1$ contains exactly one set from the chain.
Proof 3.12.
Let $X$ be a finite set and $f$ a symmetric submodular function. Let $\{A_{1},A_{2},\dots,A_{m}\}$ be a chain of order $k+1$ on the connectivity system $(X,f)$, and let $U$ be an ultrafilter of order $k+1$ on $(X,f)$. By the properties of the chain, we have $A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{m}$ with $f(A_{i})\leq k$ for each $i$. By the properties of the ultrafilter, for each set $A_{i}$, either $A_{i}\in U$ or $X\setminus A_{i}\in U$. Since the sets in the chain are nested, the ultrafilter cannot contain two different sets $A_{i}$ and $A_{j}$ with $i\neq j$ without violating the ultrafilter condition. Therefore, there must be a unique $A_{j}$ such that $A_{j}\in U$ and for all $i<j$, $A_{i}\notin U$. This proof is completed.
Theorem 3.13.
Let $(X,f)$ be a connectivity system with $X$ as a finite set and $f$ as a symmetric submodular function. The following holds:
  1. 1.
    Chain Extension: Every chain in the set of subsets of $X$, where $f(A)\leq k$ for all $A$ in the chain, can be extended to an ultrafilter on the connectivity system $(X,f)$.
  2. 2.
    Ultrafilter Inducing a Chain: Conversely, every ultrafilter on the connectivity system $(X,f)$ induces a maximal chain in the set of subsets of $X$.
Proof 3.14.
  1. 1.
    Chain Extension to Ultrafilter: Let $X$ be a finite set and $f$ a symmetric submodular function. Given a chain $C=\{A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{m}\}$ where $f(A_{i})\leq k$ for each $i$, we want to show that this chain can be extended to an ultrafilter on the connectivity system $(X,f)$. Consider the filter $F$ generated by the chain $C$. This filter includes all supersets of elements in $C$ that satisfy $f(A)\leq k$. The symmetric submodular condition ensures that the intersection of any two sets in $F$ also belongs to $F$, maintaining the filter structure. By Theorem 3.2, this filter $F$ can be extended to a maximal filter, which is an ultrafilter $U$. This ultrafilter $U$ satisfies the condition that for any set $B\subseteq X$, either $B\in U$ or $X\setminus B\in U$, ensuring that $f(B)\leq k$ for all $B\in U$. Therefore, every chain $C$ in the set of subsets of $X$ can be extended to an ultrafilter $U$ on the connectivity system $(X,f)$.
  2. 2.
    Ultrafilter Inducing a Maximal Chain: Let $X$ be a finite set and $f$ a symmetric submodular function. Given an ultrafilter $U$ on the connectivity system $(X,f)$, we aim to show that $U$ induces a maximal chain in the set of subsets of $X$. For any set $A\in U$, consider the collection of subsets of $A$ that are also in $U$. This collection forms a chain because $U$ is a maximal filter, meaning that for any subset $A\subseteq B\subseteq X$, $B\in U$ if $A\in U$ and $f(B)\leq k$. The ultrafilter $U$ ensures that for any set $B\subseteq X$, either $B\in U$ or $X\setminus B\in U$, which means that the chain induced by $U$ cannot be extended further within $U$. This guarantees that the chain is maximal.
This proof is completed.
Theorem 3.15.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a chain of order $k+1$ on a connectivity system $(X,f)$, every ultrafilter of order $k$ does not contain a maximal set from the chain of order $k+1$.
Proof 3.16.
Consider a chain $\{A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{m}\}$ in the connectivity system $(X,f)$, where $f(A_{i})\leq k$ for all $i$. This chain is said to be of order $k+1$ if $f(A_{i})\leq k$ for all sets in the chain. An ultrafilter of order $k$ on $(X,f)$ is a maximal filter where $f(A)<k$ for all $A$ in the ultrafilter. Suppose there exists an ultrafilter $U$ of order $k$ that contains the maximal set $A_{m}$ from the chain $\{A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{m}\}$. Since $f$ is symmetric and submodular, $f(A_{m})\leq k$ holds because the chain is of order $k+1$. However, if $f(A_{m})=k$, then $A_{m}$ cannot belong to the ultrafilter $U$ of order $k$, because by definition, $U$ only contains sets where $f(A)<k$. This contradiction implies that the ultrafilter $U$ of order $k$ cannot contain the maximal set $A_{m}$. Therefore, every ultrafilter of order $k$ does not contain a maximal set from a chain of order $k+1$.
Theorem 3.17.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a chain of order $k$ on a connectivity system $(X,f)$, if no chain of order $k+1$ exists, then no ultrafilter of order $k+1$ exists.
Proof 3.18.
Let $\{A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{m}\}$ be a chain of order $k$ in the connectivity system $(X,f)$, where $f(A_{i})\leq k$ for all $i$. Suppose no chain of order $k+1$ exists, meaning there is no set $A\subseteq X$ such that $f(A)\leq k$. Assume there exists an ultrafilter $U$ of order $k+1$ on $(X,f)$. By definition, an ultrafilter of order $k+1$ contains sets $B\subseteq X$ where $f(B)\leq k$. If such a set $B$ exists in $U$, then $B$ should be part of a chain of order $k+1$. However, since we assumed that no chain of order $k+1$ exists, no set $B\subseteq X$ can satisfy $f(B)\leq k$. This proof is completed.
In the future, we will consider the relationship between a chain on a connectivity system and branch-width on a connectivity system.
3.4 Operation and Single-element-chain on a Connectivity System
In the future, we will consider Operation and Single-element-chain.
Definition 3.19.
Given a chain $\{A_{1},A_{2},\dots,A_{m}\}$ of order $k+1$ on a connectivity system $(X,f)$, where $A_{1}\subseteq A_{2}\subseteq\dots\subseteq A_{m}$ and for each $i$, $f(A_{i})\leq k$:
  1. 1.
    Extension of the Chain: The extension of the chain by adding a new subset $A_{m+1}$ is defined as follows: Let $A_{m+1}\subseteq X$ be a subset such that $A_{m}\subseteq A_{m+1}$ and $f(A_{m+1})\leq k$. The resulting extended chain is $\{A_{1},A_{2},\dots,A_{m},A_{m+1}\}$.
  2. 2.
    Deletion of a Subset: The deletion of a subset $A_{i}$ from the chain is defined as follows: Remove $A_{i}$ from the chain, resulting in the chain $\{A_{1},\dots,A_{i-1},A_{i+1},\dots,A_{m}\}$. The condition that $f(A_{j})\leq k$ for all $j$ in the modified chain must still hold.
Definition 3.20.
A single-element-chain is a special case of a chain where each subset $A_{i}$ in the chain differs from the previous subset $A_{i-1}$ by exactly one element. Formally, a single-element-chain of order $k+1$ is defined as a sequence of subsets $\{A_{1},A_{2},\dots,A_{m}\}$ such that:
  1. 1.
    $A_{1}\subseteq A_{2}\subseteq\dots\subseteq A_{m}$.
  2. 2.
    For each $i$ $(1\leq i\leq m)$, $f(A_{i})\leq k$.
  3. 3.
    $|A_{i}\setminus A_{i-1}|=1$ for each $i$ $(2\leq i\leq m)$.
Definition 3.21.
The single-element-extension of a chain $\{A_{1},A_{2},\dots,A_{m}\}$ involves adding a new subset $A_{m+1}$ such that $A_{m+1}=A_{m}\cup\{e\}$, where $e\in X\setminus A_{m}$ and $f(A_{m+1})\leq k$. The new chain is $\{A_{1},A_{2},\dots,A_{m},A_{m+1}\}$.
Definition 3.22.
The single-element-deletion of a subset $A_{i}$ from the chain involves removing one element $e$ from $A_{i}$, resulting in $A_{i}^{\prime}=A_{i}\setminus\{e\}$. The new chain, after performing the deletion, must still satisfy the condition $f(A_{j})\leq k$ for all $j$. The resulting chain is $\{A_{1},\dots,A_{i-1},A_{i}^{\prime},A_{i+1},\dots,A_{m}\}$.
4 Prefilter and Filter Subbase on a Connectivity System
We explain Prefilter and Filter Subbase on a connectivity system $(X,f)$. Prefilters and filter subbases, like ultrafilters, are crucial in set theory and topology for studying limits, convergence, and compactness. Prefilters are non-empty collections of sets closed under finite intersections, forming the basis for constructing filters. Filter subbases are non-empty collections of sets that generate filters by taking finite intersections, aiding in the study of convergence, limits, and compactness [673, 550, 261]. In this section, we add the condition of symmetric submodularity to prefilters and filter subbases and then perform verification.
4.1 Prefilter on a Connectivity System
First, we explain the definition of a prefilter and ultra-prefilter on a connectivity system $(X,f)$.
Definition 4.1.
Let $X$ be a finite set and $f$ be a symmetric submodular function. A prefilter of order $k+1$ on a connectivity system $(X,f)$ is a non-empty proper set family $P\subseteq 2^{X}$ that satisfies:
  1. (P1)
    $\emptyset\notin P$ (Proper).
  2. (P2)
    For any $A\in P$, $f(A)\leq k$ (k-efficiency).
  3. (P3)
    For any $B,C\in P$, there exists some $A\in P$ such that $A\subseteq B\cap C$ and $f(A)\leq k$ (Downward Directed under k-efficiency).
Definition 4.2.
Let $X$ be a finite set and $f$ be a symmetric submodular function. An ultra-prefilter $P\subseteq 2^{X}$ on a connectivity system $(X,f)$ is a prefilter of order $k+1$ that satisfies the following additional property:
  1. (P4)
    $\forall A\subseteq X,f(A)\leq k\Rightarrow\exists B\in P$ such that $B\subseteq A$ or $B\subseteq X\setminus A$ and $f(B)\leq k$.
Next, we demonstrate that a filter on a connectivity system can be considered a prefilter on that system. This concept extends the notions of prefilters and ultra-prefilters from set theory by incorporating the condition of symmetric submodularity.
Theorem 4.3.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Any filter of order $k+1$ on a connectivity system $(X,f)$ is also a prefilter of order $k+1$ on the same connectivity system $(X,f)$.
Proof 4.4.
To prove that a filter $Q$ on a connectivity system is also a prefilter, we need to show that $Q$ satisfies all the conditions of a prefilter.
  1. 1.
    Axiom (P1): By definition, a filter $Q$ is non-empty because it contains subsets of $X$ that satisfy the given conditions. And by Axiom (Q3), $\emptyset\notin Q$. Therefore, $Q$ is a proper set family.
  2. 2.
    Axiom (P2): By Axiom (Q0), $\forall A\in Q,f(A)\leq k$. Therefore, every element in $Q$ satisfies the k-efficiency condition.
  3. 3.
    Axiom (P3): For any $B,C\in Q$ such that $f(B)\leq k$ and $f(C)\leq k$, we need to show there exists some $A\in Q$ such that $A\subseteq B\cap C$ and $f(A)\leq k$.
    • Given $B,C\in Q$, by Axiom (Q1), $f(B\cap C)\leq k\Rightarrow B\cap C\in Q$.
    • Since $B\cap C\in Q$ and by the definition of $Q$, $f(B\cap C)\leq k$, we can choose $A=B\cap C$.
    • Therefore, $A\in Q$, $A\subseteq B\cap C$, and $f(A)\leq k$.
Hence, $Q$ satisfies all the conditions of a prefilter. This proof is completed.
Next, we show that an ultrafilter on a connectivity system is also an ultra-prefilter on a connectivity system.
Theorem 4.5.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Any ultrafilter of order $k+1$ on a connectivity system $(X,f)$ is also an ultra-prefilter of order $k+1$ on the same connectivity system $(X,f)$.
Proof 4.6.
Let $X$ be a finite set and $f$ be a symmetric submodular function. To prove that an ultrafilter $Q$ on a connectivity system is also an ultra-prefilter of order $k+1$ on a connectivity system, we need to show that $Q$ satisfies all the conditions of an ultra-prefilter of order $k+1$. Axioms (P1)-(P3) hold obviously. So we show that axiom (P4) holds.
For any $A\subseteq X$ such that $f(A)\leq k$, we need to show there exists some $B\in Q$ such that $B\subseteq A$ or $B\subseteq X\setminus A$ and $f(B)\leq k$. Given $A\subseteq X$ such that $f(A)\leq k$, by Axiom (Q4), either $A\in Q$ or $X\setminus A\in Q$. If $A\in Q$, then we can choose $B=A$, and $B\subseteq A$ with $f(B)\leq k$. If $X\setminus A\in Q$, then we can choose $B=X\setminus A$, and $B\subseteq X\setminus A$ with $f(B)\leq k$.
Hence, $Q$ satisfies all the conditions of an ultra-prefilter. This proof is completed.
4.2 Filter Subbase on a Connectivity System
Next, we introduce the definitions of filter subbase and ultrafilter subbase. These concepts extend the traditional notions of filter subbases and ultrafilter subbases from set theory by incorporating the condition of symmetric submodularity.
Definition 4.7.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a connectivity system $(X,f)$, a set family $S\subseteq 2^{X}$ is called a filter subbase of order $k+1$ if it satisfies the following conditions:
  1. (SB1)
    $S\neq\emptyset$.
  2. (SB2)
    $\emptyset\notin S$.
  3. (SB3)
    $\forall A\in S$, $f(A)\leq k$.
Definition 4.8.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a connectivity system $(X,f)$, a set family $S\subseteq 2^{X}$ is called an ultrafilter subbase of order $k+1$ if it satisfies the following conditions:
  1. (SB1)
    $S\neq\emptyset$.
  2. (SB2)
    $\emptyset\notin S$.
  3. (SB3)
    $\forall A\in S$, $f(A)\leq k$.
  4. (SB4)
    $\forall A\subseteq X$ such that $f(A)\leq k$, there exists $B\in S$ such that either $B\subseteq A$ or $B\subseteq X\setminus A$ and $f(B)\leq k$.
We now explore the relationship between a subbase on a connectivity system and a filter on a connectivity system. The following theorem establishes this relationship.
Theorem 4.9.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Given a filter subbase $S$ of order $k+1$ on a connectivity system $(X,f)$, we can generate a filter $Q$ of order $k+1$ as follows: The filter $Q$ of order $k+1$ on a connectivity system $(X,f)$ generated by $S$ is the set of all subsets of $X$ that can be formed by finite intersections of elements of $S$. Formally,
$Q=\{A\subseteq X\mid\exists A_{1},A_{2},\dots,A_{n}\in S\text{ such that }A=A_{1}\cap A_{2}\cap\dots\cap A_{n}\text{ and }f(A)\leq k\}.$
Proof 4.10.
To prove that $Q$ is a filter of order $k+1$, we need to show that $Q$ satisfies the conditions of a filter of order $k+1$.
Axiom (Q0): This axiom obviously holds. By construction, every element in $Q$ is formed by finite intersections of elements in $S$. Since $S$ is a filter subbase of order $k+1$, all elements $A_{i}$ in $S$ satisfy $f(A_{i})\leq k$. By the submodularity of $f$, for any intersection $A=A_{1}\cap A_{2}\cap\dots\cap A_{n}$, we have $f(A)=f(A_{1}\cap A_{2}\cap\dots\cap A_{n})\leq\min(f(A_{1}),f(A_{2}),\dots,f(A_{n}))\leq k$.
Axiom (Q3): Since $S$ is a filter subbase of order $k+1$ on a connectivity system $(X,f)$, it is non-empty. Suppose $S=\{A_{1},A_{2},\dots,A_{m}\}$. Consider any single element $A_{i}\in S$. Since $S$ is non-empty, $A_{i}$ exists, and $f(A_{i})\leq k$. Therefore, $A_{i}\in Q$, implying $Q$ is non-empty. And by the definition of a filter subbase, $\emptyset\notin S$. Since $Q$ is generated by finite intersections of elements of $S$, and $\emptyset$ cannot be formed by any finite intersection of non-empty sets, $\emptyset\notin Q$. Therefore, $Q$ is proper.
Axiom (Q1): By construction, $Q$ consists of all subsets of $X$ that can be formed by finite intersections of elements of $S$ and satisfy $f(A)\leq k$. Thus, by definition, $Q$ satisfies the k-efficiency condition. Let $A,B\in Q$. Then there exist $A_{1},A_{2},\dots,A_{n}\in S$ and $B_{1},B_{2},\dots,B_{m}\in S$ such that $A=A_{1}\cap A_{2}\cap\dots\cap A_{n}$ and $f(A)\leq k$, $B=B_{1}\cap B_{2}\cap\dots\cap B_{m}$ and $f(B)\leq k$. Consider the intersection $A\cap B$:
$A\cap B=(A_{1}\cap A_{2}\cap\dots\cap A_{n})\cap(B_{1}\cap B_{2}\cap\dots\cap B_{m}).$
Since $S$ is a filter subbase and contains $A_{i}$ and $B_{j}$ for all $i,j$, $A\cap B$ is formed by the finite intersection of elements of $S$. Moreover, because $f$ is submodular and symmetric, $f(A\cap B)\leq\min(f(A),f(B))\leq k$. Thus, $A\cap B\in Q$.
Axiom (Q2): Let $A\in Q$ and $A\subseteq B\subseteq X$. By definition, there exist $A_{1},A_{2},\dots,A_{n}\in S$ such that $A=A_{1}\cap A_{2}\cap\dots\cap A_{n}$ and $f(A)\leq k$. If $f(B)\leq k$, we need to show $B\in Q$. Since $A\subseteq B$ and $f(B)\leq k$, $B$ can be considered as a superset satisfying the condition for being in $Q$.
Therefore, $Q$ is a filter of order $k+1$ generated by the filter subbase $S$. This proof is completed.
Theorem 4.11.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Given a filter subbase $S$ of order $k+1$ on a connectivity system $(X,f)$, we can generate a prefilter $Q$ of order $k+1$ as follows:
$Q=\{A\subseteq X\mid\exists A_{1},A_{2},\dots,A_{n}\in S\text{ such that }A=A_{1}\cap A_{2}\cap\dots\cap A_{n}\text{ and }f(A)\leq k\}.$
Proof 4.12.
To demonstrate that $Q$ is a prefilter of order $k+1$, we need to verify that $Q$ satisfies the conditions of a prefilter as defined by axioms (P1) to (P3).
Axiom (P1): Since $S$ is a non-empty filter subbase, by definition, $S$ does not contain $\emptyset$. $Q$ is formed by finite intersections of elements in $S$, and because $\emptyset$ cannot be formed by such intersections, $\emptyset\notin Q$. Therefore, $Q$ is proper and non-empty, satisfying axiom (P1).
Axiom (P2): By construction, each element $A\in Q$ is the intersection of a finite number of sets $A_{1},A_{2},\dots,A_{n}\in S$ such that $f(A)\leq k$. Since $S$ is a filter subbase of order $k+1$, all elements $A_{i}\in S$ satisfy $f(A_{i})\leq k$. The submodularity of function $f$ ensures that $f(A_{1}\cap A_{2}\cap\dots\cap A_{n})\leq\min(f(A_{1}),f(A_{2}),\dots,f(A_{n}))\leq k$. Therefore, every element $A\in Q$ satisfies $f(A)\leq k$, ensuring k-efficiency as required by axiom (P2).
Axiom (P3): For any $B,C\in Q$, there exist sets $B_{1},B_{2},\dots,B_{m}\in S$ and $C_{1},C_{2},\dots,C_{n}\in S$ such that $B=B_{1}\cap B_{2}\cap\dots\cap B_{m}$ and $C=C_{1}\cap C_{2}\cap\dots\cap C_{n}$, with $f(B)\leq k$ and $f(C)\leq k$. Consider the intersection $A=B\cap C$. Since $A=(B_{1}\cap B_{2}\cap\dots\cap B_{m})\cap(C_{1}\cap C_{2}\cap\dots\cap C_{n})$, $A$ is formed by a finite intersection of elements in $S$. The submodularity and symmetry of function $f$ imply that $f(A)=f(B\cap C)\leq\min(f(B),f(C))\leq k$. Therefore, $A\in Q$, satisfying axiom (P3).
Given that $Q$ satisfies all the conditions (P1)-(P3) of a prefilter, we conclude that $Q$ is indeed a prefilter of order $k+1$ generated by the filter subbase $S$. This completes the proof.
Theorem 4.13.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Given an ultrafilter subbase $S$ of order $k+1$ on a connectivity system $(X,f)$, we can generate an ultrafilter $Q$ of order $k+1$ as follows: The filter $Q$ of order $k+1$ on a connectivity system $(X,f)$ generated by $S$ is the set of all subsets of $X$ that can be formed by finite intersections of elements of $S$. Formally,
$Q=\{A\subseteq X\mid\exists A_{1},A_{2},\dots,A_{n}\in S\text{ such that }A=A_{1}\cap A_{2}\cap\dots\cap A_{n}\text{ and }f(A)\leq k\}.$
Proof 4.14.
The proof can be established using a method nearly identical to that used in Theorem 4.9. This completes the proof.
Theorem 4.15.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Given an ultrafilter subbase $S$ of order $k+1$ on a connectivity system $(X,f)$, we can generate an ultra-prefilter $Q$ of order $k+1$ as follows:
$Q=\{A\subseteq X\mid\exists A_{1},A_{2},\dots,A_{n}\in S\text{ such that }A=A_{1}\cap A_{2}\cap\dots\cap A_{n}\text{ and }f(A)\leq k\}.$
Proof 4.16.
The proof can be established using a method nearly identical to that used in Theorem 4.11. This completes the proof.
Based on the above, the relationships shown in the following diagram become clear.
Figure 2: Relation about generating a filter
5 Ultrafilter on Infinite Connectivity System and Countable Connectivity System
5.1 Ultrafilter on Infinite Connectivity System
Until now, we have considered ultrafilters on finite sets, but now we turn our attention to ultrafilters on infinite sets. First, let’s define what an infinite connectivity system is.
Definition 5.1 ([127, 237, 164]).
Consider an infinite set $X$ equipped with a symmetric submodular function $f$ mapping from the set of subsets of $X$ to ${}\mathbb{N}\cup\{\infty\}$. This function satisfies the following conditions for all subsets $A$ and $B$ of $X$:
  1. 1.
    Symmetry: $f(A)=f(X\setminus A)$.
  2. 2.
    Submodularity: $f(A)+f(B)\geq f(A\cup B)+f(A\cap B)$.
In the context of infinite sets, we also require $f$ to be $k$-limit-closed. Specifically, if $k\in\mathbb{N}$ and $(A_{i})_{i\in I}$ is a chain of subsets of $X$ each with connectivity at most $k$, then the union $\bigcup_{i\in I}A_{i}$ also has connectivity at most $k$. It is important to note that the ground set of a (possibly infinite) matroid, along with its connectivity function, constitutes a connectivity system, and the connectivity function is defined to be $k$-limit-closed. A pair $(X,f)$ is called an infinite connectivity system.
Given the known lemma, it is reasonable to consider a deep relationship between finite sets and infinite sets.
Lemma 5.2 ([127]).
Let $X$ be an infinite set and $f$ be a symmetric submodular function. Let $k\in\mathbb{N}$, and let $A\subseteq X$ be a set such that all finite subsets have connectivity at most $k$. Then also $f(A)\leq k$.
Following the model of ultrafilters on finite sets, we define ultrafilters on infinite sets as follows [304].
Definition 5.3 (cf. [304]).
Let $X$ be an infinite set and $f$ be a symmetric submodular function. In an infinite connectivity system $(X,f)$, the set family $F\subseteq 2^{X}$ is called a filter of order $k+1$ if the following axioms hold true:
  1. (Q0)
    $\forall A\in F,f(A)\leq k$.
  2. (Q1)
    $A\in F,B\in F,f(A\cap B)\leq k\Rightarrow A\cap B\in F$.
  3. (Q2)
    $A\in F,A\subseteq B\subseteq X,f(B)\leq k\Rightarrow B\in F$.
  4. (Q3)
    $\emptyset\notin F$.
An ultrafilter of order $k+1$ on an infinite connectivity system $(X,f)$ is a filter of order $k+1$ on an infinite connectivity system $(X,f)$ which satisfies the following additional axiom:
  1. (Q4)
    $\forall A\subseteq X,f(A)\leq k\Rightarrow\text{either }A\in F\text{ or }X\setminus A\in F$.
A filter of order $k+1$ on an infinite connectivity system $(X,f)$ is principal if the filter satisfies the following axiom [304]:
  1. (QP5)
    $A\in F$ for all $A\subseteq X$ with $|A|=1$ and $f(A)\leq k$.
A filter of order $k+1$ on an infinite connectivity system $(X,f)$ is non-principal if the filter satisfies the following axiom [304]:
  1. (Q5)
    $A\notin F$ for all $A\subseteq X$ with $|A|=1$.
Non-principal refers to a filter or ideal that does not contain any singletons (i.e., sets with exactly one element).
If a filter on an infinite connectivity system is weak [288, 475, 35], the following axiom holds instead of axiom (Q1):
  1. (QW1’)
    $A\in F,B\in F,f(A\cap B)\leq k\Rightarrow A\cap B\neq\emptyset$.
Note that weak filters aim at interpreting defaults via a generalized ‘most’ quantifier in first-order logic [475, 35, 622, 276]. A weak filter on an infinite connectivity system is a co-Weak Ideal on an infinite connectivity system.
If a filter on an infinite connectivity system is quasi [293, 143], the following axiom holds instead of axiom (Q1):
  1. (QQ1’)
    $A\subseteq X,B\subseteq X,A\notin F,B\notin F\Rightarrow A\cup B\notin F$.
Note that a quasi-Ultrafilter is used to perform an axiomatic analysis of incomplete social judgments [143].
The following also holds for infinite connectivity systems, just as it does for finite connectivity systems.
Theorem 5.4.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Note that Limit-Closed means that for any $k\in\mathbb{N}$ and any chain $(A_{i})_{i\in I}$ of subsets of $X$, each with connectivity at most $k$, the union $\bigcup_{i\in I}A_{i}$ also has connectivity at most $k$. An ultrafilter of order $k+1$ on an infinite connectivity system $(X,f)$ is also an ultrafilter of order $k$ on an infinite connectivity system $(X,f)$.
Proof 5.5.
By using the symmetric submodular function, an ultrafilter of order $k+1$ on an infinite connectivity system $(X,f)$ obviously holds the axioms of an ultrafilter of order $k$ on an infinite connectivity system $(X,f)$. This proof is completed.
Here are some properties that hold for filters and ultrafilters of order $k+1$ on an infinite connectivity system.
Lemma 5.6.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. If a filter of order $k+1$ on an infinite connectivity system $(X,f)$ is maximal, it satisfies the following additional axiom:
  • (Q4)
    $\forall A\subseteq X$, if $f(A)\leq k$, then either $A\in F$ or $X\setminus A\in F$.
Proof 5.7.
Assume that there exists a maximal filter $F$ of order $k+1$ on $(X,f)$ that does not satisfy axiom (Q4). That is, there exists a subset $A\subseteq X$ such that $f(A)\leq k$, but neither $A\in F$ nor $X\setminus A\in F$. Since $F$ is maximal, the addition of $A$ or $X\setminus A$ to $F$ would result in a filter that still respects the conditions of $f(A)\leq k$. Therefore, $F\cup\{A\}$ or $F\cup\{X\setminus A\}$ would also be a valid filter, contradicting the maximality of $F$. Thus, the assumption that $F$ does not satisfy (Q4) leads to a contradiction. Therefore, any maximal filter $F$ of order $k+1$ must satisfy axiom (Q4). This proof is completed.
Definition 5.8 (cf. [256]).
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. We will say that a family $F$ of order $k+1$ on an infinite connectivity system is nice if it is nonempty, closed under intersections (i.e., if $A,B\in F$ and $f(A\cap B)\leq k$, then $A\cap B\in F$), and $\emptyset\notin F$.
Lemma 5.9.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. A family of subsets of $X$ is maximal nice of order $k+1$ on an infinite connectivity system if and only if it is an ultrafilter of order $k+1$ on an infinite connectivity system.
Proof 5.10.
Let’s first recall that a family $F$ of order $k+1$ on an infinite connectivity system is called nice if it is nonempty, closed under intersections (i.e., if $A,B\in F$ and $f(A\cap B)\leq k$, then $A\cap B\in F$), and $\emptyset\notin F$.
Assume that $F$ is a maximal nice family of order $k+1$. We want to show that $F$ is an ultrafilter of order $k+1$, i.e., it satisfies axiom (Q4): $\forall A\subseteq X$, if $f(A)\leq k$, then either $A\in F$ or $X\setminus A\in F$. Suppose $F$ is maximal nice, but there exists a subset $A\subseteq X$ such that $f(A)\leq k$, and neither $A\in F$ nor $X\setminus A\in F$. Since $F$ is maximal, we can extend $F$ by adding either $A$ or $X\setminus A$ to form a new family $F^{\prime}$. But this contradicts the maximality of $F$, as it can still be extended. Therefore, $F$ must satisfy (Q4), and hence it is an ultrafilter of order $k+1$.
Now assume $F$ is an ultrafilter of order $k+1$. We need to show that $F$ is maximal nice. Suppose $F$ is not maximal nice. Then there exists a family $F^{\prime}$ such that $F\subsetneq F^{\prime}$ and $F^{\prime}$ is nice. But since $F$ is an ultrafilter, it cannot be extended by adding more sets without violating the conditions for an ultrafilter. This contradicts the assumption that $F$ is not maximal nice. Hence, $F$ must be maximal nice. This proof is completed.
In this paper, let $\text{Fr}=\{A\subseteq X\mid f(A)\leq k,X\setminus A\text{ is finite}\}$ be the Frechet filter of order $k+1$ on an infinite connectivity system $(X,f)$. We consider the following lemma.
Lemma 5.11.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. If there exists a Frechet filter of order $k+1$ on an infinite connectivity system $(X,f)$, then there exists a nonprincipal ultrafilter of order $k+1$ on an infinite connectivity system $(X,f)$.
Proof 5.12.
Assume that there does not exist a nonprincipal ultrafilter of order $k+1$ on $(X,f)$. This implies that every ultrafilter on $(X,f)$ must be principal. If Fr can be extended to an ultrafilter, this ultrafilter would be principal. A principal ultrafilter contains a set $\{x\}$ for some $x\in X$ with $f(\{x\})\leq k$. However, by definition, in a Frechet filter, $X\setminus\{x\}$ is finite, and thus $X$ cannot have only principal ultrafilters if it contains elements for which $f(\{x\})\leq k$. This contradicts the assumption that every ultrafilter is principal. Thus, there must exist a nonprincipal ultrafilter of order $k+1$ on $(X,f)$. This proof is completed.
We consider the following extended Ramsey’s Theorem on an infinite connectivity system. Note that Ramsey’s Theorem states that for any given positive integers $k$ and $r$, there is a minimum number $R(k,r)$ such that any graph with $R(k,r)$ vertices, colored with $r$ colors, contains a monochromatic clique of size $k$. Note that a monochromatic clique is a subset of vertices where all edges connecting them are the same color. This theorem can also be considered in terms of partitions, ensuring any partition of edges or vertices leads to a monochromatic subset.
Theorem 5.13.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. Then for any infinite partition $P=\{A_{1},A_{2},\dots\}$ of $X$, there exists an infinite subset $Y\subseteq X$ such that all subsets of $Y$ are contained within a single block of the partition and have connectivity at most $k$.
Proof 5.14.
Let $X$ be an infinite set with the symmetric submodular function $f$ as defined. Let $P=\{A_{1},A_{2},\dots\}$ be a partition of $X$. We want to find an infinite subset $Y\subseteq X$ such that all subsets of $Y$ are contained within a single block of $P$ and have connectivity at most $k$.
Consider an increasing chain of subsets $(A_{i})_{i\in I}$ of $X$, where each $A_{i}$ has connectivity at most $k$. By the limit-closed property of function $f$, the union $\bigcup_{i\in I}A_{i}$ also has connectivity at most $k$.
Using the classical form of Ramsey’s Theorem, we know that for any infinite set $X$ and any partition of $X$ into finitely many pieces, there exists an infinite subset $Y\subseteq X$ such that all pairs in $Y$ lie within the same piece of the partition. To apply Ramsey’s Theorem in the context of the connectivity system $(X,f)$, we need to ensure that the infinite subset $Y$ not only lies within a single block of the partition but also maintains the connectivity condition $f(Y)\leq k$.
Start with an infinite subset $Y_{0}\subseteq X$. Partition $Y_{0}$ into blocks $A_{i}$ according to the given partition $P$. Select a block $A_{j}$ that contains an infinite subset $Y_{1}\subseteq A_{j}$. By the symmetry and submodularity of function $f$, ensure $f(Y_{1})\leq k$.
By iterating the process, we can construct a nested sequence of infinite subsets $Y_{0}\supseteq Y_{1}\supseteq Y_{2}\supseteq\dots$, each lying within a single block of the partition and maintaining the connectivity condition $f(Y_{i})\leq k$.
Let $Y=\bigcap_{i=0}^{\infty}Y_{i}$. Since each $Y_{i}$ is infinite and lies within a single block of the partition, $Y$ is also infinite and lies within a single block of the partition. Moreover, by the limit-closed property of the infinite connectivity system, we obtain $f(Y)\leq k$.
Thus, we have shown that for any infinite partition $P$ of $X$, there exists an infinite subset $Y\subseteq X$ such that all subsets of $Y$ lie within a single block of the partition and have connectivity at most $k$. This completes the proof of the extended Ramsey’s Theorem on an infinite connectivity system.
Theorem 5.15.
Let $(X,f)$ be an infinite connectivity system with $f$ a symmetric submodular function satisfying $k$-Limit-Closed. A set family $U\subseteq 2^{X}$ is a weak ultrafilter of order $k+1$ if and only if it is a maximal weak filter of order $k+1$ on $(X,f)$.
Proof 5.16.
Assume that $U$ satisfies the definition of a weak ultrafilter but is not maximal. This means there exists a weak filter $G$ such that $U\subset G$ and $G$ contains an additional set $A\subset X$ where $A\in G$ but $A\notin U$. Since $U$ is a weak ultrafilter, by its definition, for any subset $A\subseteq X$ with $f(A)\leq k$, it must hold that either $A\in U$ or $X\setminus A\in U$.
Now, because $U\subset G$, every element of $U$ is also an element of $G$. Specifically, since $X\setminus A\in U$ and $U\subset G$, it follows that $X\setminus A$ must also be in $G$. However, $G$ already contains $A$ by assumption, and now it also contains $X\setminus A$. For $G$ to be a consistent weak filter, it cannot contain both $A$ and $X\setminus A$ because that would violate the condition of non-empty intersection. This leads to a contradiction because if $G$ contains both $A$ and $X\setminus A$, it would mean that $G$ is no longer a valid weak filter due to inconsistency, which contradicts the assumption that $G$ is a weak filter.
Hence, $U$ must be maximal because extending $U$ to any larger set $G$ results in an inconsistency, proving that $U$ cannot be properly contained in any larger weak filter.
We need to prove that $U$ is a weak ultrafilter of order $k+1$. Specifically, for any subset $A\subset X$ with $f(A)\leq k$ and $A\notin U$, we must show $X\setminus A\in U$.
Let $G$ be the upward closure of $U\cup\{A\}$, meaning $G$ includes all sets in $U$ as well as any sets that can be formed by taking unions of elements from $U$ and $A$.
By the assumption that $U$ is maximal, $G$ cannot form a weak filter. If $G$ were a weak filter, it would contradict the maximality of $U$.
The fact that $G$ cannot be a weak filter means that there must exist some set $Z\subseteq X$ in $G$ such that $Z\supset A$, and for some $Y\in U$, we have $Z\cap Y=\emptyset$. This condition arises because if $Z$ and $Y$ were both in $G$, it would imply that $Z$ and $Y$ must intersect, which they do not in this scenario.
The existence of such a $Z$ means that the complement $X\setminus A$ must be in $U$ to maintain the weak ultrafilter property (since $U$ must contain the complement if it does not contain the original set).
This confirms that $U$ must contain $X\setminus A$, proving that $U$ behaves like a weak ultrafilter of order $k+1$. This proof is completed.
In the future, we plan to further investigate the theorems related to ultrafilters on infinite connectivity systems. For example, we will consider the following definitions of Selective and Weakly Selective Ultrafilters on a Connectivity System (cf. [79, 77, 270, 136]).
Definition 5.17.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. Assume that for all $x\in X$, $f(\{x\})\leq k$. An ultrafilter $F$ on $(X,f)$ is called selective if for every function $g:X\to X$, there exists a subset $A\in F$ such that $g\mid A$ is either:
  • Constant: $g(x)=c$ for all $x\in A$, or
  • One-to-One: $g(x_{1})\neq g(x_{2})$ for all distinct $x_{1},x_{2}\in A$.
Definition 5.18.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. An ultrafilter $F$ on $(X,f)$ is called quasi-selective if for every function $g:X\to X$ with $g(x)\leq x$ for all $x\in X$ where $f(\{x\})\leq k$, there exists a subset $A\in F$ such that $g\mid A$ is a non-decreasing function.
5.2 Ultrafilter on Countable Connectivity System
We consider ultrafilters on countable (infinite) connectivity systems. Let $X$ be a countable set, which means it can be put into a one-to-one correspondence with the set of natural numbers $\mathbb{N}$.
In mathematics, some researchers focus on the study of games used to model situations with multiple adversaries who have conflicting interests (e.g., [145, 165, 687, 470, 236, 100, 276, 574]). We consider the investigation of the Ultrafilter Game [120, 173, 51] on a countable connectivity system. This game is an analogy of the well-known ultrafilter game in the context of countable set theory.
Game 5.19 (Ultrafilter Game on a Countable Connectivity System).
Let $X$ be a finite set and $D\subseteq X^{\mathbb{N}}$. We define a two-player game $G(D)$ as follows:
  1. 1.
    Players and Moves: Player I and Player II take turns playing elements from $X$.
    • Player I starts by playing $a_{0}\in X$.
    • Player II responds with $a_{1}\in X$.
    • Player I then plays $a_{2}\in X$, and the game continues in this alternating fashion.
  2. 2.
    Rounds: The game is played for countably many rounds, resulting in a sequence $a=(a_{0},a_{1},a_{2},\ldots)\in X^{\mathbb{N}}$.
  3. 3.
    Winning Condition:
    • Player I wins if the sequence $a$ belongs to $D$ and satisfies the condition $f(\{a\})\leq k$.
    • Player II wins if either the sequence $a$ does not belong to $D$ or $f(\{a\})>k$.
  4. 4.
    Strategies:
    • Strategy for Player I: A strategy for Player I is a rule determining Player I’s moves based on the history of the game. A winning strategy guarantees that Player I can always produce a sequence $a$ such that $a\in D$ and $f(\{a\})\leq k$, regardless of Player II’s moves.
    • Strategy for Player II: Similarly, a strategy for Player II is a rule determining Player II’s responses based on the history of the game. A winning strategy for Player II ensures that the resulting sequence $a$ either does not belong to $D$ or violates the condition $f(\{a\})\leq k$.
  5. 5.
    Determination: We say that $D\subseteq X^{\mathbb{N}}$ is determined if one of the two players has a winning strategy for the game $G(D)$. This means there exists a strategy such that either Player I or Player II can always secure a win, ensuring the game is resolved definitively in favor of one player.
We now consider the winning strategy and determination.
Lemma 5.20.
Let $X$ be a finite set and $f$ be a symmetric submodular function. For the ultrafilter game on a countable connectivity system $(X,f)$, Player I has a winning strategy by following a linear decomposition strategy.
Proof 5.21.
Consider a caterpillar tree $C$ as a path $(l_{1},b_{2},b_{3},\ldots,b_{n-1},l_{n})$, where the subgraph of $C$ induced by $\{b_{i-1},b_{i},b_{i+1}\}$ forms a connectivity system.
If $X$ has only one element, $e_{1}$, Player I starts by choosing $a_{0}=e_{1}$. Since there are no other elements, $f(\{e_{1}\})\leq k$. Player I wins because the sequence $(e_{1})$ belongs to the set $D$ and satisfies $f((e_{1}))\leq k$.
Assume that for a caterpillar tree with $n-1$ elements, Player I has a winning strategy by choosing elements according to the linear decomposition. For a caterpillar tree with $n$ elements, let the linear decomposition partition $X$ into $\{e_{1},e_{2},\ldots,e_{n}\}$.
At each turn $i$, Player I chooses $a_{i}=e_{i+1}$. This ensures that the partial sequence $(a_{0},a_{1},\ldots,a_{i})$ satisfies $f((a_{0},a_{1},\ldots,a_{i}))\leq k$.
Since the linear decomposition ensures the sequence adheres to the symmetric submodular condition, and Player I can choose elements to keep $f((a_{0},a_{1},\ldots,a_{i}))\leq k$, the sequence $a$ will belong to the set $D$.
By induction, Player I has a winning strategy for any $n$ elements by following the linear decomposition. Therefore, Player I can systematically choose elements according to the linear decomposition strategy, ensuring the sequence meets the conditions required for victory. This proof is completed.
Lemma 5.22.
Let $X$ be a finite set and $f$ be a symmetric submodular function. For the ultrafilter game on a countable connectivity system $(X,f)$, a non-principal ultrafilter does not result in a determined game.
Proof 5.23.
If $X$ has only one element, it cannot form a non-principal ultrafilter since the ultrafilter would include singletons, which contradicts the non-principal property.
Assume that for any set of size $n-1$, the non-principal ultrafilter does not lead to a determined game. For a set $X$ with $n$ elements, the non-principal ultrafilter excludes all singletons, making it impossible for Player I to use single elements to form a winning strategy.
Player II also cannot formulate a winning strategy because the non-principal nature prevents clear counteractions based on singletons or small subsets.
The lack of singletons in the ultrafilter means that neither player can systematically form a winning sequence based on single elements or small subsets.
Both players are left without a structured method to guarantee victory, leading to a non-determined game.
By induction, the non-principal ultrafilter does not lead to a determined game for any finite $n$. The exclusion of singletons and the lack of structured strategies ensure that neither player can develop a definitive winning strategy. This proof is completed.
5.3 Elimination Games with an Invisible Robber on a Countable Connectivity System
We consider another game called "elimination games" (cf. [255]). We propose the definition of the elimination games with an invisible robber on a countable connectivity system.
Game 5.24 (Elimination Game with an Invisible Robber).
[277] Let $(X,f)$ be a countable connectivity system equipped with a symmetric submodular function $f$, and let $k$ be an integer. The elimination game on this system is a pursuit-evasion game between a cop and a robber, with the following rules:
  1. 1.
    Players and Pieces:
    • The cop controls one or more tokens (representing cops).
    • The robber controls a single token (representing the robber).
    • Both the cops and the robber move on the elements of $X$.
  2. 2.
    Initial Setup:
    • Initially, no tokens occupy any elements of $X$.
    • The game begins with the cop placing some tokens on the elements of $X$ according to the function $f$.
    • The robber then chooses an initial position $r\in X$ without being seen by the cop, ensuring that $f(\{r\})\leq k$.
  3. 3.
    Moves:
    • In each round, the cop announces a new position $C_{i}$ by moving the tokens to a set of elements $C_{i+1}\subseteq X$, ensuring that $f(C_{i+1})\leq k$.
    • The robber, while invisible to the cop, can move from their current position $r$ along or against a directed path not containing any element of $C_{i+1}$ to a new position $r^{\prime}$, ensuring $f(\{r^{\prime}\})\leq k$.
  4. 4.
    Movement Rules:
    • The robber can move to any element $r^{\prime}$ within the connected component of $X\setminus C_{i+1}$ containing $r$, under the condition that $f(\{r^{\prime}\})\leq k$.
    • The cop cannot see the robber’s moves, so their strategy must account for all possible positions the robber might occupy, taking into account the submodularity condition: $f(A)+f(B)\geq f(A\cup B)+f(A\cap B)$ for any subsets $A,B\subseteq X$.
  5. 5.
    Winning Conditions:
    • The cop wins if they place a token on the element currently occupied by the robber, and the robber is unable to escape to a safe place within the connectivity constraints $f$.
    • The robber wins if they can always elude capture, continuously finding a path $r^{\prime}$ that satisfies the connectivity condition $f(\{r^{\prime}\})\leq k$ within the connected components of $X\setminus C_{i+1}$.
We can represent a cop’s strategy on a connectivity system $(X,f)$ by a finite or infinite sequence $S$ of cop positions: $S:=\{C_{0},C_{1},C_{2},\ldots\}$ where $C_{i}$ denotes the set of elements occupied by the cops at the $i$-th move.
Conjecture 5.25.
The robber’s winning strategy can be characterized as a (non-principal) single ultrafilter if the robber can always find a path to escape from the cops’ positions within the connected components of $X\setminus C_{i+1}$.
In the future, we plan to further investigate the theorems related to ultrafilters on countable connectivity systems.
5.4 Ultrafilter on Maximum-Connectivity System
We introduce a property similar to submodularity, which supports the development of corresponding theories for connectivity functions.
Definition 5.26 (cf. [262]).
A set function $\kappa$ on a finite underlying set $X$ is defined as maximum-submodular if, for all subsets $A,B\subseteq X$,
$\max(\kappa(A),\kappa(B))\geq\max(\kappa(A\cap B),\kappa(A\cup B)).$
According to [262], a set function that is normalized, symmetric, and maximum-submodular is referred to as a maximum-submodular connectivity function. A maximum-connectivity system consists of a pair $(U,\kappa)$, where $U$ is a finite set and $\kappa$ is a maximum-submodular connectivity function.
We plan to define an ultrafilter on a maximum-connectivity system and explore its relationship with tangles on maximum-connectivity systems in future research.
6 Future Tasks
This section outlines future perspectives for this study and related research. Although still in the conceptual stage or under review for submission, I aim to outline these ideas to help advance future research on graph width parameters and ultrafilters.
6.1 New Width/Length/Depth Parameters and Obstructions
In this paper, we discussed an ultrafilter that serves as an obstruction to branch-width and linear branch-width. In the future, we will consider the following width parameters. As noted in Appendix A, various width parameters are well-known. Our goal is to examine these width parameters from multiple perspectives, including their corresponding decompositions, related length [674, 201, 183] and depth parameters [225, 182, 562] and Appendix D, linear concepts, directed graph concepts, and their extensions from graphs to connectivity systems [330] and abstract separation systems [195, 185]. For a detailed discussion on length, refer to Appendix B. Additionally, we plan to define and explore the characteristics of directed branch depth, similar to directed tree depth [479], and Directed Rank-depth, similar to Rank-depth [425, 569, 452, 426]. Furthermore, we aim to investigate the relationships with ultrafilters and related definitions.
6.1.1 Consideration about a Branch Distance Decomposition
First, let’s consider the concept of a branch distance decomposition. Intuitively, this is a branch decomposition that integrates the concept of distance. This idea originates from exploring how extending tree-distance-width [715] and path-distance-width [715] to branch decompositions might work.
Definition 6.1 (Branch Distance Decomposition).
A branch distance decomposition is a branch decomposition with an additional distance function and a root. A branch decomposition of a graph $G=(V,E)$ is a pair $B=(T,\mu)$, where $T$ is a tree and $\mu$ is a bijection from the edges of $G$ to the leaves of $T$. In a branch distance decomposition, we add a distance function $d:E(T)\rightarrow\mathbb{N}$, which assigns a distance to each edge in $T$, and a root $r\in V(T)$.
For an edge $e\in E(T)$, let $T_{1}$ and $T_{2}$ be the two subtrees obtained by deleting $e$ from $T$. Let $G_{1}$ and $G_{2}$ be the subgraphs of $G$ induced by the edges mapped by $\mu$ to the leaves of $T_{1}$ and $T_{2}$, respectively. The vertices $V(G_{1})\cap V(G_{2})$ are denoted by $\text{mid}_{B}(e)$.
The branch distance width of $B$ is $\max_{e\in E(T)}(\lvert\text{mid}_{B}(e)\rvert+d(e))$. The branch distance width of $G$, denoted by $\text{BDW}(G)$, is the minimum width over all branch distance decompositions of $G$.
A linear distance decomposition of a graph $G=(V,E)$ is a triple $(P,\mu,d)$, where:
  • $P=(v_{1},v_{2},\dots,v_{n})$ is a path graph, represented as a sequence of vertices.
  • $\mu$ is a bijection from the edges of $G$ to the vertices of $P$.
  • $d:V(P)\rightarrow\mathbb{N}$ is a function assigning a distance to each vertex in $P$.
For a vertex $v_{i}\in V(P)$, let $P_{i}$ be the subpath consisting of vertices $v_{i-1},v_{i},v_{i+1}$. Let $G_{i}$ be the subgraph of $G$ induced by the edges mapped by $\mu$ to the vertices in $P_{i}$. The set of vertices $V(G_{i})$ is denoted by $\text{mid}_{L}(v_{i})$.
The width of a linear distance decomposition is defined as:
$\text{ldw}(P,\mu,d)=\max_{v_{i}\in V(P)}(\lvert\text{mid}_{L}(v_{i})\rvert+d(v_{i})).$
The linear distance width of $G$, denoted by $\text{LDW}(G)$, is the minimum width over all linear distance decompositions of $G$:
$\text{LDW}(G)=\min_{(P,\mu,d)}\text{ldw}(P,\mu,d).$
Conjecture 6.2.
For any connected graph $G$ with $\text{bw}(G)\geq 2$, $\text{TDW}(G)=\text{BDW}(G)$. Also, for any connected graph $G$ with $\text{lw}(G)\geq 2$, $\text{PDW}(G)=\text{LDW}(G)$.
Next, we consider an algorithm for constructing a branch distance decomposition. The following algorithm has significant room for improvement, and we plan to continue refining it.
Algorithm 6.3 (Constructing Branch Distance Decomposition).
  • Input: A graph $G=(V,E)$.
  • Output: A branch distance decomposition $(T,\{X_{i}\})$.
  1. 1.
    Initialization:
    • Let $T$ be a tree with a single node $r$ (the root).
    • Initialize $X_{r}=V(G)$.
  2. 2.
    Iterative Decomposition:
    • While $\lvert X_{r}\rvert>1$:
      1. (a)
        Select a vertex $v\in X_{r}$.
      2. (b)
        Find the shortest path $P$ in $G$ that includes $v$ and other vertices in $X_{r}$.
      3. (c)
        Partition $X_{r}$ into two sets $X_{1}$ and $X_{2}$ based on the vertices on either side of $P$ such that:
        • $X_{1}\cup X_{2}=X_{r}$,
        • $X_{1}\cap X_{2}$ contains the vertices on $P$.
      4. (d)
        Add two new nodes $i$ and $j$ to $T$ as children of $r$, and set:
        • $X_{i}=X_{1}$,
        • $X_{j}=X_{2}$.
      5. (e)
        Update $T$ and the corresponding sets $X_{i}$ and $X_{j}$.
  3. 3.
    Ensuring Distance Constraints:
    • For each vertex $v\in V(G)$, ensure that if $v\in X_{i}$, then $\text{dist}_{G}(X_{r},v)=\text{dist}_{T}(r,i)$.
    • Adjust $X_{i}$ and $X_{j}$ to meet this distance constraint by reassigning vertices if necessary.
  4. 4.
    Repeat: Continue decomposing each subset $X_{i}$ recursively, applying the above steps until each $X_{i}$ contains a single vertex or a base case is met.
  5. 5.
    Output the Decomposition: Once the tree $T$ and the subsets $\{X_{i}\}$ satisfy all the constraints, output the branch distance decomposition $(T,\{X_{i}\})$.
Lemma 6.4.
Above algorithm is correct.
Proof 6.5.
  1. 1.
    The decomposition process ensures that $T$ remains a tree, as new nodes are added as children to existing nodes, maintaining the tree structure.
  2. 2.
    Each vertex $v\in V(G)$ is included in at least one subset $X_{i}$ throughout the decomposition, ensuring $\bigcup_{i\in V(T)}X_{i}=V(G)$.
  3. 3.
    For each edge $\{u,v\}\in E(G)$, the decomposition ensures there are nodes $i,j\in V(T)$ such that $u\in X_{i}$ and $v\in X_{j}$, and either $i=j$ or $\{i,j\}\in E(T)$.
  4. 4.
    The algorithm maintains the distance constraint by ensuring $\text{dist}_{G}(X_{r},v)=\text{dist}_{T}(r,i)$ for each vertex $v\in V(G)$, adjusting subsets as necessary to maintain this property.
  5. 5.
    The iterative and recursive nature of the algorithm ensures that the decomposition process will terminate, as each step reduces the size of the subsets until the base case is reached.
Therefore, the lemma is proved.
Lemma 6.6.
The total time complexity of Algorithm 6.4 is $O((V+E)\log V)$.
Proof 6.7.
  • Initialization: The initialization step takes $O(1)$ time.
  • Iterative Decomposition: Each iteration involves selecting a vertex $v\in X_{r}$ and finding a shortest path $P$. Finding a shortest path can be done in $O(V+E)$ time using breadth-first search (BFS) or depth-first search (DFS). Partitioning $X_{r}$ into $X_{1}$ and $X_{2}$ based on the shortest path takes $O(V)$ time. Adding new nodes to the tree $T$ and updating the corresponding sets $X_{i}$ and $X_{j}$ takes $O(1)$ time.
  • Ensuring Distance Constraints: Adjusting subsets $X_{i}$ and $X_{j}$ to meet distance constraints might involve checking distances for all vertices, which can be done in $O(V)$ time.
  • Recursive Decomposition: The recursive decomposition process continues until each subset $X_{i}$ contains a single vertex, resulting in a tree of height $O(\log V)$.
Total Time Complexity: The total time complexity for the entire algorithm can be expressed as a sum of the time complexities of the individual steps. Each step involves $O(V+E)$ time for finding the shortest path, and there are $O(\log V)$ levels of recursion. Therefore, the lemma is proved.
And we will consider a game called Cops and Robbers with Speed $d$. The definition of the game is as follows:
Game 6.8.
Cops and Robbers with Speed $d$
In the "cops and robbers with speed $d$" game, the graph $G=(V,E)$ is endowed with a distance function that affects movement capabilities. The players consist of a single robber and multiple cops.
  • Players: One robber and $k$ cops.
  • Positions: Defined by a tuple $(r,C)$, where $r\in V$ is the robber’s position and $C\subseteq V$ with $|C|\leq k$ represents the positions of the cops.
  • Initial Position: The robber chooses an initial position $r_{0}\in V$, and the cops collectively choose their initial positions as a set $C$ where $|C|\leq k$.
  • Movement: Each round consists of the cops moving to a new set of positions $C^{\prime}$ followed by the robber moving to a new position $r^{\prime}$. The robber can move to any vertex $r^{\prime}$ that is reachable from $r$ within distance $d$ in $G\setminus(C\cap C^{\prime})$, reflecting the robber’s speed and the obstruction caused by the cops.
  • Winning Conditions: The cops win if they can move into the robber’s position, i.e., if a situation $(r,C)$ with $r\in C$ is achieved; otherwise, the robber wins if he remains uncaught indefinitely.
Conjecture 6.9.
Branch-distance-decomposition is a winning strategy for Cops and Robbers with Speed $d$.
6.1.2 Consideration about a Directed Tree Distance Decomposition
Next, we will consider a directed tree distance decomposition[280]. This idea also originates from exploring how extending tree-distance-width [715] and path-distance-width [715] to branch decompositions might work.
Definition 6.10.
[280] Let $G=(V,E)$ be a directed graph. A directed tree-distance decomposition of $G$ is a triple $(R,\{X_{i}\mid i\in V(R)\},r)$, where:
  • $R$ is an arborescence (a directed tree with a designated root $r$ such that there is a unique directed path from $r$ to any vertex in $V(R)$).
  • $\{X_{i}\mid i\in V(R)\}$ is a family of subsets of $V(G)$ satisfying the following properties:
    • (Partition Property): $\bigcup_{i\in V(R)}X_{i}=V(G)$ and $X_{i}\cap X_{j}=\emptyset$ for $i\neq j$.
    • (Distance Property): For each vertex $v\in V(G)$, if $v\in X_{i}$, then the directed distance from $X_{r}$ (the set associated with the root $r$) to $v$ in $G$ is equal to the directed distance from $r$ to $i$ in $R$. Formally, $d_{G}(X_{r},v)=d_{R}(r,i)$.
    • (Edge Coverage Property): For each edge $(u,v)\in E(G)$, there exist vertices $i,j\in V(R)$ such that $u\in X_{i}$, $v\in X_{j}$, and either $i=j$ or $(i,j)\in E(R)$.
The width of the directed tree-distance decomposition is defined as $\max_{i\in V(R)}|X_{i}|$. The directed tree-distance width of a directed graph $G$, denoted by $\text{DTDW}(G)$, is the minimum width over all possible directed tree-distance decompositions of $G$.
A directed path-distance decomposition of a directed graph $G=(V,E)$ is a special case of the directed tree-distance decomposition where the tree $T$ is a directed path. Formally, a directed path-distance decomposition is a sequence of subsets $(X_{1},X_{2},\dots,X_{t})$ of the vertex set $V$, satisfying the following conditions:
  1. 1.
    $\bigcup_{i=1}^{t}X_{i}=V$, meaning that the union of all subsets $X_{i}$ covers the entire vertex set $V$,
  2. 2.
    For each vertex $v\in V$, if $v\in X_{i}$, then the directed distance $d_{G}(X_{1},v)$ from the root set $X_{1}$ in the directed graph $G$ to $v$ corresponds to the position of $X_{i}$ in the sequence, i.e., $d_{G}(X_{1},v)=i-1$,
  3. 3.
    For each directed edge $(u,v)\in E$, there exist indices $i$ and $j$ such that $u\in X_{i}$, $v\in X_{j}$, and either $i=j$ or $|i-j|=1$, indicating that the vertices $u$ and $v$ must either be in the same subset or in consecutive subsets.
The width of a directed path-distance decomposition $(X_{1},X_{2},\dots,X_{t})$ is defined as the maximum size of any set $X_{i}$ in the decomposition, i.e.,
$\text{width}\left((X_{1},X_{2},\dots,X_{t})\right)=\max_{1\leq i\leq t}|X_{i}|.$
The directed path-distance width (DPDW) of the directed graph $G$ is defined as the minimum width over all possible directed path-distance decompositions of $G$.
In the future, we will investigate the relationships with ultrafilters and these distance widths.
6.1.3 New Linear Width Parameter
As discussed in Appendix A, various graph width parameters have been defined and extensively studied. Our goal is to introduce new linear versions of some of these width parameters and explore their relationships with other graph parameters. Specifically, we plan to define Linear-Amalgam-Decomposition, the linear version of Amalgam-Decomposition [517], Linear-Modular-Decomposition, the linear version of Modular-Decomposition [2], Linear-Tree-Cut-Decomposition, the linear version of Tree-Cut Decomposition, and Direct Linear-Branch-Decomposition, the linear version of Direct Branch-Decomposition, and analyze their interconnections. It’s important to note that for any general width parameter, it holds that (general width parameter) $\leq$ (linear width parameter); for instance, Amalgam-width $\leq$ Linear-Amalgam-width.
These restrictions to underlying path structures are often advantageous in establishing results for general parameters. Moreover, these linear parameters provide valuable structural insights, particularly in the study of special graph classes. Additionally, the relationship between path-structured graph width parameters and game theory, specifically in the context of the "invisible cops and robbers" game, is frequently discussed [451, 181, 711]. This is a game where the cops can catch the robber even without knowing the exact location of the robber. We also aim to explore potential connections between these widths and ultrafilters, further enriching the theoretical framework.
6.1.4 Matroid Length Parameters
We will explore new matroid length parameters [675, 283]. The most well-known length parameter already established is tree-length. Tree-length measures the maximum distance between any two vertices within a single bag of a tree-decomposition, minimized over all possible decompositions [169, 205, 206]. Although still in the conceptual stage, the definitions are outlined below[283].
Definition 6.11.
[283] A tree decomposition of a matroid $M$ on the ground set $E=E(M)$ is a pair $(T,\tau)$, where $T$ is a tree and $\tau:E\to V(T)$ is a mapping from the elements of $E$ to the nodes of $T$. For each node $i\in V(T)$, let $X_{i}=\tau^{-1}(i)$ be the set of elements of $E$ mapped to $i$.
For each $X_{i}$, consider the rank function $r_{M}$ of the matroid $M$. The diameter of $X_{i}$ is defined as:
$\text{diam}_{M}(X_{i})=\max_{x,y\in X_{i}}\left(r_{M}(\{x,y\})-\min(r_{M}(\{x\}),r_{M}(\{y\}))\right).$
The tree-length of the decomposition $(T,\tau)$, denoted by $\text{tl}(T,\tau)$, is defined as the maximum diameter of $X_{i}$ over all nodes $i$ in $T$:
$\text{tl}(T,\tau)=\max_{i\in V(T)}\text{diam}_{M}(X_{i}).$
The tree-length of the matroid $M$, denoted by $\text{tl}(M)$, is the minimum tree-length over all possible tree decompositions of $M$:
$\text{tl}(M)=\min_{(T,\tau)}\text{tl}(T,\tau).$
Definition 6.12.
[283] A branch decomposition of a matroid $M=(E,I)$ is a pair $(T,\mu)$, where $T$ is a subcubic tree without degree-2 nodes, and $\mu$ is a bijection from the elements of $E$ to the leaves of $T$. For each edge $e\in E(T)$, let $T_{1}$ and $T_{2}$ be the two subtrees obtained by deleting $e$ from $T$, and let $E_{1}$ and $E_{2}$ be the sets of elements of $E$ corresponding to the leaves in $T_{1}$ and $T_{2}$, respectively. The set $E_{1}\cap E_{2}$ is denoted by $\text{mid}_{B}(e)$.
For each $\text{mid}_{B}(e)$, consider the rank function $r_{M}$. The diameter of $\text{mid}_{B}(e)$ is defined as:
$\text{diam}_{M}(\text{mid}_{B}(e))=\max_{x,y\in\text{mid}_{B}(e)}\left(r_{M}(\{x,y\})-\min(r_{M}(\{x\}),r_{M}(\{y\}))\right).$
The branch-length of the decomposition $(T,\mu)$, denoted by $\text{bl}(T,\mu)$, is defined as the maximum diameter of $\text{mid}_{B}(e)$ over all edges $e$ in $T$:
$\text{bl}(T,\mu)=\max_{e\in E(T)}\text{diam}_{M}(\text{mid}_{B}(e)).$
The branch-length of the matroid $M$, denoted by $\text{bl}(M)$, is the minimum branch-length over all possible branch decompositions of $M$:
$\text{bl}(M)=\min_{(T,\mu)}\text{bl}(T,\mu).$
6.1.5 Branch-breadth and Linear-Breadth
We will explore the concepts of branch-breadth and linear-breadth for graphs. A parameter closely related to tree-length is tree-breadth, which measures how closely the vertices in each bag of a tree-decomposition can be clustered around a central vertex within a specific radius, minimized over all possible decompositions [496, 495]. A path-based version of this parameter, known as path-breadth, is also well-established. We are currently investigating the corresponding versions for branch-decompositions[496, 495]. Although still in the conceptual stage, the definitions are outlined below [278].
Definition 6.13.
[278] A branch decomposition of a graph $G=(V,E)$ is a pair $(T,\sigma)$, where $T$ is a tree with vertices of degree at most 3, and $\sigma$ is a bijection from the set of leaves of $T$ to $E$. For an edge $e$ in $T$, consider the two subtrees $T_{1}$ and $T_{2}$ obtained by removing $e$ from $T$. Let $V_{1}$ and $V_{2}$ be the sets of vertices in $G$ that are incident to edges corresponding to leaves in $T_{1}$ and $T_{2}$, respectively.
The breadth of an edge $e$ in $T$ is defined as the minimum $\rho$ such that for each $v\in V_{1}\cap V_{2}$, the set of vertices $\{v\}\cup(V_{1}\setminus V_{2})\cup(V_{2}\setminus V_{1})$ is contained within the closed neighborhood $N_{\rho}(v)$ of some vertex $v$ in $G$. The branch-breadth of a graph $G$, denoted as $\text{bb}(G)$, is the minimum breadth over all possible branch decompositions of $G$.
Definition 6.14.
[278] A linear decomposition of a graph $G=(V,E)$ is a linear ordering of the edges $(e_{1},e_{2},\dots,e_{m})$. For each index $i$ between 1 and $m-1$, let $V_{1}$ be the set of vertices incident to edges in $\{e_{1},\dots,e_{i}\}$, and $V_{2}$ be the set of vertices incident to edges in $\{e_{i+1},\dots,e_{m}\}$.
The breadth of the linear decomposition is the minimum $\rho$ such that for each vertex $v$ in $V_{1}\cap V_{2}$, the set $V_{1}\cup V_{2}$ is contained within the closed neighborhood $N_{\rho}(v)$ of some vertex $v$ in $G$. The linear-breadth of a graph $G$, denoted as $\text{lb}(G)$, is the minimum breadth over all possible linear decompositions of $G$.
6.1.6 Examination for Directed proper-path-width
We will explore the concept of proper-path-width for graphs. Proper-path-width [655, 654, 657] measures how "tightly" a graph can be decomposed into overlapping vertex subsets, ensuring that no two subsets are identical while maintaining specific connectivity constraints. We are interested in extending this definition to directed graphs. Although still in the conceptual stage, the definition is outlined below [280].
Definition 6.15.
[280] Let $\mathcal{P}=(X_{1},X_{2},\dots,X_{r})$ be a sequence of subsets of vertices of a directed graph $G=(V,E)$. The width of $\mathcal{P}$ is defined as $\max_{1\leq i\leq r}|X_{i}|-1$. The sequence $\mathcal{P}$ is called a directed-proper-path-decomposition of $G$ if the following conditions are satisfied:
  1. 1.
    For any distinct $i$ and $j$, $X_{i}\neq X_{j}$.
  2. 2.
    $\bigcup_{i=1}^{r}X_{i}=V(G)$.
  3. 3.
    For any edge $(u,v)\in E(G)$, there exists an $i$ such that $u,v\in X_{i}$.
  4. 4.
    For all $1\leq l<m<n\leq r$, $X_{l}\cap X_{n}\subseteq X_{m}$.
  5. 5.
    For all $1\leq l<m<n\leq r$, $|X_{l}\cap X_{n}|\leq|X_{m}|-2$.
The directed-proper-path-width of $G$, denoted by $\text{dppw}(G)$, is the minimum width over all directed-proper-path-decompositions of $G$. If a sequence $\mathcal{P}$ satisfies only conditions (i)-(iv), it is called a directed-path-decomposition of $G$, and the directed-path-width of $G$, denoted by $\text{dpw}(G)$, is the minimum width over all directed-path-decompositions of $G$. A (directed-proper-)path-decomposition with width $k$ is called a $k$-(directed-proper-)path-decomposition.
6.1.7 Fuzzy Tree-width
We will delve into the concept of tree-width as it applies to fuzzy graphs. Fuzzy set theory, introduced in 1965, was developed to capture the nuances of partial truth, bridging the gap between absolute truth and falsehood [720]. Building on classical graph theory, fuzzy graphs introduce uncertainty by assigning a membership degree to each edge [616]. This rapidly advancing field offers intriguing theoretical challenges and significant practical applications [530, 547, 68, 582, 587, 204, 546]. Our goal is to extend the concept of tree-width to fuzzy graphs. The formal definition of a fuzzy graph is presented below.
Definition 6.16.
[616] A fuzzy graph $G=(\sigma,\mu)$ with $V$ as the underlying set is defined as follows:
  • $\sigma:V\rightarrow[0,1]$ is a fuzzy subset of vertices, where $\sigma(x)$ represents the membership degree of vertex $x\in V$.
  • $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$, such that $\mu(x,y)\leq\sigma(x)\wedge\sigma(y)$ for all $x,y\in V$, where $\wedge$ denotes the minimum operation.
The underlying crisp graph of $G$ is denoted by $G^{*}=(\sigma^{*},\mu^{*})$, where:
  • $\sigma^{*}=\sup p(\sigma)=\{x\in V:\sigma(x)>0\}$
  • $\mu^{*}=\sup p(\mu)=\{(x,y)\in V\times V:\mu(x,y)>0\}$
A fuzzy subgraph $H=(\sigma^{\prime},\mu^{\prime})$ of $G$ is defined as follows:
  • There exists $X\subseteq V$ such that $\sigma^{\prime}:X\rightarrow[0,1]$ is a fuzzy subset.
  • $\mu^{\prime}:X\times X\rightarrow[0,1]$ is a fuzzy relation on $\sigma^{\prime}$, satisfying $\mu^{\prime}(x,y)\leq\sigma^{\prime}(x)\wedge\sigma^{\prime}(y)$ for all $x,y\in X$.
For more detailed information about fuzzy set and fuzzy graph, please refer to lecture notes or surveys[727, 260, 727, 547].
Now, We consider about the definitions of fuzzy-tree-decomposition and fuzzy-path-decomposition[285]. This definition generalizes the classical concept of tree-width by incorporating the fuzzy nature of the graph. Although still in the conceptual stage, the definitions are outlined below.
Definition 6.17.
[285] Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing vertex membership, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$. A fuzzy tree-decomposition of $G$ is a pair $(T,\{B_{t}\}_{t\in T})$, where:
  • $T=(I,F)$ is a tree with nodes $I$ and edges $F$.
  • $\{B_{t}\}_{t\in T}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $T$ such that:
    1. 1.
      For each vertex $v\in V$, the set $\{t\in I:v\in B_{t}\}$ is connected in the tree $T$.
    2. 2.
      For each edge $(u,v)\in V\times V$ with membership degree $\mu(u,v)\leq\sigma(u)\wedge\sigma(v)$, there exists some node $t\in I$ such that both $u$ and $v$ are in $B_{t}$, and the membership degree of $u$ and $v$ in $B_{t}$ is at least $\mu(u,v)$.
The width of a fuzzy tree-decomposition $(T,\{B_{t}\}_{t\in T})$ is defined as
$\max_{t\in I}\left(\sup_{v\in B_{t}}\mu(v,B_{t})-1\right),$
where $\mu(v,B_{t})$ represents the maximum membership degree of vertex $v$ in the fuzzy set $B_{t}$. The Fuzzy-Tree-Width of the fuzzy graph $G$ is the minimum width among all possible fuzzy tree-decompositions of $G$.
Definition 6.18.
[285] Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing vertex membership, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$. A fuzzy path-decomposition of $G$ is a pair $(P,\{B_{p}\}_{p\in P})$, where:
  • $P$ is a path with nodes $I$ and edges $F$.
  • $\{B_{p}\}_{p\in P}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $P$ such that:
    1. 1.
      For each vertex $v\in V$, the set $\{p\in I:v\in B_{p}\}$ is connected in the path $P$.
    2. 2.
      For each edge $(u,v)\in V\times V$ with membership degree $\mu(u,v)\leq\sigma(u)\wedge\sigma(v)$, there exists some node $p\in I$ such that both $u$ and $v$ are in $B_{p}$, and the membership degree of $u$ and $v$ in $B_{p}$ is at least $\mu(u,v)$.
The width of a fuzzy path-decomposition $(P,\{B_{p}\}_{p\in P})$ is defined as
$\max_{p\in I}\left(\sup_{v\in B_{p}}\mu(v,B_{p})-1\right),$
where $\mu(v,B_{p})$ represents the maximum membership degree of vertex $v$ in the fuzzy set $B_{p}$. The Fuzzy-Path-Width of the fuzzy graph $G$ is the minimum width among all possible fuzzy path-decompositions of $G$.
And in the future, we will consider about following definitions for fuzzy-graph.
Definition 6.19 (Fuzzy Bandwidth).
Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing the membership degree of vertices, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $V$ representing the membership degree of edges. A layout of the graph $G$ is a bijection $f:V\rightarrow\{1,2,\dots,|V|\}$. The fuzzy bandwidth of the layout $f$ is defined as:
$\max_{\{u,v\}\in E}\left(\mu(u,v)\cdot|f(u)-f(v)|\right).$
The fuzzy bandwidth of the fuzzy graph $G$, denoted $\text{fbw}(G)$, is the minimum fuzzy bandwidth over all possible layouts $f$ of $G$.
Definition 6.20 (Fuzzy Path Distance-Width).
Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing the membership degree of vertices, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$. A fuzzy path-distance decomposition of $G$ is a sequence of fuzzy subsets $(B_{1},B_{2},\dots,B_{r})$ of $V$, called bags, such that:
  1. 1.
    $B_{1}\cup B_{2}\cup\dots\cup B_{r}=V$,
  2. 2.
    For each edge $(u,v)\in E$ with membership degree $\mu(u,v)>0$, there exists an index $i\leq j$ such that $u\in B_{i}$ and $v\in B_{j}$, and the membership degree of $u$ and $v$ in $B_{i}$ and $B_{j}$ is at least $\mu(u,v)$.
  3. 3.
    For all indices $1\leq i<j<\ell\leq r$, $B_{i}\cap B_{\ell}\subseteq B_{j}$.
The width of a fuzzy path-distance decomposition $(B_{1},B_{2},\dots,B_{r})$ is defined as:
$\max_{1\leq i\leq r}\sup_{v\in B_{i}}\mu(v,B_{i})-1.$
The fuzzy path-distance width of $G$, denoted by $\text{fpdw}(G)$, is the minimum width over all possible fuzzy path-distance decompositions of $G$.
Definition 6.21 (Fuzzy Tree-length).
Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing vertex membership, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$. A fuzzy tree-decomposition of $G$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a tree with nodes $I$ and edges $F$.
  • $\{B_{t}\}_{t\in I}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $T$.
The fuzzy length of a tree-decomposition $T(G)$ is defined as:
$\lambda(T)=\max_{t\in I}\max_{u,v\in B_{t}}d_{G^{*}}(u,v),$
where $d_{G^{*}}(u,v)$ denotes the distance between vertices $u$ and $v$ in the underlying crisp graph $G^{*}$. The fuzzy tree-length of $G$, denoted by $\text{ftl}(G)$, is the minimum length over all possible fuzzy tree-decompositions of $G$.
Definition 6.22 (Fuzzy Tree-Breadth).
Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing vertex membership, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$. A fuzzy tree-decomposition of $G$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a tree with nodes $I$ and edges $F$.
  • $\{B_{t}\}_{t\in I}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $T$.
The fuzzy breadth of a tree-decomposition $T(G)$ is the minimum integer $r$ such that for every $t\in I$, there exists a vertex $v_{t}\in V(G^{*})$ with $B_{t}\subseteq D_{r}(v_{t},G^{*})$, where $D_{r}(v_{t},G^{*})$ is the disk of radius $r$ centered at $v_{t}$ in the underlying crisp graph $G^{*}$. The fuzzy tree-breadth of $G$, denoted by $\text{ftb}(G)$, is the minimum breadth over all possible fuzzy tree-decompositions of $G$.
We will also consider about directed fuzzy graph[526, 519, 359].
Definition 6.23.
[519] A directed fuzzy graph is a quadruple $\xi=(V,\sigma,\mu,E)$, where:
  • $V$ is a non-empty set of vertices.
  • $\sigma:V\to[0,1]$ is the membership function for vertices, where $\sigma(a)$ represents the membership value of vertex $a$.
  • $E$ is the set of all directed edges between vertices in $V$.
  • $\mu:E\to[0,1]$ is the membership function for edges, where $\mu(a,b)$ represents the membership value of the directed edge $(a,b)$.
These functions must satisfy the following condition:
$\mu(a,b)\leq\sigma(a)\wedge\sigma(b),$
where $\wedge$ denotes the minimum operation, ensuring that the membership value of an edge cannot exceed the membership values of the vertices it connects.
Given a directed fuzzy graph $\xi=(V,\sigma,\mu,E)$, the measure of influence between two vertices $a,b\in V$ is denoted by $\phi(a,b)$ and is defined as:
$\phi(a,b)\leq|\sigma(a)-\sigma(b)|,$
where $\phi(a,b)$ quantifies the influence based on the difference in their membership values.
Although still in the conceptual stage, the definitions are outlined below.
Definition 6.24 (Fuzzy Directed Tree-Decomposition).
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed tree-decomposition of $\xi$ is a triple $(T,X,W)$, where:
  • $T=(V_{T},E_{T})$ is a directed tree,
  • $X=\{X_{e}\mid e\in E_{T}\}$ is a set of fuzzy subsets of $V$ (called bags) associated with the edges of $T$,
  • $W=\{W_{r}\mid r\in V_{T}\}$ is a set of fuzzy subsets of $V$ associated with the vertices of $T$, satisfying the following conditions:
    1. 1.
      $W=\{W_{r}\mid r\in V_{T}\}$ is a partition of $V$,
    2. 2.
      For every edge $e=(u,v)\in E_{T}$, the fuzzy set $W_{v}$ is $X_{e}$-normal, meaning $\mu(u,v)\leq\min(\sigma(u),\sigma(v))$.
The width of a fuzzy directed tree-decomposition $(T,X,W)$ is defined as:
$\max_{r\in V_{T}}\left(\sup_{v\in W_{r}\cup\bigcup_{e\sim r}X_{e}}\mu(v,W_{r})-1\right),$
where $e\sim r$ means that $r$ is one of the endpoints of the arc $e$. The fuzzy directed tree-width of $\xi$, denoted by $\text{fdtw}(\xi)$, is the minimum width over all possible fuzzy directed tree-decompositions of $\xi$.
Definition 6.25 (Fuzzy Directed Path-Decomposition).
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed path-decomposition of $\xi$ is a sequence $(X_{1},\dots,X_{r})$ of fuzzy subsets of $V$, called bags, that satisfies the following conditions:
  1. 1.
    $X_{1}\cup\dots\cup X_{r}=V$,
  2. 2.
    For each directed edge $(u,v)\in E$ with membership degree $\mu(u,v)>0$, there exists a pair $i\leq j$ such that $u\in X_{i}$ and $v\in X_{j}$, and the membership degree of $u$ and $v$ in $X_{i}$ and $X_{j}$ is at least $\mu(u,v)$.
  3. 3.
    For all indices $1\leq i<j<\ell\leq r$, it holds that $X_{i}\cap X_{\ell}\subseteq X_{j}$.
The width of a fuzzy directed path-decomposition $(X_{1},\dots,X_{r})$ is defined as:
$\max_{1\leq i\leq r}\sup_{v\in X_{i}}\mu(v,X_{i})-1.$
[…]
Definition 6.26 (Fuzzy Directed Cut-Width).
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. The fuzzy directed cut-width of $\xi$ is defined as:
[…]
where $\Phi(\xi)$ denotes the set of all linear layouts of $\xi$, $L(i,\varphi,\xi)$ represents the set of vertices mapped to positions $\leq i$ by the linear layout $\varphi$, and $R(i,\varphi,\xi)$ represents the set of vertices mapped to positions $>i$.
Definition 6.27 (Fuzzy Directed Bandwidth).
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A layout of $\xi$ is a bijection $f:V\rightarrow\{1,2,\dots,|V|\}$. The fuzzy directed bandwidth of the layout $f$ is defined as:
$\max_{\{u,v\}\in E}\left(\mu(u,v)\cdot|f(u)-f(v)|\right).$
The fuzzy directed bandwidth of the fuzzy directed graph $\xi$, denoted by $\text{fdbw}(\xi)$, is the minimum fuzzy bandwidth over all possible layouts $f$ of $\xi$.
Definition 6.28 (Fuzzy Directed Tree-Length).
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed tree-decomposition of $\xi$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a directed tree with nodes $I$ and edges $F$,
  • $\{B_{t}\}_{t\in I}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $T$.
The fuzzy directed tree-length of a tree-decomposition $T(\xi)$ is defined as:
$\lambda(T)=\max_{t\in I}\max_{u,v\in B_{t}}d_{G^{*}}(u,v),$
where $d_{G^{*}}(u,v)$ denotes the distance between vertices $u$ and $v$ in the underlying crisp graph $G^{*}$. The fuzzy directed tree-length of $\xi$, denoted by $\text{fdtl}(\xi)$, is the minimum length over all possible fuzzy directed tree-decompositions of $\xi$.
Definition 6.29 (Fuzzy Directed Tree-Breadth).
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed tree-decomposition of $\xi$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a directed tree with nodes $I$ and edges $F$,
  • $\{B_{t}\}_{t\in I}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $T$.
The fuzzy directed tree-breadth of a tree-decomposition $T(\xi)$ is defined as the minimum integer $r$ such that for every $t\in I$, there exists a vertex $v_{t}\in V(G^{*})$ with $B_{t}\subseteq D_{r}(v_{t},G^{*})$, where $D_{r}(v_{t},G^{*})$ is the disk of radius $r$ centered at $v_{t}$ in the underlying crisp graph $G^{*}$. The fuzzy directed tree-breadth of $\xi$, denoted by $\text{fdtb}(\xi)$, is the minimum breadth over all possible fuzzy directed tree-decompositions of $\xi$.
6.1.8 Graph parameters for Fuzzy HyperGraph
In fuzzy graph theory, the concept of a fuzzy hypergraph is well-known. It extends the notion of a hypergraph by incorporating the concepts of fuzzy sets and fuzzy graphs. Numerous studies have explored its applications in real-world scenarios[547, 61, 18, 512].
Definition 6.30.
A fuzzy hypergraph on a non-empty set $X$ is defined as a pair $H=(\mu,\rho)$, where:
  • $\mu=\{\mu_{1},\mu_{2},\dots,\mu_{r}\}$ is a collection of fuzzy subsets on $X$, where each $\mu_{i}:X\rightarrow[0,1]$ represents the membership degree of elements in $X$ to the fuzzy subset $\mu_{i}$. It is required that the union of the supports of all fuzzy subsets covers the entire set $X$, i.e., $\bigcup_{i=1}^{r}\text{supp}(\mu_{i})=X$.
  • $\rho$ is a fuzzy relation on the fuzzy subsets $\mu_{i}$, which assigns a membership degree $\rho(E_{i})$ to each hyperedge $E_{i}=\{x_{1},x_{2},\dots,x_{s}\}\subseteq X$. The membership degree $\rho(E_{i})$ is constrained by the condition:
    $\rho(E_{i})\leq\min\{\mu_{i}(x_{1}),\mu_{i}(x_{2}),\dots,\mu_{i}(x_{s})\},$
    for all elements $x_{1},x_{2},\dots,x_{s}\in X$ and for each fuzzy subset $\mu_{i}\in\mu$.
We are interested in exploring whether new insights can be gained by extending the graph width parameter known as hypertree-width to fuzzy hypergraphs. Given the vast range of applications for fuzzy hypergraphs, such an extension could potentially accelerate the development of real-world applications.
Definition 6.31 (Fuzzy Hypertree-width).
Let $H=(\mu,\rho)$ be a fuzzy hypergraph defined on a non-empty set $X$, where:
  • $\mu=\{\mu_{1},\mu_{2},\dots,\mu_{r}\}$ is a collection of fuzzy subsets on $X$, with each $\mu_{i}:X\rightarrow[0,1]$ representing the membership degree of elements in $X$ to the fuzzy subset $\mu_{i}$. The union of the supports of all fuzzy subsets covers the entire set $X$, i.e., $\bigcup_{i=1}^{r}\text{supp}(\mu_{i})=X$.
  • $\rho$ is a fuzzy relation on the fuzzy subsets $\mu_{i}$, assigning a membership degree $\rho(E_{i})$ to each hyperedge $E_{i}=\{x_{1},x_{2},\dots,x_{s}\}\subseteq X$. The membership degree $\rho(E_{i})$ satisfies:
    $\rho(E_{i})\leq\min\{\mu_{i}(x_{1}),\mu_{i}(x_{2}),\dots,\mu_{i}(x_{s})\},$
    for all elements $x_{1},x_{2},\dots,x_{s}\in X$ and each fuzzy subset $\mu_{i}\in\mu$.
A fuzzy hypertree decomposition of a fuzzy hypergraph $H=(\mu,\rho)$ is a triple $D=(T,\chi,\lambda)$, where:
  • $T=(V_{T},E_{T})$ is a tree.
  • $\chi:V_{T}\rightarrow 2^{X}$ is a function that assigns to each tree node $t\in V_{T}$ a fuzzy subset $\chi(t)\subseteq X$, known as a fuzzy bag.
  • $\lambda:V_{T}\rightarrow 2^{\mu}$ is a function that assigns to each tree node $t\in V_{T}$ a collection of fuzzy hyperedges, known as a fuzzy edge cover of the fuzzy bag $\chi(t)$.
The decomposition must satisfy the following conditions:
  1. 1.
    Fuzzy Edge Cover Condition: For every fuzzy hyperedge $E_{i}$ in $H$, there exists a tree node $t\in V_{T}$ such that $\rho(E_{i})\leq\min_{x\in E_{i}}\chi(t)(x)$.
  2. 2.
    Fuzzy Connectedness Condition: For every vertex $x\in X$, the set of tree nodes $t\in V_{T}$ with $\chi(t)(x)>0$ induces a connected subtree of $T$.
  3. 3.
    Fuzzy Special Condition: For any two nodes $t,t^{\prime}\in V_{T}$ where $t^{\prime}$ is a descendant of $t$ in $T$, and for every fuzzy hyperedge $E_{i}\in\lambda(t)$, it holds that $(E_{i}\setminus\chi(t))\cap\chi(t^{\prime})=\emptyset$.
The fuzzy hypertree-width of a fuzzy hypergraph $H=(\mu,\rho)$ is defined as the minimum width over all possible fuzzy hypertree decompositions of $H$, where the width of a fuzzy hypertree decomposition $D=(T,\chi,\lambda)$ is given by the maximum value of the sum of the membership degrees $\sum_{x\in\chi(t)}\chi(t)(x)$ over all tree nodes $t\in V_{T}$.
6.1.9 Directed Hypertree Width
A directed hypergraph, which is an extension of the undirected hypergraph, has been studied for its practical applications[669, 37, 603, 681, 38, 511]. This structure has proven useful in various real-world scenarios. The definition is provided below.
Definition 6.32.
[669] Given a directed hypergraph $H=(V,E)$, where $V$ represents the set of vertices and $E$ denotes the set of hyper-arcs. Each hyper-arc $e\in E$ is defined as $e=(e_{\text{Tail}},e_{\text{Head}})$, where:
  • $e_{\text{Tail}}$ is the tail of the hyper-arc $e$, and $e_{\text{Head}}$ is the head of the hyper-arc $e$.
  • The vertices associated with the hyper-arc $e$ are denoted by $\mathcal{E}=e_{\text{Tail}}\cup e_{\text{Head}}$.
  • It is required that $e_{\text{Tail}}\neq\emptyset$, $e_{\text{Head}}\neq\emptyset$, and $e_{\text{Tail}}\cap e_{\text{Head}}=\emptyset$.
The directed hypergraph $H=(V,E)$ can be represented using two incidence matrices: $H_{\text{Tail}}$ and $H_{\text{Head}}$. These incidence matrices are defined as follows:
$h_{\text{Tail}}(v,e)=\begin{cases}1&\text{if }v\in e_{\text{Tail}},\\ 0&\text{otherwise}.\end{cases}$
$h_{\text{Head}}(v,e)=\begin{cases}1&\text{if }v\in e_{\text{Head}},\\ 0&\text{otherwise}.\end{cases}$
Here, $h_{\text{Tail}}$ and $h_{\text{Head}}$ are the incidence matrices corresponding to the tail and head of the hyper-arcs, respectively, with $v$ representing a vertex and $e$ representing a hyper-arc in $H$.
We are interested in exploring whether new insights can be gained by extending the graph width parameter known as hypertree-width to directed hypergraphs. Although still in the conceptual stage, the definitions for directed hypergraph are outlined below.
Definition 6.33.
A directed hypertree decomposition of a directed hypergraph $H=(V,E)$ is a triple $D=(T,\chi,\lambda)$, where:
  • $T$ is a tree,
  • $\chi$ is a function that assigns each tree node $t\in V(T)$ a set $\chi(t)\subseteq V$, called a bag,
  • $\lambda$ is a function that assigns each tree node $t\in V(T)$ a set $\lambda(t)\subseteq E$, called an arc cover of the bag $\chi(t)$.
The decomposition must satisfy the following conditions:
  1. 1.
    Arc Cover Condition: For each hyper-arc $e\in E(H)$, there exists a tree node $t\in V(T)$ such that the vertices in $\mathcal{E}=e_{\text{Tail}}\cup e_{\text{Head}}$ are contained in the bag $\chi(t)$.
  2. 2.
    Connectedness Condition: For every vertex $v\in V(H)$, the set of tree nodes $\{t\in V(T):v\in\chi(t)\}$ induces a connected subtree of $T$.
  3. 3.
    Special Condition: For any two nodes $t,t^{\prime}\in V(T)$ such that $t^{\prime}$ is a descendant of $t$ in $T$, and for each hyper-arc $e\in\lambda(t)$, it holds that $(e\setminus\chi(t))\cap\chi(t^{\prime})=\emptyset$.
[…]
We will also explore the concept of fuzzy directed hypertree-width. Given the active research on the real-world applications of fuzzy directed hypergraphs (ex.[17, 513, 551, 553, 552]), studying fuzzy directed hypertree-width is both meaningful and timely.
We will also explore tree-width and other width parameters on symmetric matroids[429]. Symmetric matroids are closely related to hypergraphs.
6.1.10 Weighted Width
A weighted graph assigns a numerical value, known as a weight, to each edge, and sometimes to each vertex. These graphs are widely used in applications such as network design[156], routing algorithms[399], optimization problems[620], and in modeling real-world systems like transportation networks[374, 540, 373] and social networks[599, 155, 721, 561, 724, 521, 684], where the strengths or values of relationships vary. Notably, concepts such as Weighted Tree-width[432], Weighted Path-width[537], and Weighted Cut-width[90] have already been established.
Although still in the conceptual stage, the definitions for weighted graph are outlined below.
Definition 6.34.
The weighted bandwidth of a weighted graph $G=(V,E,w)$ is defined as follows:
Consider an ordering $\varphi$ of the vertices of $G$, and let $d_{G}(u,v)$ denote the distance between vertices $u$ and $v$ in $G$ under the ordering $\varphi$.
[…]
where $w(u)$ and $w(v)$ are the weights of vertices $u$ and $v$, respectively.
Definition 6.35.
Given a weighted graph $G=(V,E,w)$ where $w:V\to\mathbb{N}$ assigns a weight to each vertex, the weighted clique-width is defined as follows:
The weighted clique-width of $G$ is the minimum number of labels needed to construct the graph using the following operations:
  1. 1.
    Creation of a vertex: Create a new vertex with a specific label.
  2. 2.
    Disjoint union: Take the disjoint union of two labeled graphs.
  3. 3.
    Edge insertion: Connect all vertices with label $i$ to all vertices with label $j$ (for $i\neq j$).
  4. 4.
    Relabeling: Change the label of all vertices with label $i$ to label $j$.
[…]
Definition 6.36.
The weighted linear-clique-width of a weighted graph $G=(V,E,w)$ is defined similarly to weighted clique-width, with the following modification:
  • The operations to construct the graph must follow a linear order, meaning each step adds to a single sequence of vertices.
[…]
Definition 6.37.
Given a weighted graph $G=(V,E,w)$, where $w:V\to\mathbb{N}$ assigns a weight to each vertex, the weighted path-distance-width is defined based on a path distance decomposition of the graph.
A path distance decomposition of a graph $G$ is a tuple ($(\{X_{i}\mid i\in I\},T=(I,F),r)$, where:
  • $T=(I,F)$ is a path (a tree where each node has at most two neighbors).
  • $\{X_{i}\mid i\in I\}$ is a family of subsets of $V(G)$, called bags, such that:
    1. 1.
      $\bigcup_{i\in I}X_{i}=V(G)$ and for all $i\neq j$, $X_{i}\cap X_{j}=\emptyset$, meaning the bags partition the vertex set $V(G)$.
    2. 2.
      For each vertex $v\in V(G)$, if $v\in X_{i}$, then $d_{G}(X_{r},v)=d_{T}(r,i)$, where $X_{r}$ is the root set and $d_{G}$ and $d_{T}$ are the distances in $G$ and $T$, respectively.
    3. 3.
      For each edge $\{v,w\}\in E(G)$, there exist $i,j\in I$ such that $v\in X_{i}$, $w\in X_{j}$, and either $i=j$ or $\{i,j\}\in F$.
The width of a path distance decomposition ($(\{X_{i}\mid i\in I\},T,r)$ is defined as:
$\text{width}(T)=\max_{i\in I}\sum_{v\in X_{i}}w(v),$
where $w(v)$ is the weight of vertex $v$.
[…]
Definition 6.38.
Given a weighted graph $G=(V,E,w)$, where $w:E\to\mathbb{N}$ assigns a weight to each edge, the weighted carving-width is defined as follows:
A call routing tree (or a carving) of a graph $G$ is a tree $T$ with internal vertices of degree 3 whose leaves correspond to the vertices of $G$. The congestion of an edge $e$ in $T$ is defined as the sum of the weights of the edges in $G$ that have endpoints in different connected components of $T-e$.
[…]
We also plan to explore width parameters such as Weighted Directed Tree-width, Weighted Directed cut-width, Weighted Directed Tree-breadth, Weighted Directed Hypertree-width and Weighted Directed Path-width in the context of weighted digraphs. Given that weighted digraphs are practical and versatile graph classes with a wide range of applications[595, 67, 548, 14, 347], defining these width parameters is both meaningful and valuable for various fields.
6.1.11 Directed Linear-width
We will consider about Directed Linear-width[282]. Although still in the conceptual stage, the definitions for weighted graph are outlined below. This is the linear version of the Branch-width of a directed graph[130] and is also an extension of the Linear-width of an undirected graph[659, 664].
Definition 6.39.
[282] Let $D=(V,E)$ be a digraph. The function $f_{D}:2^{E(D)}\to\mathbb{N}$ is defined as:
$f_{D}(X)=\left|S_{V}(X)\cup S_{V}(E(D)\setminus X)\right|,$
where $S_{V}(X)$ and $S_{V}(E(D)\setminus X)$ are the directed vertex-separators corresponding to the edge subsets $X$ and $E(D)\setminus X$, respectively.
A directed branch-decomposition of $D$ is any layout of $f_{D}$ on $E(D)$ using a tree $T$ where the vertices correspond to subsets of $E(D)$. The width of this decomposition is the maximum value of $f_{D}(X)$ over all edges $e$ in $T$.
A directed linear-branch-decomposition is a specific type of directed branch-decomposition where the edges $E(D)$ are arranged in a linear ordering $(e_{1},e_{2},\dots,e_{m})$. The function $f_{D}(X)$ is then evaluated for each possible cut in this ordering, corresponding to the sets $\{e_{1},\dots,e_{i}\}$ and $\{e_{i+1},\dots,e_{m}\}$ for $11\leq i\leq m-$ 1.
The directed linear-branch width of $D$ is defined as the maximum value of $f_{D}(X)$ over all possible cuts in the linear ordering of the edges of $D$.
6.1.12 Directed Connectivity system
In the literature [597], the concepts of Directed subset and Directed Submodular Function were defined. The definitions are provided below. We intend to explore whether these concepts can be used to characterize Directed Tree-width and Directed Width parameters in the future.
Definition 6.40.
[597] Suppose that $S$ is a finite set, $s=|S|$. A directed subset $X$ of $S$ is a collection of elements of $S$, where the elements of $X$ are distinguished between forward and backward elements. We identify $X$ with an $s$-dimensional vector, its incidence vector. For an element $e\in S$, let
$X(e)=1,\quad\text{if }e\text{ is a forward element of }X;$
$X(e)=-1,\quad\text{if }e\text{ is a backward element of }X;$
$X(e)=0,\quad\text{if }e\text{ is not in }X.$
Define $-X$ by $-X(e)=-X(e)$, and $\text{abs}(X)$ by $\text{abs}(X)(e)=|X(e)|$.
Suppose that $X$ and $Y$ are two directed subsets of $S$. If all forward elements of $X$ are forward elements of $Y$ and all backward elements of $X$ are backward elements of $Y$, then we say that $X$ is contained in $Y$ and denote this fact as $X\subseteq Y$. An element $e$ is a forward element of $X\cap Y$ if and only if it is a forward element of both $X$ and $Y$. Similarly, an element $e$ is a backward element of $X\cap Y$ if and only if it is a backward element of both $X$ and $Y$. An element $e$ is a forward element of $X\cup Y$ if and only if it is a forward element of one of $X$ and $Y$ and not a backward element of another. Similarly, an element $e$ is a backward element of $X\cup Y$ if and only if it is a backward element of one of $X$ and $Y$ and not a forward element of another. Equivalently,
$(X\cup Y)(e)=\begin{cases}1,&\text{if }X(e)+Y(e)>0,\\ -1,&\text{if }X(e)+Y(e)<0,\\ 0,&\text{if }X(e)+Y(e)=0.\end{cases}$
$(X\cap Y)(e)=\begin{cases}1,&\text{if }X(e)=Y(e)=1,\\ -1,&\text{if }X(e)=Y(e)=-1,\\ 0,&\text{otherwise}.\end{cases}$
Definition 6.41.
[597] Suppose that $\mathcal{S}$ is an interesting family of directed subsets of a finite set $S$. A set-function $f:\mathcal{S}\to\mathbb{R}$ is called a directed submodular function on $\mathcal{S}$ if for all interesting directed subsets $X$ and $Y$ in $\mathcal{S}$,
[…]
We are considering whether the Directed Connectivity System, based on the above Directed Subset and Directed Submodular Function, can be used for any characterizations. Additionally, we are also exploring the possibility of further extending this framework using bisubmodular functions [402, 26] or other related concepts.
Furthermore, we are interested in exploring whether Tree-width, Path-width, Branch-width, Rank-width, Linear-width, and other related parameters can be defined on a ditroid (the directed set version of a matroid). The definition of a ditroid is provided below [597]. Also we are interested in width-parameters on fuzzy matroids [632, 500, 339, 633] or connectoids[114, 115].
Definition 6.42.
[597] Let $X$ a finite set. Two directed subsets $A$ and $B$ of $X$ are called non-cancelling if there is no $e$ in $X$ such that $A(e)=-B(e)\neq 0$. Otherwise, $A$ and $B$ are called cancelling.
Suppose that $\mathcal{M}$ is a family of directed subsets of $X$. We call $D=(X,\mathcal{M})$ a ditroid if the following conditions are satisfied:
  1. 1.
    $\emptyset\in\mathcal{M}$,
  2. 2.
    If $A\in\mathcal{M}$ and $B\subseteq A$, then $B\in\mathcal{M}$,
  3. 3.
    If $A$ and $B\in\mathcal{M}$ are non-cancelling with $|A|=|B|+1$, then there exists $e$ in $X$ and $C$ in $\mathcal{M}$ such that $B(e)=0$, $C(e)=A(e)\neq 0$, and $C(e^{\prime})=B(e^{\prime})$ for all $e^{\prime}\neq e$.
A directed subset in $\mathcal{M}$ is called an independent directed subset of $D$. A maximal independent directed subset is called a base of $D$. The rank function of $D$ is a set-function $r:D(X)\rightarrow\mathbb{R}$ defined by
$r(A)=\max\{|B|:B\subseteq A,B\in\mathcal{M}\},\quad\text{for }A\in D(X).$
The graph class known as Bidirected Graphs [110, 710] has also been a subject of study. Since it generalizes the concept of Directed Graphs, I anticipate that research in this area will continue to grow. In fact, Bidirected Tree-width has already been defined in the literature [698]. Therefore, defining parameters such as Bidirected Path-width (the path version of Bidirected Tree-width), Bidirected Cut-width, Bidirected Branch-width, Bidirected Linear-width, and Bidirected Connectivity system and exploring their relationships, would be highly meaningful. Additionally, extending these concepts to bidirected hypergraphs is another avenue worth exploring (cf.[645]).
6.2 Consideration about Property of an Ultrafilter
We will explore additional properties of an ultrafilter on a connectivity system, including game-theoretical interpretations of ultrafilters and ultraproducts within such systems.
6.2.1 Simple Game on a Connectivity System
A simple game is a mathematical model of voting systems where a coalition of agents wins if it belongs to a specified set of “winning” coalitions, adhering to certain conditions like monotonicity (cf. [274, 545, 175, 658]). These simple games model voting power and provide a generalized interpretation of the notion of what is “majority.” This game is well-known to be closely related to ultrafilters.
A simple game on a connectivity system can be modeled similarly to the simple game in the context of voting systems, but within the framework of connectivity systems, incorporating the constraint $f(A)\leq k$. In the future, we will consider a simple game on a connectivity system.
Game 6.43.
Simple Game on a Connectivity System
Let $X$ be a finite set, $f$ be a symmetric submodular function, and $W\subseteq 2^{X}$ be any collection that satisfies the following monotonicity condition (M1). The triple $(X,W,f)$ is called a simple game on a connectivity system if for every $A\in W$, $f(A)\leq k$.
Basic Definitions of a Simple Game:
  1. 1.
    Connectivity System:
    • A connectivity system is defined as a pair $(X,f)$, where $X$ is a finite set and $f:2^{X}\to\mathbb{N}$ is a symmetric submodular function.
  2. 2.
    Winning Coalition:
    • Let $X$ be a set of agents and $W\subseteq 2^{X}$ be a collection of subsets that represent the majorities or winning coalitions of $X$.
Conditions of a Simple Game:
  1. 1.
    Monotonicity Condition (M1) on a Connectivity System:
    • The collection $W$ is closed under supersets: If $A\in W$ and $A\subseteq B\subseteq X$ with $f(B)\leq k$, then $B\in W$.
  2. 2.
    Proper Simple Game on a Connectivity System:
    • A simple game is called proper if it satisfies: $A\in W$ implies $X\setminus A\notin W$, provided $f(A)\leq k$.
  3. 3.
    Strong Simple Game on a Connectivity System:
    • A simple game is called strong if it satisfies: $A\notin W$ implies $X\setminus A\in W$, provided $f(A)\leq k$.
Conjecture 6.44.
An ultrafilter on a connectivity system is a simple game on a connectivity system.
6.2.2 The Axiom of Choice in the Context of Connectivity Systems
The Axiom of Choice states that for any set of nonempty sets, there exists a function selecting one element from each set, ensuring the product is nonempty. The Axiom of Choice is widely used in real-world applications, such as in optimization, decision theory, and economics, where selecting elements from sets is essential for solving complex problems [417, 375, 416, 275, 390, 392, 128].
In the future, we will consider the Axiom of Choice within the context of connectivity systems. To consider the Axiom of Choice within a connectivity system, we need to translate the concepts into the framework of connectivity systems:
  • Set of Nonempty Subsets:
    Let $\{S_{\alpha}:\alpha\in A\}$ be a collection of nonempty subsets of $X$ within the connectivity system $(X,f)$. Each $S_{\alpha}$ is a nonempty subset of $X$, and $f(S_{\alpha})\leq k$ for some $k$.
  • Product of Sets:
    The product $\prod_{\alpha\in A}S_{\alpha}$ represents the Cartesian product of the sets $S_{\alpha}$. In the context of a connectivity system, this product can be interpreted as a selection of one element from each $S_{\alpha}$.
  • Nonemptiness of the Product:
    The Axiom of Choice asserts that there exists a function $g:A\to X$ such that $g(\alpha)\in S_{\alpha}$ for all $\alpha\in A$. This means there is a way to select one element from each subset $S_{\alpha}$ while maintaining the constraints imposed by the connectivity system.
Additionally, we will consider selective (Ramsey) ultrafilters and $P$-ultrafilters on a connectivity system [670, 538, 588, 271].
6.2.3 Ultraproducts and Ultrapowers on Connectivity Systems
Ultraproducts combine structures using ultrafilters, allowing the creation of a new structure that retains properties from the original structures. This is often used in model theory to study logical consistency [468, 388, 387, 623, 229, 415]. We will extend the construction of ultraproducts to the context of a symmetric submodular function $f$.
Although it is still in the conceptual stage, the ultraproduct on a connectivity system can be described as follows. Let $\{M_{i}\}_{i\in I}$ be a collection of sets indexed by $I$, and let $U$ be an ultrafilter on a connectivity system on $I$. We consider the direct product $\prod_{i\in I}M_{i}$, which consists of all possible combinations of elements from each $M_{i}$. Instead of simply considering the direct product, we now define a set of subsets $A\subseteq\prod_{i\in I}M_{i}$ such that $f(A)\leq k$. This ensures that the subsets considered in the product adhere to the submodular constraints.
Using the ultrafilter $U$ on the connectivity system, we select subsets of $\prod_{i\in I}M_{i}$ that are included in the ultrafilter. This step filters the product to only include elements in subsets dictated by $U$.
Finally, we take the direct limit:
$\lim_{J\in U}\left\{A\subseteq\prod_{i\in J}M_{i}\mid f(A)\leq k\right\}.$
This limit considers the stable elements in the context of the ultrafilter $U$, yielding the ultraproduct
$\left(\prod_{i\in J}M_{i}\right)/U.$
6.2.4 Algorithm of Constructing an Ultrafilter on a Connectivity System
We will consider an algorithm for constructing an ultrafilter on a connectivity system. The algorithm is as follows.
Algorithm 6.45.
  • Input: A finite set $X$ with $\lvert X\rvert=n$. A symmetric submodular function $f:2^{X}\rightarrow\mathbb{N}$.
  • Output: An ultrafilter $U$ of order $k+1$ on the connectivity system $(X,f)$.
Algorithm:
  1. 1.
    Initialization: Start with an empty set collection $U=\emptyset$. Let $L$ be the list of all subsets $A\subseteq X$ such that $f(A)\leq k$.
  2. 2.
    Step 1: Generate Candidate Sets:
    • Enumerate all subsets $A\subseteq X$ such that $f(A)\leq k$. Store these subsets in $L$.
    • Note: The condition $f(A)\leq k$ ensures that the function values are bounded, adhering to the order $k+1$.
  3. 3.
    Step 2: Construct the Filter:
    • Select the first non-empty subset $A_{0}$ from $L$ and include it in $U$, i.e., $U=\{A_{0}\}$.
    • For each subsequent subset $A_{i}\in L$, do the following:
      • Intersection Check: Check whether for all sets $B\in U$, the intersection $A_{i}\cap B$ satisfies $f(A_{i}\cap B)\leq k$.
      • If the condition is met, include $A_{i}$ in $U$; otherwise, discard $A_{i}$.
  4. 4.
    Step 3: Extend to an Ultrafilter:
    • For each subset $A\subseteq X$ not yet in $U$ with $f(A)\leq k$, perform the following:
      • If adding $A$ maintains the properties of $U$, include $A$ in $U$.
      • If adding $A$ violates the properties of $U$, include $X\setminus A$ instead, ensuring the maximality of $U$.
  5. 5.
    Termination:
    • Continue the above steps until every subset $A\subseteq X$ with $f(A)\leq k$ is either in $U$ or its complement $X\setminus A$ is in $U$.
    • The process ends when $L$ is exhausted, ensuring that $U$ is maximal and satisfies the ultrafilter conditions.
    • Return the collection $U$, which is the desired ultrafilter of order $k+1$.
Lemma 6.46.
Above algorithm is correct.
Proof 6.47.
  1. 1.
    The construction of $U$ begins with a non-empty set $A_{0}$, ensuring that $U$ is non-empty. Additionally, $f(A_{0})\leq k$ guarantees that the initial set satisfies the $k$-efficiency condition.
  2. 2.
    Every subset $A$ included in $U$ is selected based on the condition $f(A)\leq k$. Therefore, $U$ inherently satisfies the $k$-efficiency condition.
  3. 3.
    Closed under Finite Intersections: For each new set $A_{i}$ considered, it is included in $U$ only if $A_{i}\cap B$ for all $B\in U$ also satisfies $f(A_{i}\cap B)\leq k$. This ensures closure under finite intersections.
  4. 4.
    Maximality: The final step in the algorithm ensures that for every subset $A\subseteq X$, either $A$ or $X\setminus A$ is included in $U$. This guarantees the ultrafilter’s maximality, as no additional subsets can be added without violating the ultrafilter conditions. Therefore, the lemma is proved.
Lemma 6.48.
The total time complexity of above algorithm is $O(2^{2n})$.
Proof 6.49.
The algorithm must consider every possible subset of the finite set $X$ to construct the ultrafilter $U$. Since $X$ contains $n$ elements, the power set $2^{X}$ consists of $2^{n}$ subsets. Therefore, generating all subsets of $X$ requires $O(2^{n})$ operations.
For each subset $A_{i}$ in $2^{X}$, the algorithm checks its intersection with every other subset $B$ already included in $U$ to ensure that the submodular condition $f(A_{i}\cap B)\leq k$ is satisfied.
In the worst-case scenario, this operation could involve checking intersections with up to $2^{n}$ subsets, which adds another factor of $O(2^{n})$ to the time complexity. Therefore, for each subset $A_{i}$, the total time complexity for evaluating all intersection properties is $O(2^{n})\times O(2^{n})=O(2^{2n})$.
Considering both the generation of subsets and the intersection evaluation, the overall time complexity of the algorithm is $O(2^{2n})$. Therefore, the lemma is proved.
Lemma 6.50.
The space complexity of the algorithm is $O(2^{n})$.
Proof 6.51.
The filter $U$ and the list $L$ both store collections of subsets of $X$. Since there are $2^{n}$ possible subsets, the space required to store either $U$ or $L$ is $O(2^{n})$. The total space complexity is thus $O(2^{n})$. Therefore, the lemma is proved.
We believe there is still room for improvement in the aforementioned algorithm. Additionally, we plan to investigate efficient algorithms tailored specifically to certain graphs, such as planar graphs, in the future.
6.2.5 Ultrafilter Width under Small Set Connectivity Expansion Hypothesis
In computational complexity theory, a computational hardness assumption posits that a specific problem cannot be solved efficiently. One such pivotal assumption is the Small Set Expansion Hypothesis (SSEH), introduced by Raghavendra and Steurer, which plays a critical role in understanding the hardness of approximation algorithms [600]. The SSEH asserts that it is NP-hard to differentiate between small sets of vertices in a graph with low edge expansion (few edges leaving the set) and those with high edge expansion (many edges leaving the set). This hypothesis has become fundamental in analyzing the complexity of various graph algorithms. Since its inception, SSEH has led to numerous inapproximability results, including those documented in [702, 39, 714, 137, 251, 70].
We will consider the concept of Ultrafilter Width. The Ultrafilter Width of a graph $G$ is the maximum order of an ultrafilter in $G$. Similarly, the Ultrafilter Width of a connectivity system is the maximum order of an ultrafilter in the connectivity system.
Conjecture 6.52.
Under the Small Set Expansion Hypothesis (SSEH), it is NP-hard to approximate the Ultrafilter width of a graph within a constant factor in polynomial time.
We will also introduce a new notion called Small Set Connectivity Expansion (SSCE). The definition is as follows. The transformation of the SSEH to apply to connectivity systems rather than simple graphs extends its applicability to more complex and abstract structures, allowing for the examination of expansion properties beyond traditional vertex-edge models.
Definition 6.53.
Let $(X,f)$ be a connectivity system, with $X$ a finite set and $f$ a symmetric submodular function as defined. Consider the analogy of edge expansion in a graph to the concept of a “boundary” in a connectivity system.
  • Connectivity Expansion: For a subset $S\subseteq X$, define the connectivity expansion, $\Psi_{f}(S)$, as:
    $\Psi_{f}(S)=\min\{\text{vol}_{f}(S),\text{vol}_{f}(X\setminus S)\}\cdot f(S)$
    where $f(S)$ quantifies the “connectivity boundary” of $S$, and $\text{vol}_{f}(S)=\sum_{x\in S}f(\{x\})$ represents the volume of $S$ in terms of connectivity, analogous to the sum of degrees in graph theory.
  • For any given small constant $\epsilon>0$, it is computationally challenging to find a non-trivial subset $S\subseteq X$ of the connectivity system such that $\Psi_{f}(S)<\epsilon$.
Question 6.54.
Is it NP-hard to approximate the Ultrafilter width of a graph within a constant factor in polynomial time, under the Small Set Connectivity Expansion Hypothesis (SSCEH)?
6.2.6 Various Ultrafilters on a Connectivity System
Ultrafilters are of interest due to their practical applications and mathematical properties, prompting the proposal of various similar concepts across different fields. Going forward, we aim to define concepts such as Partition-filter [377], uniform-ultrafilters [557], weak normal-ultrafilter [433], gw-ultrafilters [475], subuniform ultrafilters [556], Fuzzy-ultrafilter [646],Regular-ultrafilters [502], Good-ultrafilters [166], Semi-ultrafilters [520], Linear-ultrafilters [58, 378], OK-ultrafilters [207], complete-ultrafilters [594], Ramsey-ultrafilters [589], and Selective-ultrafilters [589] within the frameworks of Undirected Graph, Directed Graph, Infinite Graph, Connectivity Systems and Infinite Connectivity Systems. Additionally, we plan to explore their relationships with graph width parameters.
6.3 Other research Themes for graph width parameters
In the field of graph width parameters, research has been conducted on the following topics. Moving forward, I hope to contribute to the advancement of graph theory by exploring these aspects.
  • Analysis of hierarchies and graph problems within restricted and extended states of specific graphs (ex.[109, 133, 364]).
    • Hierarchies and graph problems restricted to graph width parameters [13, 46].
    • Example of specific graphs: Line Graph[383, 385], planar graph[184, 350, 611, 23], AT-Free Graph[342, 584, 111], chordal graph[371, 203, 118], and sequence Graph[268, 368], etc…
  • Algorithms under constraints imposed by graph width parameters [690, 445, 489, 93].
  • Structural properties of generalized or restricted graph width parameters.
  • Interpretation of graph width parameters using game theory[474, 641, 356, 304].
  • The relationship between forbidden minors and graph width parameters[366, 654, 457, 653].
7 Acknowledgments
I humbly express my sincere gratitude to all those who have extended their invaluable support, enabling me to successfully accomplish this paper.
8 Conflict of Interest
The author declares no conflicts of interest.
Appendix A Various Width Parameters
Numerous graph width parameters have been identified and studied extensively. As illustrated, many of these parameters have significant applications in real-world scenarios and other fields of research, which suggests that the study of graph width parameters will continue to grow in importance. Examining the relationships among these parameters, such as inequalities, upper bounds, and lower bounds, is a well-established research topic (cf. Appendix C). Additionally, we aim to explore the potential connections between these width parameters and ultrafilters in future studies.
Note that unless otherwise specified, we consider any graph when comparing graph parameters. We say that a parameter $p$ upper bounds a parameter $q$ if there exists a non-decreasing function $f$ such that $f(p(G))\geq q(G)$ for all graphs $G$. Conversely, if $p$ does not upper bound $q$, then $q$ is considered unbounded with respect to $p$. Additionally, if value $a$ upper bounds value $b$ and $b$ upper bounds value $c$, then value $a$ upper bounds $c$.
A.1 Tree-width
Tree-width is a measure of how tree-like a graph is, defined by its minimum tree decomposition width [613, 84, 608, 609, 624, 396, 612]. The definitions of Tree-width on a graph are provided below.
Definition A.1.
A tree-decomposition of a graph $G=(V,E)$ is a pair ($(\{X_{i}\mid i\in I\},T=(I,F))$, where:
  • $\{X_{i}\mid i\in I\}$ is a collection of subsets of $V$, known as bags.
  • $T=(I,F)$ is a tree whose nodes correspond to the bags $X_{i}$.
The tree-decomposition must satisfy the following conditions:
  1. 1.
    $\bigcup_{i\in I}X_{i}=V$, meaning every vertex of $G$ appears in at least one bag.
  2. 2.
    For every edge $uv\in E$, there is a bag $X_{i}$ such that $u,v\in X_{i}$.
  3. 3.
    For all $i,j,k\in I$, if $j$ lies on the path from $i$ to $k$ in $T$, then $X_{i}\cap X_{k}\subseteq X_{j}$. Equivalently, for every vertex $v\in V$, the set of bags containing $v$ forms a connected subtree of $T$.
Definition A.2.
The width of a tree-decomposition ($(\{X_{i}\mid i\in I\},T=(I,F))$ is defined as:
$\text{width}(T)=\max_{i\in I}|X_{i}|-1.$
The tree-width of a graph $G$, denoted by $\text{tw}(G)$, is the minimum width over all possible tree-decompositions of $G$. A graph has tree-width 1 if and only if it is a tree.
Examples of obstructions for tree-decomposition are Tangle [613], Ultrafilter [286], and Bramble [154, 92, 91, 491]. For more detailed information, please refer to lecture notes or surveys [72, 462, 85, 126].
A.1.1 Hierarchy for treewidth
We consider about Hierarchy for treewidth.
Theorem A.3.
The relationships between tree-width and other parameters are as follows:
  • If $G$ has tree-width $k$, then $G$ is a subgraph of a chordal graph with maximum clique size $k+1$ [85].
  • Let $G$ be a connected graph with treewidth $tw(G)$ and gonality $dgon(G)$. Then $tw(G)\leq dgon(G)$ [678].
  • If $G$ has tree-width $k$, the dimension of $G$ is at most $k$ [85].
  • If $G$ has tree-width $k$, $G$ is $k$-decomposable [85].
  • If $G$ has tree-width $k-1$, $G$ has haven $k$ [625].
  • If $G$ has tree-width $k$, $G$ has no screen of thickness at least $k+2$ [85].
  • If $G$ has tree-width $k$, $k+1$ cops can search $G$ in the Seymour-Thomas search game [85].
  • If $G$ has tree-width $k$, $k+1$ cops can monotonely search $G$ in the Seymour-Thomas search game [85].
  • If $G$ has tree-width $k$, the number of searchers needed to search $G$ in the fugitive search game with an inert fugitive is at most $k+1$ [85].
  • If $G$ has tree-width $k$, the number of searchers needed to monotonically search $G$ in the fugitive search game with an inert fugitive is at most $k+1$ [85].
[…]
  • $G$ is a partial strict $k$-tree [215].
  • $qn(G)\leq pw(G)\leq(tw(G)+1)\log n$ (queue-number $qn(G)$, path-width $pw$) [328].
  • The stack-number $sn(G)\leq tw(G)+1$ [328].
  • The track number track $\text{track number}(G)\leq\text{path-width}(G)+1\leq 1+(\text{tree-width}(G)+1)\log n$ [215].
  • The tree-width of a graph is a lower bound on the positive semidefinite zero forcing number $Z^{+}(G)$ [230].  [^2] The zero forcing number is the minimum number of initially colored vertices required to color the entire graph by repeatedly applying the ”color-change rule(cf.[430, 227]).
  • $tw(G)\leq la(G)\leq tw(G)+1$ ($la(G)$: Largeur d’arborescence. Also called Two-sided tree-width) [677, 178].
  • The spanheight of a graph upper-bounds the treewidth of a graph [615].
  • Ultrafilter on a graph is an obstacle of tree-width [287].
  • Let $G$ be any non-complete graph. Then toughness $\leq tw(G)/2$ [332].
  • $\text{treewidth}(G)\leq\text{pathwidth}(G)\leq(\text{treewidth}(G)+1)\log(n)$.
  • $\text{bw}(G)\leq\text{tw}(G)+1\leq\frac{3}{2}\cdot\text{bw}(G)$ for a graph $G$ with $\text{bw}(G)\geq 2$ (bw: branch-width) [84].
  • $\text{tw}(G)\leq 2\cdot\text{tcw}(G)^{2}+3\cdot\text{tcw}(G)$ (tcw: Tree-cut-width) [323].
  • $\text{boow}(G)\leq\text{tw}(G)+1$ (boow: Boolean-width) [323].
  • If a graph $G$ has bounded tree-width, then it also has bounded sparse twin-width, bounded clique-width, and bounded twin-width [5].
  • If a graph $G$ has bounded Tree-distance-width, then it also has bounded Tree-width [715].
  • If a graph $G$ has bounded tree-width, then it also has bounded Maximum clique number, and chromatic number [726, 618].
  • If a graph $G$ has bounded tree-width, then it also has bounded Bramble Number (bn), Tangle Number, Lexicographic Tree Product Number (LTP), Linkedness Number (link), Cartesian Tree Product Number (ctp), and Well-Linked Number (wl). The converse is also true [384].
[…]
  • $\text{carw}(G)\cdot\Delta(G)\leq\text{tw}(G)\leq 2\times\text{carw}(G)$ (where $\Delta(G)$ is the maximum degree of $G$ and $\text{carw}(G)$ is carving width) [661, 69, 560].
  • Minimal upper bounds for tree-width is distance to outerplanar [84].
  • Tree-width is a minimal upper bound for book thickness (page number [60]) and acyclic chromatic number [358, 223, 219, 220].  [^3] The acyclic chromatic number of a graph is the minimum number of colors needed to color the vertices such that no two adjacent vertices share the same color and no cycle is bichromatic(cf.[134, 685]).
  • If a graph $G$ has bounded Vertex-Cover Number, then it also has bounded tree-width. [^4] The Vertex-Cover Number is the smallest number of vertices needed to cover all edges in a graph, meaning each edge has at least one endpoint in the vertex cover(cf.[686, 590]).
  • If a graph $G$ has bounded Tree-width, then it also has bounded Degeneracy, Average Degree, Minimum Degree, Domatic Number [249], and Distance to Disconnected [668]. [^5] Degeneracy is the smallest number $d$ such that every subgraph has a vertex of degree at most $d$, indicating a graph’s sparsity(cf.[566, 247]).
  • For any graph $G$, $\text{box}(G)\leq\text{tw}(G)+2$, where $\text{box}(G)$ and $\text{tw}(G)$ denote the boxicity and treewidth of $G$, respectively [149].  [^6] The boxicity of a graph is the minimum number of dimensions needed to represent the graph as an intersection graph of axis-aligned boxes in Euclidean space (cf. [246, 148]).
  • Every graph of tree-width $k$ has a balanced vertex cut of size at most $k+1$ [614].  [^7] A balanced vertex cut of size $k$ is a set of $k$ vertices whose removal divides a graph into roughly equal-sized components [614].
  • Every graph of tree width $k$ is $k$-degenerate [107].
  • Every graph of tree-width $k$ admits a 2-balanced binary tree-decomposition of width at most $23k+$ 2 [81].
  • Let $G$ be a graph. Then $\text{tw}(G)\geq$ minimum degree [333].
  • Let $G$ be a graph on $n$ vertices. Then $\text{tw}(G)\leq n-$ maximum independent set [333].  [^8] The maximum independent set is the largest set of vertices in a graph where no two vertices are adjacent, meaning no edges exist between any pair of vertices in the set(cf.[706, 631]).
  • Let $G$ be a graph and $m\geq k\geq 1$ integers. If $\text{tw}(G)\geq k+m-1$, then $G$ has a $k$-mesh of order $m$ [196].
  • If $G$ is a $k$-outerplanar graph, then $\text{tw}(G)\leq 3k-1$ [80].
Proof A.4.
Please refer to each reference.
A.1.2 Related concepts for tree-width
Related concepts include maximum order of a grid minor, maximum order of a grid-like-minor, Hadwiger number, fractional Hadwiger number, domino treewidth [88, 86], and the following concepts [358].
  • Local Tree-width [352]: A graph width parameter that associates with every natural number $r$ the maximum tree-width of an $r$-neighborhood in $G$.
  • Linear Local Tree-width: A linear variation of Local Tree-width.
  • Simple Tree-width:
    • For every $G$, we have $tw(G)\leq stw(G)\leq tw(G)+1$ [463].
    • A simple $k$-tree is a $k$-tree with the extra requirement that there is a construction sequence in which no two vertices are attached to the same $k$-clique.
  • Bag-connected Tree-width [427]: A graph width parameter that considers tree-decompositions where each cluster is connected.
  • Circuit Tree-width [605, 113]:
    • Circuit Tree-width or Expression width [419] measures, for any Boolean function, the smallest treewidth of a circuit computing the function when considering an ordered binary decision diagram (OBDD).
    • Related parameters are OBDD width [420] and SDD width [113].
  • Stacked Treewidth [705]: Width related to stacked $k$-tree.
    • A $k$-tree is called a stacked $k$-tree if during its construction no two vertices are stacked onto the same $k$-clique.
  • Edge-tree Width [518]: A graph width parameter that serves as the tree-like analogue of cut-width, or as an edge-analogue of tree-width.
  • Layered Treewidth [108, 216]: The minimum integer $k$ such that $G$ has a tree-decomposition and layering where each bag has at most $k$ vertices in each layer. $\text{The minimum degree }\leq 3\quad ltw(G)\leq 1.$ [217]. $ltw(G)\leq 2g+3,\text{ for any genus }g\text{ graph}$.
  • Weighted Tree-width: The weighted treewidth of a weighted graph is defined analogously to the (unweighted) treewidth [432].
  • 2-dimensional treewidth: Width defined as the minimum integer $k$ such that $G$ has two $k$-orthogonal tree decompositions $S$ and $T$.
[…]
  • Row Treewidth [108, 213]: The minimum integer $k$ such that $G$ is isomorphic to a subgraph of some graph $H$ with tree-width at most $k$ and for some path $P$.
  • Connected Tree-width [198, 554]: The minimum width of a tree-decomposition whose parts induce connected subgraphs.
  • Triangular Tree-width [534]: A graph width parameter known for making the computation of permanents of matrices with bounded width efficient.
  • Special Tree-width [170]: A complexity measure of graphs, positioned between path-width and tree-width.
  • Spaghetti Tree-width [480, 89]: A graph width parameter similar to Special Tree-width. Related concepts include strongly chordal tree-width and Directed Spaghetti Tree-width [89].
  • Pared Tree Width and Acceptance Width [447, 448]: Graph width parameters relevant to the computation of an alternating finite automaton.
  • Co-treewidth [210]: The tree-width of the complement of the input graph.
  • Constant Treewidth [30, 119]: A graph width parameter characteristic of control flow graphs in most programs.
  • Dynamic Treewidth [21, 472]: A graph width parameter designed to be computed for a dynamic graph $G$.
  • Project-join Tree Width [604]: A graph width parameter linked to weighted projected model counting with graded project-join trees.
  • Semantic Tree-width [259]: This parameter determines whether a query is equivalent to a query of tree-width $k$, applicable to the class of Unions of Conjunctive Regular Path Queries with two-way navigation.
  • Induced Tree Width [680]: A graph width parameter where the complexity of algorithms is dictated by the highest-dimensional factor across computations.
  • Free-connex Tree-width [222, 102]: A graph width parameter that requires a connected set of nodes, including the root, containing exactly the output variables.
  • Effective Width [245]: A graph width parameter for Partially Ordered Time, positioned between the well-known properties of width and dimension.
  • Embedded Width [693]: Embedded width is a variation on tree-width, which is restricted to surface-embedded graphs.
  • Backdoor-tree-width [226, 326]: Backdoor-tree-width combines backdoor and tree-width approaches to create a fixed-parameter tractable framework for abstract argumentation, ensuring manageable parameter values. Related concept is C-backdoor tree-width [209].
  • Dual Tree-width [639]: The Tree-width of the dual of a graph.
  • Singly-crossing treewidth [585]: Singly-crossing treewidth has been explored in algorithmic applications in [585] where the structure of a planar graph is integrated with a tree decomposition that includes clique-sums of bounded-size non-planar components.
A.1.3 Application aspect for tree-width
We consider about application aspect for tree-width. Due to the versatility of tree structures in handling various types of data, tree-width has numerous applications across different fields.
  • Protein structure
    Research on proteins is thriving in the field of medicine. Protein structure refers to the three-dimensional arrangement of amino acids in a protein. It can be modeled as a graph where nodes represent amino acids, and edges represent bonds or interactions, capturing the protein’s complex folding and connectivity(cf.[410, 596]). Applications have also been studied in the context of tree-width[709, 586, 400, 707].
  • RNA and DNA
    Research on RNA and DNA is also thriving in the field of medicine. RNA and DNA are molecules that store genetic information. DNA is double-stranded and stable, forming a double helix, while RNA is single-stranded and more flexible, often acting in protein synthesis. Applications have also been studied in the context of tree-width[362, 717, 723, 252].
  • metabolic networks
    Metabolic networks are interconnected pathways within cells that convert nutrients into energy and essential molecules, supporting life processes. Applications have also been studied in the context of tree-width[158]. In addition, the application of Tree-width is being researched from biological and medical perspectives[583, 63].
  • social network
    A social network is a structure of individuals or organizations connected by social relationships, like friendships or collaborations. Applications have also been studied in the context of tree-width[4, 360, 527].
  • Bayesian network
    A Bayesian network is a graphical model representing probabilistic relationships among variables using nodes and directed edges (cf.[592]).Tree-width relates to Bayesian networks by measuring the complexity of inference; lower tree-width allows more efficient probabilistic computations and algorithms[234, 666, 602, 488, 564, 565].
  • Neural networks
    Neural networks are computational models inspired by the human brain, consisting of interconnected nodes that process data and learn patterns. A related concept is graph neural networks [725]. Applications have also been studied in the context of tree-width[47].
  • Database
    Applications for Database have also been studied in the context of tree-width[258, 354, 346, 345].
  • Internet
    Applications for Internet have also been studied in the context of tree-width[177].
  • Markov network
    Markov networks, or Markov random fields, are undirected graphical models representing the joint distribution of variables. They are used in applications like image processing and statistical physics, emphasizing local dependencies (cf.[683]). Applications for Markov network have also been studied in the context of tree-width[644, 501, 438].
  • Tensor network
    Tensor networks are mathematical structures used to efficiently represent and compute high-dimensional data. They are applied in quantum physics, machine learning, and compressing large datasets by capturing complex relationships among variables(cf.[421]). Applications for Tensor network have also been studied in the context of tree-width[221].
For those interested in a more detailed understanding of tree-width, I recommend consulting lecture notes and surveys. Fortunately, there are many excellent books and survey papers available on the topic [83, 72, 396, 85, 462, 384, 87, 94].
A.2 Path-width
Path-width [453, 454, 372, 442] quantifies how similar a graph is to a path by determining the minimum width of its path decomposition. An example of an obstruction to low path-width is blockage [73]. Path-width is notably applied in VLSI design [571]. Path-width is also called interval-width [459, 265].
As an example, the definition of Path-width in the context of matroids is provided. For the definition of a matroid, please refer to Definition A.24 and related references.
Definition A.5.
[442] Given an ordering $(e_{1},e_{2},\dots,e_{n})$ of the elements of $M$, define the width of the ordering as:
$w_{M}(e_{1},e_{2},\dots,e_{n})=\max_{i\in[n]}f_{M}(e_{1},e_{2},\dots,e_{i}).$
For simplicity, we use $f_{M}(e_{1},e_{2},\dots,e_{i})$ instead of $f_{M}(\{e_{1},e_{2},\dots,e_{i}\})$.
The pathwidth of $M$ is defined as:
$\text{pw}(M)=\min w_{M}(e_{1},e_{2},\dots,e_{n}),$
where the minimum is taken over all orderings $(e_{1},e_{2},\dots,e_{n})$ of $E(M)$. An ordering $(e_{1},e_{2},\dots,e_{n})$ of $E(M)$ such that $w_{M}(e_{1},e_{2},\dots,e_{n})=\text{pw}(M)$ is called an optimal ordering.
A.2.1 Hierarchy for path-width
We consider about Hierarchy for path-width.
Theorem A.6.
The relationships between path-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Path-width, then it also has bounded Tree-width (Trivially).
  • The path-width of any graph is equal to one less than the smallest clique number of an interval graph [84].
  • If a graph $G$ has bounded Tree-depth, then it also has bounded Path-width [618].
  • If a graph $G$ has bounded Path-width, then it also has bounded linear-rank-width [13].
  • If a graph $G$ has bounded Path-width, then it also has bounded linear-clique-width and linear NLC-width [369].
  • If $G$ has path-width $k$, the interval thickness of $G$ is at most $k+1$ [85].
  • If $G$ has path-width $k$, the minimum progressive black pebble demand over all directives of $G$ is at most $k+1$ [85].
  • If $G$ has path-width $k$, the minimum progressive black and white pebble demand over all directives of $G$ is at most $k+1$ [85].
  • Vertex separation number is equivalent to path-width [235].
  • Node search number is also equal to path-width [460].
  • Path-width is less than or equal to cut-width [471].
  • If a graph $G$ has bounded Path-width, then it also has bounded clique-width [253].
  • If a graph $G$ has bounded Band-width, then it also has bounded path-width [5].
  • If $G$ has $T_{h}$-witness, then $pw(G)\geq h$[351].
Proof A.7.
Please refer to each reference.
A.2.2 Related concepts for path-width
Related concepts for path-width are following:
  • DAG-path-width [440]: Linear layout parameter of DAG-tree-width.
  • Layered path-width [108, 216]: Linear layout parameter of Layered tree-width. For any graph, layered treewidth (pathwidth) $\leq$ treewidth (pathwidth) (trivially). And queue number is bounded by layered pathwidth [215]. Every graph with pathwidth $\text{Every graph with pathwidth }k\text{ has layered pathwidth at most }\frac{k+1}{2}.$ Every graph with layered pathwidth $\text{Every graph with layered pathwidth }\lambda\text{ has track-number at most }3\lambda.$
  • Circuit Pathwidth: Path version of Circuit Tree-width [605, 113].
    • Related parameters are OBDD width [420] and SDD width [113].
  • Simple Pathwidth: The simple pathwidth of $G$ is the smallest $w$ such that $G$ has a $w$-simple path decomposition of width $\leq w$ [71].
  • Row path-width [108, 213]: Linear layout parameter of Row tree-width.
  • Linear Local path-width [212] : A path variation of Linear Local Tree-width.
  • D-path-width and clique preserving d-path-width [604]: A graph width parameter that has the splitting power of branching programs of bounded repetition and CNFs of bounded width.
  • Semantic path-width [259]: Linear layout parameter of Semantic tree-width.
  • Proper-path-width [656, 653, 657, 655, 654, 532, 531]: Path-width that relates to mixed search game (A pursuit-evasion game on graphs where a searcher combines edge and node searches to capture a hidden intruder).
    • For any graph $G$, $pw(G)\leq ppw(G)\leq pw(G)+1$ [655, 654].
    • For any simple graph $G$ and an integer $k$, $ppw(G)\leq k$ if and only if $G$ is a partial $k$-path [655, 654].
    • For any simple graph $G$, Mixed Search Number = Proper-path-width [655, 654].
    • edge search $\text{edge search number}(G)-1\leq\text{Proper-path-width}(G)\leq\text{edge search number}(G)$ [655, 654].
    • node search $\text{node search number}(G)-1\leq\text{Proper-path-width}(G)\leq\text{node search number}(G)$ [655, 654].
  • Connected path-width [179]: Linear layout parameter of Connected tree-width. Connected pathwidth of any graph $G$ is at most $12\times pw(G)+$ 1 [180].
  • Persistence path-width [208]: Persistence path-width refers to a graph path decomposition with width $k$, where each vertex of the graph is contained in at most $l$ nodes of the path.
  • Linked path decompositions [335]: Weaker variant of lean path decomposition (cf. [662, 53]).
  • Dual path-width [639]: The path-width of the dual of a graph (cf. [146]).
A.2.3 Application aspect for path-width
We consider about application aspect for path-width.
  • VLSI
    VLSI (Very Large Scale Integration) refers to the process of integrating thousands to millions of transistors onto a single semiconductor chip (cf.[428, 542, 40]). It’s crucial for developing modern electronic devices, enabling complex circuits and high-performance computing. Applications for VLSI have also been studied in the context of path-width [570, 543, 541].
  • Compiler
    Applications for Compiler have also been studied in the context of path-width[82, 168].
  • Blockchain
    Applications for blockchain have also been studied in the context of path-width[443, 441].
  • Natural Language
    Applications for Natural Language have also been studied in the context of path-width[473].
A.3 Cut-width
Cut-width (also called folding number[162]), referenced in sources [163, 523, 522], measures the minimum number of edges crossing any vertical cut in a linear layout of a graph’s vertices. Related concepts include page-width [358].
Definition A.8.
Let $G=(V,E)$ be an undirected graph, where $V$ is the set of vertices and $E$ is the set of edges. The cutwidth of $G$ is the smallest integer $k$ such that there exists an ordering of the vertices $v_{1},v_{2},\dots,v_{n}$ in $V$, where for every $\ell=1,2,\dots,n-1$, the number of edges $(v_{i},v_{j})\in E$ with $i\leq\ell$ and $j>\ell$ is at most $k$.
A.3.1 Hierarchy for cut-width
We consider about Hierarchy for cut-width.
Theorem A.9.
The relationships between cut-width and other graph parameters are as follows:
  • If a graph $G$ has bounded cut-width, it also has bounded carving-width [151].
  • The cut-width of a graph can provide a lower bound on another parameter, the crossing number [202]. [^9] The crossing number of a graph is the minimum number of edge crossings in any drawing of the graph in the plane(cf.[630, 469]).
  • In subcubic graphs (graphs with a maximum degree of three), the cut-width equals the path-width plus one [524].
[…]
Proof A.10.
Please refer to each reference.
A.3.2 Related concepts for cut-width
Related parameter is Weighted cut-width.
  • Weighted cut-width
    Cut-width which considers the weight of the graph[125, 124].
  • Circuit cut-width
    Circuit cut-width is a graph-width of circuits that characterizes the complexity of solving ATPG (Automatic Test Pattern Generation), demonstrating that circuits with smaller cut-widths can be solved more efficiently [593, 160].
A.3.3 Application aspect for cut-width
We consider about application aspect for cut-width.
  • VLSI
    Applications for VLSI have been studied in the context of cut-width [56].
A.4 MIM-width
MIM-width (maximum induced matching width) [412, 413] measures the largest induced matching in any bipartite graph induced by the cuts in the decomposition.
Definition A.11.
[412, 413] A branch decomposition is a pair $(T,L)$, where $T$ is a subcubic tree and $L$ is a bijection from $V(G)$ to the set of leaves of $T$. For each edge $e$ of $T$, let $T^{e}_{1}$ and $T^{e}_{2}$ be the two connected components of $T-e$. Define $(A^{e}_{1},A^{e}_{2})$ as the vertex bipartition of $G$ such that, for each $i\in\{1,2\}$, the set $A^{e}_{i}$ contains all vertices in $G$ that are mapped by $L$ to leaves within $T^{e}_{i}$.
For a vertex set $A\subseteq V(G)$, let $\text{mim}(A)$ denote the maximum size of an induced matching in $G[A,V(G)\setminus A]$. The mim-width of $(T,L)$, denoted by $\text{mimw}(T,L)$, is defined as $\max_{e\in E(T)}\text{mim}(A^{e}_{1})$. The minimum mim-width over all branch decompositions of $G$ is referred to as the mim-width of $G$. If $|V(G)|\leq 1$, then $G$ does not admit a branch decomposition, and the mim-width of $G$ is defined to be 0.
A.4.1 Hierarchy for MIM-width
Theorem A.12.
The relationships between MIM-width and other graph parameters are as follows:
  • If a graph $G$ has bounded MIM-width, then it also has bounded Sim-width [618].
  • If a graph $G$ has bounded Independent set or Dominating Number, then it also has bounded MIM-width [618].
  • If a graph $G$ has bounded Boolean-width, then it also has bounded MIM-width [412, 413].
  • If a graph $G$ has bounded Clique-width, then it also has bounded MIM-width [411].
Proof A.13.
Please refer to each reference.
A.4.2 Related concepts for MIM-width
Related concepts are following:
  • One-sided maximum induced matching-width [59]: Graph width parameters which solves maximum independent set problem. Any graph with tree-independence number $k$ has o-mim-width at most $k$ [59].
  • SIM width [434, 264]: Graph width parameters which is a more useful parameter of MIM-width. Any graph with minor-matching hypertree-width $k$ has sim-width at most $k$ [59].
A.5 Linear MIM-width
Linear MIM-width [412, 413] is the linear restriction of MIM-width. If a graph $G$ has bounded linear-MIM-width, then it also has bounded MIM-width (trivially).
A.5.1 Related concepts for Linear-MIM-width
Related concept is following:
  • Linear SIM width [434]: Linear layout parameter of SIM width. If a graph $G$ has bounded linear-SIM-width, then it also has bounded SIM-width (trivially).
A.6 Boolean-width
Boolean-width [129] measures the width of a graph decomposition based on the number of different unions of neighborhoods across cuts.
Definition A.14.
[129] A decomposition tree of a graph $G$ is a pair $(T,\delta)$, where $T$ is a tree with internal nodes of degree three, and $\delta$ is a bijection between the leaves of $T$ and the vertices of $G$. Removing an edge from $T$ yields two subtrees, and correspondingly, a cut $\{A,\bar{A}\}$ of $G$, where $A\subseteq V(G)$ corresponds to the leaves of one subtree.
Let $f:2^{V(G)}\rightarrow\mathbb{R}$ be a symmetric cut function, meaning $f(A)=f(\bar{A})$ for all $A\subseteq V(G)$. The f-width of $(T,\delta)$ is defined as the maximum value of $f(A)$ over all cuts $\{A,\bar{A}\}$ corresponding to the removal of an edge from $T$.
For rooted trees, subdivide an edge of $T$ to introduce a new root $r$, yielding a binary tree $T_{r}$. The subtree of $T_{r}$ rooted at a node $u$ corresponds to a subset $A_{u}^{r}\subseteq V(G)$, denoted simply by $A_{u}$ when the root is clear. For an edge $\{u,v\}$ with $u$ as a child of $v$ in $T_{r}$, the associated cut is denoted $\{A_{u},\bar{A}_{u}\}$.
A divide-and-conquer approach on $(T,\delta)$, following the edges of $T_{r}$ bottom-up, solves the problem recursively by combining solutions from the cuts given by edges from parent nodes to their children. In the context of independent sets, if two sets $X\subseteq A$ and $X^{\prime}\subseteq A$ have identical neighborhood unions across the cut (i.e., $N(X)\cap\bar{A}=N(X^{\prime})\cap\bar{A}$), they can be treated equivalently, suggesting the minimization of distinct neighborhood unions across cuts.
Definition A.15 (Boolean-Width).
[129] Given a graph $G$ and a subset $A\subseteq V(G)$, define the set of neighborhood unions across the cut $\{A,\bar{A}\}$ as:
$U(A)=\{Y\subseteq\bar{A}:\exists X\subseteq A\text{ such that }Y=N(X)\cap\bar{A}\}.$
The bool-dim function for $G$ is defined by:
[…]
The boolean-width of a decomposition tree $(T,\delta)$, denoted by $\text{boolw}(T,\delta)$, and the boolean-width of a graph $G$, denoted by $\text{boolw}(G)$, are defined using the bool-dim function as $f$.
The relationships between boolean-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Tree-width, then it also has bounded Boolean-width [412, 413].
A.7 Linear-Boolean-width
Linear Boolean-width [129] is the linear restriction of Boolean-width. If a graph $G$ has bounded Linear Boolean-width, then it also has bounded Boolean-width (trivially).
For any graph $G$, it holds that $\text{lboolw}(G)\leq\text{pw}(G)+1$ (Path-width) [123].
A.8 Band-width
Band-width [159, 676] measures the minimum width of a band or interval in an optimal linear arrangement of a graph’s vertices. It is a width parameter introduced from VLSI design and networking. Related concepts include split-band-width [264].
Definition A.16.
[159, 676] Let $f:V(G)\to\mathbb{Z}$ be a layout of a graph $G=(V,E)$. The bandwidth of this layout $f$ is defined as $\max_{\{u,v\}\in E(G)}|f(u)-f(v)|$. The bandwidth of the graph $G$ is the minimum bandwidth over all possible layouts of $G$, and this graph parameter is denoted by $\text{BW}(G)$.
A.8.1 Hierarchy for Band-width
We consider about hierarchy for Band-width.
Theorem A.17.
The relationships between band-width(bw) and other graph parameters are as follows:
  • $\text{ppw}(G)\leq\text{bw}(G)\leq 2\times\text{ppw}(G)$ (ppw: path-partition-width)
  • Max-leaf number is a minimal upper bound for band-width [668].
  • $\text{bw}(G)\leq 2\times\text{pdw}(G)$ (pdw: path-distance-width)
  • If a graph $G$ has bounded band-width, then it also has bounded maximum degree, $c$-closure [467], acyclic chromatic number, and $h$-index [240] [668]  [^10] The $h$-index of a graph is the maximum integer $h$ such that the graph has at least $h$ vertices, each with a degree of at least $h$(cf.[241, 74])..
  • Treespan is less than or equal to band-width [266].
  • The slope number $\text{sn}(G)$ satisfies: $\text{sn}(G)\leq\frac{1}{2}\text{bw}(G)\cdot(\text{bw}(G)+1)+1.$[218]
Proof A.18.
Please refer to each reference.
A.9 Boundary-width
Vertex-boundary-width [576] measures the minimum boundary of vertices in a discrete isoperimetric problem. Note that the discrete isoperimetric problem seeks to determine the minimum "boundary" size of a set of vertices in a graph, given a fixed number of vertices in that set, often relating to optimizing shapes in grid-like structures or networks.
Edge-boundary-width [576] measures the minimum boundary of edges in a discrete isoperimetric problem.
A.10 Carving-width
Carving-width [626, 660, 509] measures a graph’s complexity by focusing on edge cuts, similar to tree-width, but with an emphasis on edges rather than vertices.
Definition A.19.
[626, 660, 509] A tree $T$ is called a carving of a graph $G$, and the pair $(T,w)$ is referred to as a carving decomposition of $G$, where $w(e)$ is a weight function defined on the edges $e\in E(T)$. The width of a carving decomposition $(T,w)$ is the maximum weight $w(e)$ over all edges $e$ in $T$. The carving-width of $G$, denoted by $\text{cw}(G)$, is the minimum width among all possible carving decompositions of $G$. If $|V(G)|=1$, we define $\text{cw}(G)=0$.
A.10.1 Related concept for Carving-width
Related concept is following:
Dual Carving-width [510]: The carving-width of the dual of a graph.
A.10.2 Application for carving-width
Carving-width is closely related to the concept of congestion in network theory. For example, consider a real-world scenario where a network with low performance is used—this often leads to frequent delays and congestion. The level of congestion experienced in such cases is analogous to the carving-width. Extensive research has been conducted to minimize delays and congestion, making this measure particularly relevant in network optimization studies [626]. This graph width parameter is a minimal upper bound for maximum degree and tree-width [618, 54].
A.11 Branch-width
Branch-width [330, 577, 613, 95] measures the smallest width of a branch decomposition of a graph. The width is the minimum number of edges connecting subgraph pairs. This graph width parameter is known to be extendable to matroids and connectivity systems. Examples of obstructions are Tangle [330, 393], Profile [192, 193, 238, 190, 199], $k$-block [231, 244, 464], Ultrafilter [304], Maximal ideal [713], loose tangle [577], Quasi-Ultrafilter [293], Weak-Ultrafilter [288], tangle kit[407], Bramble[516, 24], (k,m)-obstacle [269]. Related concepts include sphere-cut width [75] and Rooted Branch-width[418].
A.11.1 Hierarchy for branch-width
We consider about Hierarchy for branch-width.
Theorem A.20.
The relationships between branch-width and other graph parameters are as follows:
  • For all matroids $M$, $\text{bw}(M)\leq\text{bd}(M)$ (Branch-depth) [572].
  • $\text{bw}(M)\leq\text{cdd}(M)\leq\text{cd}(M)$ and $\text{bw}(M)\leq\text{cdd}(M)\leq\text{dd}(M)$ [618] (cdd:contraction-deletion-depth, cd:contraction-depth, and dd:deletion-depth).
  • $\text{bw}(G)\leq\text{tw}(G)+1\leq\frac{3}{2}\cdot\text{bw}(G)$ for a graph $G$ with $\text{bw}(G)\geq 2$ (tree-width) [84].
  • For every graph $G$ with $\text{bw}(G)\neq 0$, $\text{boolw}(G)\leq\text{bw}(G)$ (Boolean-width) [5].
  • The branch-width of a graph is a lower bound on the zero-forcing number [248].
  • $bw(G)\leq tw(G)+1\leq Z(G)+1$ ($Z(G)$: zero-forcing number)[248].
Proof A.21.
Please refer to each reference.
A.12 Linear Branch-width
Linear Branch-width (Also called Linear-width[301] or caterpillar width[76]) is the linear restriction of Branch-width. This graph width parameter is known to be extendable to matroids and connectivity systems. Examples of obstructions are Linear Tangle [303], Single Ideal [300], Linear Loose Tangle [292], Linear-obstacle [269], Ultra Matroid [284], Ultra Antimatroid [290], Ultra Greedoid [290], Ultra Quasi-matroid [289] (cf. [444]).
A.12.1 Hierarchy for Linear-branch-width
We consider about Hierarchy for Linear-branch-width.
Theorem A.22.
The relationships between Linear Branch-width and other graph parameters are as follows:
  • $\text{pw}(G)\leq\text{lbw}(G)\leq 1+\text{pw}(G)$ (path-width) [568].
  • If a graph $G$ has bounded Linear Branch-width, then it also has bounded Branch-width (trivially).
Proof A.23.
Please refer to each reference.
A.13 Rank-width
Rank-width [404] measures the complexity of a graph based on the rank of adjacency matrices over cuts, closely linked to Matroid theory [578, 695, 699] and the Rank function. Examples of obstructions are $\rho$-Tangle [403].
The concept of a rank function is introduced from the world of matroids, and parameters such as Branch-width, Tree-width, and Tangles have been extended from graphs to matroids[330, 577, 304, 394, 395]. Therefore, we will briefly introduce the concept of matroids. For those interested in learning more about matroids, please refer to books, lecture notes, or surveys on the topic[579, 697, 607, 700, 329].
Definition A.24 (Matroid).
(cf.[579, 697]) A matroid is a collection of subsets $\mathcal{F}$ of a set $E$ that satisfies the following conditions:
  1. 1.
    $E\in\mathcal{F}$,
  2. 2.
    If $F_{1},F_{2}\in\mathcal{F}$, then $F_{1}\cap F_{2}\in\mathcal{F}$,
  3. 3.
    If $F\in\mathcal{F}$ and $\{F_{1},F_{2},\ldots,F_{k}\}$ is the set of minimal members of $\mathcal{F}$ that properly contain $F$, then the sets $F_{1}\setminus F,F_{2}\setminus F,\ldots,F_{k}\setminus F$ partition $E\setminus F$.
Matroid can also be defined as follows.
Definition A.25 (Matroid of Rank $d+1$).
(cf.[579, 697])
A matroid on $E$ of rank $d+1$ is a function $r:2^{E}\rightarrow\mathbb{Z}$ satisfying the following conditions:
  1. 1.
    $0\leq r(S)\leq|S|$ for any $S\subseteq E$,
  2. 2.
    If $S\subseteq U\subseteq E$, then $r(S)\leq r(U)$,
  3. 3.
    $r(S\cup U)+r(S\cap U)\leq r(S)+r(U)$ for any $S,U\subseteq E$,
  4. 4.
    $r(\{0,\ldots,n\})=d+1$.
Based on the above, we define Rank-width as follows.
Definition A.26.
[404] Let $G=(V,E)$ be a graph. The cut-rank function $\rho_{G}$ is defined for any subset $X\subseteq V$ as the rank of the $|X|\times|V\setminus X|$ binary matrix $A_{X}$, where the entry at position $(i,j)$ is 1 if and only if the $i$-th vertex in $X$ is adjacent to the $j$-th vertex in $V\setminus X$. We set $\rho_{G}(X)=0$ when $X=\emptyset$ or $X=V$.
A rank-decomposition of $G$ is a pair $(T,L)$, where $T$ is a subcubic tree and $L:V\to\text{Leaves}(T)$ is a bijection. For each edge $e\in E(T)$, removing $e$ partitions the leaves of $T$ into two sets $A_{e}$ and $B_{e}$. The width of $e$ is defined as $\rho_{G}(L^{-1}(A_{e}))$. The width of the rank-decomposition $(T,L)$ is the maximum width over all edges $e\in E(T)$.
The rank-width of $G$, denoted by $\text{rw}(G)$, is the minimum width over all rank-decompositions of $G$. If $|V|<2$, we define $\text{rw}(G)=0$.
A.13.1 Hierarchy for rank-width
Theorem A.27.
The relationships between rank-width and other graph parameters are as follows:
  • $\text{rw}(G)\leq\text{tw}(G)+1\leq\text{pw}(G)+1$ (tw: tree-width and pw: path-width) [661].
  • Neighborhood Diversity is greater than or equal to rank-width [314].
Proof A.28.
Please refer to each reference.
A.13.2 Related concept for rank-width
Related concept is following:
  • Bi-rank-width [435]: A graph width parameter that discusses possible extensions of the notion of rank-width to colored graphs.
  • Signed-rank-width [312, 327, 318, 322]: Rank-width used for model counting.
  • $\mathbb{Q}$-Rank-Width [408]: A modified width parameter that utilizes the rank function over the rational field, as opposed to the binary field traditionally used. The $\mathbb{Q}$-rank-width does not exceed its corresponding clique-width. Furthermore, the $\mathbb{Q}$-rank-width is consistently less than or equal to the rank-width.
And Linear Rank-width is the linear restriction of Rank-width. If a graph G has bounded Linear Rank-width, then it also has bounded Rank-width (trivially).
A.14 Clique-width
Clique-width [406] measures the complexity of a graph based on a composition mechanism using vertex labels. Note that a clique in a graph is a subset of vertices where every pair of distinct vertices is connected by an edge, forming a complete subgraph (cf.[652]).
Definition A.29 (Clique-width).
Let $k$ be a positive integer. A $k$-graph $(G,\text{lab})$ consists of a graph $G$ and a labeling function $\text{lab}:V(G)\rightarrow\{1,2,\ldots,k\}$, which assigns one of $k$ labels to each vertex in $G$. In this context, all graphs are finite, undirected, and do not contain loops or parallel edges. The label of a vertex $v$ is denoted by $\text{lab}(v)$.
We define the following operations on $k$-graphs:
  1. 1.
    For $i\in\{1,\ldots,k\}$, $\cdot_{i}$ denotes an isolated vertex labeled by $i$.
  2. 2.
    For $i,j\in\{1,2,\ldots,k\}$ with $i\neq j$, the unary operator $\eta_{i,j}$ adds edges between all vertices labeled $i$ and all vertices labeled $j$. Formally, if $(G,\text{lab})$ is a $k$-graph, then $\eta_{i,j}(G,\text{lab})=(G^{\prime},\text{lab})$, where $V(G^{\prime})=V(G)$ and $E(G^{\prime})=E(G)\cup\{vw:\text{lab}(v)=i,\text{lab}(w)=j\}$.
  3. 3.
    The unary operator $\rho_{i\rightarrow j}$ relabels all vertices labeled $i$ to $j$. Formally, if $(G,\text{lab})$ is a $k$-graph, then $\rho_{i\rightarrow j}(G,\text{lab})=(G,\text{lab}^{\prime})$, where
    $\text{lab}^{\prime}(v)=\begin{cases}j&\text{if }\text{lab}(v)=i,\\ \text{lab}(v)&\text{otherwise}.\end{cases}$
  4. 4.
    The binary operation $\oplus$ denotes the disjoint union of two $k$-graphs. If $G_{1}$ and $G_{2}$ are two graphs, then $G_{1}\oplus G_{2}$ is their disjoint union.
A well-formed expression $t$ using these operations is called a $k$-expression. The value $\text{val}(t)$ of a $k$-expression $t$ is the $k$-graph produced by performing the operations in $t$. If $\text{val}(t)=(G,\text{lab})$, we say that $t$ is a $k$-expression of $G$.
The clique-width of a graph $G$, denoted by $\text{cwd}(G)$, is the minimum $k$ such that there exists a $k$-expression for $G$.
A.14.1 Hierarchy for clique-width
Theorem A.30.
The relationships between clique-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Clique-width, then it also has bounded twin-width [309].
[…]
  • Neighborhood Diversity is greater than or equal to clique-width [314].
  • If a graph $G$ has bounded Twin-cover, then it also has bounded Rank-width/Clique-width [313].
  • For any graph $G$, For any graph $G$, $\text{sd}(G)\leq 2\text{clique-width}(G)-2$. (sd: symmetric difference). [20]
  • For any graph $G$, $\text{fun}(G)\leq 2\text{clique-width}(G)-1$. (fun: functionality). [20, 224]
Proof A.31.
Please refer to each reference.
A.14.2 Related concepts for clique-width
Related concepts are following:
  • H-Clique-Width [398]: A graph width parameter which aims to bridge the gap between classical hereditary width measures and the recently introduced graph product structure theory [380].
  • Signed Clique-width [651]: Clique-width used for model counting.
  • Multi-Clique-Width (MCW) [307]: A graph width parameter that provides a natural extension of tree-width. For every graph $G$, it holds that $\text{mcw}(G)\leq\text{tw}(G)+2$ [307] (tw:Tree-width), $\text{boolw}(G)\leq\text{mcw}(G)\leq 2^{\text{boolw}(G)}$ [307] (boolw:Boolean-width), and $\text{mcw}(G)\leq\text{cw}(G)$ [307] (cw:Clique-width).
  • Clique-Partitioned Treewidth [331]: A parameter for graphs that is similar to tree-width but additionally considers the structure (clique) of each bag.
A.15 Linear-Clique-width
Linear Clique-width is the linear restriction of Clique-width.
A.15.1 Hierarchy for Linear-Clique-width
The relationships between Linear Clique-width and other graph parameters are as follows:
  • If a graph $G$ has bounded linear clique-width, then it also has bounded clique-width (trivially).
  • If a graph $G$ has bounded shrub-depth, then it also has bounded linear clique-width [412, 413].
  • If a graph $G$ has bounded linear clique-width, then it also has bounded clique-tree-width [369].
A.16 NLC-width
NLC-width and NLCT-width [369, 689] measure the complexity of a graph based on a composition mechanism similar to clique-width.
Definition A.32.
[369, 689] Let $k$ be a positive integer. The class $\text{NLC}_{k}$ of labeled graphs is defined recursively as follows:
  1. 1.
    A single vertex graph labeled $a$, where $a\in[k]$, belongs to $\text{NLC}_{k}$.
  2. 2.
    Let $G=(V_{G},E_{G},\text{lab}_{G})$ and $J=(V_{J},E_{J},\text{lab}_{J})$ be two vertex-disjoint labeled graphs in $\text{NLC}_{k}$, and let $S\subseteq[k]\times[k]$ be a relation. The graph $G\times_{S}J$ is defined as $(V^{\prime},E^{\prime},\text{lab}^{\prime})$, where:
    $V^{\prime}:=V_{G}\cup V_{J},$
    $E^{\prime}:=E_{G}\cup E_{J}\cup\{\{u,v\}\mid u\in V_{G},v\in V_{J},(\text{lab}_{G}(u),\text{lab}_{J}(v))\in S\},$
    and the labeling function $\text{lab}^{\prime}$ is given by:
    $\text{lab}^{\prime}(u):=\begin{cases}\text{lab}_{G}(u)&\text{if }u\in V_{G},\\ \text{lab}_{J}(u)&\text{if }u\in V_{J}.\end{cases}$
    Then, $G\times_{S}J$ is in $\text{NLC}_{k}$.
  3. 3.
    Let $G=(V_{G},E_{G},\text{lab}_{G})$ be a labeled graph in $\text{NLC}_{k}$, and let $R:[k]\to[k]$ be a function. The graph $\circ R(G)$ is defined as $(V_{G},E_{G},\text{lab}^{\prime})$, where the labeling function $\text{lab}^{\prime}$ is given by:
    $\text{lab}^{\prime}(u):=R(\text{lab}_{G}(u))\quad\text{for all }u\in V_{G}.$
    Then, $\circ R(G)$ is in $\text{NLC}_{k}$.
The NLC-width of a labeled graph $G$ is the smallest integer $k$ such that $G\in\text{NLC}_{k}$.
A.16.1 Hierarchy for NLC-width
We consider about hierarchy for NLC-width.
Theorem A.33.
The relationships between NLC-width and other graph parameters are as follows:
[…]
Proof A.34.
Please refer to each reference.
A.17 Linear NLC-width
Linear NLC-width [369] is the linear restriction of NLC-width.
A.17.1 Hierarchy for Linear-NLC-width
We consider about Hierarchy for Linear-NLC-width.
Theorem A.35.
The relationships between Linear NLC-width and other graph parameters are as follows:
  • If a graph $G$ has bounded linear NLC-width, then it also has bounded NLC-width (trivially).
  • If a graph $G$ has bounded linear NLC-width, then it also has bounded NLCT-width [369].
[…]
Proof A.36.
Please refer to each reference.
A.18 Hypertree-width
Hypertree-width [11] measures the complexity of hypergraphs by extending tree-width concepts to cover hyperedges.
First, we introduce the concept of a hypergraph [122]. Hypergraphs have numerous real-world applications across various fields. For readers interested in more detailed information or exploring these applications, we recommend consulting lecture notes, surveys, or books on the subject (e.g., [122, 254, 450, 27, 547, 121, 57]).
Definition A.37 (Hypergraph).
[122] A hypergraph $H=(V,E)$ consists of a finite set $V$ of vertices and a finite family $E=\{e_{i}\}_{i\in I}$ of non-empty subsets of $V$, called hyperedges, where $I$ is a finite index set. The set $V$ is also denoted by $V(H)$, and the family $E$ is denoted by $E(H)$.
The order of the hypergraph $H$, denoted by $|V|$, is the cardinality of the vertex set $V$. The size of the hypergraph, denoted by $|E|$, is the cardinality of the hyperedge family $E$.
  • The empty hypergraph is defined as the hypergraph where:
    • $V=\emptyset$,
    • $E=\emptyset$.
  • A trivial hypergraph is defined as a hypergraph where:
    • $V\neq\emptyset$,
    • $E=\emptyset$.
Next, we introduce the concept of a hypertree-width.
Definition A.38.
[11] A hypertree decomposition of a hypergraph $H=(V,E)$ is a triple $D=(T,\chi,\lambda)$, where:
  • $T$ is a tree,
  • $\chi$ is a function that assigns each tree node $t\in V(T)$ a set $\chi(t)\subseteq V$, known as a bag,
  • $\lambda$ is a function that maps each tree node $t\in V(T)$ to a set $\lambda(t)\subseteq E$, known as an edge cover of the bag $\chi(t)$.
The decomposition must satisfy the following conditions:
  1. 1.
    Edge Cover Condition (P1): For every hyperedge $e\in E(H)$, there exists a tree node $t\in V(T)$ such that $e\subseteq\chi(t)$.
  2. 2.
    Connectedness Condition (P2): For every vertex $v\in V(H)$, the set of tree nodes $t\in V(T)$ with $v\in\chi(t)$ induces a connected subtree of $T$.
  3. 3.
    Special Condition (P3): For any two nodes $t,t^{\prime}\in V(T)$ where $t^{\prime}$ is a descendant of $t$ in $T$, and for every hyperedge $e\in\lambda(t)$, it holds that $(e\setminus\chi(t))\cap\chi(t^{\prime})=\emptyset$.
The width of a hypertree decomposition $D=(T,\chi,\lambda)$ is defined as the maximum size of any edge cover $\lambda(t)$ over all tree nodes $t\in V(T)$. The hypertree width $\text{htw}(H)$ of the hypergraph $H$ is the minimum width over all possible hypertree decompositions of $H$.
The known obstructions are Hypertangles [12] and Hyperbrambles [12]. In the future, we will consider a hyperultrafilter analogous to [304] (cf.[525]).
A.18.1 Related concepts for Hypertree-width
Related concepts are following:
  • Hyperbranch width [12]: A generalized graph width parameter that extends the concept of hypertree-width.
  • Hyper-T-width and Hyper-D-width [617]: These graph width parameters were introduced as the first stable connectivity measures for hypergraphs, offering tools for designing algorithms to address various problems in hypergraph theory.
  • Fractional hypertree-width [528]: A hypergraph measure that is analogous to tree-width and hypertree-width.
  • Hyperpath-width and Hypertree-depth [10]: path-width and tree-depth in hypergraph.
  • Closure tree-width and Closure hypertree-width [9] : This is a width parameter used for applications in databases and similar fields.
A.18.2 Applications for Hypertree-width
We consider about Applications for Hypertree-width.
  • Artificial intelligence
    Similar to tree-width, the concept of hypertree-width is applied in the techniques and concepts of artificial intelligence that utilize graphs[228, 573].
  • Database
    The relationship between queries and hypertree-width has been particularly studied in the context of databases[334, 671, 343, 349, 422].
A.19 Modular-width
Modular-width [2] is defined using modular decompositions of the graph.
Definition A.39.
[2] A graph $G$ is said to have a tree-model with $m$ colors and depth $d\geq 1$ if there exists a rooted tree $T$ of height $d$ such that:
  1. 1.
    The set of leaves of $T$ is exactly $V(G)$.
  2. 2.
    Every root-to-leaf path in $T$ has length exactly $d$.
  3. 3.
    Each leaf of $T$ is assigned one of $m$ colors (this is not a graph coloring in the traditional sense).
  4. 4.
    The existence of an edge between $u,v\in V(G)$ depends only on the colors of $u$ and $v$, and the distance between $u$ and $v$ in $T$.
The class of all graphs that have a tree-model with $m$ colors and depth $d$ is denoted by $\text{TM}_{m}(d)$.
Definition A.40.
[2] A class of graphs $\mathcal{G}$ has shrub-depth $d$ if there exists some $m$ such that $\mathcal{G}\subseteq\text{TM}_{m}(d)$, and for every natural number $m$, we have $\mathcal{G}\nsubseteq\text{TM}_{m}(d-1)$.
A.19.1 Hierarchy for Modular-width
We consider about hierarchy for Modular-width.
Theorem A.41.
The relationships between modular-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Modular-width, then it also has bounded Clique-width [308].
  • If a graph $G$ has bounded vertex cover or neighborhood diversity, then it also has bounded Modular-width [308].  [^11] Neighborhood diversity is a graph parameter that measures the number of distinct vertex neighborhoods in a graph, where vertices with identical neighborhoods are grouped together(cf.[465, 315]).
  • If a graph $G$ has bounded twin-cover, then it also has bounded Modular-width [308].  [^12] The twin-cover of a graph is the smallest set of vertices such that every edge is either covered by this set or connects two vertices with identical neighborhoods[308].
Proof A.42.
Please refer to each reference.
A.20 Submodular-width
Submodular-width [344] is defined using the submodular condition property.
A.21 Amalgam-width
Amalgam-width [517] is a new matroid width parameter based on matroid amalgamation. Note that matroid amalgamation is a process of combining two matroids on overlapping ground sets into a single matroid, preserving their independent sets while maintaining matroid properties, particularly submodularity (cf.[591, 558]).
A.22 Kelly-width
Kelly-width [535, 458] is a parameter for directed graphs, analogous to tree-width. Related concepts include Kelly-path-width [455].
Definition A.43.
[535] A Kelly-decomposition of a directed graph $G$ is a triple $D:=(D,(B_{t})_{t\in V(D)},(W_{t})_{t\in V(D)})$ that satisfies the following conditions:
  • $D$ is a directed acyclic graph (DAG) and $(B_{t})_{t\in V(D)}$ is a partition of $V(G)$.
  • For all $t\in V(D)$, $W_{t}\subseteq V(G)$ is a set of vertices that guards $B_{t}^{\downarrow}:=\bigcup_{t^{\prime}\leq_{D}t}B_{t^{\prime}}$, where $\leq_{D}$ denotes the partial order induced by $D$.
  • For each node $s\in V(D)$, there exists a linear order on its children $t_{1},\ldots,t_{p}$ such that for all $1\leq i\leq p$, $W_{t_{i}}\subseteq B_{s}\cup W_{s}\cup\bigcup_{j<i}B_{t_{j}}^{\downarrow}$. Similarly, there is a linear order on the roots such that $W_{r_{i}}\subseteq\bigcup_{j<i}B_{r_{j}}^{\downarrow}$.
The width of $D$ is defined as $\max\{|B_{t}\cup W_{t}|:t\in V(D)\}$. The Kelly-width of $G$ is the minimum width over all possible Kelly-decompositions of $G$.
A.22.1 Hierarchy for Kelly-width
We consider about hierarchy for Kelly-width.
Theorem A.44.
The relationships between Kelly-width and other graph parameters are as follows:
  • A directed graph $G$ has directed elimination width $\leq k$ if and only if $G$ has Kelly-width $\leq k+1$.[535]
  • The Kelly-width is less than or equal to the Directed Path-width plus one [320].
  • If a digraph $G$ has bounded Kelly-width, then it also has bounded Directed Tree-width [62, 25].
  • If a graph has bounded Kelly-path-width, then it has also bounded Kelly-width (trivially).
Proof A.45.
Please refer to each reference.
A.23 Monoidal Width
Monoidal Width [493] measures the complexity of morphisms in monoidal categories. Note that monoidal categories are categories equipped with a binary operation (tensor product) and an identity object, allowing the composition of objects and morphisms, similar to how multiplication works in algebra (cf.[672])  [^13] The tensor product of two vector spaces combines them into a new vector space, where each element is a pair from the original spaces, allowing multilinear maps to be expressed as linear maps(cf.[157, 640])..
A.24 Tree-cut Width
Tree-cut width [324] is a graph width parameter that serves as an analogue to tree-width, specifically for edge cuts. This concept is accompanied by related width parameters, including 0-Tree-cut width [325].
Definition A.46.
[324] A tree-cut decomposition of a graph $G$ is a pair $(T,\mathcal{X})$, where $T$ is a tree and $\mathcal{X}=\{X_{t}\subseteq V(G):t\in V(T)\}$ is a near-partition of $V(G)$. Each set $X_{t}$ is called a bag.
For an edge $e=(u,v)$ in $T$, let $T_{u}$ and $T_{v}$ be the two connected components of $T-e$ containing $u$ and $v$, respectively. The corresponding partition of $V(G)$ is $(\bigcup_{t\in T_{u}}X_{t},\bigcup_{t\in T_{v}}X_{t})$, and the set of edges with one endpoint in each part is denoted by $\text{cut}(e)$.
A tree-cut decomposition is rooted if one node $r\in V(T)$ is designated as the root. For any $t\in V(T)\setminus\{r\}$, let $e(t)$ be the unique edge on the path from $r$ to $t$. The adhesion of $t$, denoted $\text{adh}(t)$, is defined as $|\text{cut}(e(t))|$. For the root $r$, set $\text{adh}(r)=0$.
The torso of a tree-cut decomposition $(T,\mathcal{X})$ at a node $t$ is the graph $H_{t}$ obtained as follows: if $T$ has a single node, then $H_{t}=G$. Otherwise, for each component $T_{i}$ of $T-t$, define the vertex set $Z_{i}=\bigcup_{b\in V(T_{i})}X_{b}$. The torso $H_{t}$ is formed by contracting each $Z_{i}$ into a single vertex $z_{i}$, adding an edge between $z_{i}$ and any vertex $v\in V(G)\setminus Z_{i}$ that was adjacent to a vertex in $Z_{i}$. This may create parallel edges.
The 3-center of $(G,X)$ is obtained by repeatedly suppressing vertices in $V(G)\setminus X$ with degree at most 2. For a node $t$, the 3-center of its torso $H_{t}$ with respect to $X_{t}$ is denoted $\widetilde{H}_{t}$, and the torso-size is defined as $\text{tor}(t)=|\widetilde{H}_{t}|$.
The width of a tree-cut decomposition $(T,\mathcal{X})$ is defined as $\max_{t\in V(T)}\{\text{adh}(t),\text{tor}(t)\}$. The tree-cut width of $G$, denoted $\text{tcw}(G)$, is the minimum width over all tree-cut decompositions of $G$.
A.24.1 Related Concepts for tree-cut width
Several related concepts further explore the idea of tree-cut width:
  • Slim Tree-cut Width [325]: Slim tree-cut width is a graph width parameter that meets all structural and algorithmic criteria for an edge-cut-based analogue of tree-width, while being less restrictive than edge-cut width.
    • If a graph $G$ has bounded slim tree-cut width, then it also has bounded tree-cut width [325].
  • Edge-crossing Width and $\alpha$-Edge-crossing Width [153, 494]: These parameters are defined by the number of edges that cross a bag in a tree-cut decomposition.
  • Edge-cut Width [117]: Edge-cut width is an algorithmically driven analogue of tree-width that focuses on edge cuts.
    • If a graph $G$ has bounded feedback edge set number, then it also has bounded edge-cut width [153].
  • Edge-Tangle [291]: Edge-Tangle is another related concept that further refines the understanding of edge cuts in graph decompositions.
A.25 Resolution width
Resolution width [36] measures proof complexity using the existential pebble game in finite model theory.
Related concept is Linear Resolution width. This width is linear layout of Resolution width.
A.26 Twin-width
Twin-width [729, 733, 731, 732, 730, 103] is a graph width parameter that measures the complexity of graphs by evaluating their ability to avoid a fixed minor. This includes classes such as planar graphs and graphs with bounded tree-width or clique-width.
Definition A.47.
Let $G=(V,E,R)$ be a trigraph, where $V$ is the vertex set, $E$ is the set of black edges, and $R$ is the set of red edges. For any two vertices $u,v\in V$, the contraction of $u$ and $v$ in $G$, denoted $G/u,v=(V^{\prime},E^{\prime},R^{\prime})$, is defined as follows:
  • The vertex set is $V^{\prime}=(V\setminus\{u,v\})\cup\{w\}$, where $w$ is the new vertex representing the contracted pair $u,v$.
  • The edges of $G/u,v$ are defined such that:
    $wx\in E^{\prime}\quad\text{if and only if}\quad ux\in E\ \text{and}\ vx\in E,$
    $wx\notin E^{\prime}\cup R^{\prime}\quad\text{if and only if}\quad ux\notin E\cup R\ \text{and}\ vx\notin E\cup R,$
    $wx\in R^{\prime}\quad\text{otherwise}.$
In other words, contracting $u$ and $v$ results in red edges for any vertex $x$ that is not consistently adjacent to both $u$ and $v$. The trigraph $G/u,v$ is called a d-contraction if both $G$ and $G/u,v$ are $d$-trigraphs.
A trigraph $G$ on $n$ vertices is said to be d-collapsible if there exists a sequence of $d$-contractions that reduces $G$ to a single vertex. Formally, this means there is a sequence of trigraphs $G=G_{n},G_{n-1},\dots,G_{2},G_{1}$, where each $G_{i-1}$ is a contraction of $G_{i}$, and $G_{1}$ is a single vertex graph.
The minimum $d$ for which $G$ is $d$-collapsible is called the twin-width of $G$, denoted by $\text{tww}(G)$.
A.26.1 Hierarchy for Twin-width
We consider about hierarchy for Twin-width.
Theorem A.48.
The relationships between Twin-width and other graph parameters are as follows:
  • Every graph class of bounded Genus has bounded Twin-width [668]. [^14] The genus of a graph is the minimum number of ”holes” in a surface required to embed the graph without edge crossings. It quantifies the graph’s topological complexity (cf.[712, 636]).
  • Every graph of bounded rank-width, also has bounded twin-width[105].
  • Every graph of bounded stack number or bounded queue number also has bounded twin-width[104]
  • Every graph class of bounded Distance to Planar, Sparse-Twin-width, or Feedback Edge Set has bounded Twin-width [668]  [^15] The Feedback Edge Set of a graph is the smallest set of edges whose removal makes the graph acyclic, eliminating all cycles (cf.[15, 376])..
  • If a graph $G$ has bounded Clique-width, then it also has bounded Twin-width [309].
  • If a graph $G$ has bounded Twin-width, then it also has bounded monadically dependent [309].
Proof A.49.
Please refer to each reference.
A.26.2 Related concepts for Twin-width
Related concepts for Twin-width are the following:
  • Sparse Twin-width [310, 103]: A graph width parameter that generalizes results for classes with bounded linear clique-width and clique-width. If a graph $G$ has bounded Sparse Twin-width, then it also has bounded nowhere dense [309].
  • Signed Twin-width [319]: A graph width parameter specifically applied to CNF formulas, demonstrating that BWMC (Bounded Width Monotone Circuits) is fixed-parameter tractable.
A.27 Decomposition width
Decomposition width [482] is related to matroid theory and second-order logic.
A.28 Minor-matching hypertree width
Minor-matching hypertree width [718] measures the complexity of a tree decomposition for graphs and hypergraphs, ensuring polynomially-sized independent sets in each decomposition bag.
A.29 Tree-clique width
Tree-clique width [31] extends the algorithmic benefits of tree-width to more structured and dense graphs.
A.30 Tree-distance-width
Tree distance width (TDW) [715] measures the width from a distance perspective, restricted by tree-width. Related concepts are Rooted Tree-distance-width (RTDW) [715] and Connected Tree-distance-width [506, 575].
Definition A.50.
[715] A tree distance decomposition of a graph $G=(V,E)$ is a triple ($(\{X_{i}\}_{i\in I},T=(I,F),r)$ where:
  • $\bigcup_{i\in I}X_{i}=V(G)$ and $X_{i}\cap X_{j}=\emptyset$ for all $i\neq j$.
  • For each vertex $v\in V$, if $v\in X_{i}$, then the distance from $v$ to the root set $X_{r}$ in $G$ is equal to the distance from $i$ to the root $r$ in the tree $T$.
  • For each edge $\{v,w\}\in E$, there exist $i,j\in I$ such that $v\in X_{i}$, $w\in X_{j}$, and either $i=j$ or $\{i,j\}\in F$.
The node $r$ is called the root of the tree $T$, and $X_{r}$ is called the root set of the tree distance decomposition.
The width of a tree distance decomposition ($(\{X_{i}\}_{i\in I},T,r)$ is defined as $\max_{i\in I}|X_{i}|$. The tree distance width of a graph $G$ is the minimum width over all possible tree distance decompositions of $G$, denoted by $\text{TDW}(G)$.
A rooted tree distance decomposition of $G$ is a tree distance decomposition where the root set $X_{r}$ contains exactly one vertex (i.e., $|X_{r}|=1$). The rooted tree distance width of a graph $G$ is the minimum width over all rooted tree distance decompositions of $G$, denoted by $\text{RTDW}(G)$.
A.30.1 Hierarchy for Tree-distance-width
We consider about Hierarchy for Tree-distance-width
Theorem A.51.
The relationships between Tree-distance-width and other graph parameters are as follows:
[…]
  • $\text{TDW}(G)\leq\text{RTDW}(G)$  (RTDW: Rooted Tree-distance width).
  • $\text{TDW}(G)\leq\text{PDW}(G)\leq\text{RPDW}(G)$  (PDW: Path-distance-width and RPDW: Rooted Path Distance width).
Proof A.52.
Please refer to each reference.
A.31 Path-distance-width
Path distance width (PDW) [715] measures the width from a distance perspective, constrained by bandwidth. Related concepts include Rooted Path Distance width (RPDW) [715], Connected Path-distance-width, and c-connected Path-distance-width [506, 575]. Distance width is particularly useful in problems like Graph Isomorphism, where the goal is to determine if two graphs are identical by checking if there is a one-to-one correspondence between their vertices and edges that preserves the connections (cf.[606, 355, 499]).
Definition A.53.
[715] A (rooted) path distance decomposition of a graph $G=(V,E)$ is defined similarly to the (rooted) tree distance decomposition and (rooted) tree distance width, but with the requirement that the tree $T$ is a path.
For simplicity, a (rooted) path distance decomposition is denoted by $(X_{1},X_{2},\dots,X_{t})$, where $X_{1}$ is the root set of the decomposition.
The corresponding graph parameters are the path distance width (PDW) and rooted path distance width (RPDW).
A.31.1 Hierarchy for Path-distance-width
We consider about Hierarchy for Path-distance-width.
Theorem A.54.
The relationships between Path-distance-width and other graph parameters are as follows:
  • $\text{bandwidth}(G)\leq 2\times\text{Path-distance-width}(G)-1$
  • If a graph $G$ has bounded Path distance width, then it also has bounded Tree distance width (trivially).
  • If a graph $G$ has bounded Rooted Path distance width, then it also has bounded Rooted Tree distance width (trivially).
Proof A.55.
Please refer to each reference.
A.32 Tree-partition-width
Tree-partition-width [701] is the minimum width of a tree-partition of a graph. This width is also called 0-quasi-tree-partitions-width [504].
Definition A.56.
[701] Let $G$ be a graph. A graph $H$ is called a partition of $G$ if the following conditions hold:
  1. 1.
    Each vertex of $H$ is a subset of the vertices of $G$, referred to as a bag.
  2. 2.
    Every vertex of $G$ belongs to exactly one bag in $H$.
  3. 3.
    Two bags $A$ and $B$ in $H$ are adjacent if and only if there is an edge in $G$ with one endpoint in $A$ and the other endpoint in $B$.
The width of a partition is defined as the maximum number of vertices in any bag.
Informally, the graph $H$ is obtained from a proper partition of $V(G)$ by identifying the vertices in each bag, deleting loops, and replacing parallel edges with a single edge.
If a forest $T$ is a partition of a graph $G$, then $T$ is called a tree-partition of $G$. The tree-partition width of $G$, denoted by $\text{tpw}(G)$, is the minimum width of a tree-partition of $G$.
A.32.1 Hierarchy for Tree-partition-width
We consider about Hierarchy for Tree-partition-width .
Theorem A.57.
The relationships between Tree-partition-width and other graph parameters are as follows:
[…]
  • 2×Tree partition $12\times\text{Tree partition width}(G)\geq\text{Tree-width}(G)+$ 1 [701].
Proof A.58.
Please refer to each reference.
Related concepts are Directed Tree partition width [414], which considers Tree partition width in Digraphs.
A.33 Path-partition-width
Path-partition-width is the minimum width of a path-partition of a graph. This width is also called 0-quasi-path-partitions-width [504]. Related concepts are Directed Path partition width [414], which considers Path partition width in Digraphs.
A.33.1 Hierarchy for Path-partition-width
We consider about Hierarchy for Path-partition-width.
Theorem A.59.
  • For any graph $G$: $\frac{1}{2}(\text{bw}(G)+1)\leq\text{ppw}(G)\leq\text{bw}(G).$ ($bw$:Band-width)[218]
  • If a graph $G$ has bounded Path partition width, then it also has bounded Tree partition width (trivially).
Proof A.60.
Please refer to each reference.
A.34 CV-width
CV-width [580] is specific to CNFs and dominates the tree-width of the CNF incidence graph. CNFs, or Conjunctive Normal Forms, are Boolean formulas expressed as a conjunction (AND) of clauses, where each clause is a disjunction (OR). And linear CV width is the linear restriction of CV width.
A.35 Dominating-set-width
Dominating-set-width [536] measures the number of different dominating sets on a subgraph.
A.36 Point width
Point width [138] for hypergraphs provides a condition for the tractability of Max-CSPs, generalizing bounded MIM-width and $\beta$-acyclicity.
A.37 Neighbourhood width
Neighbourhood width [363, 361] considers all neighborhoods of the vertices. Neighbourhood width is less than path-width plus one[363, 361].
Definition A.61.
[363, 361] A linear layout for a graph $G=(V,E)$ is a bijection $\varphi:V\rightarrow\{1,\dots,|V|\}$. We denote the set of all linear layouts for $G$ by $\Phi(G)$. For a given $\varphi\in\Phi(G)$ and $1\leq i\leq|V|$, we define the left and right sets as follows:
$L(i,\varphi,G)=\{u\in V\mid\varphi(u)\leq i\}\quad\text{and}\quad R(i,\varphi,G)=\{u\in V\mid\varphi(u)>i\}.$
Let $G=(V,E)$ be a graph, and let $U,W\subseteq V$ be two disjoint vertex sets. The restricted neighborhood of a vertex $u$ in $W$ is denoted by $N_{W}(u)=\{v\in W\mid\{u,v\}\in E\}$. The set of all neighborhoods of the vertices in $U$ with respect to $W$ is denoted by $N(U,W)=\{N_{W}(u)\mid u\in U\}$.
This allows us to define the neighborhood-width of a graph $G$, denoted by $\text{nw}(G)$, as follows:
$\text{nw}(G)=\min_{\varphi\in\Phi(G)}\max_{1\leq i\leq|V|-1}|N(L(i,\varphi,G),R(i,\varphi,G))|.$
A.38 Fusion-width
Fusion-width [306] generalizes tree-width and encompasses graphs of bounded tree-width and clique-width.
Definition A.62.
[306] A k-fusion-tree expression is an expression constructed from the following elements:
  • $i^{(m)}$, which creates a graph consisting of $m$ isolated vertices labeled $i$, where $i$ is an element of the set $\{1,\dots,k\}$.
  • $\eta_{i,j}$, a unary operation that creates an edge between every vertex labeled $i$ and every vertex labeled $j$ for $i\neq j$.
  • $\rho_{i\rightarrow j}$, a unary operation that changes all labels $i$ to $j$.
  • $\theta_{i}$, a unary operation that merges all vertices labeled $i$ into a single vertex. The new vertex is labeled $i$ and is adjacent to every vertex not labeled $i$ that was adjacent to any vertex labeled $i$ before the operation.
  • $\oplus$, a binary operation that forms the disjoint union of two graphs.
Finally, the generated graph is obtained by deleting all labels.
Definition A.63.
[306] The fusion-width $\text{fw}(G)$ of a graph $G$ is the smallest $k$ such that there exists a $k$-fusion-tree expression that generates $G$.
A.38.1 Hierarchy for Fusion-width
We consider about Hierarchy for Fusion-width.
Theorem A.64.
The relationships between Fusion-width and other graph parameters are as follows:
  • Fusion-width(G) $\leq$ Tree-width(G) + 2 [306],
  • Fusion-width(G) $\leq$ Clique-width(G) [306].
A.39 Directed NLC width
Directed NLC width [370] is NLC width on directed graphs.
Definition A.65 (Directed NLC-width).
[370] Let $k$ be a positive integer. And the new operation $\otimes(\overrightarrow{S},\overleftarrow{S})$ utilizes two relations $\overrightarrow{S}$ and $\overleftarrow{S}$ for label pairs, where $\overrightarrow{S}$ defines the arcs directed from left to right and $\overleftarrow{S}$ defines the arcs directed from right to left. The class $\text{dNLC}_{k}$ of labeled digraphs is defined recursively as follows:
  1. 1.
    The single vertex digraph $\bullet_{a}$ for some $a\in[k]$ is in $\text{dNLC}_{k}$.
  2. 2.
    Let $G=(V_{G},E_{G},\text{lab}_{G})\in\text{dNLC}_{k}$ and $J=(V_{J},E_{J},\text{lab}_{J})\in\text{dNLC}_{k}$ be two vertex-disjoint labeled digraphs, and let $\overrightarrow{S},\overleftarrow{S}\subseteq[k]^{2}$ be two relations. Then the operation $G\otimes(\overrightarrow{S},\overleftarrow{S})J:=(V^{\prime},E^{\prime},\text{lab}^{\prime})$ is defined by:
    $V^{\prime}:=V_{G}\cup V_{J},$
    $E^{\prime}:=$ $\,E_{G}\cup E_{J}\cup\{(u,v)\mid u\in V_{G},v\in V_{J},(\text{lab}_{G}(u),\text{lab}_{J}(v))\in\overrightarrow{S}\}$
    $\,\cup\{(v,u)\mid u\in V_{G},v\in V_{J},(\text{lab}_{G}(u),\text{lab}_{J}(v))\in\overleftarrow{S}\},$
    $\text{lab}^{\prime}(u):=\begin{cases}\text{lab}_{G}(u)&\text{if }u\in V_{G},\\ \text{lab}_{J}(u)&\text{if }u\in V_{J},\end{cases}$
    for every $u\in V^{\prime}$, which results in a digraph in $\text{dNLC}_{k}$.
  3. 3.
    Let $G=(V_{G},E_{G},\text{lab}_{G})\in\text{dNLC}_{k}$ and $R:[k]\rightarrow[k]$ be a function. Then the operation $\circ_{R}(G):=(V_{G},E_{G},\text{lab}^{\prime})$ is defined by:
    $\text{lab}^{\prime}(u):=R(\text{lab}_{G}(u))$
    for every $u\in V_{G}$, resulting in a digraph in $\text{dNLC}_{k}$.
[…]
A.40 Directed Tree-width
Directed Tree-width [423, 8] is Tree-width on directed graphs. Examples of obstructions are directed tangle [336] and directed ultrafilter [299].
Definition A.66 (Directed Tree-width, [423, 8] ).
A (arboreal) tree-decomposition of a digraph $G=(V_{G},E_{G})$ is a triple $(T,X,W)$, where:
  • $T=(V_{T},E_{T})$ is an out-tree,
  • $X=\{X_{e}\mid e\in E_{T}\}$ and $W=\{W_{r}\mid r\in V_{T}\}$ are sets of subsets of $V_{G}$,
such that the following conditions hold:
  1. (dtw-1)
    $W=\{W_{r}\mid r\in V_{T}\}$ is a partition of $V_{G}$ into non-empty subsets.
  2. (dtw-2)
    For every $(u,v)\in E_{T}$, the set $\bigcup_{r\in V_{T},v\leq r}W_{r}$ is $X_{(u,v)}$-normal.
The width of an arboreal tree-decomposition $(T,X,W)$ is defined as:
$\max_{r\in V_{T}}\left\{|W_{r}\cup\bigcup_{e\sim r}X_{e}|\right\}-1,$
where $e\sim r$ means that $r$ is one of the endpoints of the arc $e$.
[…]
A.40.1 Hierarchy for Directed Tree-width
We consider about Hierarchy for Directed Tree-width .
Theorem A.67.
The relationships between Directed Tree-width and other graph parameters are as follows:
  • If a digraph $G$ has bounded DAG-width, then it also has bounded Directed Tree-width [62].
  • If a digraph $G$ has bounded Kelly-width, then it also has bounded Directed Tree-width [62, 25].
  • For every digraph $D$, it holds that $\text{dtw}(D)\geq\text{cycle-degeneracy}(D)$.
Proof A.68.
Please refer to each reference.
A.41 Directed path-width
Directed Path-width (d-pw) [48, 367] is Path-width on directed graphs.
Definition A.69 (Directed Path-width).
[48, 367] Let $G=(V,E)$ be a digraph. A directed path-decomposition of $G$ is a sequence $(X_{1},\dots,X_{r})$ of subsets of $V$, called bags, that satisfies the following conditions:
  1. 1.
    $X_{1}\cup\dots\cup X_{r}=V$,
  2. 2.
    For each edge $(u,v)\in E$, there exists a pair $i\leq j$ such that $u\in X_{i}$ and $v\in X_{j}$,
  3. 3.
    For all indices $1\leq i<j<\ell\leq r$, it holds that $X_{i}\cap X_{\ell}\subseteq X_{j}$.
The width of a directed path-decomposition $X=(X_{1},\dots,X_{r})$ is defined as:
$\max_{1\leq i\leq r}|X_{i}|-1.$
[…]
A.41.1 Hierarchy for Directed Path-width
We consider about Hierarchy for Directed Path-width .
Theorem A.70.
[281] Let $D$ be a directed graph. The Directed Path-width $\text{dpw}(D)$ has the following relationships with various other graph parameters:
  1. 1.
    Directed Vertex Separation Number: The Directed Path-width is equal to the Directed Vertex Separation Number [82].
    $\text{dpw}(D)=\text{Directed Vertex Separation Number}(D)$
  2. 2.
    Kelly-width: The Kelly-width is less than or equal to the Directed Path-width plus one [320].
[…]
  3. 3.
    Entanglement: The entanglement of $D$ is less than or equal to the Directed Path-width [598] (cf.[139]).
    $\text{entanglement}(D)\leq\text{dpw}(D)$
  4. 4.
    Cycle Rank: The Directed Path-width is less than or equal to the Cycle Rank of $D$ [357].
    $\text{dpw}(D)\leq\text{Cycle Rank}(D)$
  5. 5.
    Kelly Path-width: The Kelly Path-width of $D$ is equal to its Directed Path-width [456].
[…]
  6. 6.
    DAG-path-width: The Directed Path-width is less than or equal to the DAG-path-width [441].
    $\text{dpw}(D)\leq\text{DAG-path-width}(D)$
  7. 7.
    Diblockage: A directed graph has Directed Path-width $\geq k-1$ if and only if it has a diblockage of order $\geq k$ [243].
    $\text{dpw}(D)\geq k-1\iff\text{diblockage order}(D)\geq k$
  8. 8.
    Directed Path-Distance-Width: The Directed Path-width is bounded above by twice the Directed Path-Distance-Width minus one [281].
    $\text{dpw}(D)\leq 2\times\text{DPDW}(D)-1$
  9. 9.
    Directed Proper-Path-Width: The Directed Proper-Path-Width is bounded by the Directed Path-width and is at most one greater [281].
    $\text{dpw}(D)\leq\text{dppw}(D)\leq\text{dpw}(D)+1$
  10. 10.
    Directed Tree-width: The Directed Path-width is greater than or equal to the Directed Tree-width [365].
    $\text{dtw}(D)\leq\text{dpw}(D)$
  11. 11.
    Directed Cut-width: The Directed Path-width is less than or equal to the Directed Cut-width [365].
    $\text{dpw}(D)\leq\text{dcw}(D)$
  12. 12.
    Linear Width Parameters: For any directed graph $G$, the Directed Path-width is related to various linear width parameters as follows [365]:
    $\text{dpw}(G)\leq\min(\Delta^{-}(G),\Delta^{+}(G))\cdot\text{d-nw}(G),$
    $\text{dpw}(G)\leq\min(\Delta^{-}(G),\Delta^{+}(G))\cdot\text{d-lnlcw}(G),$
    $\text{dpw}(G)\leq\min(\Delta^{-}(G),\Delta^{+}(G))\cdot\text{d-lcw}(G),$
[…]
Proof A.71.
Please refer to each reference.
A.42 Directed Branch-width
Directed Branch-width [131] is Branch-width on directed graphs.
A.43 Directed Cut-width
[…]
Definition A.72.
[161] The directed cut-width of a digraph $G=(V,E)$ is defined as:
[…]
where $\Phi(G)$ denotes the set of all linear layouts of $G$, $L(i,\varphi,G)$ represents the set of vertices mapped to positions $\leq i$ by the linear layout $\varphi$, and $R(i,\varphi,G)$ represents the set of vertices mapped to positions $>i$.
A.44 Directed Clique-width
Directed Clique-width [171, 370] is clique width on directed graphs. And Directed Linear-Clique-width (d-lcw) [370] is linear-concept of clique width on directed graphs. We introduce about the definitions of directed clique-width and directed linear-clique-width.
Definition A.73.
[370] Directed Clique-width. Let $k$ be a positive integer. The class $\text{dCW}_{k}$ of labeled digraphs is defined recursively as follows:
  1. 1.
    A single vertex digraph $\bullet_{a}$ for some $a\in[k]$ is in $\text{dCW}_{k}$.
  2. 2.
    Let $G=(V_{G},E_{G},\text{lab}_{G})\in\text{dCW}_{k}$ and $J=(V_{J},E_{J},\text{lab}_{J})\in\text{dCW}_{k}$ be two vertex-disjoint labeled digraphs. Then the disjoint union $G\oplus J:=(V^{\prime},E^{\prime},\text{lab}^{\prime})$ defined by:
    $V^{\prime}:=V_{G}\cup V_{J},$
    $E^{\prime}:=E_{G}\cup E_{J},$
    $\text{lab}^{\prime}(u):=\begin{cases}\text{lab}_{G}(u)&\text{if }u\in V_{G},\\ \text{lab}_{J}(u)&\text{if }u\in V_{J},\end{cases}$
    for every $u\in V^{\prime}$, is in $\text{dCW}_{k}$.
  3. 3.
    Let $a,b\in[k]$ be two distinct integers, and let $G=(V_{G},E_{G},\text{lab}_{G})\in\text{dCW}_{k}$ be a labeled digraph.
    1. (a)
      The operation $\rho_{a\rightarrow b}(G):=(V_{G},E_{G},\text{lab}^{\prime})$ defined by:
      $\text{lab}^{\prime}(u):=\begin{cases}\text{lab}_{G}(u)&\text{if }\text{lab}_{G}(u)\neq a,\\ b&\text{if }\text{lab}_{G}(u)=a,\end{cases}$
      for every $u\in V_{G}$, results in a digraph in $\text{dCW}_{k}$.
    2. (b)
      The operation $\alpha_{a,b}(G):=(V_{G},E^{\prime},\text{lab}_{G})$ defined by:
      $E^{\prime}:=E_{G}\cup\{(u,v)\mid u,v\in V_{G},u\neq v,\text{lab}(u)=a,\text{lab}(v)=b\},$
      results in a digraph in $\text{dCW}_{k}$.
[…]
Definition A.74 (Directed Linear Clique-width).
[…]
  1. 1.
    Creation of a new vertex with label $a$ (denoted by $\bullet_{a}$).
  2. 2.
    Disjoint union of a labeled digraph $G$ and a single vertex labeled $a$ (denoted by $G\oplus\bullet_{a}$).
  3. 3.
    Inserting an arc from every vertex with label $a$ to every vertex with label $b$ (where $a\neq b$, denoted by $\alpha_{a,b}$).
  4. 4.
    Changing label $a$ to label $b$ (denoted by $\rho_{a\rightarrow b}$).
A.45 Directed Linear-NLC-width
Directed Linear NLC-width (d-lnlcw) [370, 364] is Linear NLC width on directed graphs.
A.45.1 Hierarchy for Linear-NLC-width
We consider about Hierarchy for Linear-NLC-width.
Theorem A.75.
The relationships between Directed Linear NLC-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Directed Linear NLC-width, then it also has bounded Directed NLC-width (trivially).
  • For every digraph $G$, we have d-lnlcw($G$) $\leq$ d-lcw($G$) $\leq$ d-lnlcw($G$) + 1 (Directed Linear-Clique-width) [364].
A.46 Directed Neighbourhood width
Directed Neighbourhood width (d-nw) [364] is Neighbourhood width on directed graphs.
Definition A.76.
[364] Let $G=(V,E)$ be a digraph, and let $U,W\subseteq V$ be two disjoint vertex sets. The set of all out-neighbors of $u$ into the set $W$ is defined by $N^{+}_{W}(u)=\{v\in W\mid(u,v)\in E\}$, and the set of all in-neighbors of $u$ from the set $W$ is defined by $N^{-}_{W}(u)=\{v\in W\mid(v,u)\in E\}$. The directed neighborhood of a vertex $u$ into the set $W$ is given by $N_{W}(u)=(N^{+}_{W}(u),N^{-}_{W}(u))$. The set of all directed neighborhoods of the vertices in the set $U$ into the set $W$ is denoted by $N(U,W)=\{N_{W}(u)\mid u\in U\}$.
For a layout $\varphi\in\Phi(G)$, we define
[…]
The directed neighborhood-width of a digraph $G$ is given by
[…]
A.46.1 Hierarchy for Neighbourhood width
We consider about Hierarchy for Neighbourhood width.
Theorem A.77.
For every digraph $G$, we have:
  • d-nw($G$) $\leq$ d-lnlcw($G$) $\leq$ d-nw($G$) + 1 (Directed Linear NLC-width) [364],
  • d-nw($G$) $\leq$ d-lcw($G$) $\leq$ d-nw($G$) + 1 (Directed Linear-Clique-width) [364].
Proof A.78.
Please refer to each reference.
A.47 Directed Rank-width
Directed Rank-width [436] is Rank-width on directed graphs.
A.48 Directed Linear-Rank-width
Directed Linear Rank-width (d-lrw) [364] is Linear rank width on directed graphs.
Definition A.79.
[364] Let $G=(V,E)$ be a digraph, and let $V_{1},V_{2}\subseteq V$ be a disjoint partition of the vertex set of $G$. Define the adjacency matrix $M^{V_{1}}_{V_{2}}=(m_{ij})$ over the four-element field $\mathbb{GF}(4)$ for the partition $V_{1}\cup V_{2}$, where
$m_{ij}=\begin{cases}0&\text{if }(v_{i},v_{j})\notin E\text{ and }(v_{j},v_{i})\notin E,\\ a&\text{if }(v_{i},v_{j})\in E\text{ and }(v_{j},v_{i})\notin E,\\ a^{2}&\text{if }(v_{i},v_{j})\notin E\text{ and }(v_{j},v_{i})\in E,\\ 1&\text{if }(v_{i},v_{j})\in E\text{ and }(v_{j},v_{i})\in E.\end{cases}$
In $\mathbb{GF}(4)$, the field elements are $\{0,1,a,a^{2}\}$, with the properties $01+a+a^{2}=$ 0 and $a^{3}=1$.
[…]
A.48.1 Hierarchy for Directed Linear-Rank-width
We consider about Hierarchy for Directed Linear-Rank-width.
Theorem A.80.
The relationships between Directed Linear Rank-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Directed Linear Rank-width, then it also has bounded Directed Rank-width (trivially).
  • For every digraph $G$, we have d-lrw($G$) $\leq$ d-nw($G$) [364] (d-nw: Directed Neighbourhood width).
Proof A.81.
Please refer to each reference.
A.49 Elimination width
Elimination width os a graph width parameter to address issues with acyclic digraphs [255].
A.50 Flip-width
The flip-width [101] parameters are defined using variants of the Cops and Robber game. Related concepts include radius-one flip-width [667].
A.50.1 Hierarchy for Flip-width
We consider about Hierarchy for Flip-width.
Theorem A.82.
The relationships between Flip-width (or radius-one flip-width) and other graph parameters are as follows:
  • Every class of bounded twin-width has bounded flip-width [667].
  • If a graph $G$ has bounded radius-one flip-width, then it also has bounded symmetric difference, functionality, and VC-dimension [667].
Proof A.83.
Please refer to each reference.
A.51 Stretch-width
Stretch-width [728] lies strictly between clique-width and twin-width.
A.51.1 Hierarchy for Stretch-width
We consider about Hierarchy for Stretch-width.
Theorem A.84.
The relationships between Stretch-width and other graph parameters are as follows:
  • The stretch-width of any graph is at most twice its clique-width.
  • There is a constant $c$ such that for every graph $G$, $\text{tw}(G)\leq c\Delta(G)^{4}\text{stw}(G)^{2}\log|V(G)|$. (Tree-width and maximal degree)
Proof A.85.
Please refer to each reference.
A.52 Cops-width
The Cop-width [101, 391, 1] parameters are defined using variants of the Cops and Robber game.
A.52.1 Hierarchy for cops-width
We consider about Hierarchy for cops-width.
Theorem A.86.
The relationships between (radius-)cop-width and other graph parameters are as follows:
[…]
Proof A.87.
Please refer to each reference.
A.52.2 Related concepts for cops-width
Related concepts are the following:
  • Marshal width and Monotone marshal width [6]: Graph width parameters related to a winning strategy in the Cops and Robber game.
  • Game-width [624, 239]: A graph width parameter of the Cops and Robber game. The game width of a graph equals its tree-width plus 1 [624].
A.53 Perfect matching width
Perfect matching width [337] is a width parameter for matching covered graphs based on a branch decomposition.
Definition A.88 (Matching-Porosity).
[386] Let $G$ be a matching covered graph and $X\subseteq V(G)$. We define the matching-porosity of $\partial(X)$ as follows:
$\text{mp}(\partial(X)):=\max_{M\in\mathcal{M}(G)}|M\cap\partial(X)|.$
A perfect matching $M\in\mathcal{M}(G)$ is maximal with respect to a cut $\partial(X)$ if there is no perfect matching $M^{\prime}\in\mathcal{M}(G)$ such that $\partial(X)\cap M\subset\partial(X)\cap M^{\prime}$.
A perfect matching $M\in\mathcal{M}(G)$ maximizes a cut $\partial(X)$ if $\text{mp}(\partial(X))=|M\cap\partial(X)|$.
Definition A.89 (Perfect Matching Width).
[386] Let $G$ be a matching covered graph. A perfect matching decomposition of $G$ is a tuple $(T,\delta)$, where $T$ is a cubic tree and $\delta:L(T)\to V(G)$ is a bijection. The width of $(T,\delta)$ is given by $\max_{e\in E(T)}\text{mp}(e)$ and the perfect matching width of $G$ is then defined as
$\text{pmw}(G):=\min_{\begin{subarray}{c}(T,\delta)\\ \text{perfect matching}\\ \text{decomposition of }G\end{subarray}}\max_{e\in E(T)}\text{mp}(\partial(e)).$
A.54 Bisection-width
The Bisection problem seeks to partition the vertices of a graph into two equally sized sets while minimizing the cut size. The Bisection Width is a width parameter closely related to this Bisection problem [64, 643, 66]. If a graph G has bounded Band-width, then it also has bounded bisection-width [668].
Definition A.90.
[66]. A bisection $\xi$ of an undirected graph $G=(V,E)$, where $|V|=n$, is a partition $V=V_{0}\cup V_{1}$ such that $|V_{0}|,|V_{1}|\in\left\{\left\lfloor\frac{n}{2}\right\rfloor,\left\lceil\frac{n}{2}\right\lceil\right\}$. For simplicity, we assume that $n$ is even throughout this paper. The number of edges between $V_{0}$ and $V_{1}$ for a given bisection $\xi$ is called the cut size of $\xi$. The minimum cut size among all possible bisections of $G$ is called the bisection width of $G$ and is denoted by $\text{bw}(G)$. Formally, this is defined as
$\text{bw}(G):=\min\{|\{\{v,w\}\in E\mid\xi(v)\neq\xi(w)\}|\mid\xi\text{ is a bisection of }G\}.$
A.55 Maximum matching width
Maximum matching width is a graph width parameter used to consider fast algorithms for the dominating set problem [418].
Definition A.91.
A rooted branch decomposition is a branch decomposition $(T,\delta)$ where an edge of $T$ is subdivided to create a root $r$. For an internal vertex $v\in V(T)$, $\delta(v)$ is the union of $\delta(l)$ for all leaves $l$ that have $v$ as an ancestor.
Given a symmetric function $f:2^{X}\to\mathbb{R}$, the $f$-value of an edge $e$ in $(T,\delta)$ is $f(A)=f(B)$, where $A$ and $B$ are the parts of the partition induced by $e$. The $f$-width of $(T,\delta)$ is the maximum $f$-value over all edges of $T$, denoted by $f(T,\delta)$. The $f$-width of $X$ is the minimum $f$-width over all branch decompositions of $X$. If $|X|\leq 1$, the $f$-width is defined as $f(\emptyset)$.
For a graph $G$ and a subset $S\subseteq E(G)$, the branchwidth $\text{bw}(G)$ is the $f$-width of $E(G)$ for $f(S)$, defined as the number of vertices incident with edges in both $S$ and $E(G)\setminus S$.
The maximum matching-width of a graph $G$, denoted by $\text{mmw}(G)$, is defined as the $f$-width of $V(G)$ for $f=\text{mm}$, where $\text{mm}(S)$ is the size of the maximum matching in $G[S,V(G)\setminus S]$.
A.55.1 Hierarchy for Maximum matching width
We consider about Hierarchy for Maximum matching width.
Theorem A.92.
The relationships between Maximum matching width and other graph parameters are as follows:
  • If a graph $G$ has bounded Maximum matching width, then it also has bounded tree-depth [618].
  • If a graph $G$ has bounded Maximum matching width, then it also has bounded Maximum induced matching-width [618].
  • For any graph $G$, we have $\text{mmw}(G)\leq\text{branch-width}(G)\leq\text{tree-width}(G)+1\leq 3\cdot\text{mmw}(G).$
Proof A.93.
Please refer to each reference.
A.56 Linear Maximum Matching width
Linear Maximum Matching width is the linear restriction of Maximum Matching width.
A.56.1 Hierarchy for Linear-Maximum matching width
We consider about Hierarchy for Linear-Maximum matching width.
Theorem A.94.
The relationships between Linear MIM-width and other graph parameters are as follows:
  • $\text{lmmw}(G)\leq\text{pw}(G)\leq 2\times\text{lmmw}(G)$. (Path width) [568].
  • If a graph $G$ has bounded Linear MIM-width, then it also has bounded MIM-width (trivially).
Proof A.95.
Please refer to each reference.
A.57 Clustering-width
Clustering-width is graph width parameter of the smallest number of variables whose deletion results in a variable-disjoint union of hitting formulas[567].
A.58 Query-width
Query-width is graph width parameter of a common generalization of both acyclicity and tree-width [96, 263].
A.59 Universal width
Universal width [446, 381] is a width parameter for an alternating finite automaton (AFA). Note that an alternating finite automaton (AFA) is a computational model where states can make simultaneous existential or universal transitions, allowing parallel computation paths.
A.59.1 Related concepts for Universal width
Related concepts are the following:
  • Maximal universal width [446]: Maximal Graph width parameter of universal width.
  • Combined width [449]: Width parameter for an alternating finite automaton (AFA).
  • Maximal combined width [449]: Maximal Graph width parameter of combined width.
  • Split-width [16, 174]: Width parameter for an alternating finite automaton (AFA).
A.60 Degree-width
Degree-width [176] is graph width parameter for tournaments that we can be seen as a measure of how far is the tournament from being acyclic.
A.61 Median-width
Median-width [647, 648] is a graph width parameter used for median graphs (A median graph is an undirected graph where any three vertices have a unique median vertex that lies on the shortest paths between each pair [466, 461].) as the underlying graph of the decomposition. For any graph $G$, Median-width(G) $\leq$ Tree-width(G) + 1 [648].
A.61.1 Related concepts for Median-width
Lattice-width [648] refers to a parameter that constrains the median graph of a decomposition to be isometrically embeddable into the Cartesian product of $i$ paths.
A.62 Directed Modular-width
Directed Modular-width [649] is Modular-width on directed graphs.
A.63 Cycle-width
Cycle-width [386] measures graph complexity, focusing on cycle structures and highlighting differences between directed and undirected tree-width by emphasizing local graph properties.
A.64 Arc-width
Arc-width [49, 32] is the maximum number of arcs.
Definition A.96.
[49, 32] The arc-representation of a graph $G$ is a mapping $\phi$ from the vertex set $V(G)$ to the set of arcs on a base circle, such that adjacent vertices of $G$ are mapped to intersecting arcs. The width (in a representation) of a point $P$ on the base circle is the number of representing arcs containing $P$. The width of $\phi$ is the maximum width of the points on the base circle. The arc-width of a graph $G$ is the minimal possible width of such arc-representations, denoted as $\text{aw}(G)$.
A.64.1 Related concepts for arc-width
Vortex-width [610, 663] measures how closely a graph can be "nearly drawn" on a surface with respect to a given cyclic ordering of vertices. Vortex-width is equal to Arc-width.
A.65 Match width and braid width
Match width and braid width [132] are parameters used to measure structural complexity in XML documents; they help determine XPath satisfiability based on document depth.
A.66 Clique cover width
The clique cover width of $G$, denoted by $ccw(G)$, is the minimum value of the bandwidth of all graphs that are obtained by contracting the cliques in a clique cover of G into a single vertex [628, 629, 627].
A.67 Questionable-width
Questionable-width [514, 515] relates to graph width measurements concerning questionable-width. A “question” identifies the initial divergence between two sequences of elements from orders or sets, indexed ordinally. In a questionable representation with finite width of an order $O$, comparisons are resolved by examining the “question” that evaluates elements within the finite order $O$ [514, 515].
A.68 plane-width
The plane-width of a graph is defined as the minimum diameter of the image of the graph’s vertex set [431, 431].
Appendix B Various Length Parameters
In this section, we consider about various length parameters.
B.1 Tree-length and Branch-length
We begin by discussing Tree-length and other well-known parameters. Tree-length measures the maximum distance between any two vertices within a single bag of a tree-decomposition, minimized over all possible decompositions[169, 205, 206]. The formal definition is provided below.
Definition B.1.
[169, 205, 206] The length of a tree-decomposition $T(G)=(\{X_{i}\mid i\in I\},T=(I,F))$ of a graph $G$ is defined as:
$\lambda(T)=\max_{i\in I}\max_{u,v\in X_{i}}d_{G}(u,v),$
where $d_{G}(u,v)$ denotes the distance between vertices $u$ and $v$ in the graph $G$. The tree-length of a graph $G$, denoted by $\text{tl}(G)$, is the minimum length over all possible tree-decompositions of $G$. Chordal graphs have tree-length 1.
A parameter closely related to tree-length is tree-breadth. Tree-breadth measures how tightly the vertices in each bag of a tree-decomposition can be clustered around a central vertex within a specific radius, minimized over all decompositions[496, 495]. The formal definition is provided below.
Definition B.2.
[496, 495] The breadth of a tree-decomposition $T(G)=(\{X_{i}\mid i\in I\},T=(I,F))$ of a graph $G$ is the minimum integer $r$ such that for every $i\in I$, there exists a vertex $v_{i}\in V(G)$ with $X_{i}\subseteq D_{r}(v_{i},G)$, where $D_{r}(v_{i},G)=\{u\in V(G)\mid d_{G}(u,v_{i})\leq r\}$ is the disk of radius $r$ centered at $v_{i}$ in $G$. The tree-breadth of $G$, denoted by $\text{tb}(G)$, is the minimum breadth over all possible tree-decompositions of $G$.
It is known that for any graph $G$:
$1\leq\text{tb}(G)\leq\text{tl}(G)\leq 2\cdot\text{tb}(G).$
A similar concept known as Branch-length is equivalent to Tree-length [675]. Additionally, linear-length is equal to path-length [675].
Definition B.3.
[675] Let $B=(T,\mu)$ be a branch decomposition of a graph $G$. The branch-length of $B$, denoted by $\text{bl}(B)$, is defined as
$\text{bl}(B)=\max_{e\in E(T)}\text{diam}_{G}(\text{mid}_{B}(e)).$
The branch-length of $G$, denoted by $\text{bl}(G)$, is defined as
$\text{bl}(G)=\min_{B}\text{bl}(B),$
where the minimum is taken over all branch decompositions of $G$.
B.2 Future tasks for length parameters
We will consider the following length parameters. These are merely suggestions, and we plan to explore them further in the future. Mathematical verification of these parameters is forthcoming.
  • Directed tree length: The length of a directed tree decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The directed tree length of a graph $G$ is the minimum length over all directed tree decompositions of $G$.
  • Directed Path length: The length of a directed path decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The directed path length of a graph $G$ is the minimum length over all directed path decompositions of $G$.
  • Rank length: The length of a rank decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The rank length of a graph $G$ is the minimum length over all rank decompositions of $G$.
  • Linear Rank length: The length of a linear rank decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The linear rank length of a graph $G$ is the minimum length over all linear rank decompositions of $G$.
  • Boolean length: The length of a boolean decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The boolean length of a graph $G$ is the minimum length over all boolean decompositions of $G$.
  • Linear Boolean length: The length of a linear boolean decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The linear boolean length of a graph $G$ is the minimum length over all linear boolean decompositions of $G$.
  • Hypertree length: The length of a hypertree decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The hypertree length of a graph $G$ is the minimum length over all hypertree decompositions of $G$.
  • Carving length: The length of a carving decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The carving length of a graph $G$ is the minimum length over all carving decompositions of $G$.
  • Cut length: The length of a cut decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The cut length of a graph $G$ is the minimum length over all cut decompositions of $G$.
  • Tree-cut length: The length of a tree-cut decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The tree-cut length of a graph $G$ is the minimum length over all tree-cut decompositions of $G$.
  • Tree distance length: The length of a tree distance decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The tree distance length of a graph $G$ is the minimum length over all tree distance decompositions of $G$.
  • Path distance length: The length of a path distance decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The path distance length of a graph $G$ is the minimum length over all path distance decompositions of $G$.
  • Clique length: The length of a clique decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The clique length of a graph $G$ is the minimum length over all clique decompositions of $G$.
  • Linear Clique length: The length of a linear clique decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The linear clique length of a graph $G$ is the minimum length over all linear clique decompositions of $G$.
  • NLC length: The length of a NLC decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The NLC length of a graph $G$ is the minimum length over all NLC decompositions of $G$.
  • Linear NLC length: The length of a linear NLC decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The linear NLC length of a graph $G$ is the minimum length over all linear NLC decompositions of $G$.
  • Modular length: The length of a modular decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The modular length of a graph $G$ is the minimum length over all modular decompositions of $G$.
Appendix C Comparing Various Graph Parameters (Over 70 Parameters)
We examine the relationships among various graph parameters. Please refer to the supplemental file “Supplemental Figure: Comparing Graph Width Parameters” for more details. Understanding these relationships is highly beneficial for developing algorithms.
Note C.1.
Note that unless otherwise specified, we consider any graph when comparing graph parameters. We say that a parameter $p$ upper bounds a parameter $q$ if there exists a non-decreasing function $f$ such that $f(p(G))\geq q(G)$ for all graphs $G$. Conversely, if $p$ does not upper bound $q$, then $q$ is considered unbounded with respect to $p$. Additionally, if value $a$ upper bounds value $b$ and $b$ upper bounds value $c$, then value $a$ upper bounds $c$.
Note C.2.
Comparing graph width parameters mathematically means evaluating how different measures of graph complexity, such as tree-width, path-width, and clique-width, relate to each other. These comparisons help in determining which parameter is more restrictive or more general, and thus, influence the choice of techniques and tools in algorithm design.
Appendix D Various Graph Depth Parameters
Many graph depth parameters are known. Examining the relationships among these depth parameters, such as inequalities, upper bounds, and lower bounds, is a well-known research topic.
D.1 Tree-depth
Tree-depth is a graph depth parameter introduced under several names as a measure of the sparsity of a graph [225, 562, 353, 397, 147].
Definition D.1.
The tree-depth of a graph $G$, denoted by $\text{td}(G)$, is defined based on the concept of a rooted forest. A rooted forest is a collection of trees where each tree has a designated root and all edges are directed away from the root. The closure of a rooted forest is the undirected graph obtained by connecting two vertices if there is a directed path between them in the forest. The height of a rooted forest is the length of the longest directed path from any root. The tree-depth of $G$ is the minimum height of a rooted forest whose closure contains $G$ as a subgraph.
D.1.1 Hierarchy for Tree Depth
We coinsider about hierarchy for Tree Depth.
Theorem D.2.
The relationships between Tree-depth and other graph parameters are as follows:
  • Cluster vertex deletion number is greater than or equal to Tree-depth [123].
  • Bridge-depth (a graph width parameter which measures the kernelization complexity of structural parameterizations of the Vertex Cover problem) is less than or equal to Tree-depth [112].
  • If a graph $G$ has bounded Tree-depth, then it also has bounded Path-width [618].
[…]
Proof D.3.
Please refer to each reference.
D.2 Branch-depth
Branch-depth is a graph depth parameter generalizing tree-depth of graphs [182]. Note that the radius of a tree is the minimum integer $r$ such that there exists a node in the tree with a distance of at most $r$ from every other node.
Definition D.4.
[182] A decomposition of a connectivity function on a set $V$ is a pair $(T,\phi)$, where $T$ is a tree with at least one internal node, and $\phi$ is a bijection from $V$ to the set of leaves of $T$. The radius of a decomposition $(T,\phi)$ is defined as the radius of the tree $T$. For an internal node $v$ of $T$, the components of the graph induced by $v$ result in a partition $\{V_{i}\}$ of $V$ by removing $v$. The width of $v$ is defined as $\max_{i}|V_{i}|$. The width of the decomposition $(T,\phi)$ is the maximum width among all internal nodes of $T$. We say that a decomposition $(T,\phi)$ is an $(r,d)$-decomposition of $V$ if its width is at most $d$ and its radius is at most $r$. The branch-depth of $V$ is the minimum integer $r$ such that there exists an $(r,d)$-decomposition of $V$.
D.2.1 Hierarchy for branch-depth
We coinsider about hierarchy for branch-depth.
Theorem D.5.
The relationships between branch-depth and other graph parameters are as follows:
  • The branch-width is less than or equal to its branch-depth [182].
  • The branch-depth of a connected graph is less than or equal to its tree-depth [182].
  • Let $G$ be a graph, $k$ be its branch-depth, and $t$ be its tree-depth. Then $k-1\leq t\leq\max(2k^{2}-k+1,2)$ [182].
  • For all matroids $M$, the branch-depth of $M$ is less than or equal to contraction-deletion depth of $M$. And
[…]
    [182].
[…]
Proof D.6.
Please refer to each reference.
D.2.2 Related Parameters for Branch-depth
Deletion-depth [182], Contraction-depth [182], Contraction-deletion Depth [182], Contraction*-depth [437], and Contraction-deletion Depth [437] are graph depth parameters analogous to operations on matroids.
D.3 Rank-depth
Rank depth [182] is a graph depth parameter defined by the Rank depth-decomposition. Note that Rank depth-decomposition of a graph $G$ is a branch-depth-decomposition of the cut-rank function [425, 569, 452, 426]. A class of simple graphs has bounded rank-depth if and only if it has bounded shrub-depth [182].
Definition D.7.
[182] The cut-rank function of a simple graph $G=(V,E)$ is a function $\rho_{G}$ defined on the subsets of $V$. For a subset $X\subseteq V$, $\rho_{G}(X)$ is the rank of the $X\times(V\setminus X)$ matrix $A_{X}=(a_{ij})_{i\in X,j\in V\setminus X}$ over the binary field. Here, $a_{ij}=1$ if and only if vertices $i$ and $j$ are adjacent, and $a_{ij}=0$ otherwise. This cut-rank function serves as an instance of a connectivity function on the vertex set of a simple graph, as discussed in a paper by Oum and Seymour [30]. We define the rank-depth of a simple graph $G$, denoted by $rd(G)$, as the branch-depth of $\rho_{G}$.
D.4 Shrub-depth
Shrub-depth is a graph depth parameter that captures the height of dense graphs [425, 317]. A related parameter is SC-depth [425, 317].
D.4.1 Hierarchy for Shrub-depth
We consider about Hierarchy for Shrub-depth.
Theorem D.8.
The relationships between Shrub-depth and other graph parameters are as follows:
  • If $G$ is of tree-depth $\leq k$, then $G$ is of shrub-depth $\leq k$ [316].
  • If $G$ is of bounded shrub-depth, then $G$ is of bounded clique-width [316].
  • If $G$ is of bounded twin-cover, then $G$ is of bounded Shrub-depth [308].
Proof D.9.
Please refer to each reference.
D.5 Directed Tree-depth
Directed Tree-depth is tree-depth on directed graphs [479].
Related parameter is DAG-depth. DAG-depth is a structural depth measure of directed graphs, which naturally extends the tree-depth of ordinary graphs [65, 321]. Definition of DAG-depth is following.
Definition D.10 (DAG-depth).
[65, 321] Let $D$ be a digraph, and let $R_{1},\ldots,R_{p}$ be the reachable fragments of $D$. The DAG-depth $\text{dd}_{p}(D)$ is defined inductively as follows:
$\text{dd}_{p}(D)=\begin{cases}1&\text{if }|V(D)|=1,\\ 1+\min_{v\in V(D)}(\text{dd}_{p}(D-v))&\text{if }p=1\text{ and }|V(D)|>1,\\ \max_{1\leq i\leq p}\text{dd}_{p}(R_{i})&\text{otherwise}.\end{cases}$