### v8

# Various Properties of Various Ultrafilters, Various Graph Width Parameters, and Various Connectivity Systems (with Survey)

Abstract
This paper investigates ultrafilters in the context of connectivity systems, defined as pairs $(X,f)$ where $X$ is a finite set and $f$ is a symmetric submodular function. Ultrafilters, essential in topology and set theory, are extended to these systems, with a focus on their relationship to graph width parameters, which help analyze graph complexity.
We demonstrate theorems for ultrafilters on connectivity systems and explore related concepts such as prefilters, ultra-prefilters, and subbases. New parameters for width, length, and depth are introduced, providing further insight into graph width. The study also includes a comparison of various graph width parameters and their related concepts, offering a foundation for future research in graph theory and computational complexity. Additionally, we explore connections to other mathematical disciplines, including set theory, lattice theory, and matroid theory, expanding the scope of ultrafilters and graph width. (It also includes information similar to that found in surveys, aiming to promote future research on graph width parameters.)
keywords:
Ultrafilter, Filter, Connectivity System, Prefilter, Branch-width
category:
[…]
1 Introduction
1.1 Filter and Graph width parameters
Filters are essential collections of sets in topology and set theory, characterized by their closure under supersets and finite intersections. They can be viewed as a way to "focus" on certain subsets of a space, much like a lens that sharpens specific details while filtering out others. An ultrafilter, being a maximal filter on a set, is particularly significant for addressing fundamental concepts such as limits, convergence, and compactness. Its unique properties render it indispensable in various fields, including non-standard analysis, model theory, social choice theory (social judgments), group theory, boolean algebra, geometry, probability theory, vector theory, semigroup theory, abstract algebra, topology, set theory, infinite combinatorics, fuzzy theory, graph theory, matroid theory, lattice theory, and first-order logic, where it provides powerful tools for both mathematical and logical applications[529, 526, 1240, 657, 490, 493, 6, 298, 579, 795, 1109, 156, 255, 912, 514, 253, 517, 354].
Graph theory, a key branch of mathematics, delves into the study of networks consisting of nodes and edges, with a focus on their paths, structures, and properties [1207, 338, 183]. Among the critical metrics in this field is the "graph width parameter," which measures the maximum width across all cuts or layers within a hierarchical decomposition of the graph. This metric is crucial for analyzing a graph’s complexity and structure, serving as a primary factor in transforming computationally hard graph problems into more tractable ones when the graph class is restricted to having bounded width. Various width parameters have been rigorously explored and are widely recognized in the literature [565, 966, 410, 427, 529, 48, 1038, 814, 618, 711, 500, 813, 496, 499, 510, 683, 515, 1153, 518]. For further information, see Appendix A, “Various Width Parameters,” and Appendix C, “Comparing Various Graph Parameters (Over 70 Parameters).”
Branch-width is an important graph width parameter, defined through branch decomposition where the leaves correspond to the graph’s edges [565, 966, 164]. It is closely related to tree-width, another significant graph width parameter. Tree-width is determined by a tree decomposition, which represents the graph as a tree structure, grouping vertices into "bags," with the tree-width being the size of the largest bag minus one [1038, 1033]. Conversely, branch-width involves a branch decomposition where the graph is segmented into a tree-like structure, with the branch-width representing the maximum size of the minimum cut between two parts of the graph. The relationship between branch-width (bw(G)) and tree-width (tw(G)) is given by the inequalities: bw(G) $\leq$ tw(G)+1 $\leq$ $\frac{3}{2}$cdot bw(G) for a graph $G$ with bw(G) $\geq$ 2 [164]. This relationship shows that both parameters are linked, with tree-width generally being larger, but not excessively so. Extending tree-width to the connectivity system framework leads to the concept of branch-width, highlighting their conceptual connection.
Graph width parameters offer several significant advantages in both theoretical and practical contexts:
  1. 1.
    Foundational Role in Theoretical Graph Theory: Graph width parameters play a critical role in the Graph Minors project serving as a fundamental combinatorial tool [565, 966]. This project, which is pivotal in the structure theory of graphs, uses graph width parameters to explore graph properties and their relationships through a series of influential theorems and algorithms.
  2. 2.
    Algorithmic Efficiency: Graph width parameters are conducive to algorithmic applications, particularly in the field of fixed-parameter tractable (FPT) algorithms [470, 531, 452, 1098]. Algorithms that are parameterized by Graph width parameters often demonstrate superior efficiency and are widely utilized in computational graph theory. This aspect is particularly valuable in optimizing complex computations and in the development of algorithms that can efficiently solve problems considered intractable by other means.
  3. 3.
    Practical Applicability: In real-world applications, graphs derived from various domains such as programming language [248, 1149], road networks [1246, 1202], and organizational structures [946] often exhibit small width. This characteristic simplifies complex problems, making them more manageable and allowing for the application of advanced graph algorithms. As a result, studying graph width parameters can directly impact the effectiveness and efficiency of practical solutions in engineering, software development, and logistics [1196, 265, 1225, 1085, 248, 337, 740, 340].
A pair $(X,f)$, consisting of a finite set $X$ and a symmetric submodular function $f$, is recognized as a connectivity system [565]. This concept is widely utilized in the analysis of graph structures, particularly in relation to graph width parameters such as branch-width and tree-width [565, 966, 529]. Exploring the duality (the min-max theorem) within these parameters enhances our understanding of the relationships among various graph decompositions and measures of graph complexity. In this context, "duality" refers to a theorem or relationship where the presence (or absence) of one entity implies the absence (or presence) of its corresponding dual entity, often referred to as a minmax theorem [1072]. Ultrafilters on connectivity systems are known to exhibit such a dual relationship with branch-width [529]. Concepts like ultrafilters, often termed as obstructions, play a crucial role in determining the values of graph width parameters.
1.2 Our Contribution
This paper outlines our contributions as follows:
  • Section 2: We primarily explain the basic concepts of ultrafilters on connectivity systems and graph-width parameters, along with previously known concepts. We discuss the relationship between ultrafilters on connectivity systems and well-known concepts like Tangle from Graph Minor theory and Matroid commonly used in optimization theory.
  • Section 3: We delve into ultrafilters on connectivity systems, considering Tukey’s Lemma for these systems. Tukey’s Lemma asserts that every non-empty collection of sets, closed under supersets, contains a maximal element. This fundamental result in set theory is often used to prove the existence of ultrafilters and related concepts [1240, 657, 6]. Additionally, we explore chains and antichains on connectivity systems.
  • Section 4: We discuss prefilters, ultra-prefilters, and subbases on connectivity systems. Prefilters, ultra-prefilters, and subbases are known concepts used to generate filters and are studied in various fields. This exploration enhances our understanding of these concepts in Set Theory and their significance in different mathematical and logical contexts.
  • Section 5: We examine the connectivity system discussed for finite sets from the perspective of infinite connectivity systems and countable connectivity systems. We analyze the properties of ultrafilters on these connectivity systems.
  • Section 6: We investigate new many parameters such as width, length, and depth, as well as the Ultrafilter game on connectivity systems, ultraproduct on connectivity systems, and the Axiom of Choice on connectivity systems. We also consider a Small Set Expansion [1217, 93, 1232, 241, 443, 142].
  • Appendix: This section presents an investigation and comparison of various graph width parameters and their related parameters. It also includes information similar to that found in surveys, aiming to promote future research on graph width parameters.
We aim to make new contributions to the study of Graph Width Parameters and Graph Algorithms by interpreting key concepts like Tree-decomposition, Branch-decomposition, Tangles, and their related graph parameters through the lens of various mathematical disciplines. For instance, in this paper, we explore concepts from Set Theory (Ultrafilter, prefilter), Lattice Theory (chain, antichain), Model Theory (Ultraproduct), Hypergraph Theory (fuzzy hypergraph, clutter), Matroid Theory (Single-element-extension), Topology (P-point, Q-point), Fuzzy Theory (Fuzzy Tree-width, bipolar fuzzy graphs), First-order Logic (weak ultrafilter), Social Judgment (Voting system, Quasi Ultrafilter), Game Theory (cops and robbers, simple game, monotone search game), and Sperner Theory (Sperner system and trace), among others.
2 Definitions and Notations in this paper
This section provides mathematical definitions for each concept. Before delving into specific definitions, let’s outline the basic mathematical concepts used in this text.
2.1 Notation in this paper
We explain the notation used in this paper.
2.1.1 Notation for Set theory
In set theory, a set is a collection of distinct elements or objects, considered as an entity and often denoted with curly braces. A subset is a set where all elements are also contained within another set. Boolean algebra $(X,\cup,\cap)$ is a mathematical structure with a set $X$, union ()($\cup$)( ∪ ), and intersection ()($\cap$)( ∩ ), satisfying specific axioms for operations. In this paper, we consider finite sets (except in Section 5).
Note 2.1.
In this paper, we use expressions like $A\subseteq X$ to indicate that $A$ is a subset of $X$, $A\cup B$ to represent the union of two subsets $A$ and $B$ (both of which are subsets of $X$), and $A=\emptyset$ to signify an empty set. Specifically, $A\cap B$ denotes the intersection of subsets $A$ and $B$. Similarly, $A\setminus B$ represents the difference between subsets $A$ and $B$.
Definition 2.2.
The powerset of a set $A$, denoted as $2^{A}$, is the set of all possible subsets of $A$, including the empty set and $A$ itself.
Example 2.3 (Powerset of a Set).
Let $A=\{1,2\}$. The powerset of $A$, denoted as $2^{A}$, is the set of all possible subsets of $A$. This includes the empty set, the individual elements, and the set $A$ itself. Therefore, the powerset of $A$ is:
$2^{A}=\{\emptyset,\{1\},\{2\},\{1,2\}\}.$
Similarly, for a set $B=\{a,b,c\}$, the powerset $2^{B}$ is:
$2^{B}=\{\emptyset,\{a\},\{b\},\{c\},\{a,b\},\{a,c\},\{b,c\},\{a,b,c\}\}.$
In this case, there are $82^{3}=$ 8 subsets, as the size of the powerset is $2^{|B|}$, where $|B|$ is the number of elements in the original set.
Definition 2.4.
In set theory, a partition of a set is a way of dividing the set into non-overlapping, non-empty subsets, such that every element of the original set is included in exactly one subset. These subsets are called the "blocks" or "parts" of the partition (cf.[942]).
Example 2.5 (Partition of a Set).
Consider the set $X=\{1,2,3,4,5,6\}$. A partition of $X$ is a collection of non-overlapping, non-empty subsets whose union equals $X$. For instance, the following subsets form a partition of $X$:
$P=\{\{1,2\},\{3,4\},\{5,6\}\}.$
Here, each element of $X$ belongs to exactly one subset, and the subsets do not overlap. Hence, $P$ is a valid partition of $X$.
Definition 2.6.
In set theory, a separation of a set $X$ is a partition of $X$ into two disjoint, non-empty subsets $A$ and $B$ such that their union equals $X$ (i.e., $A\cup B=X$ and $A\cap B=\emptyset$). The subsets $A$ and $B$ are often referred to as the separated components of $X$.
Definition 2.7.
In set theory, the cardinality of a set refers to the number of elements in the set. For finite sets, the cardinality is simply the count of elements in the set.
Formally, the cardinality of a set $A$ is denoted by $|A|$. If there exists a bijection (a one-to-one correspondence) between two sets $A$ and $B$, they are said to have the same cardinality, i.e., $|A|=|B|$.
For more detailed information on set theory, please refer to surveys and lecture notes (ex.[829, 1174, 477, 811]).
2.1.2 Notation for undirected graph
Except appendix and section 6, we consider a simple undirected graph $G$, where the vertex set is denoted by $V(G)$ and the edge set by $E(G)$. For simplicity, we will often write $G=(V,E)$, with $V$ representing the vertices and $E$ the edges. If $X$ is a subset of the vertices $V(G)$ (or the edges $E(G)$), then $X^{c}$ represents the complement set $V(G)\setminus X$ (or $E(G)\setminus X$, respectively), which includes all elements not in $X$.
Definition 2.8.
The degree of a vertex refers to the number of edges connected to it. For example, if a vertex has three edges connected to it, its degree is three.
Definition 2.9.
(cf.[218]) A path graph is a type of tree that consists of a sequence of vertices where each vertex is connected to exactly two others, except for the two endpoints, which are connected to only one other vertex. A linear ordering of a graph’s vertices is an arrangement of the vertices in a linear sequence in such a way that each vertex comes before all vertices to which it is connected by edges in the sequence, except possibly the previous vertex.
Example 2.10.
Consider the path graph $P_{4}$ consisting of four vertices $v_{1},v_{2},v_{3},v_{4}$ and three edges $\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\}$. The vertices can be arranged linearly as $v_{1},v_{2},v_{3},v_{4}$, which respects the edge connections, making it a valid path graph. The endpoints $v_{1}$ and $v_{4}$ are connected to only one other vertex, while the intermediate vertices $v_{2}$ and $v_{3}$ are each connected to two vertices.
Definition 2.11 (Connectedness).
(cf.[1175]) A graph $G=(V,E)$ is said to be connected if for every pair of vertices $u,v\in V$, there exists a path $P\subseteq G$ that connects $u$ and $v$. Formally, $G$ is connected if:
$\forall u,v\in V,\exists P\subseteq G\text{ such that }P\text{ is a path from }u\text{ to }v.$
Definition 2.12.
(cf.[203]) A tree is a connected, acyclic graph where any two vertices are connected by exactly one path, symbolizing hierarchical relationships. The root of a tree is the topmost node from which all other nodes descend. In a tree, vertices with a degree of 1 are called leaves, while all other vertices are referred to as inner vertices. A ternary tree is a specific type of tree where all inner vertices have a degree of 3.
Example 2.13.
Consider a tree with 7 vertices, labeled $v_{1}$ to $v_{7}$, where $v_{1}$ is the root. The edges of the tree are $\{v_{1},v_{2}\}$, $\{v_{1},v_{3}\}$, $\{v_{1},v_{4}\}$, $\{v_{2},v_{5}\}$, $\{v_{2},v_{6}\}$, and $\{v_{3},v_{7}\}$. This forms a ternary tree, as the root $v_{1}$ has three children $v_{2}$, $v_{3}$, and $v_{4}$, making it an inner vertex with degree 3. The vertices $v_{5}$, $v_{6}$, $v_{7}$, and $v_{4}$ are leaves with a degree of 1.
Definition 2.14.
A subgraph is a graph formed from a subset of a graph’s vertices and edges. A subpath is a continuous segment of a path, consisting of consecutive edges and vertices from the original path, and is useful for analyzing specific portions of larger paths.
Example 2.15.
Consider the original graph $G$ with vertices $V(G)=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ and edges $E(G)=\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\},\{v_{4},v_{5}\}\}$. A subgraph $H$ can be formed by selecting the subset of vertices $V(H)=\{v_{2},v_{3},v_{4}\}$ and the corresponding edges $E(H)=\{\{v_{2},v_{3}\},\{v_{3},v_{4}\}\}$. This subgraph $H$ represents a subpath of the original path in $G$, highlighting a specific segment of the larger path.
Definition 2.16.
The distance between two vertices $u$ and $v$ in a graph $G=(V,E)$ is defined as the length of the shortest path connecting $u$ and $v$ in $G$. Formally, the distance $d_{G}(u,v)$ is given by:
$d_{G}(u,v)=\min\{\ell(P)\mid P\text{ is a path in }G\text{ connecting }u\text{and }v\}$
where $\ell(P)$ denotes the length of the path $P$, which is the number of edges in $P$.
Example 2.17.
Consider the graph $G$ with vertices $V(G)=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ and edges $E(G)=\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\},\{v_{4},v_{5}\},\{v_{1},v_{3}\}\}$.
To find the distance $d_{G}(v_{1},v_{4})$, observe that there are two paths connecting $v_{1}$ and $v_{4}$:
  • Path 1: $v_{1}\to v_{2}\to v_{3}\to v_{4}$, with a length of 3 edges.
  • Path 2: $v_{1}\to v_{3}\to v_{4}$, with a length of 2 edges.
The shortest path is Path 2, so the distance $d_{G}(v_{1},v_{4})$ is 2.
Definition 2.18.
A clique in an undirected graph $G=(V,E)$ is a subset $C\subseteq V$ of vertices such that every pair of distinct vertices in $C$ is connected by an edge in $G$. Formally, for all $u,v\in C$ with $u\neq v$, the edge $(u,v)\in E$.
Example 2.19.
Consider a graph $G$ with vertex set $V(G)=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ and edge set
$E(G)=\{\{v_{1},v_{2}\},\{v_{1},v_{3}\},\{v_{1},v_{4}\},\{v_{2},v_{3}\},\{v_{2},v_{4}\},\{v_{3},v_{4}\},\{v_{4},v_{5}\}\}.$
The subset $C=\{v_{1},v_{2},v_{3},v_{4}\}$ forms a clique because every pair of distinct vertices in $C$ is connected by an edge:
  • $\{v_{1},v_{2}\},\{v_{1},v_{3}\},\{v_{1},v_{4}\}\in E(G)$,
  • $\{v_{2},v_{3}\},\{v_{2},v_{4}\}\in E(G)$,
  • $\{v_{3},v_{4}\}\in E(G)$.
However, the subset $\{v_{3},v_{4},v_{5}\}$ is not a clique, as $\{v_{3},v_{5}\}\notin E(G)$.
Definition 2.20 (union).
(cf.[357, 503]) The union of two graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ is a graph $G=(V,E)$ where:
  • The vertex set $V$ is the union of the vertex sets of $G_{1}$ and $G_{2}$:
    $V=V_{1}\cup V_{2}.$
  • The edge set $E$ is the union of the edge sets of $G_{1}$ and $G_{2}$:
    $E=E_{1}\cup E_{2}.$
Thus, the union of $G_{1}$ and $G_{2}$ combines the vertices and edges of both graphs, without duplicating any elements.
Definition 2.21 (Intersection).
(cf.[885, 289]) The intersection of two graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ is a graph $G=(V,E)$ where:
  • The vertex set $V$ is the intersection of the vertex sets of $G_{1}$ and $G_{2}$:
    $V=V_{1}\cap V_{2}.$
  • The edge set $E$ is the intersection of the edge sets of $G_{1}$ and $G_{2}$:
    $E=E_{1}\cap E_{2}.$
Thus, the intersection of $G_{1}$ and $G_{2}$ consists of only the vertices and edges that are common to both graphs.
For more detailed information on graphs, please refer to surveys and lecture notes (ex.[1013, 182, 339, 1183, 181]).
2.1.3 Notation for directed graph
We define a directed graph $D$, where the vertex set is denoted by $V(D)$ and the edge set by $E(D)$. For clarity, we will use the notation $D=(V,E)$, where $V$ represents the vertices and $E$ represents the directed edges. If $X$ is a subset of either the vertex set $V(D)$ or the edge set $E(D)$, then $X^{c}$ denotes the complement, $V(D)\setminus X$ or $E(D)\setminus X$, respectively, consisting of all elements not contained in $X$.
Definition 2.22.
The degree of a vertex in a directed graph is defined by the number of directed edges incident to it, separated into in-degree and out-degree. For example, a vertex with three incoming edges has an in-degree of three.
Definition 2.23.
A path in a directed graph is a sequence of vertices where each vertex is connected to the next by a directed edge. A linear ordering of the vertices of a directed graph arranges them in a sequence such that, for each directed edge, the source vertex precedes the target vertex in the sequence.
Definition 2.24.
A tree in the context of directed graphs is a directed acyclic graph (DAG) that is connected and contains no directed cycles. A directed tree with $n$ vertices always has $n-1$ directed edges. A leaf in a directed tree is a vertex with exactly one incoming or outgoing edge, making it a terminal vertex. In this context, a node is any vertex, which can be either a leaf or an internal vertex depending on its degree. A subcubic directed tree is one in which every vertex has a total degree (in-degree plus out-degree) of at most three, meaning no vertex is connected by more than three directed edges.
Example 2.25 (Directed Tree with Specific Values).
Consider the directed tree $T=(V,E)$, where the set of vertices is $V=\{1,2,3,4,5\}$ and the set of directed edges is $E=\{(1,2),(1,3),(2,4),(2,5)\}$. The directed tree can be illustrated as follows: $1\rightarrow 2\rightarrow 4$, $1\rightarrow 2\rightarrow 5$, $1\rightarrow 3$.
In this example:
  • The root of the tree is vertex $1$, as it has no incoming edges.
  • The vertices $4$ and $5$ are leaves, as they have exactly one incoming edge and no outgoing edges.
  • The tree has $5$ vertices and $4$ directed edges, satisfying the condition that a directed tree with $n$ vertices has $n-1$ edges.
  • The tree is connected, acyclic, and contains no directed cycles, making it a valid directed tree.
This is also an example of a subcubic directed tree, as no vertex has a degree greater than 3. For instance, vertex $2$ has an in-degree of 1 and an out-degree of 2, giving it a total degree of 3.
For more detailed information on directed graphs, please refer to surveys and lecture notes (ex.[100, 99]).
2.2 Symmetric Submodular Function and Connectivity System
The definition of a symmetric submodular function is presented below. This concept is extensively used and discussed in numerous scholarly articles [803, 687, 97]. While symmetric submodular functions are generally defined over real numbers, this paper specifically considers those restricted to natural numbers. This submodular function is also sometimes referred to as the connectivity function [565]. Additionally, a variant of submodular function known as the submodular partition function is also well-known [67, 1261]. Also, $k$-Submodular [675, 1200, 1199], two-dimensional submodular [887], monotone-submodular [442], and maximum-submodular [460] are also known.
Definition 2.26.
Let $X$ be a finite set. A function $f:2^{X}\rightarrow\mathbb{N}$ is called symmetric submodular if it satisfies the following conditions:
  • $\forall A\subseteq X,\ f(A)=f(X\setminus A)$.
  • $\forall A,B\subseteq X,\ f(A)+f(B)\geq f(A\cap B)+f(A\cup B)$.
In this paper, a pair $(X,f)$ of a finite set $X$ and a symmetric submodular function $f$ is called a connectivity system. This concept is frequently used in discussions of graph width parameters, such as branch-width and tree-width, to analyze graph structures (e.g., [346, 511, 505, 565]).
The following is an example illustrating the concept of a symmetric submodular function.
Example 2.27.
Consider a simple undirected graph $G=(V,E)$ where $V$ is the set of vertices and $E$ is the set of edges. Let $V=\{1,2,3,4\}$ and the edges $E=\{(1,2),(2,3),(3,4),(4,1)\}$, forming a cycle. Define a function $f:2^{V}\rightarrow\mathbb{N}$ as follows:
$f(A)=|E(A,V\setminus A)|$
where $E(A,V\setminus A)$ is the set of edges with one endpoint in $A$ and the other endpoint in $V\setminus A$. The symmetric submodularity condition is satisfied.
Example 2.28.
For a set of random variables $X_{1},X_{2},\dots,X_{n}$, the entropy function $H(S)$ for a subset $S\subseteq\{X_{1},X_{2},\dots,X_{n}\}$ is defined as the joint entropy of the variables in $S$. This function is known to be submodular and symmetric. Generally, the entropy function $H(S)$ itself is not symmetric, as it depends on the specific set of variables $S$. However, the mutual information between sets of variables, which is related to entropy, is symmetric. And note that entropy, in information theory, measures the uncertainty or randomness of a random variable’s outcomes. It quantifies the average amount of information produced by a stochastic source of data (cf.[1090, 592]).
Example 2.29.
In the context of network flow, the cut-value function $f(S)$ for a subset $S\subseteq V$ (where $V$ is the set of vertices) is the total capacity of edges crossing from $S$ to $V\setminus S$. This is an example of a symmetric submodular function.
Example 2.30.
Let $\Omega=\{v_{1},v_{2},\dots,v_{n}\}$ be the set of vertices of a directed graph. For any set of vertices $S\subseteq\Omega$, define the function $f(S)$ to denote the number of edges $e=(u,v)$ such that $u\in S$ and $v\in\Omega\setminus S$. This function counts the number of edges leaving the set $S$.
The function $f(S)$ is submodular, meaning that for any two sets $A$ and $B$ with $A\subseteq B\subseteq\Omega$ and any vertex $x\in\Omega\setminus B$, it holds that:
$f(A\cup\{x\})-f(A)\geq f(B\cup\{x\})-f(B).$
However, $f(S)$ is not symmetric. Specifically, $f(S)$ depends on the direction of the edges, so in general, $f(S)\neq f(\Omega\setminus S)$. This asymmetry arises because the number of edges leaving $S$ is typically different from the number of edges entering $S$.
Moreover, this function can be generalized by assigning non-negative weights to the directed edges. In this case, $f(S)$ would represent the total weight of the edges leaving the set $S$, which still maintains the submodular property without being symmetric.
It is known that a symmetric submodular function $f$ satisfies the following useful properties:
Lemma 2.31.
[565] Let $X$ be a finite set. A symmetric submodular function $f$ satisfies:
  1. 1.
    $\forall A\subseteq X,\ f(A)\geq f(\emptyset)=f(X)=0$.
  2. 2.
    $\forall A,B\subseteq X,\ f(A)+f(B)\geq f(A\setminus B)+f(B\setminus A)$.
Proof 2.32.
The following results can be obtained:
  1. 1.
    $f(A)+f(A)=f(A)+f(A)\geq f(A\cup A)+f(A\cap A)=f(X)+f(\emptyset)=f(\emptyset)+f(\emptyset)$.
  2. 2.
    $f(A)+f(B)=f(A)+f(B)\geq f(A\cup B)+f(A\cap B)=f(B\setminus A)+f(A\setminus B)$.
Thus, the proof is completed.
2.3 Ultrafilter for set theory
We consider about ultrafilter on Connectivity System. First, let’s introduce the general concept of an Ultrafilter in set theory. The definition of an Ultrafilter in Set Theory is described as follows.
Definition 2.33.
Let $X$ be a set. A collection $\mathcal{F}\subseteq 2^{X}$ is called a filter on $X$ if it satisfies the following conditions:
  1. (BF1)
    If $A$ and $B$ are both in $\mathcal{F}$, then their intersection $A\cap B$ is also in $\mathcal{F}$.
  2. (BF2)
    If $A$ is in $\mathcal{F}$ and $A\subseteq B\subseteq X$, then $B$ is also in $\mathcal{F}$.
  3. (BF3)
    The empty set $\emptyset$ is not in $\mathcal{F}$.
Definition 2.34.
A maximal filter, which cannot be extended any further while still being a filter, is called an ultrafilter. An ultrafilter satisfies an additional condition:
  1. (BT1)
    For any subset $A\subseteq X$, either $A$ is in $\mathcal{F}$ or its complement $X\setminus A$ is in $\mathcal{F}$, but not both.
Definition 2.35.
(cf.[745, 50]) Given a filter $\mathcal{F}$ on a set $X$, we say a subset $A\subseteq X$ is $\mathcal{F}$-stationary if, for all $B\in\mathcal{F}$, we have $A\cap B\neq\emptyset$.
Example 2.36.
Consider the set $X=\{1,2,3\}$. The collection =$\mathcal{F}=\{\{1,2,3\},\{1,2\},\{2,3\},\{2\}\}$ is a filter on $X$. It satisfies all the conditions of a filter:
  • The intersection of any two sets in $\mathcal{F}$ is also in $\mathcal{F}$.
  • If $A$ is in $\mathcal{F}$ and $A\subseteq B\subseteq X$, then $B$ is also in $\mathcal{F}$.
  • The empty set is not in $\mathcal{F}$.
Example 2.37.
(cf.[562, 713]) Example of a filter is the Fréchet filter on an infinite set $X$. The Fréchet filter consists of all cofinite subsets of $X$ (i.e., subsets whose complements are finite). This collection forms a filter because:
  • The intersection of two cofinite sets is also cofinite.
  • Any superset of a cofinite set is cofinite.
  • The empty set is not cofinite, so it is not in the filter.
Example 2.38 (neighbourhood filter).
(cf.[435, 692]) Let $X$ be a topological space and let $x\in X$ be a point. The collection $\mathcal{N}(x)$, defined as the set of all neighbourhoods of $x$, forms a filter on $X$. This is because:
  1. 1.
    For any neighbourhoods $N_{1},N_{2}\in\mathcal{N}(x)$, their intersection $N_{1}\cap N_{2}$ is also a neighbourhood of $x$, thus $N_{1}\cap N_{2}\in\mathcal{N}(x)$
  2. 2.
    If $N\in\mathcal{N}(x)$ and $N\subseteq N^{\prime}$, where $N^{\prime}$ is an open set containing $x$, then $N^{\prime}\in\mathcal{N}(x)$
  3. 3.
    $x$ belongs to every neighbourhood in $\mathcal{N}(x)$, ensuring that $\mathcal{N}(x)$ is non-empty.
Example 2.39.
Consider the set $X=\mathbb{R}$, the set of all real numbers. The collection =$\mathcal{F}=\{A\subseteq\mathbb{R}\mid\exists M\in\mathbb{R},\ A\supseteq[M,\infty)\}$ is a filter on $\mathbb{R}$. This collection forms a filter because:
  • The intersection of any two sets that contain some $[M,\infty)$ also contains a set of the form $[M^{\prime},\infty)$ for some $M^{\prime}\geq\max\{M_{1},M_{2}\}$, so it is in $\mathcal{F}$.
  • If $A\in\mathcal{F}$ and $A\subseteq B\subseteq\mathbb{R}$, then $B\in\mathcal{F}$.
  • The empty set does not contain any interval of the form $[M,\infty)$, so it is not in the filter.
Example 2.40.
Consider a topological space $(X,\tau)$. The collection of all dense open sets  [^1] Dense open sets are subsets of a topological space that are both open and dense, meaning they intersect every non-empty open set, covering the entire space in a ”thick” way(cf.[916, 593]). in $X$ forms a filter. This is because the intersection of two dense open sets is also a dense open set, and any superset of a dense open set is dense and open. The empty set is not dense, so it is not in the filter.
Example 2.41.
In a measure space $(X,\mathcal{M},\mu)$, where $\mu$ is a measure on the sigma-algebra $\mathcal{M}$  [^2] A sigma-algebra is a collection of subsets of a given set that is closed under complements, countable unions, and countable intersections. It provides the framework for defining measures in measure theory., the collection of all subsets of $X$ that have positive measure (i.e., $\mu(A)>0$) forms a filter. This filter includes all sets with non-zero measure, and satisfies the filter conditions.
Example 2.42 (Voting system).
 [^3] A voting system is a method by which voters select candidates or policies, usually involving preferences or choices aggregated to determine the final outcome, often using mechanisms like majority, plurality, or ranked voting. (cf.[602, 95, 292, 861])
Consider a set $X$ that represents a collection of voters $X=\{v_{1},v_{2},\dots,v_{n}\}$. Each voter has a preference for a particular candidate in an election. We define a collection $\mathcal{F}$ of subsets of $X$ as follows :
  • A subset $A\subseteq X$ belongs to $\mathcal{F}$ if the majority of voters in $A$ support a particular candidate, say Candidate $C$.
This collection $\mathcal{F}$ satisfies the properties of an ultrafilter:
  1. 1.
    Non-emptiness and Completeness: The set $\mathcal{F}$ is non-empty because there exists at least one subset of voters (e.g., the entire set $X$) that supports Candidate $C$.
  2. 2.
    Closure under Intersection: If two subsets $A$ and $B$ are in $\mathcal{F}$, meaning that the majority in both $A$ and $B$ supports Candidate $C$, then their intersection $A\cap B$ also belongs to $\mathcal{F}$ because the majority in $A\cap B$ still supports Candidate $C$.
  3. 3.
    Closure under Supersets: If $A$ is in $\mathcal{F}$ and $A\subseteq B\subseteq X$, then $B$ is also in $\mathcal{F}$. This is because if the majority of $A$ supports Candidate $C$, and $B$ includes all of $A$, then the majority in $B$ still supports $C$.
  4. 4.
    Decisiveness: For any subset $A\subseteq X$, either $A$ or its complement $X\setminus A$ is in $\mathcal{F}$. This corresponds to the idea that, in any group of voters, either the majority supports $C$ or the majority does not, with no undecided or neutral outcomes.
This voting system example illustrates an ultrafilter by capturing the idea of a decisive majority in a set of voters. In this context, the ultrafilter $\mathcal{F}$ helps identify the subsets of voters that consistently support a particular candidate across various possible groupings.
Several related concepts to Filters and Ultrafilters have been proposed. Here, we introduce just one: the Weak Filter. This concept, discussed in the field of logic, is used to interpret defaults through a generalized "most" quantifier in first-order logic [796, 798, 1066, 1064, 797, 1063, 1062, 86, 799, 252]. The definition and an example are provided below.
Definition 2.43.
[796] Let $A,B$ be a set. A weak filter over a non-empty set $X$ is a collection $\mathcal{F}_{w}\subseteq 2^{X}$ that satisfies the following conditions:
  1. (WFB0)w
    If $A\in\mathcal{F}_{w}$ and $A\subseteq B\subseteq X$, then $B\in\mathcal{F}_{w}$.
  2. (WBF1)w
    If $A,B\in\mathcal{F}_{w}$, then $A\cap B\neq\emptyset$.
  3. (WFB2)w
    $X\in\mathcal{F}_{w}$.
Definition 2.44.
[796] Let $A,B$ be a set. A weak ultrafilter over a non-empty set $X$ is a complete weak filter, meaning it additionally satisfies the following condition:
  1. (WFB3)w
    For every $A\subseteq X$, $A\notin\mathcal{F}_{w}$ if and only if $X\setminus A\in\mathcal{F}_{w}$.
Example 2.45.
Let $I=\{1,2,3\}$. Consider the collection $\mathcal{F}_{w}=\{\{1,2,3\},\{1,2\},\{2,3\},\{1,3\}\}$. This collection $\mathcal{F}_{w}$ is a weak filter over $X$ because it satisfies the following:
  • For any $A\in\mathcal{F}_{w}$ and any $A\subseteq B\subseteq X$, we have $B\in\mathcal{F}_{w}$.
  • The intersection of any two sets in $\mathcal{F}_{w}$ is non-empty.
  • The entire set $X=\{1,2,3\}$ is in $\mathcal{F}_{w}$.
Several properties of ultrafilters in set theory are listed below. As will be explained later, in a connectivity system, the focus is on sets where $f(A)\leq k$. Therefore, the properties of ultrafilters in set theory do not necessarily hold in a connectivity system.
Definition 2.46 (Finite intersection property).
(cf.[673]) Let $\mathcal{C}$ be a collection of subsets of $X$. We say that $\mathcal{C}$ has the finite intersection property (FIP) if, for any $D_{1},\dots,D_{n}\in\mathcal{C}$, we have $D_{1}\cap\cdots\cap D_{n}\neq\emptyset$.
Theorem 2.47.
Let $X$ be a finite set, and let $\mathcal{F}$ be a filter on $X$ with $A\subseteq X$. Then, $\mathcal{F}\cup\{A\}$ has the finite intersection property (FIP) if and only if $X\setminus A\notin\mathcal{F}$.
Proof 2.48.
Let $X$ be a finite set, $\mathcal{F}$ be a filter on $X$, and $A\subseteq X$. We aim to prove that $\mathcal{F}\cup\{A\}$ has the finite intersection property (FIP) if and only if $X\setminus A\notin\mathcal{F}$.
Assume that $\mathcal{F}\cup\{A\}$ has the finite intersection property. By definition, this means that for any finite collection of sets $D_{1},D_{2},\dots,D_{n}\in\mathcal{F}\cup\{A\}$, we have $D_{1}\cap D_{2}\cap\dots\cap D_{n}\neq\emptyset$. Now, suppose for contradiction that $X\setminus A\in\mathcal{F}$. Consider the two sets $A$ and $X\setminus A$. Clearly, $A\cap(X\setminus A)=\emptyset$. Since $A\in\mathcal{F}\cup\{A\}$ and $X\setminus A\in\mathcal{F}$, their intersection should be non-empty if $\mathcal{F}\cup\{A\}$ has the FIP. But $A\cap(X\setminus A)=\emptyset$, which contradicts the assumption that $\mathcal{F}\cup\{A\}$ has the FIP. Therefore, it must be that $X\setminus A\notin\mathcal{F}$.
Now assume that $X\setminus A\notin\mathcal{F}$. We need to show that $\mathcal{F}\cup\{A\}$ has the finite intersection property. Consider any finite collection of sets $D_{1},D_{2},\dots,D_{n}\in\mathcal{F}\cup\{A\}$. If none of these sets are equal to $A$, then they all belong to $\mathcal{F}$. Since $\mathcal{F}$ is a filter, we know that $D_{1}\cap D_{2}\cap\dots\cap D_{n}\neq\emptyset$ because filters have the FIP by definition. If one of the sets is $A$, then the intersection becomes $A\cap D_{1}\cap D_{2}\cap\dots\cap D_{n}$, where all other sets are in $\mathcal{F}$. Since $X\setminus A\notin\mathcal{F}$, $A\cap D_{i}\neq\emptyset$ for all $D_{i}\in\mathcal{F}$, meaning that $A\cap D_{1}\cap D_{2}\cap\dots\cap D_{n}\neq\emptyset$. Thus, $\mathcal{F}\cup\{A\}$ has the FIP.
We have shown that $\mathcal{F}\cup\{A\}$ has the finite intersection property if and only if $X\setminus A\notin\mathcal{F}$.
Definition 2.49 (uniform ultrafilter).
(cf.[937]) An ultrafilter $\mathcal{U}$ on $Y$ is called uniform if $|A|=|Y|$ for every $A\in\mathcal{U}$.
Theorem 2.50.
Let $\mathcal{U}$ be an ultrafilter on $Y$, and suppose that $Z\in\mathcal{U}$ has minimal cardinality among the sets in $\mathcal{U}$. Then $\mathcal{U}\cap Z$ is a uniform ultrafilter on $Z$.
Proof 2.51.
We need to show that $\mathcal{U}\cap Z$ is a uniform ultrafilter on $Z$.
Since $\mathcal{U}$ is an ultrafilter on $Y$ and $Z\in\mathcal{U}$, the family $\mathcal{U}\cap Z$ consists of subsets of $Z$ that are in $\mathcal{U}$. To verify that $\mathcal{U}\cap Z$ is an ultrafilter on $Z$, we must confirm that it satisfies the properties of an ultrafilter: - For any set $A\subseteq Z$, either $A\in\mathcal{U}\cap Z$ or $Z\setminus A\in\mathcal{U}\cap Z$. This holds because $Z\in\mathcal{U}$, and $\mathcal{U}$ is an ultrafilter on $Y$, meaning for any subset of $Z$, either it or its complement (within $Z$) is in $\mathcal{U}$. - $\mathcal{U}\cap Z$ contains $Z$ itself, which ensures non-emptiness.
Thus, $\mathcal{U}\cap Z$ is an ultrafilter on $Z$.
By assumption, $Z$ has minimal cardinality among the sets in $\mathcal{U}$. For $\mathcal{U}\cap Z$ to be a uniform ultrafilter on $Z$, we need to show that $|A|=|Z|$ for every $A\in\mathcal{U}\cap Z$.
Since $Z$ is of minimal cardinality, any subset $A\in\mathcal{U}\cap Z$ must be such that $|A|=|Z|$. If $|A|<|Z|$ for some $A\in\mathcal{U}\cap Z$, then $A$ would have a smaller cardinality than $Z$, contradicting the assumption that $Z$ has minimal cardinality in $\mathcal{U}$.
Thus, every set in $\mathcal{U}\cap Z$ has the same cardinality as $Z$, meaning that $\mathcal{U}\cap Z$ is a uniform ultrafilter on $Z$.
We have shown that $\mathcal{U}\cap Z$ is both an ultrafilter and uniform on $Z$, completing the proof.
Theorem 2.52 (ultrafilter on $A$ induced by $\mathcal{U}$).
Let $\mathcal{U}$ be an ultrafilter on $X$ and let $A\in\mathcal{U}$. Define the set
$A\cap\mathcal{U}:=\{A\cap B:B\in\mathcal{U}\}.$
Then $A\cap\mathcal{U}$ is an ultrafilter on $A$.
Proof 2.53.
Since $\mathcal{U}$ is an ultrafilter on $X$, it is non-empty and closed under intersections. For any $B_{1},B_{2}\in\mathcal{U}$, we have $A\cap B_{1}\in A\cap\mathcal{U}$ and $A\cap B_{2}\in A\cap\mathcal{U}$. Therefore,
$(A\cap B_{1})\cap(A\cap B_{2})=A\cap(B_{1}\cap B_{2})\in A\cap\mathcal{U}.$
Thus, $A\cap\mathcal{U}$ is non-empty and closed under intersections.
If $A\cap B\in A\cap\mathcal{U}$ and $A\cap B\subseteq A\cap C$ for some $C\subseteq X$, then $B\subseteq C$ since $\mathcal{U}$ is an ultrafilter on $X$. Thus, $A\cap C\in A\cap\mathcal{U}$, satisfying the superset condition.
For any subset $C\subseteq A$, we need to show that either $C\in A\cap\mathcal{U}$ or $A\setminus C\in A\cap\mathcal{U}$. Since $\mathcal{U}$ is an ultrafilter on $X$, for any subset $C\subseteq A$, either $C\in\mathcal{U}$ or $X\setminus C\in\mathcal{U}$. - If $C\in\mathcal{U}$, then $A\cap C\in A\cap\mathcal{U}$. - If $X\setminus C\in\mathcal{U}$, then $A\cap(X\setminus C)=A\setminus C\in A\cap\mathcal{U}$.
Thus, either $C\in A\cap\mathcal{U}$ or $A\setminus C\in A\cap\mathcal{U}$, satisfying the maximality condition.
We have shown that $A\cap\mathcal{U}$ is non-empty, closed under intersections, contains supersets, and satisfies the maximality condition. Therefore, $A\cap\mathcal{U}$ is an ultrafilter on $A$, called the ultrafilter on $A$ induced by $\mathcal{U}$.
For more detailed information for ultrafilter, please refer to surveys and lecture notes (ex.[579, 1109]).
2.4 Ultrafilter on Connectivity system
We introduce some properties of the ultrafilter on the connectivity system $(X,f)$ as an extension of the Ultrafilter on Boolean Algebras.
First, we consider ultrafilters on graphs. This concept extends the idea of ultrafilters from set theory to graphs and has a dual relationship with tree-width[495]. Refer to Appendix A for the definition of tree-width.
Definition 2.54.
[495] Let $G$ be a graph. A G-Ultrafilter of order $k$ is a family $\mathcal{F}$ of separations of $G$ satisfying the following conditions:
  • (FG0)
    The order of all separations $(A,B)\in\mathcal{F}$ is less than $k$.
  • (FG1)
    For all separations $(A,B)$ of $G$ with order less than $k$, either $(A,B)\in\mathcal{F}$ or $(B,A)\in\mathcal{F}$.
  • (FG2)
    If $(A_{1},B_{1})\in\mathcal{F}$, $A_{1}\subseteq A_{2}$, and $(A_{2},B_{2})$ is a separation of $G$ with order less than $k$, then $(A_{2},B_{2})\in\mathcal{F}$.
  • (FG3)
    If $(A_{1},B_{1})\in\mathcal{F}$ and $(A_{2},B_{2})\in\mathcal{F}$, and $(A_{1}\cap A_{2},B_{1}\cup B_{2})$ is a separation of $G$ with order less than $k$, then $(A_{1}\cap A_{2},B_{1}\cup B_{2})\in\mathcal{F}$.
  • (FG4)
    If $V(A)=V(G)$, then $(A,B)\in\mathcal{F}$.
Next, we introduce definitions of a filter and an ultrafilter on a connectivity system [529]. These concepts extend the traditional notions of filter and ultrafilter from set theory by incorporating the condition of symmetric submodularity. Also, an ultrafilter on the connectivity system $(X,f)$ is co-Maximal ideal [1231] on the connectivity system $(X,f)$.
Definition 2.55 ([529]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a connectivity system, the set family $F\subseteq 2^{X}$ is called a filter of order $k+1$ if the following axioms hold true:  [^4] In studies involving connectivity systems, some papers treat the order as $k$, while others handle it as $k+1$.
  • (Q0)
    $\forall A\in F,\ f(A)\leq k$
  • (Q1)
    $A\in F,\ B\in F,\ f(A\cap B)\leq k\Rightarrow A\cap B\in F$
  • (Q2)
    $A\in F,\ A\subseteq B\subseteq X,\ f(B)\leq k\Rightarrow B\in F$
  • (Q3)
    $\emptyset\notin F$
An Ultrafilter on a connectivity system is a filter on a connectivity system that satisfies the following condition $(Q4)$. The Ultrafilter on a Connectivity System has a dual relationship with branch-width, a graph width parameter [529]. Note that an ultrafilter on a connectivity system is a non-empty and proper (i.e., $\emptyset\notin F$) family.
  • (Q4)
    $\forall A\subseteq X,\ f(A)\leq k\Rightarrow\text{either }A\in F\text{ or }X\setminus A\in F$
Definition 2.56.
[529] A filter is principal if the filter satisfies the following axiom $(QP5)$ [529]:
  • (QP5)
    $A\in F$ for all $A\subseteq X$ with $|A|=1$ and $f(A)\leq k$.
A filter is non-principal if the filter satisfies the following axiom $(Q5)$ [529]:
  • (Q5)
    $A\notin F$ for all $A\subseteq X$ with $|A|=1$.
Non-principal refers to a filter or ideal that does not contain any singletons (i.e., sets with exactly one element). It is not generated by any finite set.
If a filter is weak [497, 795, 87, 506], the following axiom $(QW1^{\prime})$ holds instead of axiom $(Q1)$. Note that weak filter aims at interpreting defaults via a generalized ‘most’ quantifier in first-order logic [795, 87, 1065, 482]. A weak filter on a connectivity system is co-Weak Ideal on a connectivity system:
  • (QW1’)
    $A\in F,\ B\in F,\ f(A\cap B)\leq k\Rightarrow A\cap B\neq\emptyset$
If a filter is quasi [512, 254, 506], the following axiom $(QQ1^{\prime})$ holds instead of axiom $(Q1)$. Note that a quasi-Ultrafilter is used to do an axiomatic analysis of incomplete social judgments [254]:
  • (QQ1’)
    $A\subseteq X,\ B\subseteq X,\ A\notin F,\ B\notin F\Rightarrow A\cup B\notin F$
[…]
  • (QS1)
    For any $A\in F$, $e\in X$, if $f(\{e\})\leq k$ and $f(A\cap(X\setminus\{e\}))\leq k$, then $A\cap(X\setminus\{e\})\in F$.
[…]
  • (QSD1)
    For $A\in F$, $e\in X$, if $f(A\setminus\{e\})\leq k$, then $A\setminus\{e\}\in F$.
If a filter on a connectivity system is connected, the following holds true [507].
  • (CF4)
    For any $A,A^{\prime}\in\mathcal{F}$ with $A\cup A^{\prime}\neq X$ , $f(A\cap A^{\prime})\leq k$ and $f(A\cup A^{\prime})\leq k$, also $A\cap A^{\prime}$ is an element of $\mathcal{F}$.
  • (CF5)
    $\{e\}\notin\mathcal{F}$ for any $e\in X$ such that $f(\{e\})\leq k$ and $\emptyset\notin\mathcal{F}$.
An ultrafilter $\mathcal{U}$ on connectivity system $(X,f)$ is called uniform if $|A|=|X|$ for every $A\in\mathcal{U}$.
Note that an equivalent form of a given ultrafilter $U\subseteq 2^{X}$ on a connectivity system is manifested as a two-valued morphism. This relationship is defined through a function $m$ where $m(A)=1$ if $A$ is an element of $U$, and $m(A)=0$ otherwise. Additionally, the function $m$ holds the value 1 for subsets where the value of the submodular function is at most $k$. This suggests that ultrafilters are well-suited for use in two-player games [84, 1059], such as Cops and Robbers. Indeed, in literature [529], an ultrafilter on a connectivity system is employed as a winning strategy.
2.5 Other concepts related to ultrafilters on a connectivity system
In this subsection, we consider other concepts related to ultrafilters on a connectivity system.
2.5.1 Relation between Tangle and Ultrafilter
A concept closely related to ultrafilters on a connectivity system is the tangle. Tangles are known to be dual to the concepts of tree-decomposition and branch-decomposition, and they are extensively used in the study of graph width parameters [347, 412, 422, 1139, 1050, 411, 565, 966, 1038, 529, 1039, 413, 419, 1143, 340, 526, 47, 342]. Tangles have played a significant role in algorithms across various fields, including graph minors, width parameters, and graph isomorphism problems. It is known that tangles on graphs and ultrafilters on graphs are complementarily equivalent concepts [495]. The definition of a tangle in a graph is provided below [1039].
Definition 2.57.
[1039] A tangle in a graph $G$ of order $k$ is a set $T$ of separations of $G$, each of order less than $k$, such that:
  1. (T1)
    For every separation $(A,B)$ of $G$ of order less than $k$, one of $(A,B)$ or $(B,A)$ is an element of $T$.
  2. (T2)
    If $(A_{1},B_{1}),(A_{2},B_{2}),(A_{3},B_{3})\in T$, then $A_{1}\cup A_{2}\cup A_{3}\neq G$.
  3. (T3)
    If $(A,B)\in T$, then $V(A)\neq V(G)$.
Several related concepts to tangles have been proposed, including Directed Tangle (obstruction of Directeded Tree-width) [573], Hypertangle[16], Connected Tangle (obstruction of Connected Tree-width) [491], Distance-Tangle (obstruction of tree-distance-width) [487], Edge-Tangle[837], Matching Tangle[471], and Linear Tangle (Obstruction of Linear-width)[527].
Subsequently, the concept of a tangle in graphs was extended to matroids and connectivity systems[565]. The definition of a tangle on a connectivity system is provided below.
Definition 2.58 ([565]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. A family $T\subseteq 2^{X}$ is a tangle of order $k+1$ on a connectivity system $(X,f)$ if $T$ satisfies the following axioms:
  • (T1)
    $\forall A\in T,\ f(A)\leq k$,
  • (T2)
    $A\subseteq X,\ f(A)\leq k\Rightarrow\text{either }(A\in T)\text{ or }(X\setminus A\in T)$,
  • (T3)
    $A,B,C\in T\Rightarrow A\cup B\cup C\neq X$,
  • (T4)
    $\forall e\in X,\ X\setminus\{e\}\notin T$.
A tangle of order $k+1$ is also a tangle of order $k$. Moreover, if $k$ is too large, a tangle cannot exist. The maximum $k$ for which a tangle can exist is called the tangle number, which coincides with the branch-width. The same principle applies to ultrafilters. (cf.[513, 565]) Actually, in an ultrafilter on a connectivity system, the following properties hold.
Theorem 2.59.
Let $(X,f)$ be a connectivity system. An ultrafilter of order $k+1$ on a connectivity system $(X,f)$ is also an ultrafilter of order $k$ on a connectivity system $(X,f)$.
Proof 2.60.
An ultrafilter of order $k+1$ on a connectivity system $(X,f)$ obviously holds axioms of an ultrafilter of order $k$ on a connectivity system $(X,f)$.
Given a set $Y\subseteq X$, the collection $\{A\subseteq X\mid A\subseteq Y,f(A)\leq k\}$ forms a tangle. This tangle is referred to as a principal tangle generated by the set $Y$ [598]. The set $Y$ is called the generator of this tangle [343].
A family $F$ on a connectivity system $(X,f)$ is a co-tangle if there is a tangle $T$ such that $A\in F$ if and only if $X\setminus A\in T$. Filter on a connectivity system is a co-tangle on the connectivity system. A filter on a connectivity system also satisfies the following conditions [529]:
  • (FT1)
    $A,B,C\in F\Rightarrow A\cap B\cap C\neq\emptyset$.
A Single-Filter on a connectivity system is a co-linear tangle on that system. It is important to note that a linear tangle is a concept dual to linear-width, which will be discussed later. The main difference between a linear tangle and a co-linear tangle lies in one of the sets in axiom (T2), which is replaced by a singleton set $\{e\}$. A linear tangle of order $k+1$ is also a linear tangle of order $k$. Furthermore, if $k$ is too large, a linear tangle cannot exist. In this paper, the largest possible $k$ for which a linear tangle can exist is referred to as the linear tangle number.
Single-Filter on a connectivity system also satisfies the following conditions [505]:
  • (FLT1)
    $A,B\in F,\ e\in X,\ f(\{e\})\leq k\Rightarrow A\cap B\cap(X\setminus\{e\})\neq\emptyset$.
[…]
  • (FLTR1)
    $A,B\in F,\ |A|\neq|B|,\ e\in X,\ f(\{e\})\leq k\Rightarrow A\cap B\cap(X\setminus\{e\})\neq\emptyset$.
Additionally, concepts such as Directed Ultrafilter [523] , defined on directed graphs, Edge-Ultrafilter [508], and Ultrafilter Tangle [813, 812, 815, 414], on the Abstract separation system, which abstracts the Connectivity system, have also been studied. It is fascinating to see the connection between seemingly unrelated concepts of Ultrafilter and Tangle when the condition of symmetric submodularity is applied.
2.5.2 Bramble and Ultrafilter
It is known that a Bramble can be constructed using a concept essentially equivalent to an Ultrafilter on separation [346]. A Bramble (cf. Strict Bramble[822], Line bramble[641, 640], tight bramble[896]) is recognized in graph game theory, particularly in the Cops and Robbers game, as an obstruction to well-known graph width parameters such as escape routes and tree-width [642, 72].
For reference, the definition of a bramble in a graph is introduced below[1071, 173].
Definition 2.61.
[1071, 173] Let $G=(V,E)$ be a graph. Two subsets $W_{1},W_{2}\subseteq V$ are said to touch if they share at least one vertex or if there is an edge in $E$ connecting them (i.e., $W_{1}\cap W_{2}\neq\emptyset$ or there is an edge $\{w_{1},w_{2}\}\in E$ with $w_{1}\in W_{1}$ and $w_{2}\in W_{2}$). A set $B$ of mutually touching connected vertex sets is called a bramble. A subset of $V$ is said to cover $B$ if it is a hitting set for $B$ (i.e., a set that intersects every element of $B$). The order of a bramble $B$ is the minimum size of a hitting set for $B$. The bramble number of $G$ is the maximum order of all brambles in $G$.
Theorem 2.62.
[1071] Let $k$ be a non-negative integer. A graph has treewidth $k$ if and only if it has bramble number $k+1$.
2.5.3 Ultrafilter, $\pi$-system, and superfilter on a Connectivity System
As stated in the introduction, one of the aims of this paper is to explore Filters and Graph width parameters from various perspectives to deepen understanding. In this section, we examine the relationships between different concepts and filters. First, we consider about $\pi$-system, $\lambda$-system, superfilter on a connectivity system.
  • $\pi$-system: A collection of subsets closed under intersection, ensuring that the intersection of any two subsets remains within the system[279, 722, 398].
  • $\lambda$-system ($d$-system): A collection of subsets closed under complement and countable unions, often used in measure theory to prove the uniqueness of measures[606, 1176, 722].
  • Superfilter: A collection of subsets containing all supersets of its elements and closed under finite unions, generalizing the concept of a filter[1056, 1158, 1159].
Definition 2.63.
Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. A $\pi$-system of order $k+1$ on $(X,f)$ is a collection $P\subseteq 2^{X}$ of subsets of $X$ such that:
  1. 1.
    $P$ is non-empty.
  2. 2.
    If $A,B\in P$ and $f(A\cap B)\leq k$, then $A\cap B\in P$, where $k$ is a fixed integer.
Theorem 2.64.
Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. A filter of order $k+1$ on $(X,f)$ is a $\pi$-system of order $k+1$ on $(X,f)$.
Proof 2.65.
This is evident.
Definition 2.66.
Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. A $\lambda$-system of order $k+1$ on $(X,f)$ is a collection $D\subseteq 2^{X}$ of subsets of $X$ satisfying the following conditions:
  1. 1.
    $X\in D$ and $f(X)\leq k$.
  2. 2.
    If $A\in D$ and $f(X\setminus A)\leq k$, then $X\setminus A\in D$.
  3. 3.
    If $\{A_{n}\}$ is a sequence of pairwise disjoint subsets in $D$ and $f\left(\bigcup_{n=1}^{\infty}A_{n}\right)\leq k$, then $\bigcup_{n=1}^{\infty}A_{n}\in D$.
Theorem 2.67.
Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. A $\pi$-system of order $k+1$ is not a $\lambda$-system of order $k+1$ on $(X,f)$.
Proof 2.68.
This is evident.
Corollary 2.69.
Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. A filter of order $k+1$ on $(X,f)$ is not a $\lambda$-system of order $k+1$ on $(X,f)$.
Definition 2.70 (Superfilter on a Connectivity System).
Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. A nonempty family $S\subseteq 2^{X}$ is called a superfilter on the connectivity system $(X,f)$ if it satisfies the following conditions:
  • (SUF1)
    $\forall A\in S,\ f(A)\leq k$.
  • (SUF2)
    If $A\in S$ and $B\supseteq A$ with $f(B)\leq k$, then $B\in S$.
  • (SUF3)
    If $A\cup B\in S$ with $f(A)\leq k$ and $f(B)\leq k$, then $A\in S$ or $B\in S$.
Theorem 2.71.
Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. An ultrafilter of order $k+1$ on $(X,f)$ is a Superfilter of order $k+1$ on $(X,f)$.
Proof 2.72.
Let $F$ be an ultrafilter of order $k+1$ on the connectivity system $(X,f)$. We need to show that $F$ satisfies the conditions (SUF1), (SUF2), and (SUF3) for a superfilter.
  1. 1.
    Condition (SUF1): By definition of an ultrafilter (Q0), for every $A\in F$, we have $f(A)\leq k$. This directly satisfies (SUF1).
  2. 2.
    Condition (SUF2): Suppose $A\in F$ and $B\supseteq A$ with $f(B)\leq k$. By the condition (Q2) of the ultrafilter, since $A\subseteq B$ and $f(B)\leq k$, it follows that $B\in F$. This satisfies (SUF2).
  3. 3.
    Condition (SUF3): We prove this by contradiction. Assume $A\cup B\in F$, with $f(A)\leq k$ and $f(B)\leq k$, but that $A\notin F$ and $B\notin F$.
    Since $F$ is an ultrafilter, by condition (Q4), for any subset $C\subseteq X$ with $f(C)\leq k$, either $C\in F$ or $X\setminus C\in F$. Thus, if $A\notin F$, then $X\setminus A\in F$. Similarly, since $B\notin F$, $X\setminus B\in F$.
    Now, consider the intersection of the complements: $(X\setminus A)\cap(X\setminus B)=X\setminus(A\cup B).$ By the intersection property of filters (Q1), we have: $X\setminus(A\cup B)\in F.$ However, $A\cup B\in F$ and $X\setminus(A\cup B)\in F$ cannot both be true in an ultrafilter, as this would contradict the maximality condition (Q4).
    Therefore, our assumption that $A\notin F$ and $B\notin F$ must be false. Thus, at least one of $A\in F$ or $B\in F$ must hold, satisfying (SUF3).
2.5.4 Ultrafilter and Grill on a Connectivity System
It is also known that Filters are equivalent to concepts like Grills and Primals, or their complementary equivalents, within a Connectivity system[507]. Note that Grills [1148, 879, 680, 1046, 737] and Primals [363], which are related to filters, serve similar purposes but under different constraints and applications, particularly in topology. And it is also known that Weak-Filters are equivalent to concepts like weak-Grills and weak-Primals, or their complementary equivalents, within a Connectivity system[507].
2.5.5 Ultrafilter and independent set on a Connectivity System
A filter of order $k+1$ on a connectivity system is a co-independent set[935] (also known as a co-abstract simplicial complex[779] or co-stable set[356, 330]) of order $k+1$ on a connectivity system. The concept of an independent set is closely related to matroids, while an abstract simplicial complex is deeply connected to commutative algebra. The definitions are provided below.
Definition 2.73.
Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. Let $I\subseteq 2^{X}$ be a collection of subsets of $X$. A pair $(X,I)$ is called an independence system of order $k+1$ (also called abstract simplicial complex or stable set of order $k+1$) on a a connectivity system if it satisfies the following properties ($I$ is called independent set of order $k+1$):
  • (IN1)
    The empty set is independent, i.e., $\emptyset\in I$.
  • (IN2)
    For every $X\in I$ and for every $Y\subseteq X$ such that $f(Y)\leq k$, we have $Y\in I$.
2.5.6 Ultrafilter and $\sigma$-filter on connectiviy system
And following $\sigma$-filter of order $k+1$ on the connectivity system is a filter on connectivity system. A $\sigma$-filter is a mathematical structure in measure theory, dual to a $\sigma$-ideal. It is closed under supersets and countable intersections, and contains the entire set[112]. Originally, a $\sigma$-filter is considered on a measurable space, but here, we extend the concept by introducing submodular conditions as axioms for general sets.
Definition 2.74.
Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. A subset $F\subseteq 2^{X}$ is called a $\sigma$-filter of order $k+1$ on the connectivity system if it satisfies the following properties:
  • (SIF1)
    $X\in F$.
  • (SIF2)
    If $A,B\subseteq X$ such that $A\in F$ and $A\subseteq B$, and if $f(B)\leq k$, then $B\in F$.
  • (SIF3)
    If $\{A_{n}\}_{n\in\mathbb{N}}\subseteq F$ and $f\left(\bigcap_{n\in\mathbb{N}}A_{n}\right)\leq k$, then $\bigcap_{n\in\mathbb{N}}A_{n}\in F$.
2.5.7 Ultrafilter and Game concepts on a Connectivity System
In the context of lattices and games, closure systems are frequently discussed[251, 7]. When extending to the concept of a closure system of order $k+1$ on a connectivity system, an ultrafilter on a connectivity system can be considered a closure system. The term "closure system" is often referred to by other names, such as Moore family[250, 332, 296], intersection ring (of sets)[1123], protopology[1235], topped intersection structure[949], or intersection semilattice[229]. The definition is provided as follows.
Definition 2.75.
(cf.[251]) Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. A family $\mathcal{F}$ on $(X,f)$ is a closure system of order $k+1$ on a connectivity system if it is $\cap$-stable and contains $X$, i.e., it satisfies the following conditions:
  1. (CL1)
    $F_{1},F_{2}\in\mathcal{F},f(F_{1}\cap F_{2})\leq k\Rightarrow F_{1}\cap F_{2}\in\mathcal{F}$.
  2. (CL2)
    $X\in\mathcal{F}$.
If $F\in\mathcal{F}$, $F$ is said to be $\mathcal{F}$-closed or simply closed. $\mathcal{K}$ will denote the set of all closure systems on $S$.
A related concept is the intersection-closed system on a connectivity system (cf. [114]). While the closure system axiom (CL1) holds, instead of (CL2), we have $\emptyset\in\mathcal{F}$. Furthermore, we can also define the union-closed system on a connectivity system, extending the set-theoretic notion of a union-closed system [1166].
Definition 2.76 (Union-Closed System on a Connectivity System).
(cf. [1166]) Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. A family $\mathcal{F}$ on $X$ is called a union-closed system of order $k+1$ on a connectivity system if it is $\cup$-stable and contains the empty set, i.e., it satisfies the following conditions:
  1. (UC1)
    $F_{1},F_{2}\in\mathcal{F},f(F_{1}\cup F_{2})\leq k\Rightarrow F_{1}\cup F_{2}\in\mathcal{F}$.
  2. (UC2)
    $\emptyset\in\mathcal{F}$, $X\in\mathcal{F}$
This concept is used in the study of cooperative games (cf.[90, 981, 929, 1, 930]), where players form coalitions to maximize their total payoff through cooperation. The payoffs are distributed among the players according to predetermined rules, such as the Shapley value.
Additionally, cooperative games on accessible union-stable systems[56], antimatroids[55, 53]  [^5] Antimatroids are combinatorial structures that model feasible sets with a hereditary and accessible property[1004, 621, 353]. Also called convex geometry[56], simplicial complexes[876], regular set system[819], $k$-regular set systems[1223], building set ($=$principal filter)[791], crossing set-system[739], communication feasible set system[1002], coalition structure[1253], augmenting system[149], Monotone Set Systems[731], S-extremal set systems[897], Chain quasi-building systems[1124], Intersection-closed quasi-building systems[1124], Union stable quasi-building systems[1124], and union-stable systems[52, 54] have been studied. However, it should be noted that a filter on a connectivity system is neither an intersection-closed system nor a union-closed system on a connectivity system.
Question 2.77.
By extending these systems to a connectivity system, could they have any relation to branch-width and ultrafilter, or can they be characterized in some way?
Question 2.78.
Could defining cooperative games on filters reveal any distinct characteristics?
Note that an ultrafilter can also be considered a perfect voting system [355] on a connectivity system. Also we introduce about a simple voting game. A simple voting game (SVG) is a mathematical model used to represent decision-making in a voting system[94].
Definition 2.79.
[94] It consists of a set of voters $X$ and a collection of coalitions $F$ (subsets of $X$) where each coalition either passes or rejects a proposal. The properties of an SVG are:
  1. 1.
    $\varnothing\notin F$ (if no voters support the bill, it is rejected),
  2. 2.
    $A\in F$ (if all voters support the bill, it is passed),
  3. 3.
    If $A\subseteq A^{\prime}$ and $A\in F$, then $A^{\prime}\in F$ (if a coalition supports the bill, any larger coalition also supports it).
The coalitions in $F$ are called winning coalitions. And a proper winning coalition is defined as follows:
A simple voting game $F$ is called proper if for every pair of winning coalitions $A,A^{\prime}\in F$, their intersection is non-empty, i.e., $A\cap A^{\prime}\neq\varnothing$.
We consider extending the concepts of a simple voting game, winning coalitions, and proper winning coalitions to a connectivity system by incorporating the conditions of a submodular function. In this case, an ultrafilter on the connectivity system can be interpreted as the set of winning coalitions and proper winning coalitions within the connectivity system framework.
Other graph width parameters and various related games are also well known. Further details are provided in Appendix F.
2.5.8 Ultrafilter and Majority system on connectivity system
Next, we consider about majority system on connectivity system. A Majority Space is a set system where, for any subset $A$, either $A$ or its complement is a majority. Ultrafilters in set theory are examples of majority systems[1049, 972, 971].
Definition 2.80.
(cf.[1049, 972, 971]) Let $(X,f)$ be a connectivity system, where $X$ is a set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. We will call any set $M\subseteq 2^{X}$ a majority system of order $k+1$ on a connectivity system if it satisfies the following properties:
  1. (MA1)
    If $A\subseteq X,f(A)\leq k$ then either $A\in M$ or $X\setminus A\in M$.
  2. (MA2)
    If $A\in M$, $B\in M$, and $A\cap B=\emptyset$, then $B=X\setminus A$.
  3. (MA3)
    Suppose that $A\in M$ and $F\subseteq A$ is any finite set. If $G$ is any set such that $G\cap A=\emptyset$ and $|F|\leq|G|,f((A\setminus F)\cup G)\leq k$ then $(A\setminus F)\cup G\in M$.
The pair $\langle X,M\rangle$ is called a weak majority space. Given a set $X$, a subset $A\subseteq X$ is called a strict majority (with respect to $M$) if $A\in M$ and $X\setminus A\notin M$. $A$ is called a weak majority if $A\in M$ and $X\setminus A\in M$. Any set $A\in M$ is called a majority set.
Conjecture 2.81.
Let $(X,f)$ be a connectivity system, where $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function, and let $\mathcal{U}$ be a non-principal ultrafilter of order $k+1$ on $(X,f)$. Then $\mathcal{U}$ is a majority space of order $k+1$ on $(X,f)$.
The results discussed so far, along with related findings, can be summarized as follows. In a later section, we will also explore the relationships between Ultrafilters and concepts such as prefilters, subbases, chains, antichains, and sequences.
Theorem 2.82.
If $\mathcal{U}$ is a non-principal ultrafilter of order $k+1$ on a connectivity system, then
  • $\mathcal{U}$ is co-tangle[529] of order $k+1$ on a connectivity system.
  • $\mathcal{U}$ is co-tangle-kit of order $k+1$ on a connectivity system[685].  [^6] The ”Tangle kit” serves as a tool for algorithm development and holds an equivalent relationship with the tangle[685].
  • $\mathcal{U}$ is co-loose-tangle of order $k+1$ on a connectivity system[517]  [^7] In [517], the concept of a ”loose tangle,” a relaxed version of a tangle, was introduced. A maximal loose tangle is equivalent to, or complementary to, the concepts of a tangle and an ultrafilter. Additionally, the ”loose tangle kit” serves as a tool for algorithm development and holds an equivalent relationship with the loose tangle..
  • $\mathcal{U}$ is co-loose-tangle-kit of order $k+1$ on a connectivity system[517].
  • $\mathcal{U}$ is maximal ideal of order $k+1$ on a connectivity system[1231].
  • $\mathcal{U}$ is superfilter of order $k+1$ on a connectivity system.
  • $\mathcal{U}$ is a closure system of order $k+1$ on a connectivity system.
  • $\mathcal{U}$ is $\sigma$-filter of order $k+1$ on the connectivity system.
  • $\mathcal{U}$ is co-independence system of order $k+1$ (also called co-abstract simplicial complex or co-stable set of order $k+1$) on a a connectivity system.
  • $\mathcal{U}$ is $\pi$-system of order $k+1$ on a connectivity system.
  • $\mathcal{U}$ is Grills of order $k+1$ on a connectivity system[507].
  • $\mathcal{U}$ is co-Primals of order $k+1$ on a connectivity system[507].
  • $\mathcal{U}$ is weak-ultrafilter of order $k+1$ on a connectivity system.
  • $\mathcal{U}$ is winning coalition and proper winning coalition of simple voting game on a connectivity system.
  • $\mathcal{U}$ is robber’s winning strategy for monotone search game on a connectivity system in [529].  [^8] A monotone search game involves cops and a robber, where cops aim to capture the robber by progressively choosing subsets while the robber tries to evade capture[529].
  • $\mathcal{U}$ is robber’s winning strategy or open monotone search game on a connectivity system in [529].  [^9] An open monotone search game involves cops and a robber on a set $X$, where the robber reveals a strategy, and cops attempt to capture the robber by selecting subsets of $X$ [529].
  • $\mathcal{U}$ is co-profile of order $k+1$ on a connectivity system[513].  [^10] A profile in separation systems is a collection of consistent, nested separations that satisfy specific properties, often used to distinguish tangles or analyze the structure of graphs in abstract graph theory. Profiles, like tangles, have been the subject of numerous studies [420, 246, 351, 428, 412, 677].
2.6 Branch width on Connectivity System
In this section, we consider about branch width on connectivity system.
2.6.1 Branch width
Branch width is a significant graph width parameter that involves a branch decomposition where the leaves correspond to the graph’s edges. Each edge is assigned a value from a symmetric submodular function, measuring connectivity. Branch width generalizes the width of symmetric submodular functions on graphs, making it crucial in graph theory (e.g., [565, 529, 566]).
First, we introduce the definition of a branch decomposition for undirected graphs [177].
Definition 2.83.
[177] A branch decomposition of a graph $G$ is a pair $(T,\tau)$, where $T$ is a tree whose vertices have degree 1 or 3, and $\tau$ is a bijection from the set of leaves of $T$ to the set of edges $E(G)$. The order of an edge $e$ in $T$ is defined as the number of vertices $v\in V(G)$ such that there exist leaves $\tau_{1},\tau_{2}$ in $T$ lying in different components of $T-e$, with both $\tau(\tau_{1})$ and $\tau(\tau_{2})$ incident to $v$ in $G$. The width of the branch decomposition $(T,\tau)$ is the maximum order of any edge in $T$. The branchwidth of $G$ is the minimum width over all possible branch decompositions of $G$.
In cases where $|E(G)|=1$, the branch-width of $G$ is defined to be 0; if $|E(G)|=0$, then $G$ has no branch decomposition. For a graph $G$ with $|E(G)|=1$, the branch decomposition consists of a tree with a single vertex, and the width of this branch decomposition is also considered to be 0.
Note that tree-width[1039] is a measure that is based on a tree decomposition of the graph, which involves grouping vertices into "bags" and examining how these bags connect. But branch-width is related to a branch decomposition of the graph, which involves splitting the graph into parts and looking at how the edges connect across these parts. Both concepts are related to the idea of how "complicated" the graph’s structure is, but they approach this complexity from slightly different angles: tree-width from the perspective of vertices and their connectivity, and branch-width from the perspective of edges.
The definition of branch decomposition generalized to a connectivity system is provided below[966].
Definition 2.84 ([966]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. Let $(X,f)$ be a connectivity system. The pair $(T,\mu)$ is a branch decomposition tree of $(X,f)$ if $T$ is a ternary tree such that $|L(T)|=|X|$ and $\mu$ is a bijection from $L(T)$ to $X$, where $L(T)$ denotes the leaves in $T$. For each $e\in E(T)$, we define $\text{bw}(T,\mu,e)$ as $f\left(\bigcup_{v\in L(T_{1})}\mu(v)\right)$, where $T_{1}$ is a tree obtained by removing $e$ from $T$ (taking into account the symmetry property of $f$). The width of $(T,\mu)$ is defined as the maximum value among $\text{bw}(T,\mu,e)$ for all $e\in E(T)$. The branch-width of $X$, denoted by $\text{bw}(X)$, is defined as the minimum width among all possible branch decomposition trees of $X$.
Example 2.85.
Consider the connectivity system $(X,f)$ where $X=\{a,b,c,d\}$ and $f$ is defined as follows: for any subset $S\subseteq X$, $f(S)$ equals the number of elements in $S$. We construct a branch decomposition tree $T$ as follows:
  • $T$ is a ternary tree with leaves $L(T)=\{v_{a},v_{b},v_{c},v_{d}\}$.
  • The bijection $\mu$ maps $v_{a}$ to $a$, $v_{b}$ to $b$, $v_{c}$ to $c$, and $v_{d}$ to $d$.
To define the width of this branch decomposition tree, we consider each edge $e\in E(T)$ and the corresponding tree $T_{1}$ obtained by removing $e$ from $T$. For instance, remove an edge $e$ that splits $T$ into subtrees $T_{1}$ and $T_{2}$ where $T_{1}$ contains leaves $\{v_{a},v_{b}\}$ and $T_{2}$ contains leaves $\{v_{c},v_{d}\}$. In this case, $\mu(\{v_{a},v_{b}\})=\{a,b\}$ and $\mu(\{v_{c},v_{d}\})=\{c,d\}$. The width $\text{bw}(T,\mu,e)$ is then calculated as:
$\text{bw}(T,\mu,e)=f(\{a,b\})=2$
Repeating this process for all edges $e$ in $T$, we determine the maximum width among all these values. In this simple example, the width $\text{bw}(T,\mu)$ is the maximum value of $f(S)$ for any partition $S$ of $X$, which is 2. Finally, the branch-width $\text{bw}(X)$ of the connectivity system $(X,f)$ is the minimum width among all possible branch decomposition trees of $X$. Here, the minimum width is achieved by our constructed tree $T$, and thus $\text{bw}(X)=2$.
In graph theory, the duality theorem for width parameters, such as tree-width and branch-width, is discussed, highlighting their dual concepts like tangles and branch decompositions[565, 966, 208, 505, 529, 48]. Additionally, obstructions are minimal structures or subgraphs that prevent a graph from having a width parameter below a certain threshold, providing insight into the graph’s complexity. The following duality theorem is known for the branch-width of the connectivity system $(X,f)$ and the ultrafilter of the connectivity system $(X,f)$.
Theorem 2.86 ([529]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. The branch-width of the connectivity system $(X,f)$ is at most $k$ if and only if no (non-principal) ultrafilter of order $k+1$ exists.
Also, the following duality theorem is known for the branch-width of the connectivity system $(X,f)$ and the tangle of the connectivity system $(X,f)$ [565]. The following duality theorem is also known for the branch-width of the connectivity system $(X,f)$ and the tangle of the connectivity system $(X,f)$ [565]. Additionally, a concept similar to the tangle, known as a loose tangle and loose tangle kit, is also recognized to have a duality relationship [966]. It is known that a loose tangle has a complementary equivalence relationship with a filter [517].
Theorem 2.87 ([565]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. The branch-width of the connectivity system $(X,f)$ is at most $k$ if and only if no tangle of order $k+1$ exists.
2.6.2 Linear Branch width
Here, we introduce the concept of linear decomposition, which is a linear version of branch decomposition. Like branch decomposition, linear branch decomposition has been the subject of extensive research [617, 527, 524]. Focusing on linear structures often facilitates deriving results for both general width parameters and linear width parameters.
Definition 2.88.
(cf.[525]) Let $X$ be a finite set and $f$ be a symmetric submodular function. Let $C$ be a caterpillar, defined as a tree where interior vertices have a degree of 3 and leaves have a degree of 1. Consider $C$ as the path $(l_{1},b_{2},b_{3},\ldots,b_{n-1},l_{n})$. For $12\leq i\leq n-$ 1, the subgraph of $C$ induced by $\{b_{i-1},b_{i},b_{i+1}\}$ forms a connectivity system $(X,f)$. The Linear Decomposition of $C$ is a process that partitions the elements of $X$ into the sets $\{e_{1}\},\{e_{2}\},\ldots,\{e_{n-1}\},\{e_{n}\}$. For each $11\leq i\leq n-$ 1, define $w_{i}:=f(\{e_{1},\ldots,e_{i}\})$. The width of the Linear Decomposition is given by $\max\{w_{1},\ldots,w_{n-1},f(e_{1}),\ldots,f(e_{n-1}),f(e_{n})\}$. The linear width of $(X,f)$ is the smallest width among all Linear Decompositions of $(X,f)$.
Example 2.89.
Consider a finite set $X=\{a,b,c,d\}$ and a symmetric submodular function $f$ defined on $2^{X}$. Suppose $C$ is a caterpillar tree, where the path is represented by vertices $l_{1},b_{2},b_{3},l_{n}$ with edges between them. Let’s assume the caterpillar has 4 vertices with the structure $l_{1}-b_{2}-b_{3}-l_{n}$, where $l_{1}$ and $l_{n}$ are leaves, and $b_{2}$ and $b_{3}$ are interior vertices. In this scenario, we define a Linear Decomposition of $X$ by partitioning it into individual elements $\{a\},\{b\},\{c\},\{d\}$. For each partition step $i$, where $1\leq i\leq 3$, the width $w_{i}$ is calculated using the submodular function $f$.
  • For example, $w_{1}=f(\{a\})$, $w_{2}=f(\{a,b\})$, $w_{3}=f(\{a,b,c\})$.
  • If the subgraph of $C$ induced by $\{b_{2},b_{3}\}$ forms the connectivity system with the function $f(\{b_{2}\})=2$, $f(\{b_{3}\})=3$, and $f(\{b_{2},b_{3}\})=2$, then the width of this Linear Decomposition would be the maximum value among these: $\max\{w_{1},w_{2},w_{3},f(\{b_{2}\}),f(\{b_{3}\})\}$.
Finally, the linear branch width of $(X,f)$ is the smallest width obtained by considering all possible Linear Decompositions of $(X,f)$. In this case, the linear width would be determined by minimizing the maximum width $w_{i}$ across different decompositions of the caterpillar tree.
A single filter on a connectivity system is known to have a dual relationship with linear-width[514]. Similarly, both linear tangles[528] and linear loose tangles[511] are also known to exhibit this duality.
Theorem 2.90 ([505]).
Let $X$ be a finite set and $f$ be a symmetric submodular function. The linear-branch-width of the connectivity system $(X,f)$ is at most $k$ if and only if no (non-principal) Single-Ultrafilter of order $k+1$ exists.
3 Tukey’s Lemma and Chain for Connectivity Systems
In this section, we explain some properties of ultrafilters on connectivity systems.
3.1 Chain for Connectivity Systems
First, let’s consider a chain on a connectivity system. Note that a chain in mathematics is crucial because it represents a totally ordered sequence of elements, which is fundamental in studying hierarchical structures, lattice theory, and optimization problems [113, 845, 382, 846, 639, 313, 577, 478].
Definition 3.1.
Let $X$ be a finite set and $f$ be a symmetric submodular function. A chain of order $k+1$ on a connectivity system is a sequence of subsets $\{A_{1},A_{2},\dots,A_{m}\}$ of $X$ such that:  [^11] It is unclear whether there is a direct connection, but in game theory, monotone strategies are sometimes examined under conditions similar to those used for chains (cf.[472, 147]). 121212 There are various concepts similar to that of a chain. In particular, the notion of a filtration has been introduced in several areas of mathematics. This will be introduced in a footnote. Let $X$ be a finite set and $f$ be a symmetric submodular function. The following is an introduction to this concept, with the condition of symmetric submodularity added. Although not listed below, other concepts such as filtered probability spaces, Helly family, stopped $\sigma$-algebra, stopped $\sigma$-field, and $D$-generic are also well-known [1003, 180, 773, 733]. A filtered algebra over a field $X_{A}$ of order $k+1$ is an algebra $(A,\cdot)$ with an increasing sequence of subspaces(cf.[1114, 21]): $\{0\}\subseteq A_{0}\subseteq A_{1}\subseteq\dots\subseteq A_{i}\subseteq\dots\subseteq A,$ such that: $A=\bigcup_{i\in\mathbb{N}}A_{i}$ and for each $i$ $(1\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$. In an algebraic structure, a filtration $\mathcal{F}=(A_{i})_{i\in I}$ of order $k+1$ is an indexed family of a given algebraic structure $A$, where the index $i\in I$ runs over a totally ordered set. The filtration must satisfy (cf.[458, 322]): $i\leq j,f(A_{i})\leq k,f(A_{j})\leq k\implies A_{i}\subseteq A_{j}\quad\text{for all}\quad i,j\in I.$ For a group $G$, filtrations of order $k+1$ are often indexed by $\mathbb{N}$. A filtration of order $k+1$ is a nested sequence of order $k+1$ (descending filtration) $G_{i}$ of normal subgroups, such that(cf.[98, 1211]): $G_{i+1}\subseteq G_{i}\quad\text{for all}\quad i\in\mathbb{N}$ such that for each $i$ $(1\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $G_{i}$ satisfies $f(G_{i})\leq k$. In linear algebra, a flag of order $k+1$ is an sequence of subspaces of a finite-dimensional vector space $X_{V}$, where each subspace is a proper subset of the next (cf.[316, 1122, 1074]): $\{0\}=A_{0}\subset A_{1}\subset A_{2}\subset\dots\subset A_{k}=X_{V}.$ such that for each $i$ $(1\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$.
  1. 1.
    $A_{1}\subseteq A_{2}\subseteq\dots\subseteq A_{m}$.
  2. 2.
    For each $i$ $(1\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$.
3.2 Tukey’s Lemma for Connectivity System
We demonstrate the theorem (Tukey’s Lemma for Connectivity Systems) by using a chain of order $k+1$ on a connectivity system. Tukey’s Lemma states that in any non-empty collection of sets closed under taking supersets, there exists a maximal set within that collection. Intuitively, this means you can always find the largest possible set that cannot be expanded further while staying within the collection. Additionally, in set theory, Tukey’s Lemma is known to be closely related to Zorn’s Lemma (e.g., [830, 238, 117]). These lemmas are also widely recognized for their broad range of applications.
Theorem 3.2.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Any filter on a connectivity system $(X,f)$, where $f$ is a symmetric submodular function and $X$ is a non-empty finite set, can be extended to an ultrafilter. Specifically, for any $k$-efficient subset $A\subseteq X$ (i.e., $f(A)\leq k$), there exists an ultrafilter containing the filter.
Proof 3.3.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Here, we consider chains composed of $k$-efficient subsets.
First, we define the initial filter. Let $Q_{0}$ be a filter on the connectivity system $(X,f)$ where $f$ is a symmetric submodular function and $X$ is a non-empty finite set that satisfies the axioms $Q_{0}$ through $Q_{3}$.
Next, define the collection $\mathcal{F}$ as the set of all filters on $(X,f)$ that contain $Q_{0}$. This means $\mathcal{F}$ is the set of all sets $Q$ such that:
  • $Q_{0}\subseteq Q$
  • $Q$ satisfies axioms $Q_{0}$ through $Q_{3}$
$\mathcal{F}$ is partially ordered by inclusion. We need to show that every chain on $(X,f)$ (totally ordered subset on $(X,f)$) in $\mathcal{F}$ has an upper bound in $\mathcal{F}$.
Let $C\subseteq\mathcal{F}$ be a chain in $\mathcal{F}$, where each element of $C$ is a filter that respects the $k$-efficiency condition, meaning $f(A)\leq k$ for every $A$ in the filter.
Define $Q_{C}=\bigcup C$. We need to show that $Q_{C}$ is a filter and an element of $\mathcal{F}$.
Axiom (Q0): This axiom obviously holds.
Axiom (Q2): For any $A,B\in Q_{C}$, there exist $Q_{A},Q_{B}\in C$ such that $A\in Q_{A}$ and $B\in Q_{B}$. Since $C$ is a chain on the connectivity system, one of these filters contains the other; assume without loss of generality that $Q_{A}\subseteq Q_{B}$. Thus, $A,B\in Q_{B}$, and because $Q_{B}$ is a filter, $A\cap B\in Q_{B}\subseteq Q_{C}$. Additionally, since both $A$ and $B$ are in $Q_{C}$, $f(A)\leq k$ and $f(B)\leq k$ by Axiom (Q0). Therefore, $f(A\cap B)\leq k$ must hold because if it did not, $A\cap B$ would not be in $Q_{B}$, contradicting the closure under intersection in Axiom (Q1).
Axiom (Q1): For any $A\in Q_{C}$ and $A\subseteq B\subseteq X$, there exists $Q_{A}\in C$ such that $A\in Q_{A}$. Since $Q_{A}$ is a filter and $f(B)\leq k$, $B\in Q_{A}\subseteq Q_{C}$.
Axiom (Q3): Since $\emptyset\notin Q$ for any $Q\in C$, we have $\emptyset\notin Q_{C}$.
Thus, $Q_{C}$ is a filter on the connectivity system and $Q_{C}\in\mathcal{F}$.
$\mathcal{F}$ contains a maximal element $Q$. This maximal filter $Q$ satisfies axioms (Q0) through (Q3).
To show $Q$ is an ultrafilter, we need to verify Axiom (Q4): For any $A\subseteq X$ such that $f(A)\leq k$, suppose $A\notin Q$. We must show $X\setminus A\in Q$.
Assume $A\notin Q$. If $X\setminus A\notin Q$, then we will show that this leads to a contradiction, meaning $Q$ must contain either $A$ or $X\setminus A$.
  1. 1.
    Constructing $Q^{\prime}=Q\cup\{A\}$: If $A\notin Q$, consider adding $A$ to $Q$. This new set $Q^{\prime}=Q\cup\{A\}$ must be checked against the filter axioms.
  2. 2.
    Closed under Intersection: For $Q^{\prime}$, consider any $B\in Q$:
    • If $A$ were added to $Q$, $B\cap A$ might violate the $k$-efficiency condition. Specifically, since $A\notin Q$ and $Q$ is maximal, there could be an element $B\in Q$ such that $B\cap A$ does not satisfy $f(B\cap A)\leq k$. Since $B\cap A\notin Q$, this would violate Axiom (Q1), meaning $Q^{\prime}$ cannot be a filter.
  3. 3.
    Constructing $Q^{\prime\prime}=Q\cup\{X\setminus A\}$: If $X\setminus A\notin Q$, consider adding $X\setminus A$ to $Q$. This new set $Q^{\prime\prime}=Q\cup\{X\setminus A\}$ must also be checked against the filter axioms.
  4. 4.
    Closed under Intersection: For $Q^{\prime\prime}$, consider any $B\in Q$:
    • If $X\setminus A\notin Q$, then there must exist some $B\in Q$ such that $B\cap(X\setminus A)$ does not satisfy $f(B\cap(X\setminus A))\leq k$. Since $Q$ is maximal, adding $X\setminus A$ would create a set that fails to be closed under intersection, as $B\cap(X\setminus A)\notin Q$. This violates Axiom (Q1), meaning $Q^{\prime\prime}$ cannot be a filter.
Therefore, the assumption that both $A\notin Q$ and $X\setminus A\notin Q$ leads to contradictions. Hence, $Q$ must contain either $A$ or $X\setminus A$. This proof is completed.
We gain the following property of an ultrafilter on a connectivity system.
Theorem 3.4.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In any finite non-empty connectivity system $(X,f)$ where $f$ is a symmetric submodular function, there exists at least one ultrafilter.
Proof 3.5.
Consider the collection $\mathcal{F}$ of all filters on $(X,f)$. This collection is non-empty because the trivial filter $\{X\}$ exists. By Tukey’s Lemma for Connectivity Systems, every filter can be extended to an ultrafilter. Thus, there exists at least one maximal element in $\mathcal{F}$ under inclusion that satisfies all the filter axioms and the ultrafilter condition $(Q4)$. This proof is completed.
3.3 Antichain for Connectivity Systems: Relationship among Ultrafilter, Chain, and Antichain
We consider an antichain on a connectivity system. An antichain is a collection of elements in a partially ordered set where no element is comparable to another, meaning no element in the set precedes or follows any other element in the ordering (cf. [113, 845, 382, 846, 639, 577, 478]).
Definition 3.6.
Let $X$ be a finite set and $f$ a symmetric submodular function. An antichain of order $k+1$ on a connectivity system is defined as a collection of subsets $\{A_{1},A_{2},\dots,A_{m}\}$ of $X$ such that:
  1. 1.
    Antichain Condition: For any two distinct subsets $A_{i}$ and $A_{j}$ in the collection, neither is a subset of the other, i.e., $A_{i}\not\subseteq A_{j}$ and $A_{j}\not\subseteq A_{i}$ for all $1\leq i,j\leq m$ with $i\neq j$.
  2. 2.
    Submodular Condition: For each $i$ $(1\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$.
The relationship between chains and antichains is established by the following theorem. In this paper, the following theorem is called Dilworth’s Theorem [970, 120, 540, 903] on a connectivity system.
Theorem 3.7 (Dilworth’s Theorem for Connectivity Systems).
Let $(X,f)$ be a connectivity system with a symmetric submodular function $f$ and let $\mathcal{A}\subseteq 2^{X}$ be a family of subsets of $X$ such that $f(A)\leq k$ for all $A\in\mathcal{A}$. Then, the size of the largest antichain in $\mathcal{A}$ is equal to the minimum number of chains of order $k+1$ needed to cover $\mathcal{A}$.
Proof 3.8.
Let $X$ be a finite set and $f$ a symmetric submodular function. We proceed by induction on the number of elements in $X$.
For $|X|=1$, the statement trivially holds because any non-empty subset $\mathcal{A}$ of $X$ is both a chain and an antichain.
Assume the statement holds for all connectivity systems with less than $|X|$ elements. Consider a connectivity system $(X,f)$ with $|X|$ elements. Let $\mathcal{A}$ be a family of subsets of $X$ such that $f(A)\leq k$ for all $A\in\mathcal{A}$. We need to prove that the size of the largest antichain in $\mathcal{A}$ is equal to the minimum number of chains required to cover $\mathcal{A}$.
  • Let $\mathcal{A}$ be an antichain such that no element in $\mathcal{A}$ is a subset of any other. Due to the submodularity condition, the maximum value of $f(A\cap B)$ for any $A,B\in\mathcal{A}$ is less than or equal to $k$, ensuring the sets in the antichain are maximally disconnected.
  • Given any partition of $X$ into chains, by the pigeonhole principle, there is at least one element $A$ in $\mathcal{A}$ that intersects every chain. Each of these chains has an associated subset $A$ that satisfies $f(A)\leq k$, ensuring that the number of chains required to cover $\mathcal{A}$ is equal to the largest antichain.
By induction, the theorem holds for all finite connectivity systems. This proof is completed.
The relationship between ultrafilters and antichains is established by the following theorem.
Theorem 3.9.
Let $X$ be a finite set and $f$ be a symmetric submodular function. A maximal antichain on a connectivity system $(X,f)$ intersects with every ultrafilter of order $k+1$.
Proof 3.10.
Let $X$ be a finite set and $f$ a symmetric submodular function. Suppose $\{B_{1},B_{2},\dots,B_{n}\}$ is a maximal antichain of order $k+1$ on the connectivity system $(X,f)$. Let $U$ be an ultrafilter of order $k+1$ on $(X,f)$. Note that by definition, an ultrafilter $U$ satisfies: For any $A\subseteq X$ such that $f(A)\leq k$, either $A\in U$ or $X\setminus A\in U$ (this is the defining property of an ultrafilter on a connectivity system).
Now, consider each set $B_{i}$ in the antichain. Since $B_{i}$ are mutually incomparable and $f(B_{i})\leq k$ for all $i$, at least one of the sets $B_{i}$ must belong to the ultrafilter $U$, otherwise, the union $X\setminus\bigcup_{i=1}^{n}B_{i}$ would belong to $U$, which contradicts the maximality of the antichain (since $U$ would then not intersect with any $B_{i}$).
Hence, every ultrafilter of order $k+1$ intersects with the maximal antichain $\{B_{1},B_{2},\dots,B_{n}\}$. This proof is completed.
The relationship between ultrafilters and chains is established by the following theorem.
Theorem 3.11.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a chain of order $k+1$ on a connectivity system $(X,f)$, every ultrafilter of order $k+1$ contains exactly one set from the chain.
Proof 3.12.
Let $X$ be a finite set and $f$ a symmetric submodular function. Let $\{A_{1},A_{2},\dots,A_{m}\}$ be a chain of order $k+1$ on the connectivity system $(X,f)$, and let $U$ be an ultrafilter of order $k+1$ on $(X,f)$. By the properties of the chain, we have $A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{m}$ with $f(A_{i})\leq k$ for each $i$. By the properties of the ultrafilter, for each set $A_{i}$, either $A_{i}\in U$ or $X\setminus A_{i}\in U$. Since the sets in the chain are nested, the ultrafilter cannot contain two different sets $A_{i}$ and $A_{j}$ with $i\neq j$ without violating the ultrafilter condition. Therefore, there must be a unique $A_{j}$ such that $A_{j}\in U$ and for all $i<j$, $A_{i}\notin U$. This proof is completed.
Theorem 3.13.
Let $(X,f)$ be a connectivity system with $X$ as a finite set and $f$ as a symmetric submodular function. The following holds:
  1. 1.
    Chain Extension: Every chain in the set of subsets of $X$, where $f(A)\leq k$ for all $A$ in the chain, can be extended to an ultrafilter on the connectivity system $(X,f)$.
  2. 2.
    Ultrafilter Inducing a Chain: Conversely, every ultrafilter on the connectivity system $(X,f)$ induces a maximal chain in the set of subsets of $X$.
Proof 3.14.
  1. 1.
    Chain Extension to Ultrafilter: Let $X$ be a finite set and $f$ a symmetric submodular function. Given a chain $C=\{A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{m}\}$ where $f(A_{i})\leq k$ for each $i$, we want to show that this chain can be extended to an ultrafilter on the connectivity system $(X,f)$. Consider the filter $F$ generated by the chain $C$. This filter includes all supersets of elements in $C$ that satisfy $f(A)\leq k$. The symmetric submodular condition ensures that the intersection of any two sets in $F$ also belongs to $F$, maintaining the filter structure. By Theorem 3.2, this filter $F$ can be extended to a maximal filter, which is an ultrafilter $U$. This ultrafilter $U$ satisfies the condition that for any set $B\subseteq X$, either $B\in U$ or $X\setminus B\in U$, ensuring that $f(B)\leq k$ for all $B\in U$. Therefore, every chain $C$ in the set of subsets of $X$ can be extended to an ultrafilter $U$ on the connectivity system $(X,f)$.
  2. 2.
    Ultrafilter Inducing a Maximal Chain: Let $X$ be a finite set and $f$ a symmetric submodular function. Given an ultrafilter $U$ on the connectivity system $(X,f)$, we aim to show that $U$ induces a maximal chain in the set of subsets of $X$. For any set $A\in U$, consider the collection of subsets of $A$ that are also in $U$. This collection forms a chain because $U$ is a maximal filter, meaning that for any subset $A\subseteq B\subseteq X$, $B\in U$ if $A\in U$ and $f(B)\leq k$. The ultrafilter $U$ ensures that for any set $B\subseteq X$, either $B\in U$ or $X\setminus B\in U$, which means that the chain induced by $U$ cannot be extended further within $U$. This guarantees that the chain is maximal.
This proof is completed.
Theorem 3.15.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a chain of order $k+1$ on a connectivity system $(X,f)$, every ultrafilter of order $k$ does not contain a maximal set from the chain of order $k+1$.
Proof 3.16.
Consider a chain $\{A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{m}\}$ in the connectivity system $(X,f)$, where $f(A_{i})\leq k$ for all $i$. This chain is said to be of order $k+1$ if $f(A_{i})\leq k$ for all sets in the chain. An ultrafilter of order $k$ on $(X,f)$ is a maximal filter where $f(A)<k$ for all $A$ in the ultrafilter. Suppose there exists an ultrafilter $U$ of order $k$ that contains the maximal set $A_{m}$ from the chain $\{A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{m}\}$. Since $f$ is symmetric and submodular, $f(A_{m})\leq k$ holds because the chain is of order $k+1$. However, if $f(A_{m})=k$, then $A_{m}$ cannot belong to the ultrafilter $U$ of order $k$, because by definition, $U$ only contains sets where $f(A)<k$. This contradiction implies that the ultrafilter $U$ of order $k$ cannot contain the maximal set $A_{m}$. Therefore, every ultrafilter of order $k$ does not contain a maximal set from a chain of order $k+1$.
Theorem 3.17.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a chain of order $k$ on a connectivity system $(X,f)$, if no chain of order $k+1$ exists, then no ultrafilter of order $k+1$ exists.
Proof 3.18.
Let $\{A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{m}\}$ be a chain of order $k$ in the connectivity system $(X,f)$, where $f(A_{i})\leq k$ for all $i$. Suppose no chain of order $k+1$ exists, meaning there is no set $A\subseteq X$ such that $f(A)\leq k$. Assume there exists an ultrafilter $U$ of order $k+1$ on $(X,f)$. By definition, an ultrafilter of order $k+1$ contains sets $B\subseteq X$ where $f(B)\leq k$. If such a set $B$ exists in $U$, then $B$ should be part of a chain of order $k+1$. However, since we assumed that no chain of order $k+1$ exists, no set $B\subseteq X$ can satisfy $f(B)\leq k$. This proof is completed.
In the future, we will consider the relationship between a chain on a connectivity system and branch-width on a connectivity system.
3.4 Sequence of connectivity system
Next, we consider about sequence chain (strong chain).
Definition 3.19.
Let $X$ be a finite set and $f$ be a symmetric submodular function. A sequence chain of order $k+1$ on a connectivity system $(X,f)$ is a sequence chain of subsets $\{A_{0},A_{1},\dots,A_{m}\}$ of $X$ such that:
  1. 1.
    $A_{0}=\emptyset$ and $A_{m}=X$.
  2. 2.
    $A_{0}\subseteq A_{1}\subseteq\dots\subseteq A_{m}$.
  3. 3.
    For each $i$ $(0\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$.
Theorem 3.20.
Given a finite set $X$ and a symmetric submodular function $f:2^{X}\rightarrow\mathbb{N}$, if there exists a sequence chain $\{A_{0},A_{1},\dots,A_{m}\}$ of order $k+1$ on the connectivity system $(X,f)$, then no antichain of order $k+1$ can exist on the same system.
Proof 3.21.
Assume that there exists a sequence chain $\{A_{0},A_{1},\dots,A_{m}\}$ of order $k+1$ on the connectivity system $(X,f)$. By the definition of a sequence chain:
  1. 1.
    The sequence chain is a nested chain of subsets:
    $A_{0}\subseteq A_{1}\subseteq\dots\subseteq A_{m},$
    where $A_{0}=\emptyset$, $A_{m}=X$, and for each $i$, $f(A_{i})\leq k$.
Now, suppose, for the sake of contradiction, that an antichain $\{B_{1},B_{2},\dots,B_{n}\}$ of order $k+1$ also exists on the same connectivity system. By the definition of an antichain:
  1. 1.
    The antichain is a collection of subsets where no two distinct subsets are comparable by inclusion:
    $B_{i}\not\subseteq B_{j}\text{ and }B_{j}\not\subseteq B_{i}\quad\text{for all}1\leq i,j\leq n,\ i\neq j,$
    and for each $i$, $f(B_{i})\leq k$.
Since $\{A_{0},A_{1},\dots,A_{m}\}$ is a sequence chain, for each $B_{i}$ in the antichain, it must relate to the sets in the sequence chain in the following way:
  • Comparability by Inclusion: Since the sequence chain $\{A_{0},A_{1},\dots,A_{m}\}$ is a chain, each $B_{i}$ in the antichain must be comparable by inclusion with each set in the sequence chain. Specifically, for each $B_{i}$, there must exist some $A_{j}$ in the sequence chain such that either: $B_{i}\subseteq A_{j}\quad\text{or}\quad A_{j}\subseteq B_{i}.$
    • If $B_{i}\subseteq A_{j}$: Then $B_{i}$ is contained within $A_{j}$, and since $A_{j}$ is part of the sequence chain, this violates the antichain condition because $B_{i}$ is now comparable to $A_{j}$ by inclusion.
    • If $A_{j}\subseteq B_{i}$: Then $A_{j}$ is contained within $B_{i}$, and similarly, this violates the antichain condition because $B_{i}$ is now comparable to $A_{j}$ by inclusion.
Given that the sequence chain imposes a strict order by inclusion on the subsets, any subset $B_{i}$ in the antichain would necessarily have to be either a subset of some $A_{j}$ or a superset of some $A_{j}$. This directly contradicts the requirement that no two distinct subsets in an antichain are comparable by inclusion.
Theorem 3.22.
Given a finite set $X$ and a symmetric submodular function $f:2^{X}\rightarrow\mathbb{N}$, if there exists a sequence chain $\{A_{0},A_{1},\dots,A_{m}\}$ of order $k+1$ on the connectivity system $(X,f)$, then no non-principal ultrafilter of order $k+1$ can exist on the same system.
Proof 3.23.
Since $F$ is a non-principal ultrafilter, it does not contain any singleton sets. Consider the following two cases:
  • Case 1: $A_{i}\in F$ for some $A_{i}$ in the sequence chain.
    By the ultrafilter condition (Q4), since $A_{0}\subseteq A_{1}\subseteq\dots\subseteq A_{m}$ and $A_{i}\in F$, all supersets of $A_{i}$ must also be in $F$ (including $A_{m}=X$). However, $X$ cannot be in $F$ because $f(X)\leq k$ would imply that the entire set $X$ is part of the filter, which contradicts the non-principal nature of $F$, as non-principal filters do not include all of $X$ unless $X$ is trivial.
  • Case 2: $A_{i}\notin F$ for all $A_{i}$ in the sequence chain.
    Since $A_{0}=\emptyset$ and $A_{m}=X$, by the ultrafilter condition (Q4), if $A_{i}\notin F$, then $X\setminus A_{i}\in F$. However, because the sequence chain is nested, if $A_{i}\notin F$, then $X\setminus A_{i}$ must be included in $F$. This contradicts the chain structure since, for some $j>i$, $A_{j}\subseteq X\setminus A_{i}$, and hence $A_{j}$ should be in $F$, contradicting the assumption that $A_{j}\notin F$.
Theorem 3.24.
Let $X$ be a finite set, and let $f:2^{X}\to\mathbb{N}$ be a symmetric submodular function. If there exists a sequence chain $\{A_{0},A_{1},\dots,A_{m}\}$ of order $k$ on the connectivity system $(X,f)$, then there exists a branch decomposition of the connectivity system $(X,f)$ with width at most $k$.
Proof 3.25.
Given the existence of a sequence chain $\{A_{0},A_{1},\dots,A_{m}\}$ of order $k$, we will construct a branch decomposition of $(X,f)$.
Start with a binary tree structure corresponding to the nested sequence chain $\{A_{0},A_{1},\dots,A_{m}\}$. For each subset $A_{i}$, assign a node $t_{i}$ in the tree such that $t_{0}$ corresponds to $A_{0}=\emptyset$, $t_{m}$ corresponds to $A_{m}=X$, and for all other $t_{i}$, $t_{i}$ corresponds to $A_{i}$.
Construct a tree $T$ such that each node $t_{i}$ is associated with a subset $A_{i}$ from the sequence chain. The leaves of $T$ are bijectively mapped to the elements of $X$. For every edge $e$ in $T$, define its width as the connectivity function $f$ evaluated on the set of vertices corresponding to the subtree induced by removing $e$ from $T$.
Since $f(A_{i})\leq k$ for each $i$, the width of each edge $e$ in the tree $T$ is at most $k$. Therefore, the branch-width of the constructed tree $(T,\mu)$ is at most $k$.
Theorem 3.26.
If there exists a sequence chain $\{A_{0},A_{1},\dots,A_{m}\}$ of order $k+1$ on the connectivity system $(X,f)$, then
  • No antichain of order $k+1$ can exist.
  • No non-principal ultrafilter of order $k+1$ can exist on the same system.
  • No maximal ideal of order $k+1$ can exist on the same system [1231].
  • No tangle of order $k+1$ can exist on the same system [529].
  • No (Maximal) loose tangle of order $k+1$ can exist on the same system [517].
  • There exists a branch decomposition of the connectivity system $(X,f)$ with width at most $k$.
3.5 Other chain concepts on connectivity system
3.5.1 Separation chain on connectivity system
In graph theory, the concept of a separation chain is also utilized, particularly in proving graph width parameters such as path-width and cut-width[471, 990, 771]. We define the separation chain and separation sequence chain within a connectivity system. These concepts are essentially equivalent to their non-separation counterparts but incorporate the perspective of separations.
Definition 3.27.
Let $X$ be a finite set, and let $f:2^{X}\rightarrow\mathbb{N}$ be a symmetric submodular function. A separation chain of order $k+1$ on a connectivity system $(X,f)$ is a sequence of separations $\{(A_{1},B_{1}),(A_{2},B_{2}),\dots,(A_{m},B_{m})\}$ where $A_{i}\subseteq X$ and $B_{i}=X\setminus A_{i}$ for each $i$, such that:
  1. 1.
    $A_{1}\subseteq A_{2}\subseteq\dots\subseteq A_{m}$ and $B_{1}\supseteq B_{2}\supseteq\dots\supseteq B_{m}$.
  2. 2.
    For each $i$ $(1\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$.
  3. 3.
    The width of the separation chain is defined as $\max_{1\leq i\leq m}|A_{i}\cap B_{i-1}|-1$.
Definition 3.28.
Let $X$ be a finite set, and let $f:2^{X}\rightarrow\mathbb{N}$ be a symmetric submodular function. A separation sequence chain of order $k+1$ on a connectivity system $(X,f)$ is a sequence of separations $\{(A_{0},B_{0}),(A_{1},B_{1}),\dots,(A_{m},B_{m})\}$ where $A_{i}\subseteq X$ and $B_{i}=X\setminus A_{i}$ for each $i$, such that:
  1. 1.
    $A_{0}=\emptyset$, $B_{0}=X$, $A_{m}=X$, and $B_{m}=\emptyset$.
  2. 2.
    $A_{0}\subseteq A_{1}\subseteq\dots\subseteq A_{m}$ and $B_{0}\supseteq B_{1}\supseteq\dots\supseteq B_{m}$.
  3. 3.
    For each $i$ $(0\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$.
  4. 4.
    The width of the separation sequence chain is defined as $\max_{1\leq i\leq m}|A_{i}\cap B_{i-1}|-1$.
3.5.2 Sperner system on connectivity system
Concepts closely related to chains and antichains include the Sperner system and trace. We aim to explore the concept of a Sperner system within a connectivity system. A Sperner system is a collection of subsets where no subset is entirely contained within another. It is significant in combinatorics and other mathematical fields for studying set systems and their properties, such as in Boolean lattices and set theory [886, 589, 978, 530, 431, 917]. In set theory, a trace refers to the intersection of a set system with a subset, capturing how the system "traces" the structure of that subset [1081, 1060, 1173, 978]. Although still in the conceptual stage, we outline the related definitions below.
Definition 3.29 (Sperner System on a Connectivity System).
Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function.
A Sperner system on $(X,f)$ is a collection $\mathcal{F}\subseteq 2^{X}$ such that there do not exist subsets $A,B\in\mathcal{F}$ with $A\subsetneq B$ and $f(A)\leq k$ and $f(B)\leq k$ for some $k$.
Definition 3.30 ($l$-Chain of Order $k+1$ on a Connectivity System).
An $l$-chain of order $k+1$ on a connectivity system $(X,f)$ is a subcollection $\{A_{1},A_{2},\dots,A_{l}\}\subseteq 2^{X}$ such that:
  1. 1.
    $A_{1}\subsetneq A_{2}\subsetneq\dots\subsetneq A_{l}$.
  2. 2.
    For each $i$ $(1\leq i\leq l)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$.
Definition 3.31 ($l$-Sperner System on a Connectivity System).
An $l$-Sperner system on a connectivity system $(X,f)$ is a collection $\mathcal{F}\subseteq 2^{X}$ such that no $l$-chain of order $k+1$ exists within $\mathcal{F}$.
We say that an $l$-Sperner system on $(X,f)$ is saturated if, for every set $S\in 2^{X}\setminus\mathcal{F}$, the collection $\mathcal{F}\cup\{S\}$ contains an $l$-chain of order $k+1$ on $(X,f)$.
Definition 3.32 (Trace (Restriction) on a Connectivity System of Order $k+1$).
[978] Let $(X,f)$ be a connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. For a set $F\subseteq 2^{X}$ and a subset $Y\subseteq X$, the trace (Restriction) of $F$ on $Y$ of order $k+1$ on a connectivity system is defined as:
$F|Y=\{F\cap Y:F\in F\text{ and }f(F\cap Y)\leq k\},$
where $k$ is a fixed integer. This trace represents the collection of subsets of $Y$ formed by intersecting $Y$ with each subset in $F$, ensuring that the submodular function evaluated on these intersections remains within the desired bound.  [^13] Although not directly related, similar operational concepts include Shattering and VC-dimension(cf.[1181, 434, 898]). Let $X$ be a finite set, and let $f:2^{X}\rightarrow\mathbb{N}$ be a symmetric submodular function. A subset $B\subseteq X$ is said to be shattered of order $k+1$ by $S$ if $\{A\cap B:f(A\cap B)\leq k,A\in S\}\in 2^{X}.$ The VC-dimension of order $k+1$ on the connectivity system $(X,f)$ is the cardinality of the largest subset $B\subseteq X$ to be shattered by $S$. A related concept known as fat-shattering is well-known in the literature [108]. 141414 Although the trace is a concept from topology, there are several related concepts in other fields. We will introduce a few definitions that extend the trace concept to a connectivity system. Please note that the following is at the conceptual stage. Let $X$ be a set, and let $f:2^{X}\rightarrow\mathbb{N}$ be a symmetric submodular function. The downward closure of $\mathcal{B}$ of order $k+1$ on a connectivity system (cf.[1151]) is defined as: :=$\mathcal{B}^{\downarrow}:=\{S\subseteq B\mid f(S)\leq k,B\in\mathcal{B}\}$ The kernel of a family $\mathcal{B}$ of order $k+1$ on a connectivity system is defined as the intersection of all sets in $\mathcal{B}$, subject to the condition that a function $f$ evaluated on this intersection is less than or equal to $k$ (cf.[364, 1156]). Formally: $\textit{kernel}\quad\mathcal{B}:=\bigcap_{B\in\mathcal{B}}B\quad\text{such that}\quad f\left(\bigcap_{B\in\mathcal{B}}B\right)\leq k.$ A family $\mathcal{B}$ is said to be countably deep of order $k+1$ on a connectivity system if, for every countable subset $\mathcal{C}\subseteq\mathcal{B}$, the kernel for $\mathcal{C}$ of order $k+1$ on a connectivity system belongs to $\mathcal{B}$ (cf.[365, 925]). Formally, we write: $\text{A family }\mathcal{B}\text{ is countably deep if }\forall\mathcal{C}\subseteq\mathcal{B},\,|\mathcal{C}|\leq\aleph_{0}\implies\textit{kernel}\quad\mathcal{C}\in\mathcal{B},$ where $\aleph_{0}$ denotes the cardinality of countable sets. A mesh of order $k+1$ on a connectivity system between two families of sets $\mathcal{B}$ and $\mathcal{C}$ occurs if every set in $\mathcal{B}$ intersects every set in $\mathcal{C}$ (cf.[362]). If $\mathcal{B}$ and $\mathcal{C}$ do not mesh (of order $k+1$ on a connectivity system), they are said to be dissociated (of order $k+1$ on a connectivity system). Formally, we write: #$\mathcal{B}\#\mathcal{C}\quad\text{if and only if}\quad B\cap C\neq\varnothing\quad\text{for all}\quad B\in\mathcal{B},C\in\mathcal{C},and\quad f(B\cap C)\leq k.$ The elementwise intersection (cf.[757, 1210, 204]) of $\mathcal{B}$ and $\mathcal{C}$ of order $k+1$ on a connectivity system is: $\mathcal{B}(\cap)\mathcal{C}=\{B\cap C\mid f(B\cap C)\leq k,B\in\mathcal{B}\text{ and }C\in\mathcal{C}\},$ The elementwise union (cf.[757, 1210, 204]) of $\mathcal{B}$ and $\mathcal{C}$ of order $k+1$ on a connectivity system is: $\mathcal{B}(\cup)\mathcal{C}=\{B\cup C\mid f(B\cup C)\leq k,B\in\mathcal{B}\text{ and }C\in\mathcal{C}\},$ The elementwise subtraction (cf.[757, 1210, 204]) of $\mathcal{B}$ and $\mathcal{C}$ of order $k+1$ on a connectivity system is: $\mathcal{B}(\setminus)\mathcal{C}=\{B\setminus C\mid f(B\setminus C)\leq k,B\in\mathcal{B}\text{ and }C\in\mathcal{C}\},$ where $\mathcal{B}\setminus\mathcal{C}$ denotes the usual subtraction of order $k+1$ on a connectivity system. Additionally, related concepts include a $T_{1}$ space (accessible space) [78], a Hausdorff space[444, 995, 1112], a $T_{0}$ space[736], a Compact space[1014, 862], and an $R_{0}$ space (symmetric space) [78].
Definition 3.33 (Strong Trace (Strong Restriction) on a Connectivity System of Order $k+1$).
[978] A set $F\subseteq 2^{X}$ strongly traces a subset $Y\subseteq X$ of order $k+1$ if there exists a set $B\subseteq Y$ such that for any subset $Z\subseteq Y$, we have $B\cup Z\in F$ and $f(B\cup Z)\leq k$. The set $B$ is called the support of $Y$ by $F$, and the set of all such supports is denoted by $S_{F}(Y)$.
A concept related to the Sperner system is the clutter[701]. A clutter is a family of subsets (or hyperedges) in a hypergraph where no subset is contained within another, and it is frequently used in combinatorial optimization.
3.5.3 Chain-Decomposition on connectivity system
In the study of chains, concepts such as chain decomposition[1150, 708], antichain decomposition[902, 403], and symmetric chains[73, 1206, 1178] have been extensively researched. These concepts are commonly utilized in partially ordered sets theory.
By adapting and defining these ideas within the context of connectivity systems, we aim to explore whether any interesting properties emerge. A chain decomposition breaks down a collection of subsets into disjoint chains. A symmetric chain extends the chain concept to include subsets of every size within a specified range, and a symmetric chain decomposition divides a collection into disjoint symmetric chains. An antichain decomposition partitions a collection into disjoint antichains. While still in the conceptual stage, the definitions are provided below.
Definition 3.34 (Disjoint Chains on a Connectivity System).
Two chains $\{A_{1},A_{2},\dots,A_{m}\}$ and $\{B_{1},B_{2},\dots,B_{n}\}$ on a connectivity system $(X,f)$ are said to be disjoint if no subset in one chain is a subset of any subset in the other chain. Formally, for all $i$ and $j$, $A_{i}\cap B_{j}=\emptyset$ or $A_{i}\not\subseteq B_{j}$ and $B_{j}\not\subseteq A_{i}$.
Definition 3.35 (Chain Decomposition on a Connectivity System).
A chain decomposition on a connectivity system $(X,f)$ is a partition of a collection of subsets $F\subseteq 2^{X}$ into disjoint chains of order $k+1$, where each chain satisfies the conditions specified in the chain definition.
Definition 3.36 (Symmetric Chain on a Connectivity System).
A symmetric chain on a connectivity system $(X,f)$ is a chain $\{A_{1},A_{2},\dots,A_{m}\}$ such that:
  1. 1.
    The chain contains subsets of every cardinality $i$ for $i\in\{k,k+1,\dots,|X|-k\}$, where $20\leq k\leq|X|/$ 2.
  2. 2.
    For each $i$ in the chain, the symmetric submodular function $f(A_{i})$ satisfies $f(A_{i})\leq k$.
Definition 3.37 (Symmetric Chain Decomposition on a Connectivity System).
A symmetric chain decomposition of a collection $F\subseteq 2^{X}$ on a connectivity system $(X,f)$ is a partition of $F$ into disjoint symmetric chains, where each symmetric chain satisfies the conditions specified in the symmetric chain definition.
Question 3.38.
How does the size of a symmetric chain relate to the order on a Connectivity System?
Definition 3.39 (Disjoint Antichains on a Connectivity System).
Two antichains $\{A_{1},A_{2},\dots,A_{m}\}$ and $\{B_{1},B_{2},\dots,B_{n}\}$ on a connectivity system $(X,f)$ are said to be disjoint if no subset in one antichain is a subset of any subset in the other antichain. Formally, for all $i$ and $j$, $A_{i}\cap B_{j}=\emptyset$ or $A_{i}\not\subseteq B_{j}$ and $B_{j}\not\subseteq A_{i}$.
Definition 3.40 (Antichain Decomposition on a Connectivity System).
An antichain decomposition on a connectivity system $(X,f)$ is a partition of a collection of subsets $F\subseteq 2^{X}$ into disjoint antichains of order $k+1$, where each antichain satisfies the conditions specified in the antichain definition.
3.5.4 Operation and Single-element-chain on a Connectivity System
In the future, we will delve into the concepts of operations and single-element-chains. We anticipate that the relationships observed between general chains, sequences, and ultrafilters will similarly apply to single-element structures, such as single-element-chains, single-element-sequences, single ultrafilters, and linear-width. The single-element extension (cf.[304, 300, 1097, 1121, 927]) on a connectivity system is also referred to as a one-point extension (cf.[793, 794]) or one-element lifting(cf.[538, 840, 539]) on the connectivity system. We also consider about single-element-coextension(cf.[920, 762, 1110]) on a connectivity system.
Definition 3.41.
Given a chain $\{A_{1},A_{2},\dots,A_{m}\}$ of order $k+1$ on a connectivity system $(X,f)$, where $A_{1}\subseteq A_{2}\subseteq\dots\subseteq A_{m}$ and for each $i$, $f(A_{i})\leq k$:
  1. 1.
    Extension of the Chain: The extension of the chain by adding a new subset $A_{m+1}$ is defined as follows: Let $A_{m+1}\subseteq X$ be a subset such that $A_{m}\subseteq A_{m+1}$ and $f(A_{m+1})\leq k$. The resulting extended chain is $\{A_{1},A_{2},\dots,A_{m},A_{m+1}\}$.
  2. 2.
    Deletion of a Subset: The deletion of a subset $A_{i}$ from the chain is defined as follows: Remove $A_{i}$ from the chain, resulting in the chain $\{A_{1},\dots,A_{i-1},A_{i+1},\dots,A_{m}\}$. The condition that $f(A_{j})\leq k$ for all $j$ in the modified chain must still hold.
Definition 3.42.
A single-element-chain is a special case of a chain where each subset $A_{i}$ in the chain differs from the previous subset $A_{i-1}$ by exactly one element. Formally, a single-element-chain of order $k+1$ is defined as a sequence of subsets $\{A_{1},A_{2},\dots,A_{m}\}$ such that:
  1. 1.
    $A_{1}\subseteq A_{2}\subseteq\dots\subseteq A_{m}$.
  2. 2.
    For each $i$ $(1\leq i\leq m)$, $f(A_{i})\leq k$.
  3. 3.
    $|A_{i}\setminus A_{i-1}|=1$ for each $i$ $(2\leq i\leq m)$.
Definition 3.43.
A single-element-sequence chain is a specific type of sequence chain on a connectivity system $(X,f)$, where $f$ is a symmetric submodular function, and $X$ is a finite set. A single-element-sequence chain of order $k+1$ is defined as a sequence chain of subsets $\{A_{0},A_{1},\dots,A_{m}\}$ of $X$ such that:
  1. 1.
    $A_{0}=\emptyset$ and $A_{m}=X$.
  2. 2.
    $A_{0}\subseteq A_{1}\subseteq\dots\subseteq A_{m}$.
  3. 3.
    For each $i$ $(0\leq i\leq m)$, the symmetric submodular function $f$ evaluated at $A_{i}$ satisfies $f(A_{i})\leq k$.
  4. 4.
    For each $i$ $(1\leq i\leq m)$, $|A_{i}\setminus A_{i-1}|=1$, meaning that each subset $A_{i}$ in the sequence chain differs from the previous subset $A_{i-1}$ by exactly one element.
Definition 3.44.
The single-element-extension of a chain $\{A_{1},A_{2},\dots,A_{m}\}$ involves adding a new subset $A_{m+1}$ such that $A_{m+1}=A_{m}\cup\{e\}$, where $e\in X\setminus A_{m}$ and $f(A_{m+1})\leq k$. The new chain is $\{A_{1},A_{2},\dots,A_{m},A_{m+1}\}$.
Definition 3.45.
The single-element-deletion of a subset $A_{i}$ from the chain involves removing one element $e$ from $A_{i}$, resulting in $A_{i}^{\prime}=A_{i}\setminus\{e\}$. The new chain, after performing the deletion, must still satisfy the condition $f(A_{j})\leq k$ for all $j$. The resulting chain is $\{A_{1},\dots,A_{i-1},A_{i}^{\prime},A_{i+1},\dots,A_{m}\}$.
Definition 3.46.
Let $(X,f)$ be a connectivity system, and let ${A_{1},A_{2},\dots,A_{m}}$ be a chain of order $k+1$ within this system. Consider a set $X\cup{e}$, where $e\notin X$. The connectivity system $(X\cup{e},f^{\prime})$ is a single-element coextension of $(X,f)$ if for every subset $A\subseteq X$, the function $f^{\prime}$ satisfies $f^{\prime}(A)=f(A\cup{e})$. In this way, the function $f^{\prime}$ on the coextended system $(X\cup{e},f^{\prime})$ reduces to $f$ on the original system $(X,f)$ when the element $e$ is contracted.
4 Prefilter and Filter Subbase on a Connectivity System
We explain Prefilter and Filter Subbase on a connectivity system $(X,f)$. Prefilters and filter subbases, like ultrafilters, are crucial in set theory and topology for studying limits, convergence, and compactness. Prefilters are non-empty collections of sets closed under finite intersections, forming the basis for constructing filters. Filter subbases are non-empty collections of sets that generate filters by taking finite intersections, aiding in the study of convergence, limits, and compactness [1162, 921, 459]. In this section, we add the condition of symmetric submodularity to prefilters and filter subbases and then perform verification.
4.1 Prefilter on a Connectivity System
First, we explain the definition of a prefilter and ultra-prefilter on a connectivity system $(X,f)$.
Definition 4.1.
Let $X$ be a finite set and $f$ be a symmetric submodular function. A prefilter of order $k+1$ on a connectivity system $(X,f)$ is a non-empty proper set family $P\subseteq 2^{X}$ that satisfies:
  1. (P1)
    $\emptyset\notin P$ (Proper).
  2. (P2)
    For any $A\in P$, $f(A)\leq k$ (k-efficiency).
  3. (P3)
    For any $B,C\in P$, there exists some $A\in P$ such that $A\subseteq B\cap C$ and $f(A)\leq k$ (Downward Directed under k-efficiency).
Definition 4.2.
Let $X$ be a finite set and $f$ be a symmetric submodular function. An ultra-prefilter $P\subseteq 2^{X}$ on a connectivity system $(X,f)$ is a prefilter of order $k+1$ that satisfies the following additional property:
  1. (P4)
    $\forall A\subseteq X,f(A)\leq k\Rightarrow\exists B\in P$ such that $B\subseteq A$ or $B\subseteq X\setminus A$ and $f(B)\leq k$.
Next, we demonstrate that a filter on a connectivity system can be considered a prefilter on that system. This concept extends the notions of prefilters and ultra-prefilters from set theory by incorporating the condition of symmetric submodularity.
Theorem 4.3.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Any filter of order $k+1$ on a connectivity system $(X,f)$ is also a prefilter of order $k+1$ on the same connectivity system $(X,f)$.
Proof 4.4.
To prove that a filter $Q$ on a connectivity system is also a prefilter, we need to show that $Q$ satisfies all the conditions of a prefilter.
  1. 1.
    Axiom (P1): By definition, a filter $Q$ is non-empty because it contains subsets of $X$ that satisfy the given conditions. And by Axiom (Q3), $\emptyset\notin Q$. Therefore, $Q$ is a proper set family.
  2. 2.
    Axiom (P2): By Axiom (Q0), $\forall A\in Q,f(A)\leq k$. Therefore, every element in $Q$ satisfies the k-efficiency condition.
  3. 3.
    Axiom (P3): For any $B,C\in Q$ such that $f(B)\leq k$ and $f(C)\leq k$, we need to show there exists some $A\in Q$ such that $A\subseteq B\cap C$ and $f(A)\leq k$.
    • Given $B,C\in Q$, by Axiom (Q1), $f(B\cap C)\leq k\Rightarrow B\cap C\in Q$.
    • Since $B\cap C\in Q$ and by the definition of $Q$, $f(B\cap C)\leq k$, we can choose $A=B\cap C$.
    • Therefore, $A\in Q$, $A\subseteq B\cap C$, and $f(A)\leq k$.
Hence, $Q$ satisfies all the conditions of a prefilter. This proof is completed.
Next, we show that an ultrafilter on a connectivity system is also an ultra-prefilter on a connectivity system.
Theorem 4.5.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Any ultrafilter of order $k+1$ on a connectivity system $(X,f)$ is also an ultra-prefilter of order $k+1$ on the same connectivity system $(X,f)$.
Proof 4.6.
Let $X$ be a finite set and $f$ be a symmetric submodular function. To prove that an ultrafilter $Q$ on a connectivity system is also an ultra-prefilter of order $k+1$ on a connectivity system, we need to show that $Q$ satisfies all the conditions of an ultra-prefilter of order $k+1$. Axioms (P1)-(P3) hold obviously. So we show that axiom (P4) holds.
For any $A\subseteq X$ such that $f(A)\leq k$, we need to show there exists some $B\in Q$ such that $B\subseteq A$ or $B\subseteq X\setminus A$ and $f(B)\leq k$. Given $A\subseteq X$ such that $f(A)\leq k$, by Axiom (Q4), either $A\in Q$ or $X\setminus A\in Q$. If $A\in Q$, then we can choose $B=A$, and $B\subseteq A$ with $f(B)\leq k$. If $X\setminus A\in Q$, then we can choose $B=X\setminus A$, and $B\subseteq X\setminus A$ with $f(B)\leq k$.
Hence, $Q$ satisfies all the conditions of an ultra-prefilter. This proof is completed.
4.2 Filter Subbase on a Connectivity System
Next, we introduce the definitions of filter subbase and ultrafilter subbase. These concepts extend the traditional notions of filter subbases and ultrafilter subbases from set theory by incorporating the condition of symmetric submodularity.
Definition 4.7.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a connectivity system $(X,f)$, a set family $S\subseteq 2^{X}$ is called a filter subbase of order $k+1$ if it satisfies the following conditions:
  1. (SB1)
    $S\neq\emptyset$.
  2. (SB2)
    $\emptyset\notin S$.
  3. (SB3)
    $\forall A\in S$, $f(A)\leq k$.
Definition 4.8.
Let $X$ be a finite set and $f$ be a symmetric submodular function. In a connectivity system $(X,f)$, a set family $S\subseteq 2^{X}$ is called an ultrafilter subbase of order $k+1$ if it satisfies the following conditions:
  1. (SB1)
    $S\neq\emptyset$.
  2. (SB2)
    $\emptyset\notin S$.
  3. (SB3)
    $\forall A\in S$, $f(A)\leq k$.
  4. (SB4)
    $\forall A\subseteq X$ such that $f(A)\leq k$, there exists $B\in S$ such that either $B\subseteq A$ or $B\subseteq X\setminus A$ and $f(B)\leq k$.
We now explore the relationship between a subbase on a connectivity system and a filter on a connectivity system. The following theorem establishes this relationship.
Theorem 4.9.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Given a filter subbase $S$ of order $k+1$ on a connectivity system $(X,f)$, we can generate a filter $Q$ of order $k+1$ as follows: The filter $Q$ of order $k+1$ on a connectivity system $(X,f)$ generated by $S$ is the set of all subsets of $X$ that can be formed by finite intersections of elements of $S$. Formally,
$Q=\{A\subseteq X\mid\exists A_{1},A_{2},\dots,A_{n}\in S\text{ such that }A=A_{1}\cap A_{2}\cap\dots\cap A_{n}\text{ and }f(A)\leq k\}.$
Proof 4.10.
To prove that $Q$ is a filter of order $k+1$, we need to show that $Q$ satisfies the conditions of a filter of order $k+1$.
Axiom (Q0): This axiom obviously holds. By construction, every element in $Q$ is formed by finite intersections of elements in $S$. Since $S$ is a filter subbase of order $k+1$, all elements $A_{i}$ in $S$ satisfy $f(A_{i})\leq k$. By the submodularity of $f$, for any intersection $A=A_{1}\cap A_{2}\cap\dots\cap A_{n}$, we have $f(A)=f(A_{1}\cap A_{2}\cap\dots\cap A_{n})\leq\min(f(A_{1}),f(A_{2}),\dots,f(A_{n}))\leq k$.
Axiom (Q3): Since $S$ is a filter subbase of order $k+1$ on a connectivity system $(X,f)$, it is non-empty. Suppose $S=\{A_{1},A_{2},\dots,A_{m}\}$. Consider any single element $A_{i}\in S$. Since $S$ is non-empty, $A_{i}$ exists, and $f(A_{i})\leq k$. Therefore, $A_{i}\in Q$, implying $Q$ is non-empty. And by the definition of a filter subbase, $\emptyset\notin S$. Since $Q$ is generated by finite intersections of elements of $S$, and $\emptyset$ cannot be formed by any finite intersection of non-empty sets, $\emptyset\notin Q$. Therefore, $Q$ is proper.
Axiom (Q1): By construction, $Q$ consists of all subsets of $X$ that can be formed by finite intersections of elements of $S$ and satisfy $f(A)\leq k$. Thus, by definition, $Q$ satisfies the k-efficiency condition. Let $A,B\in Q$. Then there exist $A_{1},A_{2},\dots,A_{n}\in S$ and $B_{1},B_{2},\dots,B_{m}\in S$ such that $A=A_{1}\cap A_{2}\cap\dots\cap A_{n}$ and $f(A)\leq k$, $B=B_{1}\cap B_{2}\cap\dots\cap B_{m}$ and $f(B)\leq k$. Consider the intersection $A\cap B$:
$A\cap B=(A_{1}\cap A_{2}\cap\dots\cap A_{n})\cap(B_{1}\cap B_{2}\cap\dots\cap B_{m}).$
Since $S$ is a filter subbase and contains $A_{i}$ and $B_{j}$ for all $i,j$, $A\cap B$ is formed by the finite intersection of elements of $S$. Moreover, because $f$ is submodular and symmetric, $f(A\cap B)\leq\min(f(A),f(B))\leq k$. Thus, $A\cap B\in Q$.
Axiom (Q2): Let $A\in Q$ and $A\subseteq B\subseteq X$. By definition, there exist $A_{1},A_{2},\dots,A_{n}\in S$ such that $A=A_{1}\cap A_{2}\cap\dots\cap A_{n}$ and $f(A)\leq k$. If $f(B)\leq k$, we need to show $B\in Q$. Since $A\subseteq B$ and $f(B)\leq k$, $B$ can be considered as a superset satisfying the condition for being in $Q$.
Therefore, $Q$ is a filter of order $k+1$ generated by the filter subbase $S$. This proof is completed.
Theorem 4.11.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Given a filter subbase $S$ of order $k+1$ on a connectivity system $(X,f)$, we can generate a prefilter $Q$ of order $k+1$ as follows:
$Q=\{A\subseteq X\mid\exists A_{1},A_{2},\dots,A_{n}\in S\text{ such that }A=A_{1}\cap A_{2}\cap\dots\cap A_{n}\text{ and }f(A)\leq k\}.$
Proof 4.12.
To demonstrate that $Q$ is a prefilter of order $k+1$, we need to verify that $Q$ satisfies the conditions of a prefilter as defined by axioms (P1) to (P3).
Axiom (P1): Since $S$ is a non-empty filter subbase, by definition, $S$ does not contain $\emptyset$. $Q$ is formed by finite intersections of elements in $S$, and because $\emptyset$ cannot be formed by such intersections, $\emptyset\notin Q$. Therefore, $Q$ is proper and non-empty, satisfying axiom (P1).
Axiom (P2): By construction, each element $A\in Q$ is the intersection of a finite number of sets $A_{1},A_{2},\dots,A_{n}\in S$ such that $f(A)\leq k$. Since $S$ is a filter subbase of order $k+1$, all elements $A_{i}\in S$ satisfy $f(A_{i})\leq k$. The submodularity of function $f$ ensures that $f(A_{1}\cap A_{2}\cap\dots\cap A_{n})\leq\min(f(A_{1}),f(A_{2}),\dots,f(A_{n}))\leq k$. Therefore, every element $A\in Q$ satisfies $f(A)\leq k$, ensuring k-efficiency as required by axiom (P2).
Axiom (P3): For any $B,C\in Q$, there exist sets $B_{1},B_{2},\dots,B_{m}\in S$ and $C_{1},C_{2},\dots,C_{n}\in S$ such that $B=B_{1}\cap B_{2}\cap\dots\cap B_{m}$ and $C=C_{1}\cap C_{2}\cap\dots\cap C_{n}$, with $f(B)\leq k$ and $f(C)\leq k$. Consider the intersection $A=B\cap C$. Since $A=(B_{1}\cap B_{2}\cap\dots\cap B_{m})\cap(C_{1}\cap C_{2}\cap\dots\cap C_{n})$, $A$ is formed by a finite intersection of elements in $S$. The submodularity and symmetry of function $f$ imply that $f(A)=f(B\cap C)\leq\min(f(B),f(C))\leq k$. Therefore, $A\in Q$, satisfying axiom (P3).
Given that $Q$ satisfies all the conditions (P1)-(P3) of a prefilter, we conclude that $Q$ is indeed a prefilter of order $k+1$ generated by the filter subbase $S$. This completes the proof.
Theorem 4.13.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Given an ultrafilter subbase $S$ of order $k+1$ on a connectivity system $(X,f)$, we can generate an ultrafilter $Q$ of order $k+1$ as follows: The filter $Q$ of order $k+1$ on a connectivity system $(X,f)$ generated by $S$ is the set of all subsets of $X$ that can be formed by finite intersections of elements of $S$. Formally,
$Q=\{A\subseteq X\mid\exists A_{1},A_{2},\dots,A_{n}\in S\text{ such that }A=A_{1}\cap A_{2}\cap\dots\cap A_{n}\text{ and }f(A)\leq k\}.$
Proof 4.14.
The proof can be established using a method nearly identical to that used in Theorem 4.9. This completes the proof.
Theorem 4.15.
Let $X$ be a finite set and $f$ be a symmetric submodular function. Given an ultrafilter subbase $S$ of order $k+1$ on a connectivity system $(X,f)$, we can generate an ultra-prefilter $Q$ of order $k+1$ as follows:
$Q=\{A\subseteq X\mid\exists A_{1},A_{2},\dots,A_{n}\in S\text{ such that }A=A_{1}\cap A_{2}\cap\dots\cap A_{n}\text{ and }f(A)\leq k\}.$
Proof 4.16.
The proof can be established using a method nearly identical to that used in Theorem 4.11. This completes the proof.
4.3 Ultrafilter Number
In set theory and various mathematical fields, the concept of the Ultrafilter Number (Cardinality) is frequently discussed (ex.[564, 1011, 239]). We plan to explore whether this can be extended to Connectivity Systems and, if so, what characteristics can be derived from such an extension.
Definition 4.17 (Ultrafilter Number on a Connectivity System).
Let $(X,f)$ be a finite connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. The Ultrafilter Number $u(X,f)$ on this connectivity system is defined as the minimum size of a set $B\subseteq X$ such that $B$ forms a ultra prefilter for a non-principal ultrafilter on $(X,f)$. Formally,
$u(X,f)=\min\{|B|:B\text{ is a ultra prefilter for a non-principal ultrafilter on }(X,f)\}.$
Question 4.18.
How does the Ultrafilter Number relate to the order on a Connectivity System?
5 Ultrafilter on Infinite Connectivity System and Countable Connectivity System
5.1 Ultrafilter on Infinite Connectivity System
Until now, we have considered ultrafilters on finite sets, but now we turn our attention to ultrafilters on infinite sets. First, let’s define what an infinite connectivity system is.
Definition 5.1 ([227, 418, 293]).
Consider an infinite set $X$ equipped with a symmetric submodular function $f$ mapping from the set of subsets of $X$ to ${}\mathbb{N}\cup\{\infty\}$. This function satisfies the following conditions for all subsets $A$ and $B$ of $X$:
  1. 1.
    Symmetry: $f(A)=f(X\setminus A)$.
  2. 2.
    Submodularity: $f(A)+f(B)\geq f(A\cup B)+f(A\cap B)$.
In the context of infinite sets, we also require $f$ to be $k$-limit-closed. Specifically, if $k\in\mathbb{N}$ and $(A_{i})_{i\in I}$ is a chain of subsets of $X$ each with connectivity at most $k$, then the union $\bigcup_{i\in I}A_{i}$ also has connectivity at most $k$. It is important to note that the ground set of a (possibly infinite) matroid, along with its connectivity function, constitutes a connectivity system, and the connectivity function is defined to be $k$-limit-closed. A pair $(X,f)$ is called an infinite connectivity system.
Given the known lemma, it is reasonable to consider a deep relationship between finite sets and infinite sets.
Lemma 5.2 ([227]).
Let $X$ be an infinite set and $f$ be a symmetric submodular function. Let $k\in\mathbb{N}$, and let $A\subseteq X$ be a set such that all finite subsets have connectivity at most $k$. Then also $f(A)\leq k$.
Following the model of ultrafilters on finite sets, we define ultrafilters on infinite sets as follows [529].
Definition 5.3 (cf. [529]).
Let $X$ be an infinite set and $f$ be a symmetric submodular function. In an infinite connectivity system $(X,f)$, the set family $F\subseteq 2^{X}$ is called a filter of order $k+1$ if the following axioms hold true:
  1. (Q0)
    $\forall A\in F,f(A)\leq k$.
  2. (Q1)
    $A\in F,B\in F,f(A\cap B)\leq k\Rightarrow A\cap B\in F$.
  3. (Q2)
    $A\in F,A\subseteq B\subseteq X,f(B)\leq k\Rightarrow B\in F$.
  4. (Q3)
    $\emptyset\notin F$.
An ultrafilter of order $k+1$ on an infinite connectivity system $(X,f)$ is a filter of order $k+1$ on an infinite connectivity system $(X,f)$ which satisfies the following additional axiom:
  1. (Q4)
    $\forall A\subseteq X,f(A)\leq k\Rightarrow\text{either }A\in F\text{ or }X\setminus A\in F$.
A filter of order $k+1$ on an infinite connectivity system $(X,f)$ is principal if the filter satisfies the following axiom [529]:
  1. (QP5)
    $A\in F$ for all $A\subseteq X$ with $|A|=1$ and $f(A)\leq k$.
A filter of order $k+1$ on an infinite connectivity system $(X,f)$ is non-principal if the filter satisfies the following axiom [529]:
  1. (Q5)
    $A\notin F$ for all $A\subseteq X$ with $|A|=1$.
Non-principal refers to a filter or ideal that does not contain any singletons (i.e., sets with exactly one element).
If a filter on an infinite connectivity system is weak [497, 795, 87], the following axiom holds instead of axiom (Q1):
  1. (QW1’)
    $A\in F,B\in F,f(A\cap B)\leq k\Rightarrow A\cap B\neq\emptyset$.
Note that weak filters aim at interpreting defaults via a generalized ‘most’ quantifier in first-order logic [795, 87, 1065, 482]. A weak filter on an infinite connectivity system is a co-Weak Ideal on an infinite connectivity system.
If a filter on an infinite connectivity system is quasi [512, 254], the following axiom holds instead of axiom (Q1):
  1. (QQ1’)
    $A\subseteq X,B\subseteq X,A\notin F,B\notin F\Rightarrow A\cup B\notin F$.
Note that a quasi-Ultrafilter is used to perform an axiomatic analysis of incomplete social judgments [254].
The following also holds for infinite connectivity systems, just as it does for finite connectivity systems.
Theorem 5.4.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Note that Limit-Closed means that for any $k\in\mathbb{N}$ and any chain $(A_{i})_{i\in I}$ of subsets of $X$, each with connectivity at most $k$, the union $\bigcup_{i\in I}A_{i}$ also has connectivity at most $k$. An ultrafilter of order $k+1$ on an infinite connectivity system $(X,f)$ is also an ultrafilter of order $k$ on an infinite connectivity system $(X,f)$.
Proof 5.5.
By using the symmetric submodular function, an ultrafilter of order $k+1$ on an infinite connectivity system $(X,f)$ obviously holds the axioms of an ultrafilter of order $k$ on an infinite connectivity system $(X,f)$. This proof is completed.
Here are some properties that hold for filters and ultrafilters of order $k+1$ on an infinite connectivity system.
Lemma 5.6.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. If a filter of order $k+1$ on an infinite connectivity system $(X,f)$ is maximal, it satisfies the following additional axiom:
  • (Q4)
    $\forall A\subseteq X$, if $f(A)\leq k$, then either $A\in F$ or $X\setminus A\in F$.
Proof 5.7.
Assume that there exists a maximal filter $F$ of order $k+1$ on $(X,f)$ that does not satisfy axiom (Q4). That is, there exists a subset $A\subseteq X$ such that $f(A)\leq k$, but neither $A\in F$ nor $X\setminus A\in F$. Since $F$ is maximal, the addition of $A$ or $X\setminus A$ to $F$ would result in a filter that still respects the conditions of $f(A)\leq k$. Therefore, $F\cup\{A\}$ or $F\cup\{X\setminus A\}$ would also be a valid filter, contradicting the maximality of $F$. Thus, the assumption that $F$ does not satisfy (Q4) leads to a contradiction. Therefore, any maximal filter $F$ of order $k+1$ must satisfy axiom (Q4). This proof is completed.
Definition 5.8 (cf. [451]).
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. We will say that a family $F$ of order $k+1$ on an infinite connectivity system is nice if it is nonempty, closed under intersections (i.e., if $A,B\in F$ and $f(A\cap B)\leq k$, then $A\cap B\in F$), and $\emptyset\notin F$.
Lemma 5.9.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. A family of subsets of $X$ is maximal nice of order $k+1$ on an infinite connectivity system if and only if it is an ultrafilter of order $k+1$ on an infinite connectivity system.
Proof 5.10.
Let’s first recall that a family $F$ of order $k+1$ on an infinite connectivity system is called nice if it is nonempty, closed under intersections (i.e., if $A,B\in F$ and $f(A\cap B)\leq k$, then $A\cap B\in F$), and $\emptyset\notin F$.
Assume that $F$ is a maximal nice family of order $k+1$. We want to show that $F$ is an ultrafilter of order $k+1$, i.e., it satisfies axiom (Q4): $\forall A\subseteq X$, if $f(A)\leq k$, then either $A\in F$ or $X\setminus A\in F$. Suppose $F$ is maximal nice, but there exists a subset $A\subseteq X$ such that $f(A)\leq k$, and neither $A\in F$ nor $X\setminus A\in F$. Since $F$ is maximal, we can extend $F$ by adding either $A$ or $X\setminus A$ to form a new family $F^{\prime}$. But this contradicts the maximality of $F$, as it can still be extended. Therefore, $F$ must satisfy (Q4), and hence it is an ultrafilter of order $k+1$.
Now assume $F$ is an ultrafilter of order $k+1$. We need to show that $F$ is maximal nice. Suppose $F$ is not maximal nice. Then there exists a family $F^{\prime}$ such that $F\subsetneq F^{\prime}$ and $F^{\prime}$ is nice. But since $F$ is an ultrafilter, it cannot be extended by adding more sets without violating the conditions for an ultrafilter. This contradicts the assumption that $F$ is not maximal nice. Hence, $F$ must be maximal nice. This proof is completed.
In this paper, let $\text{Fr}=\{A\subseteq X\mid f(A)\leq k,X\setminus A\text{ is finite}\}$ be the Frechet filter of order $k+1$ on an infinite connectivity system $(X,f)$. We consider the following lemma.
Lemma 5.11.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. If there exists a Frechet filter of order $k+1$ on an infinite connectivity system $(X,f)$, then there exists a nonprincipal ultrafilter of order $k+1$ on an infinite connectivity system $(X,f)$.
Proof 5.12.
Assume that there does not exist a nonprincipal ultrafilter of order $k+1$ on $(X,f)$. This implies that every ultrafilter on $(X,f)$ must be principal. If Fr can be extended to an ultrafilter, this ultrafilter would be principal. A principal ultrafilter contains a set $\{x\}$ for some $x\in X$ with $f(\{x\})\leq k$. However, by definition, in a Frechet filter, $X\setminus\{x\}$ is finite, and thus $X$ cannot have only principal ultrafilters if it contains elements for which $f(\{x\})\leq k$. This contradicts the assumption that every ultrafilter is principal. Thus, there must exist a nonprincipal ultrafilter of order $k+1$ on $(X,f)$. This proof is completed.
We consider the following extended Ramsey’s Theorem on an infinite connectivity system. Note that Ramsey’s Theorem states that for any given positive integers $k$ and $r$, there is a minimum number $R(k,r)$ such that any graph with $R(k,r)$ vertices, colored with $r$ colors, contains a monochromatic clique of size $k$. Note that a monochromatic clique is a subset of vertices where all edges connecting them are the same color. This theorem can also be considered in terms of partitions, ensuring any partition of edges or vertices leads to a monochromatic subset.
Theorem 5.13.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. Then for any infinite partition $P=\{A_{1},A_{2},\dots\}$ of $X$, there exists an infinite subset $Y\subseteq X$ such that all subsets of $Y$ are contained within a single block of the partition and have connectivity at most $k$.
Proof 5.14.
Let $X$ be an infinite set with the symmetric submodular function $f$ as defined. Let $P=\{A_{1},A_{2},\dots\}$ be a partition of $X$. We want to find an infinite subset $Y\subseteq X$ such that all subsets of $Y$ are contained within a single block of $P$ and have connectivity at most $k$.
Consider an increasing chain of subsets $(A_{i})_{i\in I}$ of $X$, where each $A_{i}$ has connectivity at most $k$. By the limit-closed property of function $f$, the union $\bigcup_{i\in I}A_{i}$ also has connectivity at most $k$.
Using the classical form of Ramsey’s Theorem, we know that for any infinite set $X$ and any partition of $X$ into finitely many pieces, there exists an infinite subset $Y\subseteq X$ such that all pairs in $Y$ lie within the same piece of the partition. To apply Ramsey’s Theorem in the context of the connectivity system $(X,f)$, we need to ensure that the infinite subset $Y$ not only lies within a single block of the partition but also maintains the connectivity condition $f(Y)\leq k$.
Start with an infinite subset $Y_{0}\subseteq X$. Partition $Y_{0}$ into blocks $A_{i}$ according to the given partition $P$. Select a block $A_{j}$ that contains an infinite subset $Y_{1}\subseteq A_{j}$. By the symmetry and submodularity of function $f$, ensure $f(Y_{1})\leq k$.
By iterating the process, we can construct a nested sequence of infinite subsets $Y_{0}\supseteq Y_{1}\supseteq Y_{2}\supseteq\dots$, each lying within a single block of the partition and maintaining the connectivity condition $f(Y_{i})\leq k$.
Let $Y=\bigcap_{i=0}^{\infty}Y_{i}$. Since each $Y_{i}$ is infinite and lies within a single block of the partition, $Y$ is also infinite and lies within a single block of the partition. Moreover, by the limit-closed property of the infinite connectivity system, we obtain $f(Y)\leq k$.
Thus, we have shown that for any infinite partition $P$ of $X$, there exists an infinite subset $Y\subseteq X$ such that all subsets of $Y$ lie within a single block of the partition and have connectivity at most $k$. This completes the proof of the extended Ramsey’s Theorem on an infinite connectivity system.
Theorem 5.15.
Let $(X,f)$ be an infinite connectivity system with $f$ a symmetric submodular function satisfying $k$-Limit-Closed. A set family $U\subseteq 2^{X}$ is a weak ultrafilter of order $k+1$ if and only if it is a maximal weak filter of order $k+1$ on $(X,f)$.
Proof 5.16.
Assume that $U$ satisfies the definition of a weak ultrafilter but is not maximal. This means there exists a weak filter $G$ such that $U\subset G$ and $G$ contains an additional set $A\subset X$ where $A\in G$ but $A\notin U$. Since $U$ is a weak ultrafilter, by its definition, for any subset $A\subseteq X$ with $f(A)\leq k$, it must hold that either $A\in U$ or $X\setminus A\in U$.
Now, because $U\subset G$, every element of $U$ is also an element of $G$. Specifically, since $X\setminus A\in U$ and $U\subset G$, it follows that $X\setminus A$ must also be in $G$. However, $G$ already contains $A$ by assumption, and now it also contains $X\setminus A$. For $G$ to be a consistent weak filter, it cannot contain both $A$ and $X\setminus A$ because that would violate the condition of non-empty intersection. This leads to a contradiction because if $G$ contains both $A$ and $X\setminus A$, it would mean that $G$ is no longer a valid weak filter due to inconsistency, which contradicts the assumption that $G$ is a weak filter.
Hence, $U$ must be maximal because extending $U$ to any larger set $G$ results in an inconsistency, proving that $U$ cannot be properly contained in any larger weak filter.
We need to prove that $U$ is a weak ultrafilter of order $k+1$. Specifically, for any subset $A\subset X$ with $f(A)\leq k$ and $A\notin U$, we must show $X\setminus A\in U$.
Let $G$ be the upward closure of $U\cup\{A\}$, meaning $G$ includes all sets in $U$ as well as any sets that can be formed by taking unions of elements from $U$ and $A$.
By the assumption that $U$ is maximal, $G$ cannot form a weak filter. If $G$ were a weak filter, it would contradict the maximality of $U$.
The fact that $G$ cannot be a weak filter means that there must exist some set $Z\subseteq X$ in $G$ such that $Z\supset A$, and for some $Y\in U$, we have $Z\cap Y=\emptyset$. This condition arises because if $Z$ and $Y$ were both in $G$, it would imply that $Z$ and $Y$ must intersect, which they do not in this scenario.
The existence of such a $Z$ means that the complement $X\setminus A$ must be in $U$ to maintain the weak ultrafilter property (since $U$ must contain the complement if it does not contain the original set).
This confirms that $U$ must contain $X\setminus A$, proving that $U$ behaves like a weak ultrafilter of order $k+1$. This proof is completed.
In the future, we plan to further investigate the theorems related to ultrafilters on infinite connectivity systems. For example, we will consider the following definitions of Selective and Weakly Selective Ultrafilters on a Connectivity System (cf. [157, 154, 474, 240]).
Definition 5.17.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. Assume that for all $x\in X$, $f(\{x\})\leq k$. An ultrafilter $F$ on $(X,f)$ is called selective if for every function $g:X\to X$, there exists a subset $A\in F$ such that $g\mid A$ is either:
  • Constant: $g(x)=c$ for all $x\in A$, or
  • One-to-One: $g(x_{1})\neq g(x_{2})$ for all distinct $x_{1},x_{2}\in A$.
Definition 5.18.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. An ultrafilter $F$ on $(X,f)$ is called quasi-selective if for every function $g:X\to X$ with $g(x)\leq x$ for all $x\in X$ where $f(\{x\})\leq k$, there exists a subset $A\in F$ such that $g\mid A$ is a non-decreasing function.
5.1.1 P-point and Q-point on an infinite connectivity system
A P-Point is an ultrafilter such that for any partition of a set into subsets not in the ultrafilter, there exists a subset with finite intersections. A Q-Point is an ultrafilter where, for any partition of a set into finite pieces, there exists a subset with intersections of at most one element. P-Points and Q-Points are studied in fields such as topology [155, 305, 987, 880]. We explore P-Points and Q-Points within the context of an infinite connectivity system.
Definition 5.19.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. Additionally, let $\{X_{n}\}_{n<\omega}$ be a symmetric submodular partition of $X$ such that $f(X_{n})\leq k$ for all $n$.
Ramsey (Selective) Point of Order $k+1$: A subset $A\subseteq X$ is called a Ramsey Point of order $k+1$ if for every symmetric submodular partition $\{X_{n}\}_{n<\omega}$ of $X$ such that $f(X_{n})\leq k$ and $X_{n}\notin F$ for all $n$, there exists an $A\in F$ such that:
$|A\cap X_{n}|\leq 1,\quad f(A\cap X_{n})\leq k,\quad\text{and}\quad f(A)\leq k\quad\text{for all }n<\omega.$
P-Point (Weakly Selective Point) of Order $k+1$: A subset $A\subseteq X$ is called a P-Point of order $k+1$ if for every symmetric submodular partition $\{X_{n}\}_{n<\omega}$ of $X$ such that $f(X_{n})\leq k$ and $X_{n}\notin F$ for all $n$, there exists an $A\in F$ such that:
$|A\cap X_{n}|<\omega,\quad f(A\cap X_{n})\leq k,\quad\text{and}\quad f(A)\leq k\quad\text{for all }n<\omega.$
Q-Point of Order $k+1$: A subset $A\subseteq X$ is called a Q-Point of order $k+1$ if for every symmetric submodular partition $\{X_{n}\}_{n<\omega}$ of $X$ into finite sets, such that $f(X_{n})\leq k$ and $X_{n}\notin F$ for all $n$, there exists an $A\in F$ such that:
$|A\cap X_{n}|\leq 1,\quad f(A\cap X_{n})\leq k,\quad\text{and}\quad f(A)\leq k\quad\text{for all }n<\omega.$
Semi-Q-Point (Rapid Point) of Order $k+1$: A subset $A\subseteq X$ is called a Semi-Q-Point of order $k+1$ if for every symmetric submodular partition $\{X_{n}\}_{n<\omega}$ of $X$ into finite sets, such that $f(X_{n})\leq k$ and $X_{n}\notin F$ for all $n$, there exists an $A\in F$ such that:
$|A\cap X_{n}|\leq n,\quad f(A\cap X_{n})\leq k,\quad\text{and}\quad f(A)\leq k\quad\text{for all }n<\omega.$
Theorem 5.20.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. An ultrafilter $F$ on an infinite connectivity system $(X,f)$ is Ramsey (Selective) of order $k+1$ if and only if it is both a P-Point and a Q-Point of order $k+1$.
Proof 5.21.
Assume $F$ is a Ramsey ultrafilter of order $k+1$. By definition:
  1. 1.
    Since $F$ is Ramsey, for every symmetric submodular partition $\{X_{n}\}_{n<\omega}$ of $X$, there exists an $A\in F$ such that $|A\cap X_{n}|\leq 1$ and $f(A\cap X_{n})\leq k$ for all $n<\omega$.
    • This directly implies that $A$ also satisfies the conditions for a Q-Point, as it meets the finite intersection condition $|A\cap X_{n}|\leq 1$.
  2. 2.
    Similarly, the condition $|A\cap X_{n}|\leq 1$ (from being Ramsey) trivially implies $|A\cap X_{n}|<\omega$, satisfying the requirement for a P-Point.
Thus, every Ramsey ultrafilter of order $k+1$ is also a P-Point and a Q-Point of the same order.
Now, assume $F$ is both a P-Point and a Q-Point of order $k+1$. We need to show that $F$ is Ramsey of order $k+1$.
Let $\{X_{n}\}_{n<\omega}$ be a symmetric submodular partition of $X$ such that $X_{n}\notin F$ and $f(X_{n})\leq k$ for all $n$.
  1. 1.
    Since $F$ is a P-Point of order $k+1$, we can choose a set $A_{0}\in F$ such that $|A_{0}\cap X_{n}|<\omega$ and $f(A_{0}\cap X_{n})\leq k$ for all $n<\omega$.
  2. 2.
    Enumerate the elements of $X\setminus\bigcup_{n<\omega}(A_{0}\cap X_{n})=\{a_{n}\}_{n<\omega}$ and define new sets $Y_{2n}:=A_{0}\cap X_{n}$ and $Y_{2n+1}:=\{a_{n}\}$.
  3. 3.
    Since $F$ is a Q-Point of order $k+1$, we can choose a set $A_{1}\in F$ such that $|A_{1}\cap Y_{n}|\leq 1$ and $f(A_{1}\cap Y_{n})\leq k$ for all $n<\omega$.
  4. 4.
    Define $A:=A_{0}\cap A_{1}$. Then $A\in F$, and for each $n<\omega$, we have:
    $|A\cap X_{n}|=|(A_{0}\cap X_{n})\cap A_{1}|=|Y_{2n}\cap A_{1}|\leq 1$
    and $f(A\cap X_{n})\leq k$.
Since $A$ satisfies the condition $|A\cap X_{n}|\leq 1$ for all $n$, and $f(A)\leq k$, $F$ is a Ramsey ultrafilter of order $k+1$.
5.2 Ultrafilter on Countable Connectivity System
We consider ultrafilters on countable (infinite) connectivity systems. Let $X$ be a countable set, which means it can be put into a one-to-one correspondence with the set of natural numbers $\mathbb{N}$.
In mathematics, some researchers focus on the study of games used to model situations with multiple adversaries who have conflicting interests (e.g., [256, 295, 1191, 786, 417, 184, 482, 963]). We consider the investigation of the Ultrafilter Game [212, 306, 115] on a countable connectivity system. This game is an analogy of the well-known ultrafilter game in the context of countable set theory.
Game 5.22 (Ultrafilter Game on a Countable Connectivity System).
Let $X$ be a finite set and $D\subseteq X^{\mathbb{N}}$. We define a two-player game $G(D)$ as follows:
  1. 1.
    Players and Moves: Player I and Player II take turns playing elements from $X$.
    • Player I starts by playing $a_{0}\in X$.
    • Player II responds with $a_{1}\in X$.
    • Player I then plays $a_{2}\in X$, and the game continues in this alternating fashion.
  2. 2.
    Rounds: The game is played for countably many rounds, resulting in a sequence $a=(a_{0},a_{1},a_{2},\ldots)\in X^{\mathbb{N}}$.
  3. 3.
    Winning Condition:
    • Player I wins if the sequence $a$ belongs to $D$ and satisfies the condition $f(\{a\})\leq k$.
    • Player II wins if either the sequence $a$ does not belong to $D$ or $f(\{a\})>k$.
  4. 4.
    Strategies:
    • Strategy for Player I: A strategy for Player I is a rule determining Player I’s moves based on the history of the game. A winning strategy guarantees that Player I can always produce a sequence $a$ such that $a\in D$ and $f(\{a\})\leq k$, regardless of Player II’s moves.
    • Strategy for Player II: Similarly, a strategy for Player II is a rule determining Player II’s responses based on the history of the game. A winning strategy for Player II ensures that the resulting sequence $a$ either does not belong to $D$ or violates the condition $f(\{a\})\leq k$.
  5. 5.
    Determination: We say that $D\subseteq X^{\mathbb{N}}$ is determined if one of the two players has a winning strategy for the game $G(D)$. This means there exists a strategy such that either Player I or Player II can always secure a win, ensuring the game is resolved definitively in favor of one player.
We now consider the winning strategy and determination.
Lemma 5.23.
Let $X$ be a finite set and $f$ be a symmetric submodular function. For the ultrafilter game on a countable connectivity system $(X,f)$, Player I has a winning strategy by following a linear decomposition strategy.
Proof 5.24.
Consider a caterpillar tree $C$ as a path $(l_{1},b_{2},b_{3},\ldots,b_{n-1},l_{n})$, where the subgraph of $C$ induced by $\{b_{i-1},b_{i},b_{i+1}\}$ forms a connectivity system.
If $X$ has only one element, $e_{1}$, Player I starts by choosing $a_{0}=e_{1}$. Since there are no other elements, $f(\{e_{1}\})\leq k$. Player I wins because the sequence $(e_{1})$ belongs to the set $D$ and satisfies $f((e_{1}))\leq k$.
Assume that for a caterpillar tree with $n-1$ elements, Player I has a winning strategy by choosing elements according to the linear decomposition. For a caterpillar tree with $n$ elements, let the linear decomposition partition $X$ into $\{e_{1},e_{2},\ldots,e_{n}\}$.
At each turn $i$, Player I chooses $a_{i}=e_{i+1}$. This ensures that the partial sequence $(a_{0},a_{1},\ldots,a_{i})$ satisfies $f((a_{0},a_{1},\ldots,a_{i}))\leq k$.
Since the linear decomposition ensures the sequence adheres to the symmetric submodular condition, and Player I can choose elements to keep $f((a_{0},a_{1},\ldots,a_{i}))\leq k$, the sequence $a$ will belong to the set $D$.
By induction, Player I has a winning strategy for any $n$ elements by following the linear decomposition. Therefore, Player I can systematically choose elements according to the linear decomposition strategy, ensuring the sequence meets the conditions required for victory. This proof is completed.
Lemma 5.25.
Let $X$ be a finite set and $f$ be a symmetric submodular function. For the ultrafilter game on a countable connectivity system $(X,f)$, a non-principal ultrafilter does not result in a determined game.
Proof 5.26.
If $X$ has only one element, it cannot form a non-principal ultrafilter since the ultrafilter would include singletons, which contradicts the non-principal property.
Assume that for any set of size $n-1$, the non-principal ultrafilter does not lead to a determined game. For a set $X$ with $n$ elements, the non-principal ultrafilter excludes all singletons, making it impossible for Player I to use single elements to form a winning strategy.
Player II also cannot formulate a winning strategy because the non-principal nature prevents clear counteractions based on singletons or small subsets.
The lack of singletons in the ultrafilter means that neither player can systematically form a winning sequence based on single elements or small subsets.
Both players are left without a structured method to guarantee victory, leading to a non-determined game.
By induction, the non-principal ultrafilter does not lead to a determined game for any finite $n$. The exclusion of singletons and the lack of structured strategies ensure that neither player can develop a definitive winning strategy. This proof is completed.
5.3 Elimination Games with an Invisible Robber on a Countable Connectivity System
We consider another game called "elimination games" (cf. [450]). We propose the definition of the elimination games with an invisible robber on a countable connectivity system.
Game 5.27 (Elimination Game with an Invisible Robber).
[483] Let $(X,f)$ be a countable connectivity system equipped with a symmetric submodular function $f$, and let $k$ be an integer. The elimination game on this system is a pursuit-evasion game between a cop and a robber, with the following rules:
  1. 1.
    Players and Pieces:
    • The cop controls one or more tokens (representing cops).
    • The robber controls a single token (representing the robber).
    • Both the cops and the robber move on the elements of $X$.
  2. 2.
    Initial Setup:
    • Initially, no tokens occupy any elements of $X$.
    • The game begins with the cop placing some tokens on the elements of $X$ according to the function $f$.
    • The robber then chooses an initial position $r\in X$ without being seen by the cop, ensuring that $f(\{r\})\leq k$.
  3. 3.
    Moves:
    • In each round, the cop announces a new position $C_{i}$ by moving the tokens to a set of elements $C_{i+1}\subseteq X$, ensuring that $f(C_{i+1})\leq k$.
    • The robber, while invisible to the cop, can move from their current position $r$ along or against a directed path not containing any element of $C_{i+1}$ to a new position $r^{\prime}$, ensuring $f(\{r^{\prime}\})\leq k$.
  4. 4.
    Movement Rules:
    • The robber can move to any element $r^{\prime}$ within the connected component of $X\setminus C_{i+1}$ containing $r$, under the condition that $f(\{r^{\prime}\})\leq k$.
    • The cop cannot see the robber’s moves, so their strategy must account for all possible positions the robber might occupy, taking into account the submodularity condition: $f(A)+f(B)\geq f(A\cup B)+f(A\cap B)$ for any subsets $A,B\subseteq X$.
  5. 5.
    Winning Conditions:
    • The cop wins if they place a token on the element currently occupied by the robber, and the robber is unable to escape to a safe place within the connectivity constraints $f$.
    • The robber wins if they can always elude capture, continuously finding a path $r^{\prime}$ that satisfies the connectivity condition $f(\{r^{\prime}\})\leq k$ within the connected components of $X\setminus C_{i+1}$.
We can represent a cop’s strategy on a connectivity system $(X,f)$ by a finite or infinite sequence $S$ of cop positions: $S:=\{C_{0},C_{1},C_{2},\ldots\}$ where $C_{i}$ denotes the set of elements occupied by the cops at the $i$-th move.
Conjecture 5.28.
The robber’s winning strategy can be characterized as a (non-principal) single ultrafilter if the robber can always find a path to escape from the cops’ positions within the connected components of $X\setminus C_{i+1}$.
In the future, we plan to further investigate the theorems related to ultrafilters on countable connectivity systems.
5.4 Ultrafilter on Maximum-Connectivity System
We introduce a property similar to submodularity, which supports the development of corresponding theories for connectivity functions.
Definition 5.29 (cf. [460]).
A set function $\kappa$ on a finite underlying set $X$ is defined as maximum-submodular if, for all subsets $A,B\subseteq X$,
$\max(\kappa(A),\kappa(B))\geq\max(\kappa(A\cap B),\kappa(A\cup B)).$
According to [460], a set function that is normalized, symmetric, and maximum-submodular is referred to as a maximum-submodular connectivity function. A maximum-connectivity system consists of a pair $(U,\kappa)$, where $U$ is a finite set and $\kappa$ is a maximum-submodular connectivity function.
We plan to define an ultrafilter on a maximum-connectivity system and explore its relationship with tangles on maximum-connectivity systems in future research.
6 Future Tasks
This section outlines future perspectives for this study and related research. Although still in the conceptual stage or under review for submission, I aim to outline these ideas to help advance future research on graph width parameters and ultrafilters.
6.1 New Width/Length/Depth Parameters and Obstructions
In this paper, we discussed an ultrafilter that serves as an obstruction to branch-width and linear branch-width. In the future, we will consider the following width parameters. As noted in Appendix A, various width parameters are well-known. Our goal is to examine these width parameters from multiple perspectives, including their corresponding decompositions, related length [1163, 358, 334] and depth parameters [401, 328, 945] and Appendix D, linear concepts, directed graph concepts, and their extensions from graphs to connectivity systems [565] and abstract separation systems [347, 336]. For a detailed discussion on length, refer to Appendix B. Additionally, we plan to define and explore the characteristics of directed branch depth, similar to directed tree depth [804], and Directed Rank-depth, similar to Rank-depth [716, 955, 761, 717]. Furthermore, we aim to investigate the relationships with ultrafilters and related definitions.
6.1.1 Consideration about a Branch Distance Decomposition
First, let’s consider the concept of a branch distance decomposition. Intuitively, this is a branch decomposition that integrates the concept of distance. This idea originates from exploring how extending tree-distance-width [1233] and path-distance-width [1233] to branch decompositions might work.
Definition 6.1 (Branch Distance Decomposition).
A branch distance decomposition is a branch decomposition with an additional distance function and a root. A branch decomposition of a graph $G=(V,E)$ is a pair $B=(T,\mu)$, where $T$ is a tree and $\mu$ is a bijection from the edges of $G$ to the leaves of $T$. In a branch distance decomposition, we add a distance function $d:E(T)\rightarrow\mathbb{N}$, which assigns a distance to each edge in $T$, and a root $r\in V(T)$.
For an edge $e\in E(T)$, let $T_{1}$ and $T_{2}$ be the two subtrees obtained by deleting $e$ from $T$. Let $G_{1}$ and $G_{2}$ be the subgraphs of $G$ induced by the edges mapped by $\mu$ to the leaves of $T_{1}$ and $T_{2}$, respectively. The vertices $V(G_{1})\cap V(G_{2})$ are denoted by $\text{mid}_{B}(e)$.
The branch distance width of $B$ is $\max_{e\in E(T)}(\lvert\text{mid}_{B}(e)\rvert+d(e))$. The branch distance width of $G$, denoted by $\text{BDW}(G)$, is the minimum width over all branch distance decompositions of $G$.
A linear distance decomposition of a graph $G=(V,E)$ is a triple $(P,\mu,d)$, where:
  • $P=(v_{1},v_{2},\dots,v_{n})$ is a path graph, represented as a sequence of vertices.
  • $\mu$ is a bijection from the edges of $G$ to the vertices of $P$.
  • $d:V(P)\rightarrow\mathbb{N}$ is a function assigning a distance to each vertex in $P$.
For a vertex $v_{i}\in V(P)$, let $P_{i}$ be the subpath consisting of vertices $v_{i-1},v_{i},v_{i+1}$. Let $G_{i}$ be the subgraph of $G$ induced by the edges mapped by $\mu$ to the vertices in $P_{i}$. The set of vertices $V(G_{i})$ is denoted by $\text{mid}_{L}(v_{i})$.
The width of a linear distance decomposition is defined as:
$\text{ldw}(P,\mu,d)=\max_{v_{i}\in V(P)}(\lvert\text{mid}_{L}(v_{i})\rvert+d(v_{i})).$
The linear distance width of $G$, denoted by $\text{LDW}(G)$, is the minimum width over all linear distance decompositions of $G$:
$\text{LDW}(G)=\min_{(P,\mu,d)}\text{ldw}(P,\mu,d).$
Conjecture 6.2.
For any connected graph $G$ with $\text{bw}(G)\geq 2$, $\text{TDW}(G)=\text{BDW}(G)$. Also, for any connected graph $G$ with $\text{lw}(G)\geq 2$, $\text{PDW}(G)=\text{LDW}(G)$.
Next, we consider an algorithm for constructing a branch distance decomposition. The following algorithm has significant room for improvement, and we plan to continue refining it.
Algorithm 6.3 (Constructing Branch Distance Decomposition).
  • Input: A graph $G=(V,E)$.
  • Output: A branch distance decomposition $(T,\{X_{i}\})$.
  1. 1.
    Initialization:
    • Let $T$ be a tree with a single node $r$ (the root).
    • Initialize $X_{r}=V(G)$.
  2. 2.
    Iterative Decomposition:
    • While $\lvert X_{r}\rvert>1$:
      1. (a)
        Select a vertex $v\in X_{r}$.
      2. (b)
        Find the shortest path $P$ in $G$ that includes $v$ and other vertices in $X_{r}$.
      3. (c)
        Partition $X_{r}$ into two sets $X_{1}$ and $X_{2}$ based on the vertices on either side of $P$ such that:
        • $X_{1}\cup X_{2}=X_{r}$,
        • $X_{1}\cap X_{2}$ contains the vertices on $P$.
      4. (d)
        Add two new nodes $i$ and $j$ to $T$ as children of $r$, and set:
        • $X_{i}=X_{1}$,
        • $X_{j}=X_{2}$.
      5. (e)
        Update $T$ and the corresponding sets $X_{i}$ and $X_{j}$.
  3. 3.
    Ensuring Distance Constraints:
    • For each vertex $v\in V(G)$, ensure that if $v\in X_{i}$, then $\text{dist}_{G}(X_{r},v)=\text{dist}_{T}(r,i)$.
    • Adjust $X_{i}$ and $X_{j}$ to meet this distance constraint by reassigning vertices if necessary.
  4. 4.
    Repeat: Continue decomposing each subset $X_{i}$ recursively, applying the above steps until each $X_{i}$ contains a single vertex or a base case is met.
  5. 5.
    Output the Decomposition: Once the tree $T$ and the subsets $\{X_{i}\}$ satisfy all the constraints, output the branch distance decomposition $(T,\{X_{i}\})$.
Lemma 6.4.
Above algorithm is correct.
Proof 6.5.
  1. 1.
    The decomposition process ensures that $T$ remains a tree, as new nodes are added as children to existing nodes, maintaining the tree structure.
  2. 2.
    Each vertex $v\in V(G)$ is included in at least one subset $X_{i}$ throughout the decomposition, ensuring $\bigcup_{i\in V(T)}X_{i}=V(G)$.
  3. 3.
    For each edge $\{u,v\}\in E(G)$, the decomposition ensures there are nodes $i,j\in V(T)$ such that $u\in X_{i}$ and $v\in X_{j}$, and either $i=j$ or $\{i,j\}\in E(T)$.
  4. 4.
    The algorithm maintains the distance constraint by ensuring $\text{dist}_{G}(X_{r},v)=\text{dist}_{T}(r,i)$ for each vertex $v\in V(G)$, adjusting subsets as necessary to maintain this property.
  5. 5.
    The iterative and recursive nature of the algorithm ensures that the decomposition process will terminate, as each step reduces the size of the subsets until the base case is reached.
Therefore, the lemma is proved.
Lemma 6.6.
The total time complexity of Algorithm 6.4 is $O((V+E)\log V)$.
Proof 6.7.
  • Initialization: The initialization step takes $O(1)$ time.
  • Iterative Decomposition: Each iteration involves selecting a vertex $v\in X_{r}$ and finding a shortest path $P$. Finding a shortest path can be done in $O(V+E)$ time using breadth-first search (BFS) or depth-first search (DFS). Partitioning $X_{r}$ into $X_{1}$ and $X_{2}$ based on the shortest path takes $O(V)$ time. Adding new nodes to the tree $T$ and updating the corresponding sets $X_{i}$ and $X_{j}$ takes $O(1)$ time.
  • Ensuring Distance Constraints: Adjusting subsets $X_{i}$ and $X_{j}$ to meet distance constraints might involve checking distances for all vertices, which can be done in $O(V)$ time.
  • Recursive Decomposition: The recursive decomposition process continues until each subset $X_{i}$ contains a single vertex, resulting in a tree of height $O(\log V)$.
Total Time Complexity: The total time complexity for the entire algorithm can be expressed as a sum of the time complexities of the individual steps. Each step involves $O(V+E)$ time for finding the shortest path, and there are $O(\log V)$ levels of recursion. Therefore, the lemma is proved.
And we will consider a game called Cops and Robbers with Speed $d$. The definition of the game is as follows:
Game 6.8.
Cops and Robbers with Speed $d$
In the "cops and robbers with speed $d$" game, the graph $G=(V,E)$ is endowed with a distance function that affects movement capabilities. The players consist of a single robber and multiple cops.
  • Players: One robber and $k$ cops.
  • Positions: Defined by a tuple $(r,C)$, where $r\in V$ is the robber’s position and $C\subseteq V$ with $|C|\leq k$ represents the positions of the cops.
  • Initial Position: The robber chooses an initial position $r_{0}\in V$, and the cops collectively choose their initial positions as a set $C$ where $|C|\leq k$.
  • Movement: Each round consists of the cops moving to a new set of positions $C^{\prime}$ followed by the robber moving to a new position $r^{\prime}$. The robber can move to any vertex $r^{\prime}$ that is reachable from $r$ within distance $d$ in $G\setminus(C\cap C^{\prime})$, reflecting the robber’s speed and the obstruction caused by the cops.
  • Winning Conditions: The cops win if they can move into the robber’s position, i.e., if a situation $(r,C)$ with $r\in C$ is achieved; otherwise, the robber wins if he remains uncaught indefinitely.
Conjecture 6.9.
Branch-distance-decomposition is a winning strategy for Cops and Robbers with Speed $d$.
6.1.2 Consideration about a Directed Tree Distance Decomposition
Next, we will consider a directed tree distance decomposition[485]. This idea also originates from exploring how extending tree-distance-width [1233] and path-distance-width [1233] to branch decompositions might work.
Definition 6.10.
[485] Let $G=(V,E)$ be a directed graph. A directed tree-distance decomposition of $G$ is a triple $(R,\{X_{i}\mid i\in V(R)\},r)$, where:
  • $R$ is an arborescence (a directed tree with a designated root $r$ such that there is a unique directed path from $r$ to any vertex in $V(R)$).
  • $\{X_{i}\mid i\in V(R)\}$ is a family of subsets of $V(G)$ satisfying the following properties:
    • (Partition Property): $\bigcup_{i\in V(R)}X_{i}=V(G)$ and $X_{i}\cap X_{j}=\emptyset$ for $i\neq j$.
    • (Distance Property): For each vertex $v\in V(G)$, if $v\in X_{i}$, then the directed distance from $X_{r}$ (the set associated with the root $r$) to $v$ in $G$ is equal to the directed distance from $r$ to $i$ in $R$. Formally, $d_{G}(X_{r},v)=d_{R}(r,i)$.
    • (Edge Coverage Property): For each edge $(u,v)\in E(G)$, there exist vertices $i,j\in V(R)$ such that $u\in X_{i}$, $v\in X_{j}$, and either $i=j$ or $(i,j)\in E(R)$.
The width of the directed tree-distance decomposition is defined as $\max_{i\in V(R)}|X_{i}|$. The directed tree-distance width of a directed graph $G$, denoted by $\text{DTDW}(G)$, is the minimum width over all possible directed tree-distance decompositions of $G$.
A directed path-distance decomposition of a directed graph $G=(V,E)$ is a special case of the directed tree-distance decomposition where the tree $T$ is a directed path. Formally, a directed path-distance decomposition is a sequence of subsets $(X_{1},X_{2},\dots,X_{t})$ of the vertex set $V$, satisfying the following conditions:
  1. 1.
    $\bigcup_{i=1}^{t}X_{i}=V$, meaning that the union of all subsets $X_{i}$ covers the entire vertex set $V$,
  2. 2.
    For each vertex $v\in V$, if $v\in X_{i}$, then the directed distance $d_{G}(X_{1},v)$ from the root set $X_{1}$ in the directed graph $G$ to $v$ corresponds to the position of $X_{i}$ in the sequence, i.e., $d_{G}(X_{1},v)=i-1$,
  3. 3.
    For each directed edge $(u,v)\in E$, there exist indices $i$ and $j$ such that $u\in X_{i}$, $v\in X_{j}$, and either $i=j$ or $|i-j|=1$, indicating that the vertices $u$ and $v$ must either be in the same subset or in consecutive subsets.
The width of a directed path-distance decomposition $(X_{1},X_{2},\dots,X_{t})$ is defined as the maximum size of any set $X_{i}$ in the decomposition, i.e.,
$\text{width}\left((X_{1},X_{2},\dots,X_{t})\right)=\max_{1\leq i\leq t}|X_{i}|.$
The directed path-distance width (DPDW) of the directed graph $G$ is defined as the minimum width over all possible directed path-distance decompositions of $G$.
In the future, we will investigate the relationships with ultrafilters and these distance widths.
6.1.3 New Linear Width Parameter
As discussed in Appendix A, various graph width parameters have been defined and extensively studied. Our goal is to introduce new linear versions of some of these width parameters and explore their relationships with other graph parameters. Specifically, we plan to define Linear-Amalgam-Decomposition, the linear version of Amalgam-Decomposition [860], Linear-Modular-Decomposition, the linear version of Modular-Decomposition [5], Linear-Tree-Cut-Decomposition, the linear version of Tree-Cut Decomposition, and Direct Linear-Branch-Decomposition, the linear version of Direct Branch-Decomposition, and analyze their interconnections. It’s important to note that for any general width parameter, it holds that (general width parameter) $\leq$ (linear width parameter); for instance, Amalgam-width $\leq$ Linear-Amalgam-width.
These restrictions to underlying path structures are often advantageous in establishing results for general parameters. Moreover, these linear parameters provide valuable structural insights, particularly in the study of special graph classes. Additionally, the relationship between path-structured graph width parameters and game theory, specifically in the context of the "invisible cops and robbers" game, is frequently discussed [754, 325, 1228]. This is a game where the cops can catch the robber even without knowing the exact location of the robber. We also aim to explore potential connections between these widths and ultrafilters, further enriching the theoretical framework.
6.1.4 Matroid Length Parameters
We will explore new matroid length parameters [1164, 489]. The most well-known length parameter already established is tree-length. Tree-length measures the maximum distance between any two vertices within a single bag of a tree-decomposition, minimized over all possible decompositions [301, 368, 369]. Although still in the conceptual stage, the definitions are outlined below[489].
Definition 6.11.
[489] A tree decomposition of a matroid $M$ on the ground set $E=E(M)$ is a pair $(T,\tau)$, where $T$ is a tree and $\tau:E\to V(T)$ is a mapping from the elements of $E$ to the nodes of $T$. For each node $i\in V(T)$, let $X_{i}=\tau^{-1}(i)$ be the set of elements of $E$ mapped to $i$.
For each $X_{i}$, consider the rank function $r_{M}$ of the matroid $M$. The diameter of $X_{i}$ is defined as:
$\text{diam}_{M}(X_{i})=\max_{x,y\in X_{i}}\left(r_{M}(\{x,y\})-\min(r_{M}(\{x\}),r_{M}(\{y\}))\right).$
The tree-length of the decomposition $(T,\tau)$, denoted by $\text{tl}(T,\tau)$, is defined as the maximum diameter of $X_{i}$ over all nodes $i$ in $T$:
$\text{tl}(T,\tau)=\max_{i\in V(T)}\text{diam}_{M}(X_{i}).$
The tree-length of the matroid $M$, denoted by $\text{tl}(M)$, is the minimum tree-length over all possible tree decompositions of $M$:
$\text{tl}(M)=\min_{(T,\tau)}\text{tl}(T,\tau).$
Definition 6.12.
[489] A branch decomposition of a matroid $M=(E,I)$ is a pair $(T,\mu)$, where $T$ is a subcubic tree without degree-2 nodes, and $\mu$ is a bijection from the elements of $E$ to the leaves of $T$. For each edge $e\in E(T)$, let $T_{1}$ and $T_{2}$ be the two subtrees obtained by deleting $e$ from $T$, and let $E_{1}$ and $E_{2}$ be the sets of elements of $E$ corresponding to the leaves in $T_{1}$ and $T_{2}$, respectively. The set $E_{1}\cap E_{2}$ is denoted by $\text{mid}_{B}(e)$.
For each $\text{mid}_{B}(e)$, consider the rank function $r_{M}$. The diameter of $\text{mid}_{B}(e)$ is defined as:
$\text{diam}_{M}(\text{mid}_{B}(e))=\max_{x,y\in\text{mid}_{B}(e)}\left(r_{M}(\{x,y\})-\min(r_{M}(\{x\}),r_{M}(\{y\}))\right).$
The branch-length of the decomposition $(T,\mu)$, denoted by $\text{bl}(T,\mu)$, is defined as the maximum diameter of $\text{mid}_{B}(e)$ over all edges $e$ in $T$:
$\text{bl}(T,\mu)=\max_{e\in E(T)}\text{diam}_{M}(\text{mid}_{B}(e)).$
The branch-length of the matroid $M$, denoted by $\text{bl}(M)$, is the minimum branch-length over all possible branch decompositions of $M$:
$\text{bl}(M)=\min_{(T,\mu)}\text{bl}(T,\mu).$
6.1.5 Branch-breadth and Linear-Breadth
We will explore the concepts of branch-breadth and linear-breadth for graphs. A parameter closely related to tree-length is tree-breadth, which measures how closely the vertices in each bag of a tree-decomposition can be clustered around a central vertex within a specific radius, minimized over all possible decompositions [828, 827]. A path-based version of this parameter, known as path-breadth, is also well-established. We are currently investigating the corresponding versions for branch-decompositions[828, 827]. Although still in the conceptual stage, the definitions are outlined below [484].
Definition 6.13.
[484] A branch decomposition of a graph $G=(V,E)$ is a pair $(T,\sigma)$, where $T$ is a tree with vertices of degree at most 3, and $\sigma$ is a bijection from the set of leaves of $T$ to $E$. For an edge $e$ in $T$, consider the two subtrees $T_{1}$ and $T_{2}$ obtained by removing $e$ from $T$. Let $V_{1}$ and $V_{2}$ be the sets of vertices in $G$ that are incident to edges corresponding to leaves in $T_{1}$ and $T_{2}$, respectively.
The breadth of an edge $e$ in $T$ is defined as the minimum $\rho$ such that for each $v\in V_{1}\cap V_{2}$, the set of vertices $\{v\}\cup(V_{1}\setminus V_{2})\cup(V_{2}\setminus V_{1})$ is contained within the closed neighborhood $N_{\rho}(v)$ of some vertex $v$ in $G$. The branch-breadth of a graph $G$, denoted as $\text{bb}(G)$, is the minimum breadth over all possible branch decompositions of $G$.
Definition 6.14.
[484] A linear decomposition of a graph $G=(V,E)$ is a linear ordering of the edges $(e_{1},e_{2},\dots,e_{m})$. For each index $i$ between 1 and $m-1$, let $V_{1}$ be the set of vertices incident to edges in $\{e_{1},\dots,e_{i}\}$, and $V_{2}$ be the set of vertices incident to edges in $\{e_{i+1},\dots,e_{m}\}$.
The breadth of the linear decomposition is the minimum $\rho$ such that for each vertex $v$ in $V_{1}\cap V_{2}$, the set $V_{1}\cup V_{2}$ is contained within the closed neighborhood $N_{\rho}(v)$ of some vertex $v$ in $G$. The linear-breadth of a graph $G$, denoted as $\text{lb}(G)$, is the minimum breadth over all possible linear decompositions of $G$.
6.1.6 Examination for Directed proper-path-width
We will explore the concept of proper-path-width for graphs. Proper-path-width [1131, 1130, 1138, 1133] measures how "tightly" a graph can be decomposed into overlapping vertex subsets, ensuring that no two subsets are identical while maintaining specific connectivity constraints. We are interested in extending this definition to directed graphs. Although still in the conceptual stage, the definition is outlined below [485].
Definition 6.15.
[485] Let $\mathcal{P}=(X_{1},X_{2},\dots,X_{r})$ be a sequence of subsets of vertices of a directed graph $G=(V,E)$. The width of $\mathcal{P}$ is defined as $\max_{1\leq i\leq r}|X_{i}|-1$. The sequence $\mathcal{P}$ is called a directed-proper-path-decomposition of $G$ if the following conditions are satisfied:
  1. 1.
    For any distinct $i$ and $j$, $X_{i}\neq X_{j}$.
  2. 2.
    $\bigcup_{i=1}^{r}X_{i}=V(G)$.
  3. 3.
    For any edge $(u,v)\in E(G)$, there exists an $i$ such that $u,v\in X_{i}$.
  4. 4.
    For all $1\leq l<m<n\leq r$, $X_{l}\cap X_{n}\subseteq X_{m}$.
  5. 5.
    For all $1\leq l<m<n\leq r$, $|X_{l}\cap X_{n}|\leq|X_{m}|-2$.
The directed-proper-path-width of $G$, denoted by $\text{dppw}(G)$, is the minimum width over all directed-proper-path-decompositions of $G$. If a sequence $\mathcal{P}$ satisfies only conditions (i)-(iv), it is called a directed-path-decomposition of $G$, and the directed-path-width of $G$, denoted by $\text{dpw}(G)$, is the minimum width over all directed-path-decompositions of $G$. A (directed-proper-)path-decomposition with width $k$ is called a $k$-(directed-proper-)path-decomposition.
6.1.7 Fuzzy Tree-width
We will delve into the concept of tree-width as it applies to fuzzy graphs. Fuzzy set theory, introduced in 1965, was developed to capture the nuances of partial truth, bridging the gap between absolute truth and falsehood [1242]. Building on classical graph theory, fuzzy graphs introduce uncertainty by assigning a membership degree to each edge [1043]. This rapidly advancing field offers intriguing theoretical challenges and significant practical applications [882, 915, 139, 973, 983, 366, 914]. Our goal is to extend the concept of tree-width to fuzzy graphs. The formal definition of a fuzzy graph is presented below.
Definition 6.16.
[1043] A fuzzy graph $G=(\sigma,\mu)$ with $V$ as the underlying set is defined as follows:
  • $\sigma:V\rightarrow[0,1]$ is a fuzzy subset of vertices, where $\sigma(x)$ represents the membership degree of vertex $x\in V$.
  • $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$, such that $\mu(x,y)\leq\sigma(x)\wedge\sigma(y)$ for all $x,y\in V$, where $\wedge$ denotes the minimum operation.
The underlying crisp graph of $G$ is denoted by $G^{*}=(\sigma^{*},\mu^{*})$, where:
  • $\sigma^{*}=\sup p(\sigma)=\{x\in V:\sigma(x)>0\}$
  • $\mu^{*}=\sup p(\mu)=\{(x,y)\in V\times V:\mu(x,y)>0\}$
A fuzzy subgraph $H=(\sigma^{\prime},\mu^{\prime})$ of $G$ is defined as follows:
  • There exists $X\subseteq V$ such that $\sigma^{\prime}:X\rightarrow[0,1]$ is a fuzzy subset.
  • $\mu^{\prime}:X\times X\rightarrow[0,1]$ is a fuzzy relation on $\sigma^{\prime}$, satisfying $\mu^{\prime}(x,y)\leq\sigma^{\prime}(x)\wedge\sigma^{\prime}(y)$ for all $x,y\in X$.
Example 6.17.
(cf.[224, 503, 519]) Consider a fuzzy graph $G=(\sigma,\mu)$ with four vertices $V=\{v_{1},v_{2},v_{3},v_{4}\}$, as depicted in the diagram.
The membership degrees of the vertices are as follows:
$\sigma(v_{1})=0.1,\quad\sigma(v_{2})=0.3,\quad\sigma(v_{3})=0.2,\quad\sigma(v_{4})=0.4$
The fuzzy relation on the edges is defined by the values of $\mu$, where $\mu(x,y)\leq\sigma(x)\land\sigma(y)$ for all $x,y\in V$. The fuzzy membership degrees of the edges are as follows:
$\mu(v_{1},v_{2})=0.1,\quad\mu(v_{2},v_{3})=0.1,\quad\mu(v_{3},v_{4})=0.1$
$\mu(v_{4},v_{1})=0.1,\quad\mu(v_{2},v_{4})=0.3$
In this case, the fuzzy graph $G$ has the following properties:
  • Vertices $v_{1},v_{2},v_{3},v_{4}$ are connected by edges with varying membership degrees.
  • The fuzzy relations ensure that $\mu(x,y)$ for any edge $(x,y)$ does not exceed the minimum membership of the corresponding vertices.
For more detailed information about fuzzy set and fuzzy graph, please refer to lecture notes or surveys[1252, 455, 1252, 915].
Now, We consider about the definitions of fuzzy-tree-decomposition and fuzzy-path-decomposition[494]. This definition generalizes the classical concept of tree-width by incorporating the fuzzy nature of the graph. Although still in the conceptual stage, the definitions are outlined below.
Definition 6.18.
[494] Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing vertex membership, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$. A fuzzy tree-decomposition of $G$ is a pair $(T,\{B_{t}\}_{t\in T})$, where:
  • $T=(I,F)$ is a tree with nodes $I$ and edges $F$.
  • $\{B_{t}\}_{t\in T}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $T$ such that:
    1. 1.
      For each vertex $v\in V$, the set $\{t\in I:v\in B_{t}\}$ is connected in the tree $T$.
    2. 2.
      For each edge $(u,v)\in V\times V$ with membership degree $\mu(u,v)\leq\sigma(u)\wedge\sigma(v)$, there exists some node $t\in I$ such that both $u$ and $v$ are in $B_{t}$, and the membership degree of $u$ and $v$ in $B_{t}$ is at least $\mu(u,v)$.
The width of a fuzzy tree-decomposition $(T,\{B_{t}\}_{t\in T})$ is defined as
$\max_{t\in I}\left(\sup_{v\in B_{t}}\mu(v,B_{t})-1\right),$
where $\mu(v,B_{t})$ represents the maximum membership degree of vertex $v$ in the fuzzy set $B_{t}$. The Fuzzy-Tree-Width of the fuzzy graph $G$ is the minimum width among all possible fuzzy tree-decompositions of $G$.
Definition 6.19.
[494] Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing vertex membership, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$. A fuzzy path-decomposition of $G$ is a pair $(P,\{B_{p}\}_{p\in P})$, where:
  • $P$ is a path with nodes $I$ and edges $F$.
  • $\{B_{p}\}_{p\in P}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $P$ such that:
    1. 1.
      For each vertex $v\in V$, the set $\{p\in I:v\in B_{p}\}$ is connected in the path $P$.
    2. 2.
      For each edge $(u,v)\in V\times V$ with membership degree $\mu(u,v)\leq\sigma(u)\wedge\sigma(v)$, there exists some node $p\in I$ such that both $u$ and $v$ are in $B_{p}$, and the membership degree of $u$ and $v$ in $B_{p}$ is at least $\mu(u,v)$.
The width of a fuzzy path-decomposition $(P,\{B_{p}\}_{p\in P})$ is defined as
$\max_{p\in I}\left(\sup_{v\in B_{p}}\mu(v,B_{p})-1\right),$
where $\mu(v,B_{p})$ represents the maximum membership degree of vertex $v$ in the fuzzy set $B_{p}$. The Fuzzy-Path-Width of the fuzzy graph $G$ is the minimum width among all possible fuzzy path-decompositions of $G$.
And in the future, we will consider about following definitions for fuzzy-graph.
Definition 6.20 (Fuzzy Bandwidth).
Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing the membership degree of vertices, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $V$ representing the membership degree of edges. A layout of the graph $G$ is a bijection $f:V\rightarrow\{1,2,\dots,|V|\}$. The fuzzy bandwidth of the layout $f$ is defined as:
$\max_{\{u,v\}\in E}\left(\mu(u,v)\cdot|f(u)-f(v)|\right).$
The fuzzy bandwidth of the fuzzy graph $G$, denoted $\text{fbw}(G)$, is the minimum fuzzy bandwidth over all possible layouts $f$ of $G$.
Definition 6.21 (Fuzzy Path Distance-Width).
Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing the membership degree of vertices, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$. A fuzzy path-distance decomposition of $G$ is a sequence of fuzzy subsets $(B_{1},B_{2},\dots,B_{r})$ of $V$, called bags, such that:
  1. 1.
    $B_{1}\cup B_{2}\cup\dots\cup B_{r}=V$,
  2. 2.
    For each edge $(u,v)\in E$ with membership degree $\mu(u,v)>0$, there exists an index $i\leq j$ such that $u\in B_{i}$ and $v\in B_{j}$, and the membership degree of $u$ and $v$ in $B_{i}$ and $B_{j}$ is at least $\mu(u,v)$.
  3. 3.
    For all indices $1\leq i<j<\ell\leq r$, $B_{i}\cap B_{\ell}\subseteq B_{j}$.
The width of a fuzzy path-distance decomposition $(B_{1},B_{2},\dots,B_{r})$ is defined as:
$\max_{1\leq i\leq r}\sup_{v\in B_{i}}\mu(v,B_{i})-1.$
The fuzzy path-distance width of $G$, denoted by $\text{fpdw}(G)$, is the minimum width over all possible fuzzy path-distance decompositions of $G$.
Definition 6.22 (Fuzzy Tree-length).
Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing vertex membership, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$. A fuzzy tree-decomposition of $G$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a tree with nodes $I$ and edges $F$.
  • $\{B_{t}\}_{t\in I}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $T$.
The fuzzy length of a tree-decomposition $T(G)$ is defined as:
$\lambda(T)=\max_{t\in I}\max_{u,v\in B_{t}}d_{G^{*}}(u,v),$
where $d_{G^{*}}(u,v)$ denotes the distance between vertices $u$ and $v$ in the underlying crisp graph $G^{*}$. The fuzzy tree-length of $G$, denoted by $\text{ftl}(G)$, is the minimum length over all possible fuzzy tree-decompositions of $G$.
Definition 6.23 (Fuzzy Tree-Breadth).
Let $G=(\sigma,\mu)$ be a fuzzy graph, where $\sigma:V\rightarrow[0,1]$ is a fuzzy subset representing vertex membership, and $\mu:V\times V\rightarrow[0,1]$ is a fuzzy relation on $\sigma$. A fuzzy tree-decomposition of $G$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a tree with nodes $I$ and edges $F$.
  • $\{B_{t}\}_{t\in I}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $T$.
The fuzzy breadth of a tree-decomposition $T(G)$ is the minimum integer $r$ such that for every $t\in I$, there exists a vertex $v_{t}\in V(G^{*})$ with $B_{t}\subseteq D_{r}(v_{t},G^{*})$, where $D_{r}(v_{t},G^{*})$ is the disk of radius $r$ centered at $v_{t}$ in the underlying crisp graph $G^{*}$. The fuzzy tree-breadth of $G$, denoted by $\text{ftb}(G)$, is the minimum breadth over all possible fuzzy tree-decompositions of $G$.
We will also consider about directed fuzzy graph[873, 864, 607].
Definition 6.24.
[864] A directed fuzzy graph is a quadruple $\xi=(V,\sigma,\mu,E)$, where:
  • $V$ is a non-empty set of vertices.
  • $\sigma:V\to[0,1]$ is the membership function for vertices, where $\sigma(a)$ represents the membership value of vertex $a$.
  • $E$ is the set of all directed edges between vertices in $V$.
  • $\mu:E\to[0,1]$ is the membership function for edges, where $\mu(a,b)$ represents the membership value of the directed edge $(a,b)$.
These functions must satisfy the following condition:
$\mu(a,b)\leq\sigma(a)\wedge\sigma(b),$
where $\wedge$ denotes the minimum operation, ensuring that the membership value of an edge cannot exceed the membership values of the vertices it connects.
Given a directed fuzzy graph $\xi=(V,\sigma,\mu,E)$, the measure of influence between two vertices $a,b\in V$ is denoted by $\phi(a,b)$ and is defined as:
$\phi(a,b)\leq|\sigma(a)-\sigma(b)|,$
where $\phi(a,b)$ quantifies the influence based on the difference in their membership values.
Although still in the conceptual stage, the definitions are outlined below (cf.[502]).
Definition 6.25 (Fuzzy Directed Tree-Decomposition).
(cf.[502]) Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed tree-decomposition of $\xi$ is a triple $(T,X,W)$, where:
  • $T=(V_{T},E_{T})$ is a directed tree,
  • $X=\{X_{e}\mid e\in E_{T}\}$ is a set of fuzzy subsets of $V$ (called bags) associated with the edges of $T$,
  • $W=\{W_{r}\mid r\in V_{T}\}$ is a set of fuzzy subsets of $V$ associated with the vertices of $T$, satisfying the following conditions:
    1. 1.
      $W=\{W_{r}\mid r\in V_{T}\}$ is a partition of $V$,
    2. 2.
      For every edge $e=(u,v)\in E_{T}$, the fuzzy set $W_{v}$ is $X_{e}$-normal, meaning $\mu(u,v)\leq\min(\sigma(u),\sigma(v))$.
The width of a fuzzy directed tree-decomposition $(T,X,W)$ is defined as:
$\max_{r\in V_{T}}\left(\sup_{v\in W_{r}\cup\bigcup_{e\sim r}X_{e}}\mu(v,W_{r})-1\right),$
where $e\sim r$ means that $r$ is one of the endpoints of the arc $e$. The fuzzy directed tree-width of $\xi$, denoted by $\text{fdtw}(\xi)$, is the minimum width over all possible fuzzy directed tree-decompositions of $\xi$.
Definition 6.26 (Fuzzy Directed Path-Decomposition).
(cf.[502]) Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed path-decomposition of $\xi$ is a sequence $(X_{1},\dots,X_{r})$ of fuzzy subsets of $V$, called bags, that satisfies the following conditions:
  1. 1.
    $X_{1}\cup\dots\cup X_{r}=V$,
  2. 2.
    For each directed edge $(u,v)\in E$ with membership degree $\mu(u,v)>0$, there exists a pair $i\leq j$ such that $u\in X_{i}$ and $v\in X_{j}$, and the membership degree of $u$ and $v$ in $X_{i}$ and $X_{j}$ is at least $\mu(u,v)$.
  3. 3.
    For all indices $1\leq i<j<\ell\leq r$, it holds that $X_{i}\cap X_{\ell}\subseteq X_{j}$.
The width of a fuzzy directed path-decomposition $(X_{1},\dots,X_{r})$ is defined as:
$\max_{1\leq i\leq r}\sup_{v\in X_{i}}\mu(v,X_{i})-1.$
[…]
Definition 6.27 (Fuzzy Directed Cut-Width).
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. The fuzzy directed cut-width of $\xi$ is defined as:
[…]
where $\Phi(\xi)$ denotes the set of all linear layouts of $\xi$, $L(i,\varphi,\xi)$ represents the set of vertices mapped to positions $\leq i$ by the linear layout $\varphi$, and $R(i,\varphi,\xi)$ represents the set of vertices mapped to positions $>i$.
Definition 6.28 (Fuzzy Directed Bandwidth).
(cf.[502]) Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A layout of $\xi$ is a bijection $f:V\rightarrow\{1,2,\dots,|V|\}$. The fuzzy directed bandwidth of the layout $f$ is defined as:
$\max_{\{u,v\}\in E}\left(\mu(u,v)\cdot|f(u)-f(v)|\right).$
The fuzzy directed bandwidth of the fuzzy directed graph $\xi$, denoted by $\text{fdbw}(\xi)$, is the minimum fuzzy bandwidth over all possible layouts $f$ of $\xi$.
Definition 6.29 (Fuzzy Directed Tree-Length).
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed tree-decomposition of $\xi$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a directed tree with nodes $I$ and edges $F$,
  • $\{B_{t}\}_{t\in I}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $T$.
The fuzzy directed tree-length of a tree-decomposition $T(\xi)$ is defined as:
$\lambda(T)=\max_{t\in I}\max_{u,v\in B_{t}}d_{G^{*}}(u,v),$
where $d_{G^{*}}(u,v)$ denotes the distance between vertices $u$ and $v$ in the underlying crisp graph $G^{*}$. The fuzzy directed tree-length of $\xi$, denoted by $\text{fdtl}(\xi)$, is the minimum length over all possible fuzzy directed tree-decompositions of $\xi$.
Definition 6.30 (Fuzzy Directed Tree-Breadth).
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed tree-decomposition of $\xi$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a directed tree with nodes $I$ and edges $F$,
  • $\{B_{t}\}_{t\in I}$ is a collection of fuzzy subsets of $V$ (called bags) associated with the nodes of $T$.
The fuzzy directed tree-breadth of a tree-decomposition $T(\xi)$ is defined as the minimum integer $r$ such that for every $t\in I$, there exists a vertex $v_{t}\in V(G^{*})$ with $B_{t}\subseteq D_{r}(v_{t},G^{*})$, where $D_{r}(v_{t},G^{*})$ is the disk of radius $r$ centered at $v_{t}$ in the underlying crisp graph $G^{*}$. The fuzzy directed tree-breadth of $\xi$, denoted by $\text{fdtb}(\xi)$, is the minimum breadth over all possible fuzzy directed tree-decompositions of $\xi$.
6.1.8 Graph parameters for Fuzzy HyperGraph
In fuzzy graph theory, the concept of a fuzzy hypergraph is well-known. It extends the notion of a hypergraph by incorporating the concepts of fuzzy sets and fuzzy graphs. Numerous studies have explored its applications in real-world scenarios[915, 128, 35, 852, 826, 1193, 859, 854].
Definition 6.31.
A fuzzy hypergraph on a non-empty set $X$ is defined as a pair $H=(\mu,\rho)$, where:
  • $\mu=\{\mu_{1},\mu_{2},\dots,\mu_{r}\}$ is a collection of fuzzy subsets on $X$, where each $\mu_{i}:X\rightarrow[0,1]$ represents the membership degree of elements in $X$ to the fuzzy subset $\mu_{i}$. It is required that the union of the supports of all fuzzy subsets covers the entire set $X$, i.e., $\bigcup_{i=1}^{r}\text{supp}(\mu_{i})=X$.
  • $\rho$ is a fuzzy relation on the fuzzy subsets $\mu_{i}$, which assigns a membership degree $\rho(E_{i})$ to each hyperedge $E_{i}=\{x_{1},x_{2},\dots,x_{s}\}\subseteq X$. The membership degree $\rho(E_{i})$ is constrained by the condition:
    $\rho(E_{i})\leq\min\{\mu_{i}(x_{1}),\mu_{i}(x_{2}),\dots,\mu_{i}(x_{s})\},$
    for all elements $x_{1},x_{2},\dots,x_{s}\in X$ and for each fuzzy subset $\mu_{i}\in\mu$.
We are interested in exploring whether new insights can be gained by extending the graph width parameter known as hypertree-width to fuzzy hypergraphs. Given the vast range of applications for fuzzy hypergraphs, such an extension could potentially accelerate the development of real-world applications[35, 854, 1180, 128, 1192, 276]. For further information on Tree-width and Width Parameters related to other real-world uncertainties, please refer to Appendix E.
Definition 6.32 (Fuzzy Hypertree-width).
[502] Let $H=(\mu,\rho)$ be a fuzzy hypergraph defined on a non-empty set $X$, where:
  • $\mu=\{\mu_{1},\mu_{2},\dots,\mu_{r}\}$ is a collection of fuzzy subsets on $X$, with each $\mu_{i}:X\rightarrow[0,1]$ representing the membership degree of elements in $X$ to the fuzzy subset $\mu_{i}$. The union of the supports of all fuzzy subsets covers the entire set $X$, i.e., $\bigcup_{i=1}^{r}\text{supp}(\mu_{i})=X$.
  • $\rho$ is a fuzzy relation on the fuzzy subsets $\mu_{i}$, assigning a membership degree $\rho(E_{i})$ to each hyperedge $E_{i}=\{x_{1},x_{2},\dots,x_{s}\}\subseteq X$. The membership degree $\rho(E_{i})$ satisfies:
    $\rho(E_{i})\leq\min\{\mu_{i}(x_{1}),\mu_{i}(x_{2}),\dots,\mu_{i}(x_{s})\},$
    for all elements $x_{1},x_{2},\dots,x_{s}\in X$ and each fuzzy subset $\mu_{i}\in\mu$.
A fuzzy hypertree decomposition of a fuzzy hypergraph $H=(\mu,\rho)$ is a triple $D=(T,\chi,\lambda)$, where:
  • $T=(V_{T},E_{T})$ is a tree.
  • $\chi:V_{T}\rightarrow 2^{X}$ is a function that assigns to each tree node $t\in V_{T}$ a fuzzy subset $\chi(t)\subseteq X$, known as a fuzzy bag.
  • $\lambda:V_{T}\rightarrow 2^{\mu}$ is a function that assigns to each tree node $t\in V_{T}$ a collection of fuzzy hyperedges, known as a fuzzy edge cover of the fuzzy bag $\chi(t)$.
The decomposition must satisfy the following conditions:
  1. 1.
    Fuzzy Edge Cover Condition: For every fuzzy hyperedge $E_{i}$ in $H$, there exists a tree node $t\in V_{T}$ such that $\rho(E_{i})\leq\min_{x\in E_{i}}\chi(t)(x)$.
  2. 2.
    Fuzzy Connectedness Condition: For every vertex $x\in X$, the set of tree nodes $t\in V_{T}$ with $\chi(t)(x)>0$ induces a connected subtree of $T$.
  3. 3.
    Fuzzy Special Condition: For any two nodes $t,t^{\prime}\in V_{T}$ where $t^{\prime}$ is a descendant of $t$ in $T$, and for every fuzzy hyperedge $E_{i}\in\lambda(t)$, it holds that $(E_{i}\setminus\chi(t))\cap\chi(t^{\prime})=\emptyset$.
The fuzzy hypertree-width of a fuzzy hypergraph $H=(\mu,\rho)$ is defined as the minimum width over all possible fuzzy hypertree decompositions of $H$, where the width of a fuzzy hypertree decomposition $D=(T,\chi,\lambda)$ is given by the maximum value of the sum of the membership degrees $\sum_{x\in\chi(t)}\chi(t)(x)$ over all tree nodes $t\in V_{T}$.
6.1.9 Directed Hypertree Width
A directed hypergraph, which is an extension of the undirected hypergraph, has been studied for its practical applications[1155, 91, 1016, 1179, 92, 851]. This structure has proven useful in various real-world scenarios. The definition is provided below.
Definition 6.33.
[1155] Given a directed hypergraph $H=(V,E)$, where $V$ represents the set of vertices and $E$ denotes the set of hyper-arcs. Each hyper-arc $e\in E$ is defined as $e=(e_{\text{Tail}},e_{\text{Head}})$, where:
  • $e_{\text{Tail}}$ is the tail of the hyper-arc $e$, and $e_{\text{Head}}$ is the head of the hyper-arc $e$.
  • The vertices associated with the hyper-arc $e$ are denoted by $\mathcal{E}=e_{\text{Tail}}\cup e_{\text{Head}}$.
  • It is required that $e_{\text{Tail}}\neq\emptyset$, $e_{\text{Head}}\neq\emptyset$, and $e_{\text{Tail}}\cap e_{\text{Head}}=\emptyset$.
The directed hypergraph $H=(V,E)$ can be represented using two incidence matrices: $H_{\text{Tail}}$ and $H_{\text{Head}}$. These incidence matrices are defined as follows:
$h_{\text{Tail}}(v,e)=\begin{cases}1&\text{if }v\in e_{\text{Tail}},\\ 0&\text{otherwise}.\end{cases}$
$h_{\text{Head}}(v,e)=\begin{cases}1&\text{if }v\in e_{\text{Head}},\\ 0&\text{otherwise}.\end{cases}$
Here, $h_{\text{Tail}}$ and $h_{\text{Head}}$ are the incidence matrices corresponding to the tail and head of the hyper-arcs, respectively, with $v$ representing a vertex and $e$ representing a hyper-arc in $H$.
We are interested in exploring whether new insights can be gained by extending the graph width parameter known as hypertree-width to directed hypergraphs (cf.[1047]). We will also explore the concept of fuzzy directed hypertree-width. Given the active research on the real-world applications of fuzzy directed hypergraphs (ex.[33, 853, 922, 924, 923]), studying fuzzy directed hypertree-width is both meaningful and timely.
We will also explore tree-width and other width parameters on symmetric matroids[720]. Symmetric matroids are closely related to hypergraphs.
6.1.10 Weighted Width
A weighted graph assigns a numerical value, known as a weight, to each edge, and sometimes to each vertex. These graphs are widely used in applications such as network design[269], routing algorithms[672], optimization problems[1061], and in modeling real-world systems like transportation networks[623, 907, 622] and social networks[1010, 268, 1243, 944, 1249, 866, 1186], where the strengths or values of relationships vary. Notably, concepts such as Weighted Tree-width[724], Weighted Path-width[900], and Weighted Cut-width[170] have already been established.
Although still in the conceptual stage, the definitions for weighted graph are outlined below.
Definition 6.34.
The weighted bandwidth of a weighted graph $G=(V,E,w)$ is defined as follows:
Consider an ordering $\varphi$ of the vertices of $G$, and let $d_{G}(u,v)$ denote the distance between vertices $u$ and $v$ in $G$ under the ordering $\varphi$.
[…]
where $w(u)$ and $w(v)$ are the weights of vertices $u$ and $v$, respectively.
Definition 6.35.
Given a weighted graph $G=(V,E,w)$ where $w:V\to\mathbb{N}$ assigns a weight to each vertex, the weighted clique-width is defined as follows:
The weighted clique-width of $G$ is the minimum number of labels needed to construct the graph using the following operations:
  1. 1.
    Creation of a vertex: Create a new vertex with a specific label.
  2. 2.
    Disjoint union: Take the disjoint union of two labeled graphs.
  3. 3.
    Edge insertion: Connect all vertices with label $i$ to all vertices with label $j$ (for $i\neq j$).
  4. 4.
    Relabeling: Change the label of all vertices with label $i$ to label $j$.
[…]
Definition 6.36.
The weighted linear-clique-width of a weighted graph $G=(V,E,w)$ is defined similarly to weighted clique-width, with the following modification:
  • The operations to construct the graph must follow a linear order, meaning each step adds to a single sequence of vertices.
[…]
Definition 6.37.
Given a weighted graph $G=(V,E,w)$, where $w:V\to\mathbb{N}$ assigns a weight to each vertex, the weighted path-distance-width is defined based on a path distance decomposition of the graph.
A path distance decomposition of a graph $G$ is a tuple ($(\{X_{i}\mid i\in I\},T=(I,F),r)$, where:
  • $T=(I,F)$ is a path (a tree where each node has at most two neighbors).
  • $\{X_{i}\mid i\in I\}$ is a family of subsets of $V(G)$, called bags, such that:
    1. 1.
      $\bigcup_{i\in I}X_{i}=V(G)$ and for all $i\neq j$, $X_{i}\cap X_{j}=\emptyset$, meaning the bags partition the vertex set $V(G)$.
    2. 2.
      For each vertex $v\in V(G)$, if $v\in X_{i}$, then $d_{G}(X_{r},v)=d_{T}(r,i)$, where $X_{r}$ is the root set and $d_{G}$ and $d_{T}$ are the distances in $G$ and $T$, respectively.
    3. 3.
      For each edge $\{v,w\}\in E(G)$, there exist $i,j\in I$ such that $v\in X_{i}$, $w\in X_{j}$, and either $i=j$ or $\{i,j\}\in F$.
The width of a path distance decomposition ($(\{X_{i}\mid i\in I\},T,r)$ is defined as:
$\text{width}(T)=\max_{i\in I}\sum_{v\in X_{i}}w(v),$
where $w(v)$ is the weight of vertex $v$.
[…]
Definition 6.38.
Given a weighted graph $G=(V,E,w)$, where $w:E\to\mathbb{N}$ assigns a weight to each edge, the weighted carving-width is defined as follows:
A call routing tree (or a carving) of a graph $G$ is a tree $T$ with internal vertices of degree 3 whose leaves correspond to the vertices of $G$. The congestion of an edge $e$ in $T$ is defined as the sum of the weights of the edges in $G$ that have endpoints in different connected components of $T-e$.
[…]
We also plan to explore width parameters such as Weighted Directed Tree-width, Weighted Directed cut-width, Weighted Directed Tree-breadth, Weighted Directed Hypertree-width and Weighted Directed Path-width in the context of weighted digraphs. Given that weighted digraphs are practical and versatile graph classes with a wide range of applications[1001, 138, 918, 20, 591], defining these width parameters is both meaningful and valuable for various fields. Also we plan to explore about width-paremeters on Weighted hypergraph (cf.[1089, 888, 1084]).
6.1.11 Directed Linear-width
We will consider about Directed Linear-width[488]. Although still in the conceptual stage, the definitions are outlined below. This is the linear version of the Branch-width of a directed graph[231] and is also an extension of the Linear-width of an undirected graph[1141, 1147].
Definition 6.39.
[488] Let $D=(V,E)$ be a digraph. The function $f_{D}:2^{E(D)}\to\mathbb{N}$ is defined as:
$f_{D}(X)=\left|S_{V}(X)\cup S_{V}(E(D)\setminus X)\right|,$
where $S_{V}(X)$ and $S_{V}(E(D)\setminus X)$ are the directed vertex-separators corresponding to the edge subsets $X$ and $E(D)\setminus X$, respectively.
A directed branch-decomposition of $D$ is any layout of $f_{D}$ on $E(D)$ using a tree $T$ where the vertices correspond to subsets of $E(D)$. The width of this decomposition is the maximum value of $f_{D}(X)$ over all edges $e$ in $T$.
A directed linear-branch-decomposition is a specific type of directed branch-decomposition where the edges $E(D)$ are arranged in a linear ordering $(e_{1},e_{2},\dots,e_{m})$. The function $f_{D}(X)$ is then evaluated for each possible cut in this ordering, corresponding to the sets $\{e_{1},\dots,e_{i}\}$ and $\{e_{i+1},\dots,e_{m}\}$ for $11\leq i\leq m-$ 1.
The directed linear-branch width of $D$ is defined as the maximum value of $f_{D}(X)$ over all possible cuts in the linear ordering of the edges of $D$.
6.1.12 Directed Connectivity system
In the literature [1008], the concepts of Directed subset and Directed Submodular Function were defined. The definitions are provided below. We intend to explore whether these concepts can be used to characterize Directed Tree-width and Directed Width parameters in the future.
Definition 6.40.
[1008] Suppose that $S$ is a finite set, $s=|S|$. A directed subset $X$ of $S$ is a collection of elements of $S$, where the elements of $X$ are distinguished between forward and backward elements. We identify $X$ with an $s$-dimensional vector, its incidence vector. For an element $e\in S$, let
$X(e)=1,\quad\text{if }e\text{ is a forward element of }X;$
$X(e)=-1,\quad\text{if }e\text{ is a backward element of }X;$
$X(e)=0,\quad\text{if }e\text{ is not in }X.$
Define $-X$ by $-X(e)=-X(e)$, and $\text{abs}(X)$ by $\text{abs}(X)(e)=|X(e)|$.
Suppose that $X$ and $Y$ are two directed subsets of $S$. If all forward elements of $X$ are forward elements of $Y$ and all backward elements of $X$ are backward elements of $Y$, then we say that $X$ is contained in $Y$ and denote this fact as $X\subseteq Y$. An element $e$ is a forward element of $X\cap Y$ if and only if it is a forward element of both $X$ and $Y$. Similarly, an element $e$ is a backward element of $X\cap Y$ if and only if it is a backward element of both $X$ and $Y$. An element $e$ is a forward element of $X\cup Y$ if and only if it is a forward element of one of $X$ and $Y$ and not a backward element of another. Similarly, an element $e$ is a backward element of $X\cup Y$ if and only if it is a backward element of one of $X$ and $Y$ and not a forward element of another. Equivalently,
$(X\cup Y)(e)=\begin{cases}1,&\text{if }X(e)+Y(e)>0,\\ -1,&\text{if }X(e)+Y(e)<0,\\ 0,&\text{if }X(e)+Y(e)=0.\end{cases}$
$(X\cap Y)(e)=\begin{cases}1,&\text{if }X(e)=Y(e)=1,\\ -1,&\text{if }X(e)=Y(e)=-1,\\ 0,&\text{otherwise}.\end{cases}$
Definition 6.41.
[1008] Suppose that $\mathcal{S}$ is an interesting family of directed subsets of a finite set $S$. A set-function $f:\mathcal{S}\to\mathbb{R}$ is called a directed submodular function on $\mathcal{S}$ if for all interesting directed subsets $X$ and $Y$ in $\mathcal{S}$,
[…]
We are considering whether the Directed Connectivity System, based on the above Directed Subset and Directed Submodular Function, can be used for any characterizations. Additionally, we are also exploring the possibility of further extending this framework using bisubmodular functions [676, 74] or other related concepts.
Furthermore, we are interested in exploring whether Tree-width, Path-width, Branch-width, Rank-width, Linear-width, and other related parameters can be defined on a ditroid (the directed set version of a matroid). The definition of a ditroid is provided below [1008]. Also we are interested in width-parameters on fuzzy matroids [1082, 832, 578, 1083] or connectoids[206, 207].
Definition 6.42.
[1008] Let $X$ a finite set. Two directed subsets $A$ and $B$ of $X$ are called non-cancelling if there is no $e$ in $X$ such that $A(e)=-B(e)\neq 0$. Otherwise, $A$ and $B$ are called cancelling.
Suppose that $\mathcal{M}$ is a family of directed subsets of $X$. We call $D=(X,\mathcal{M})$ a ditroid if the following conditions are satisfied:
  1. 1.
    $\emptyset\in\mathcal{M}$,
  2. 2.
    If $A\in\mathcal{M}$ and $B\subseteq A$, then $B\in\mathcal{M}$,
  3. 3.
    If $A$ and $B\in\mathcal{M}$ are non-cancelling with $|A|=|B|+1$, then there exists $e$ in $X$ and $C$ in $\mathcal{M}$ such that $B(e)=0$, $C(e)=A(e)\neq 0$, and $C(e^{\prime})=B(e^{\prime})$ for all $e^{\prime}\neq e$.
A directed subset in $\mathcal{M}$ is called an independent directed subset of $D$. A maximal independent directed subset is called a base of $D$. The rank function of $D$ is a set-function $r:D(X)\rightarrow\mathbb{R}$ defined by
$r(A)=\max\{|B|:B\subseteq A,B\in\mathcal{M}\},\quad\text{for }A\in D(X).$
In the future, we aim to explore the results when extending to variants of directed graphs. Similar to undirected graphs, directed graphs have various proposed classes such as directed multigraphs[310], complete directed graphs[63], quasi-transitive digraphs [101], semicomplete multipartite digraphs[1238], and tournaments[103]. We plan to investigate how these results might apply when extended to these classes.
And the graph class known as Bidirected Graphs [200, 1227] has also been a subject of study. Since it generalizes the concept of Directed Graphs, we anticipate that research in this area will continue to grow. In fact, Bidirected Tree-width has already been defined in the literature [1209]. Therefore, defining parameters such as Bidirected Path-width (the path version of Bidirected Tree-width), Bidirected Cut-width, Bidirected Branch-width, Bidirected Linear-width, and Bidirected Connectivity system and exploring their relationships, would be highly meaningful. Additionally, extending these concepts to bidirected hypergraphs is another avenue worth exploring (cf.[1115]).
6.1.13 DAG-PathWidth and Related Parameter
A Directed Acyclic Graph (DAG) is a graph with directed edges and no cycles, commonly used in blockchains[655, 746, 742, 647]. While DAG-width is the most well-known width parameter related to DAGs[130, 102, 131, 957], the concept of DAG-pathwidth has recently been proposed[746, 742]. In light of this, we aim to explore obstructions to DAG-pathwidth and define various parameters on DAGs. Although still in the conceptual stage, the anticipated definitions are provided below.
Definition 6.43.
The DAG-cut-width of a directed acyclic graph (DAG) $D=(V,E)$ is defined as:
[…]
where:
  • $\Phi(D)$ is the set of all linear layouts (topological orderings) of $D$.
  • $L(i,\varphi,D)$ represents the set of vertices mapped to positions $\leq i$ by the linear layout $\varphi$.
  • $R(i,\varphi,D)$ represents the set of vertices mapped to positions $>i$ by the linear layout $\varphi$.
The DAG-cut-width measures the maximum number of edges that must be "cut" when partitioning the vertices in a given topological order. It quantifies the complexity of the DAG in terms of edge connectivity across different partitions of the vertex set.
Definition 6.44.
The DAG-band-width of a directed acyclic graph $D=(V,E)$ is defined as:
[…]
where:
  • $\mathcal{F}(D)$ is the set of all bijective mappings $f:V\rightarrow\{1,2,\dots,|V|\}$, representing different vertex orderings.
  • $f(u)$ and $f(v)$ are the positions of vertices $u$ and $v$ in the layout defined by $f$.
We also consider tree-width, path-width, cut-width, and other parameters on a directed acyclic hypergraph (DAH). A directed acyclic hypergraph is a generalization of a directed acyclic graph, where each hyperedge can have multiple tails and heads [994, 1070]. Since the concept of directed acyclic hypergraphs is applied in areas such as graphical modeling, extending these width parameters to DAHs is meaningful.
6.2 Consideration about Property of an Ultrafilter
We will explore additional properties of an ultrafilter on a connectivity system, including game-theoretical interpretations of ultrafilters and ultraproducts within such systems.
6.2.1 Simple Game on a Connectivity System
A simple voting game is a mathematical model of voting systems in which a coalition of agents wins if it belongs to a specified set of "winning" coalitions, subject to conditions like monotonicity (cf. [479, 913, 312, 1137]). These simple games model voting power and offer a generalized interpretation of the concept of "majority." As discussed in previous sections regarding the relationship between ultrafilters and simple voting systems, this game is well-known to be closely connected to ultrafilters.
A simple game on a connectivity system can be modeled similarly to the simple game in the context of voting systems, but within the framework of connectivity systems, incorporating the constraint $f(A)\leq k$. In the future, we will consider a simple game on a connectivity system.
Game 6.45.
Simple Game on a Connectivity System
Let $X$ be a finite set, $f$ be a symmetric submodular function, and $W\subseteq 2^{X}$ be any collection that satisfies the following monotonicity condition (M1). The triple $(X,W,f)$ is called a simple game on a connectivity system if for every $A\in W$, $f(A)\leq k$.
Basic Definitions of a Simple Game:
  1. 1.
    Connectivity System:
    • A connectivity system is defined as a pair $(X,f)$, where $X$ is a finite set and $f:2^{X}\to\mathbb{N}$ is a symmetric submodular function.
  2. 2.
    Winning Coalition:
    • Let $X$ be a set of agents and $W\subseteq 2^{X}$ be a collection of subsets that represent the majorities or winning coalitions of $X$.
Conditions of a Simple Game:
  1. 1.
    Monotonicity Condition (M1) on a Connectivity System:
    • The collection $W$ is closed under supersets: If $A\in W$ and $A\subseteq B\subseteq X$ with $f(B)\leq k$, then $B\in W$.
  2. 2.
    Proper Simple Game on a Connectivity System:
    • A simple game is called proper if it satisfies: $A\in W$ implies $X\setminus A\notin W$, provided $f(A)\leq k$.
  3. 3.
    Strong Simple Game on a Connectivity System:
    • A simple game is called strong if it satisfies: $A\notin W$ implies $X\setminus A\in W$, provided $f(A)\leq k$.
Conjecture 6.46.
An ultrafilter on a connectivity system is a simple game on a connectivity system.
6.2.2 The Axiom of Choice in the Context of Connectivity Systems
The Axiom of Choice states that for any set of nonempty sets, there exists a function selecting one element from each set, ensuring the product is nonempty. The Axiom of Choice is widely used in real-world applications, such as in optimization, decision theory, and economics, where selecting elements from sets is essential for solving complex problems [699, 626, 698, 481, 658, 660, 228].
In the future, we will consider the Axiom of Choice within the context of connectivity systems. To consider the Axiom of Choice within a connectivity system, we need to translate the concepts into the framework of connectivity systems:
  • Set of Nonempty Subsets:
    Let $\{S_{\alpha}:\alpha\in A\}$ be a collection of nonempty subsets of $X$ within the connectivity system $(X,f)$. Each $S_{\alpha}$ is a nonempty subset of $X$, and $f(S_{\alpha})\leq k$ for some $k$.
  • Product of Sets:
    The product $\prod_{\alpha\in A}S_{\alpha}$ represents the Cartesian product of the sets $S_{\alpha}$. In the context of a connectivity system, this product can be interpreted as a selection of one element from each $S_{\alpha}$.
  • Nonemptiness of the Product:
    The Axiom of Choice asserts that there exists a function $g:A\to X$ such that $g(\alpha)\in S_{\alpha}$ for all $\alpha\in A$. This means there is a way to select one element from each subset $S_{\alpha}$ while maintaining the constraints imposed by the connectivity system.
Additionally, we will consider selective (Ramsey) ultrafilters and $P$-ultrafilters on a connectivity system [1157, 901, 984, 475].
6.2.3 Ultraproducts and Ultrapowers on Connectivity Systems
Ultraproducts combine structures using ultrafilters, allowing the creation of a new structure that retains properties from the original structures. This is often used in model theory to study logical consistency [784, 656, 652, 1068, 407, 697]. We will extend the construction of ultraproducts to the context of a symmetric submodular function $f$.
Although it is still in the conceptual stage, the ultraproduct on a connectivity system can be described as follows. Let $\{M_{i}\}_{i\in I}$ be a collection of sets indexed by $I$, and let $U$ be an ultrafilter on a connectivity system on $I$. We consider the direct product $\prod_{i\in I}M_{i}$, which consists of all possible combinations of elements from each $M_{i}$. Instead of simply considering the direct product, we now define a set of subsets $A\subseteq\prod_{i\in I}M_{i}$ such that $f(A)\leq k$. This ensures that the subsets considered in the product adhere to the submodular constraints.
Using the ultrafilter $U$ on the connectivity system, we select subsets of $\prod_{i\in I}M_{i}$ that are included in the ultrafilter. This step filters the product to only include elements in subsets dictated by $U$.
Finally, we take the direct limit:
$\lim_{J\in U}\left\{A\subseteq\prod_{i\in J}M_{i}\mid f(A)\leq k\right\}.$
This limit considers the stable elements in the context of the ultrafilter $U$, yielding the ultraproduct
$\left(\prod_{i\in J}M_{i}\right)/U.$
6.2.4 Algorithm of Constructing an Ultrafilter on a Connectivity System
We will consider an algorithm for constructing an ultrafilter on a connectivity system. The algorithm is as follows.
Algorithm 6.47.
  • Input: A finite set $X$ with $\lvert X\rvert=n$. A symmetric submodular function $f:2^{X}\rightarrow\mathbb{N}$.
  • Output: An ultrafilter $U$ of order $k+1$ on the connectivity system $(X,f)$.
Algorithm:
  1. 1.
    Initialization: Start with an empty set collection $U=\emptyset$. Let $L$ be the list of all subsets $A\subseteq X$ such that $f(A)\leq k$.
  2. 2.
    Step 1: Generate Candidate Sets:
    • Enumerate all subsets $A\subseteq X$ such that $f(A)\leq k$. Store these subsets in $L$.
    • Note: The condition $f(A)\leq k$ ensures that the function values are bounded, adhering to the order $k+1$.
  3. 3.
    Step 2: Construct the Filter:
    • Select the first non-empty subset $A_{0}$ from $L$ and include it in $U$, i.e., $U=\{A_{0}\}$.
    • For each subsequent subset $A_{i}\in L$, do the following:
      • Intersection Check: Check whether for all sets $B\in U$, the intersection $A_{i}\cap B$ satisfies $f(A_{i}\cap B)\leq k$.
      • If the condition is met, include $A_{i}$ in $U$; otherwise, discard $A_{i}$.
  4. 4.
    Step 3: Extend to an Ultrafilter:
    • For each subset $A\subseteq X$ not yet in $U$ with $f(A)\leq k$, perform the following:
      • If adding $A$ maintains the properties of $U$, include $A$ in $U$.
      • If adding $A$ violates the properties of $U$, include $X\setminus A$ instead, ensuring the maximality of $U$.
  5. 5.
    Termination:
    • Continue the above steps until every subset $A\subseteq X$ with $f(A)\leq k$ is either in $U$ or its complement $X\setminus A$ is in $U$.
    • The process ends when $L$ is exhausted, ensuring that $U$ is maximal and satisfies the ultrafilter conditions.
    • Return the collection $U$, which is the desired ultrafilter of order $k+1$.
Lemma 6.48.
Above algorithm is correct.
Proof 6.49.
  1. 1.
    The construction of $U$ begins with a non-empty set $A_{0}$, ensuring that $U$ is non-empty. Additionally, $f(A_{0})\leq k$ guarantees that the initial set satisfies the $k$-efficiency condition.
  2. 2.
    Every subset $A$ included in $U$ is selected based on the condition $f(A)\leq k$. Therefore, $U$ inherently satisfies the $k$-efficiency condition.
  3. 3.
    Closed under Finite Intersections: For each new set $A_{i}$ considered, it is included in $U$ only if $A_{i}\cap B$ for all $B\in U$ also satisfies $f(A_{i}\cap B)\leq k$. This ensures closure under finite intersections.
  4. 4.
    Maximality: The final step in the algorithm ensures that for every subset $A\subseteq X$, either $A$ or $X\setminus A$ is included in $U$. This guarantees the ultrafilter’s maximality, as no additional subsets can be added without violating the ultrafilter conditions. Therefore, the lemma is proved.
Lemma 6.50.
The total time complexity of above algorithm is $O(2^{2n})$.
Proof 6.51.
The algorithm must consider every possible subset of the finite set $X$ to construct the ultrafilter $U$. Since $X$ contains $n$ elements, the power set $2^{X}$ consists of $2^{n}$ subsets. Therefore, generating all subsets of $X$ requires $O(2^{n})$ operations.
For each subset $A_{i}$ in $2^{X}$, the algorithm checks its intersection with every other subset $B$ already included in $U$ to ensure that the submodular condition $f(A_{i}\cap B)\leq k$ is satisfied.
In the worst-case scenario, this operation could involve checking intersections with up to $2^{n}$ subsets, which adds another factor of $O(2^{n})$ to the time complexity. Therefore, for each subset $A_{i}$, the total time complexity for evaluating all intersection properties is $O(2^{n})\times O(2^{n})=O(2^{2n})$.
Considering both the generation of subsets and the intersection evaluation, the overall time complexity of the algorithm is $O(2^{2n})$. Therefore, the lemma is proved.
Lemma 6.52.
The space complexity of the algorithm is $O(2^{n})$.
Proof 6.53.
The filter $U$ and the list $L$ both store collections of subsets of $X$. Since there are $2^{n}$ possible subsets, the space required to store either $U$ or $L$ is $O(2^{n})$. The total space complexity is thus $O(2^{n})$. Therefore, the lemma is proved.
We believe there is still room for improvement in the aforementioned algorithm. Additionally, we plan to investigate efficient algorithms tailored specifically to certain graphs, such as planar graphs, in the future.
6.2.5 Ultrafilter Width under Small Set Connectivity Expansion Hypothesis
We investigate ultrafilters in the context of the Small Set Connectivity Expansion Hypothesis problem within Parameterized Complexity.
Parameterized Complexity is a framework in computational complexity theory (cf.[975, 81, 1091]) that offers a more nuanced analysis of algorithmic problems than traditional complexity measures, which focus solely on input size. In this approach, each problem instance includes an additional value called a parameter—typically a non-negative integer—that captures some aspect of the input’s structure or complexity. This allows us to understand how computational difficulty depends not just on the input size but also on this specific parameter. Due to its significance, Parameterized Complexity has been extensively studied  [376, 374, 375, 317, 367].
Formally, a parameterized problem $P$ is a subset of $\Sigma^{*}\times\mathbb{N}_{0}$, where $\Sigma$ is a finite alphabet, $\Sigma^{*}$ denotes all finite strings over $\Sigma$, and $\mathbb{N}_{0}$ is the set of non-negative integers. Each instance is a pair $(x,k)$, where $x$ is the main input and $k$ is the parameter.
A problem is called fixed-parameter tractable (FPT) (cf.[1055, 373, 372]) if there exists an algorithm that decides whether $(x,k)\in P$ in time $f(k)\cdot|x|^{O(1)}$, where:
  • $f(k)$ is a computable function depending only on $k$,
  • $|x|$ is the size of the input $x$,
  • $O(1)$ indicates that the exponent is a constant independent of $k$.
In other words, for small values of $k$, the problem can be solved efficiently even if $|x|$ is large, because the computational complexity is primarily influenced by the parameter $k$.
The class XP consists of parameterized problems solvable in time $|x|^{f(k)}$, where $f(k)$ depends on $k$. For each fixed $k$, the problem is solvable in polynomial time, but the degree of the polynomial depends on $k$. Each parameterized problem can be viewed as a collection of $k$-slices, denoted $P_{k}=\{x\mid(x,k)\in P\}$, where the parameter $k$ is fixed. For problems in FPT, each $P_{k}$ can be solved in polynomial time with a degree independent of $k$. In XP, however, the degree may increase with $k$.
A problem is para-NP-hard (cf.[668]) or para-NP-complete (cf.[563]) if there exists at least one fixed $k$ such that $P_{k}$ is NP-hard or NP-complete. This means that even when $k$ is fixed, the problem remains as hard as the most difficult problems in NP. To further classify parameterized problems, the W-hierarchy (cf.[274]) is introduced, consisting of complexity classes between FPT and XP:
[…]
These inclusions are believed to be strict. Problems that are W[1]-hard are considered unlikely to be fixed-parameter tractable; that is, they probably cannot be solved in time $f(k)\cdot|x|^{O(1)}$.
In computational complexity theory, a computational hardness assumption posits that a specific problem cannot be solved efficiently. One such pivotal assumption is the Small Set Expansion Hypothesis (SSEH), introduced by Raghavendra and Steurer [1012], which plays a critical role in understanding the hardness of approximation algorithms. The SSEH asserts that it is NP-hard to distinguish between small sets of vertices in a graph with low edge expansion (few edges leaving the set) and those with high edge expansion (many edges leaving the set). This hypothesis has become fundamental in analyzing the complexity of various graph algorithms. Since its inception, the SSEH has led to numerous inapproximability results, including those documented in [694, 1217, 93, 1232, 241].
We will consider the concept of Ultrafilter Width. The Ultrafilter Width of a graph $G$ is the maximum order of an ultrafilter in $G$. Similarly, the Ultrafilter Width of a connectivity system is the maximum order of an ultrafilter in the connectivity system.
Conjecture 6.54.
Under the Small Set Expansion Hypothesis (SSEH), it is NP-hard to approximate the Ultrafilter width of a graph within a constant factor in polynomial time.
We will also introduce a new notion called Small Set Connectivity Expansion (SSCE). The definition is as follows. The transformation of the SSEH to apply to connectivity systems rather than simple graphs extends its applicability to more complex and abstract structures, allowing for the examination of expansion properties beyond traditional vertex-edge models.
Definition 6.55.
Let $(X,f)$ be a connectivity system, with $X$ a finite set and $f$ a symmetric submodular function as defined. Consider the analogy of edge expansion in a graph to the concept of a “boundary” in a connectivity system.
  • Connectivity Expansion: For a subset $S\subseteq X$, define the connectivity expansion, $\Psi_{f}(S)$, as:
    $\Psi_{f}(S)=\min\{\text{vol}_{f}(S),\text{vol}_{f}(X\setminus S)\}\cdot f(S)$
    where $f(S)$ quantifies the “connectivity boundary” of $S$, and $\text{vol}_{f}(S)=\sum_{x\in S}f(\{x\})$ represents the volume of $S$ in terms of connectivity, analogous to the sum of degrees in graph theory.
  • For any given small constant $\epsilon>0$, it is computationally challenging to find a non-trivial subset $S\subseteq X$ of the connectivity system such that $\Psi_{f}(S)<\epsilon$.
Question 6.56.
Is it NP-hard to approximate the Ultrafilter width of a graph within a constant factor in polynomial time, under the Small Set Connectivity Expansion Hypothesis (SSCEH)?
6.2.6 Various Ultrafilters on a Connectivity System
Ultrafilters are of interest due to their practical applications and mathematical properties, prompting the proposal of various similar concepts across different fields. Going forward, we aim to define concepts such as Partition-filter [629], uniform-ultrafilters [938], weak normal-ultrafilter [725], gw-ultrafilters [795], subuniform ultrafilters [936], Fuzzy-ultrafilter [1116], Regular-ultrafilters [836], Good-ultrafilters [297], Semi-ultrafilters [865], Dfin-ultrafilter [800], $I$-Ultrafilter [456], Open-Ultrafilter [245], Closed-Ultrafilter [121], Linear-ultrafilters [125, 630], OK-ultrafilters [371], Club-Ultrafilter [986], neutrosophic-ultrafilters [1048, 1241], complete-ultrafilters [1000], Ramsey-ultrafilters [985], and Selective-ultrafilters [985] within the frameworks of Undirected Graph, Directed Graph, Infinite Graph, Connectivity Systems and Infinite Connectivity Systems. Additionally, we plan to explore their relationships with graph width parameters.
6.3 Other research Themes for graph width parameters
In the field of graph width parameters, research has been conducted on the following topics. Moving forward, I hope to contribute to the advancement of graph theory by exploring these aspects.
  • Analysis of hierarchies and graph problems within restricted and extended states of specific graphs (ex.[199, 236, 612]).
    • Hierarchies and graph problems restricted to graph width parameters [18, 106].
    • Example of specific graphs: Line Graph[641, 643], planar graph[335, 596, 1036, 68], AT-Free Graph[581, 976, 201], chordal graph[619, 361, 210], and sequence Graph[471, 616], etc…
  • Algorithms under constraints imposed by graph width parameters [1198, 748, 817, 175].
  • Structural properties of generalized or restricted graph width parameters.
  • Interpretation of graph width parameters using game theory[792, 1108, 603, 529].
  • The relationship between forbidden minors and graph width parameters[614, 1130, 767, 1129].
7 Acknowledgments
I humbly express my sincere gratitude to all those who have extended their invaluable support, enabling me to successfully accomplish this paper.
8 Conflict of Interest
The author declares no conflicts of interest.
Note 8.1.
Please note that preprints and early-stage research may not have undergone peer review. Additionally, as I am an independent researcher, please understand. Sorry.
Appendix A Various Width Parameters
Numerous graph width parameters have been identified and studied extensively. As illustrated, many of these parameters have significant applications in real-world scenarios and other fields of research, which suggests that the study of graph width parameters will continue to grow in importance. Examining the relationships among these parameters, such as inequalities, upper bounds, and lower bounds, is a well-established research topic (cf. Appendix C). Additionally, we aim to explore the potential connections between these width parameters and ultrafilters in future studies.
Note that unless otherwise specified, we consider any graph when comparing graph parameters. We say that a parameter $p$ upper bounds a parameter $q$ if there exists a non-decreasing function $f$ such that $f(p(G))\geq q(G)$ for all graphs $G$. Conversely, if $p$ does not upper bound $q$, then $q$ is considered unbounded with respect to $p$. Additionally, if value $a$ upper bounds value $b$ and $b$ upper bounds value $c$, then value $a$ upper bounds $c$.
A.1 Tree-width
Tree-width is a measure of how tree-like a graph is, defined by its minimum tree decomposition width [1038, 164, 1033, 1034, 1071, 665, 1037]. Many algorithms that are NP-hard for general graphs, become easier when the treewidth is bounded by a constant. The definitions of Tree-width on a graph are provided below.
Definition A.1.
A tree-decomposition of a graph $G=(V,E)$ is a pair ($(\{X_{i}\mid i\in I\},T=(I,F))$, where:
  • $\{X_{i}\mid i\in I\}$ is a collection of subsets of $V$, known as bags.
  • $T=(I,F)$ is a tree whose nodes correspond to the bags $X_{i}$.
The tree-decomposition must satisfy the following conditions:
  1. 1.
    $\bigcup_{i\in I}X_{i}=V$, meaning every vertex of $G$ appears in at least one bag.
  2. 2.
    For every edge $uv\in E$, there is a bag $X_{i}$ such that $u,v\in X_{i}$.
  3. 3.
    For all $i,j,k\in I$, if $j$ lies on the path from $i$ to $k$ in $T$, then $X_{i}\cap X_{k}\subseteq X_{j}$. Equivalently, for every vertex $v\in V$, the set of bags containing $v$ forms a connected subtree of $T$.
Definition A.2.
The width of a tree-decomposition ($(\{X_{i}\mid i\in I\},T=(I,F))$ is defined as:
$\text{width}(T)=\max_{i\in I}|X_{i}|-1.$
The tree-width of a graph $G$, denoted by $\text{tw}(G)$, is the minimum width over all possible tree-decompositions of $G$. A graph has tree-width 1 if and only if it is a tree.
Example A.3.
(cf.[516]) Consider a complete graph $K_{n}$ with $n$ vertices. Every pair of vertices is connected by an edge, so $E(K_{n})=\{(v_{i},v_{j})\mid 1\leq i<j\leq n\}$.
  • Nodes and Bags:
    • A tree-decomposition of $K_{n}$ can consist of a single node tree $T$ with one bag containing all vertices:
      $W=\{v_{1},v_{2},\dots,v_{n}\}.$
  • Tree $T$:
    • $T$ has only one node associated with the bag $W=V(K_{n})$.
  • Properties Verification:
    • Vertex Containment (T1):
      $\bigcup_{t\in V(T)}W=V(K_{n}).$
    • Edge Containment (T2):
        For each edge $(v_{i},v_{j})\in E(K_{n})$, both $v_{i}$ and $v_{j}$ are in the bag $W$.
    • Vertex Connectivity (T3):
        There is only one bag, so this condition is trivially satisfied.
  • Width Calculation:
    • The size of the bag $W$ is $|W|=n$.
    • The width of this tree-decomposition is $n-1$.
    • Since no tree-decomposition can have a smaller width, the Tree-width of $K_{n}$ is:
      $\text{TreeWidth}(K_{n})=n-1.$
Examples of obstructions for tree-decomposition are Tangle [1038], Ultrafilter [495], and Bramble [267, 174, 172, 821]. For more detailed information, please refer to lecture notes or surveys [145, 775, 165, 219].
A.1.1 Hierarchy for treewidth
We consider about Hierarchy for treewidth.
Theorem A.4.
The relationships between tree-width and other parameters are as follows:
  • If $G$ has tree-width $k$, then $G$ is a subgraph of a chordal graph with maximum clique size $k+1$ [165].
  • Let $G$ be a connected graph with treewidth $tw(G)$ and gonality $dgon(G)$. Then $tw(G)\leq dgon(G)$ [1170].  [^15] The gonality of a graph is the minimum number of chips needed to reach any vertex using chip-firing moves without causing debt at other vertices[291].
  • If $G$ has tree-width $k$, the dimension of $G$ is at most $k$ [165].  [^16] In graph theory, the dimension of a graph refers to the minimum size of cliques formed during vertex elimination, minimized over all possible perfect elimination orderings.
  • If $G$ has tree-width $k$, $G$ is $k$-decomposable [165].
  • If $G$ has tree-width $k-1$, $G$ has haven $k$ [1072].  [^17] A haven in a graph is a function that maps each subset of vertices of size at most $k$ to a connected component, ensuring that any two components ”touch” either by intersection or adjacency.
  • If $G$ has tree-width $k$, $G$ has no screen of thickness at least $k+2$ [165].  [^18] Screen in a graph is a collection of connected subsets of vertices where each pair of subsets either intersects or contains adjacent vertices, ensuring they ”touch.”
  • If $G$ has tree-width $k$, $k+1$ cops can search $G$ in the Seymour-Thomas search game [165].
  • If $G$ has tree-width $k$, $k+1$ cops can monotonely search $G$ in the Seymour-Thomas search game [165].
  • If $G$ has tree-width $k$, the number of searchers needed to search $G$ in the fugitive search game with an inert fugitive is at most $k+1$ [165].
  • If $G$ has tree-width $k$, the number of searchers needed to monotonically search $G$ in the fugitive search game with an inert fugitive is at most $k+1$ [165].
[…]
  • $G$ is a partial strict $k$-tree [387].
  • $qn(G)\leq pw(G)\leq(tw(G)+1)\log n$ (queue-number $qn(G)$, path-width $pw$) [560].
  • The stack-number $sn(G)\leq tw(G)+1$ (stack-number is also known as book thickness or page-number) [560].
  • The track number track $\text{track number}(G)\leq\text{path-width}(G)+1\leq 1+(\text{tree-width}(G)+1)\log n$ [387].
  • The tree-width of a graph is a lower bound on the positive semidefinite zero forcing number $Z^{+}(G)$ [408].  [^19] The zero forcing number is the minimum number of initially colored vertices required to color the entire graph by repeatedly applying the ”color-change rule(cf.[721, 404]).
  • $tw(G)\leq la(G)\leq tw(G)+1$ ($la(G)$: Largeur d’arborescence. Also called Two-sided tree-width) [1169, 319].
  • The spanheight of a graph upper-bounds the treewidth of a graph [1041].
  • Ultrafilter on a graph is an obstacle of tree-width [495].
  • Let $G$ be any non-complete graph. Then toughness $\leq tw(G)/2$ [568].
  • $\text{treewidth}(G)\leq\text{pathwidth}(G)\leq(\text{treewidth}(G)+1)\log(n)$.
  • $\text{bw}(G)\leq\text{tw}(G)+1\leq\frac{3}{2}\cdot\text{bw}(G)$ for a graph $G$ with $\text{bw}(G)\geq 2$ (bw: branch-width) [164].
  • $\text{tw}(G)\leq 2\cdot\text{tcw}(G)^{2}+3\cdot\text{tcw}(G)$ (tcw: Tree-cut-width) [553].
  • $\text{boow}(G)\leq\text{tw}(G)+1$ (boow: Boolean-width) [553].
  • If a graph $G$ has bounded tree-width, then it also has bounded sparse twin-width, bounded clique-width, and bounded twin-width [9].
  • If a graph $G$ has bounded Tree-distance-width, then it also has bounded Tree-width [1233].
  • If a graph $G$ has bounded Tree-width, then it also has bounded tree-independence number [116].
  • If a graph $G$ has bounded tree-width, then it also has bounded Maximum clique number, and chromatic number [1251, 1058].
  • If a graph $G$ has bounded tree-width, then it also has bounded Bramble Number (bn), Tangle Number, Lexicographic Tree Product Number (LTP) [^20] The lexicographic tree product number of a graph $G$, denoted by $ltp(G)$, is the smallest integer $k$ such that $G$ is a minor of $T\cdot K_{k}$, where $T$ is a tree and $K_{k}$ is the complete graph on $k$ vertices[642]., Linkedness Number (link)[1030], Cartesian Tree Product Number (ctp)[1168, 320]  [^21] The Cartesian Tree Product Number $\text{ctp}(G)$ is the smallest integer $k$ such that a graph $G$ is a minor of $T^{(k)}$, where $T^{(k)}$ is the Cartesian product of a tree $T$ and $K_{k}$[642]., and Well-Linked Number (wl)[1030]. The converse is also true [642].
[…]
  • $\text{carw}(G)\cdot\Delta(G)\leq\text{tw}(G)\leq 2\times\text{carw}(G)$ (where $\Delta(G)$ is the maximum degree of $G$ and $\text{carw}(G)$ is carving width) [1144, 141, 943].
  • Minimal upper bounds for tree-width is distance to outerplanar [164].
  • Tree-width is a minimal upper bound for book thickness (page number [127]) and acyclic chromatic number [605, 397, 392, 393].  [^22] The acyclic chromatic number of a graph is the minimum number of colors needed to color the vertices such that no two adjacent vertices share the same color and no cycle is bichromatic(cf.[237, 1187]).
  • If a graph $G$ has bounded Vertex-Cover Number, then it also has bounded tree-width. [^23] The Vertex-Cover Number is the smallest number of vertices needed to cover all edges in a graph, meaning each edge has at least one endpoint in the vertex cover(cf.[1189, 992]).
  • If a graph $G$ has bounded Tree-width, then it also has bounded Degeneracy, Average Degree, Minimum Degree, Domatic Number [441], and Distance to Disconnected [1154]. [^24] Degeneracy is the smallest number $d$ such that every subgraph has a vertex of degree at most $d$, indicating a graph’s sparsity(cf.[951, 437]).
  • For any graph $G$, $\text{box}(G)\leq\text{tw}(G)+2$, where $\text{box}(G)$ and $\text{tw}(G)$ denote the boxicity and treewidth of $G$, respectively [262].  [^25] The boxicity of a graph is the minimum number of dimensions needed to represent the graph as an intersection graph of axis-aligned boxes in Euclidean space (cf. [436, 261]).
  • Every graph of tree-width $k$ has a balanced vertex cut of size at most $k+1$ [1040].  [^26] A balanced vertex cut of size $k$ is a set of $k$ vertices whose removal divides a graph into roughly equal-sized components [1040].
  • For any graph $G$, $\text{Cubicity}(G)\leq(\text{treewidth}(G)+2)\left\lfloor\log_{2}n\right\rfloor$ [260].
  • Every graph of tree width $k$ is $k$-degenerate [197].
  • Every graph of tree-width $k$ admits a 2-balanced binary tree-decomposition of width at most $23k+$ 2 [160].
  • Let $G$ be a graph. Then $\text{tw}(G)\geq$ minimum degree [569].
  • Let $G$ be a graph on $n$ vertices. Then $\text{tw}(G)\leq n-$ maximum independent set [569].  [^27] The maximum independent set is the largest set of vertices in a graph where no two vertices are adjacent, meaning no edges exist between any pair of vertices in the set(cf.[1221, 1080]).
  • Let $G$ be a graph and $m\geq k\geq 1$ integers. If $\text{tw}(G)\geq k+m-1$, then $G$ has a $k$-mesh of order $m$ [348].
  • If $G$ is a $k$-outerplanar graph, then $\text{tw}(G)\leq 3k-1$ [159].
  • Let $G$ be a graph. Then surrounding cop number is less than or equal to tree-width plus one[233, 1067].
  • For every connected graph, periodic cop number is less than or equal to tree-width plus one[1088].
  • For every connected graph, 00-domination cop number equals to tree-width plus one[467].
  • For every graph, $\infty$-colouring number equals to tree-width plus one[1167].  [^28] Also, for every graph, weak $\infty$-colouring number equals to tree-depth[1167].
[…]
  • Graphs that have a treewidth of exactly $k$ are referred to as $k$-trees, while those with a treewidth of at most $k$ are known as partial $k$-trees [162].
  • If $tw(G)\leq k$, then maximum interval-number is $k+1$[777], track-number is $k+1$[777], and bend-number is $22k-$ 2[654].  [^29] The bend-number of a graph is the minimum $k$ such that it can be represented by grid paths with at most $k$ bends each, where two vertices are adjacent if and only if their corresponding paths share a grid edge[654, 653].
  • Block number is less than or equal to treewidth + 1[1204].
  • If $G$ has treewidth $k$, it is known to have a balanced separator of size $k+1$[153].
  • Let $G$ be a graph on $n$ vertices. Then $\text{TW}(G)\leq k$ if and only if there exists a $k$-achievable set $S$ of size $n-k-1$ [850].
  • Let $G$ be a graph and $H$ a minor of $G$. Then $\text{tw}(H)\leq\text{tw}(G)$ [311].
  • Let $G$ be a graph. Then $\textit{treewidth}(G)\leq\textit{catwidth}(G)\leq\textit{pathwidth}(G)$[1140].  [^30] Catwidth is the minimum tree-decomposition width of a graph $G$ when the pathwidth of the decomposition tree $T$ is restricted to 1[1140].
  • Let $G$ be a graph. Then comparable box $\textit{comparable box dimension}\leq\textit{treewidth}(G)+1$[400].
  • Every graph $G$ has a 2-blocking partition with width at most $1350(\text{tw}(G)+1)(\Delta(G)^{2}.$ ($\Delta$: maximum degree) [359]  [^31] An $\ell$-blocking partition of a graph $G$ is a partition of the vertex set $V(G)$ into connected sets such that every path of length greater than $\ell$ in $G$ contains at least two vertices in one part of the partition[359].
  • For a connected graph $G$, we have the following inequality:
    $\text{tw}(G)<\Delta(G)\left(\text{spanning tree congestion}(G)+1\right),$
    where $\text{tw}(G)$ is the treewidth of $G$ and $\Delta(G)$ is the maximum degree of $G$[801].  [^32] The spanning tree congestion problem seeks to minimize the maximum number of detours (congestion) that pass through any edge in a spanning tree of a graph[824, 855, 801].
Proof A.5.
Please refer to each reference.
A.1.2 Related concepts for tree-width
Related concepts include maximum order of a grid minor, maximum order of a grid-like-minor[644, 605], S-functions[1027], dimension[190], Hadwiger number[259], fractional Hadwiger number[644, 605], domino treewidth [168, 166], and the following concepts [605].
  • Local Tree-width [598]: A graph width parameter that associates with every natural number $r$ the maximum tree-width of an $r$-neighborhood in $G$.
  • Linear Local Tree-width: A linear variation of Local Tree-width[388, 598].
  • Simple Tree-width: Simple Tree-width holds following property.
    • For every $G$, we have $tw(G)\leq stw(G)\leq tw(G)+1$ [776].
    • A simple $k$-tree is a $k$-tree with the extra requirement that there is a construction sequence in which no two vertices are attached to the same $k$-clique.
  • Bag-connected Tree-width [718]: A graph width parameter that considers tree-decompositions where each cluster is connected.
  • Circuit Tree-width [1026, 205]: Circuit Tree-width holds following property.
    • Circuit Tree-width or Expression width [705] measures, for any Boolean function, the smallest treewidth of a circuit computing the function when considering an ordered binary decision diagram (OBDD).
    • Related parameters are OBDD width [706] and SDD width [205].
  • Stacked Treewidth [1220]: Width related to stacked $k$-tree.
    • A $k$-tree is called a stacked $k$-tree if during its construction no two vertices are stacked onto the same $k$-clique.
  • Edge-tree Width [863]: A graph width parameter that serves as the tree-like analogue of cut-width, or as an edge-analogue of tree-width.
  • Layered Treewidth [198, 389]: The minimum integer $k$ such that $G$ has a tree-decomposition and layering where each bag has at most $k$ vertices in each layer. $\text{The minimum degree }\leq 3\quad ltw(G)\leq 1.$ [390]. $ltw(G)\leq 2g+3,\text{ for any genus }g\text{ graph}$.
  • Weighted Tree-width: The weighted treewidth of a weighted graph is defined analogously to the (unweighted) treewidth [724].
  • 2-dimensional treewidth: Width defined as the minimum integer $k$ such that $G$ has two $k$-orthogonal tree decompositions $S$ and $T$.
[…]
  • Row Treewidth [198, 385]: The minimum integer $k$ such that $G$ is isomorphic to a subgraph of some graph $H$ with tree-width at most $k$ and for some path $P$.
  • Connected Tree-width [350, 926]: The minimum width of a tree-decomposition whose parts induce connected subgraphs. Connected Tree-width equals to connected tree vertex separation number[19].
  • Triangular Tree-width [889]: A graph width parameter known for making the computation of permanents of matrices with bounded width efficient.
  • Special Tree-width [302]: A complexity measure of graphs, positioned between path-width and tree-width.
  • Spaghetti Tree-width [806, 169]: A graph width parameter similar to Special Tree-width. Related concepts include strongly chordal tree-width and Directed Spaghetti Tree-width [169].
  • Pared Tree Width and Acceptance Width [750, 751]: Graph width parameters relevant to the computation of an alternating finite automaton.
  • Co-treewidth [381]: The tree-width of the complement of the input graph.
  • Constant Treewidth [79, 211]: A graph width parameter characteristic of control flow graphs in most programs.
  • Dynamic Treewidth [59, 789]: A graph width parameter designed to be computed for a dynamic graph $G$.
  • Project-join Tree Width [1025]: A graph width parameter linked to weighted projected model counting with graded project-join trees.
  • Semantic Tree-width [454]: This parameter determines whether a query is equivalent to a query of tree-width $k$, applicable to the class of Unions of Conjunctive Regular Path Queries with two-way navigation.
  • Induced Tree Width [1177]: A graph width parameter where the complexity of algorithms is dictated by the highest-dimensional factor across computations.
  • Free-connex Tree-width [395, 186]: A graph width parameter that requires a connected set of nodes, including the root, containing exactly the output variables.
  • Effective Width [432]: A graph width parameter for Partially Ordered Time, positioned between the well-known properties of width and dimension.
  • Embedded Width [1201]: Embedded width is a variation on tree-width, which is restricted to surface-embedded graphs.
  • Backdoor-tree-width [402, 556]: Backdoor-tree-width combines backdoor and tree-width approaches to create a fixed-parameter tractable framework for abstract argumentation, ensuring manageable parameter values. Related concept is C-backdoor tree-width [380].
  • Dual Tree-width [1099]: The Tree-width of the dual of a graph.
  • Singly-crossing treewidth [979]: Singly-crossing treewidth has been explored in algorithmic applications in [979] where the structure of a planar graph is integrated with a tree decomposition that includes clique-sums of bounded-size non-planar components.
  • linear tree-width [561, 1185]: Linear layout width of tree-width.
  • VF-tree-width [14] :VF-tree-width is a graph width parameter defined using vertex-free tree decompositions, focusing on edges instead of vertices, useful in analyzing the complexity of cycle matroids[14].
  • Branched treewidth [464] : Branched treewidth is a graph width parameter defined by limiting the number of branching nodes in a tree decomposition path[464].
  • Radial tree-width [49] : Radial tree-width is a graph parameter that measures the smallest radial width in a tree-decomposition of a graph, where the radial width is defined by the largest radius of the subgraphs (bags) in the decomposition.
  • Loose tree-decomposition [896] : A relaxation of the tree-decomposition associated to treewidth. Related parameter is tight bramble[896]. And let $G$ be a graph and $k\geq 1$ be an integer. The following conditions are equivalent[896]:
    1. 1.
      $G$ has a loose tree-decomposition of width $k$[896];
    2. 2.
      $G$ is a minor of $T^{(k)}=T\boxdot K_{k}$ (i.e., $\mathrm{ctp}(G)\leq k$)[896];
    3. 3.
      Every tight bramble $\mathcal{B}$ of $G$ has order at most $k$[896];
    4. 4.
      The mixed search number against an agile and visible fugitive is at most $k$ (i.e., $\mathrm{avms}(G)\leq k$)[896];
    5. 5.
      The monotone mixed search number against an agile and visible fugitive is at most $k$ (i.e., $\mathrm{mavms}(G)\leq k$)[896].
  • Underlying treewidth [384] : The underlying treewidth of a graph class is the smallest treewidth of a graph such that any graph in the class is a subgraph of its product with a complete graph.
  • Twin-treewidth [991] : Twin-treewidth (ttw) is a graph parameter that generalizes treewidth. It measures the treewidth of the twin-class graph, which is obtained by contracting equivalence classes of true or false twin vertices in the original graph[991].
  • Threshold Treewidth [557] : Threshold treewidth is a refinement of treewidth that considers a graph partitioned into light and heavy vertices, with a constraint on the maximum number of heavy vertices in each bag of the tree decomposition[557].
  • supertree-width (of hypergraph) [841]: Supertree-width is a graph parameter that extends treewidth to hypergraphs by considering generalized hypertree decompositions and minimizing the maximum number of hyperedges in each decomposition[841].
  • Weak Tree-width : A method of decomposing a graph into subsets (bags) connected in a tree structure. It has looser constraints than standard tree decompositions, with relaxed connectivity and adjacency within the bags. Related concept is strong bramble[69].Let $G$ be a simple finite graph. We have
[…]
    [69]
A.1.3 Application aspect for tree-width
We consider about application aspect for tree-width. Due to the versatility of tree structures in handling various types of data, tree-width has numerous applications across different fields (cf.[501]).
  • Protein structure
    Research on proteins is thriving in the field of medicine. Protein structure refers to the three-dimensional arrangement of amino acids in a protein. It can be modeled as a graph where nodes represent amino acids, and edges represent bonds or interactions, capturing the protein’s complex folding and connectivity(cf.[688, 1005]). Applications have also been studied in the context of tree-width[1226, 982, 674, 1224].
  • RNA and DNA
    Research on RNA and DNA is also thriving in the field of medicine. RNA and DNA are molecules that store genetic information. DNA is double-stranded and stable, forming a double helix, while RNA is single-stranded and more flexible, often acting in protein synthesis. Applications have also been studied in the context of tree-width[610, 1237, 1247, 445].
  • metabolic networks
    Metabolic networks are interconnected pathways within cells that convert nutrients into energy and essential molecules, supporting life processes. Applications have also been studied in the context of tree-width[278]. In addition, the application of Tree-width is being researched from biological and medical perspectives[974, 134].
  • social network
    A social network is a structure of individuals or organizations connected by social relationships, like friendships or collaborations. Applications have also been studied in the context of tree-width[8, 608, 874].
  • Bayesian network
    A Bayesian network is a graphical model representing probabilistic relationships among variables using nodes and directed edges (cf.[997]).Tree-width relates to Bayesian networks by measuring the complexity of inference; lower tree-width allows more efficient probabilistic computations and algorithms[415, 1152, 1015, 816, 947, 948].
  • Neural networks
    Neural networks are computational models inspired by the human brain, consisting of interconnected nodes that process data and learn patterns. A related concept is graph neural networks [1250]. Applications have also been studied in the context of tree-width[107].
  • Database
    Applications for Database have also been studied in the context of tree-width[453, 600, 588, 584].
  • Internet
    Applications for Internet have also been studied in the context of tree-width[318].
  • Markov network
    Markov networks, or Markov random fields, are undirected graphical models representing the joint distribution of variables. They are used in applications like image processing and statistical physics, emphasizing local dependencies (cf.[1184]). Applications for Markov network have also been studied in the context of tree-width[1113, 834, 735].
  • Tensor network
    Tensor networks are mathematical structures used to efficiently represent and compute high-dimensional data. They are applied in quantum physics, machine learning, and compressing large datasets by capturing complex relationships among variables(cf.[707]). Applications for Tensor network have also been studied in the context of tree-width[394].
For those interested in a more detailed understanding of tree-width, I recommend consulting lecture notes and surveys. Fortunately, there are many excellent books and survey papers available on the topic [163, 145, 665, 165, 775, 642, 167, 176].
A.2 Path-width
Path-width [763, 764, 620, 744, 645] quantifies how similar a graph is to a path by determining the minimum width of its path decomposition. An example of an obstruction to low path-width is blockage [146]. Path-width is notably applied in VLSI design [960]. Path-width is also called interval-width [769, 465].
The definition of path decomposition for a graph is as follows.
Definition A.6.
(cf.[787]) A path decomposition of a given graph $G=(V,E)$ is a sequence of subsets of $V$, $X=(X_{1},X_{2},\dots,X_{l})$, such that the following conditions hold:
  1. 1.
    $V=\bigcup_{i=1}^{l}X_{i}$ (i.e., every vertex of $G$ appears in at least one subset $X_{i}$).
  2. 2.
    For each edge $(u,v)\in E$, there exists some $i\in\{1,2,\dots,l\}$ such that both $u$ and $v$ belong to $X_{i}$.
  3. 3.
    For each $v\in V$, there exist indices $s(v),e(v)\in\{1,2,\dots,l\}$ such that $s(v)\leq e(v)$, and $v\in X_{j}$ if and only if $j\in\{s(v),s(v)+1,\dots,e(v)\}$.
The width of the path decomposition $X$ is defined as:
$\mathrm{pw}_{X}(G)=\max\{|X_{i}|\mid i=1,2,\dots,l\}-1.$
The pathwidth of $G$, denoted $\mathrm{pw}(G)$, is the minimum width over all possible path decompositions of $G$, i.e.,
$\mathrm{pw}(G)=\min\{\mathrm{pw}_{X}(G)\mid X\text{ is a path decomposition of}G\}.$
As another example, the definition of Path-width in the context of matroids is provided. For the definition of a matroid, please refer to Definition A.29 and related references.
Definition A.7.
[744] Given an ordering $(e_{1},e_{2},\dots,e_{n})$ of the elements of $M$, define the width of the ordering as:
$w_{M}(e_{1},e_{2},\dots,e_{n})=\max_{i\in[n]}f_{M}(e_{1},e_{2},\dots,e_{i}).$
For simplicity, we use $f_{M}(e_{1},e_{2},\dots,e_{i})$ instead of $f_{M}(\{e_{1},e_{2},\dots,e_{i}\})$.
The pathwidth of $M$ is defined as:
$\text{pw}(M)=\min w_{M}(e_{1},e_{2},\dots,e_{n}),$
where the minimum is taken over all orderings $(e_{1},e_{2},\dots,e_{n})$ of $E(M)$. An ordering $(e_{1},e_{2},\dots,e_{n})$ of $E(M)$ such that $w_{M}(e_{1},e_{2},\dots,e_{n})=\text{pw}(M)$ is called an optimal ordering.
A.2.1 Hierarchy for path-width
We consider about Hierarchy for path-width.
Theorem A.8.
The relationships between path-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Path-width, then it also has bounded Tree-width (Trivially).
  • The path-width of any graph is equal to one less than the smallest clique number of an interval graph [164].
  • If a graph $G$ has bounded Tree-depth, then it also has bounded Path-width [1058].
  • If a graph $G$ has bounded Path-width, then it also has bounded linear-rank-width [18].
  • If a graph $G$ has bounded Path-width, then it also has bounded linear-clique-width and linear NLC-width [617].
  • If $G$ has path-width $k$, the interval thickness of $G$ is at most $k+1$ [165].
  • If $G$ has path-width $k$, the minimum progressive black pebble demand over all directives of $G$ is at most $k+1$ [165].
  • If $G$ has path-width $k$, the minimum progressive black and white pebble demand over all directives of $G$ is at most $k+1$ [165].
  • Vertex separation number is equivalent to path-width [416].
  • Node search number is also equal to path-width [770].
  • Path-width is less than or equal to cut-width [788].
  • If a graph $G$ has bounded Path-width, then it also has bounded clique-width [448].
  • $\text{thinness}(G)\leq\text{pathwidth}(G)+1,$[875]
  • If a graph $G$ has bounded Band-width, then it also has bounded path-width [9].
  • If $G$ has $T_{h}$-witness, then $pw(G)\geq h$[597].
  • For every graph, $\text{ds}(G)\leq\text{pw}(G)+1\leq(\Delta(G)+1)\text{ds}(G)$. ($\delta_{d}(G)$: the maximum degree of $G$. $\text{ds}(G)$:Domination search number)[805].
Proof A.9.
Please refer to each reference.
A.2.2 Related concepts for path-width
Related concepts for path-width are following:
  • DAG-path-width [741]: Linear layout parameter of DAG-tree-width.
  • Layered path-width [198, 389]: Linear layout parameter of Layered tree-width. For any graph, layered treewidth (pathwidth) $\leq$ treewidth (pathwidth) (trivially). And queue number is bounded by layered pathwidth [387]. Every graph with pathwidth $\text{Every graph with pathwidth }k\text{ has layered pathwidth at most }\frac{k+1}{2}.$ Every graph with layered pathwidth $\text{Every graph with layered pathwidth }\lambda\text{ has track-number at most }3\lambda.$
  • Circuit Pathwidth: Path version of Circuit Tree-width [1026, 205].
    • Related parameters are OBDD width [706] and SDD width [205].
  • Simple Pathwidth: The simple pathwidth of $G$ is the smallest $w$ such that $G$ has a $w$-simple path decomposition of width $\leq w$ [143].
  • Row path-width [198, 385]: Linear layout parameter of Row tree-width.
  • Linear Local path-width [383] : A path variation of Linear Local Tree-width.
  • D-path-width and clique preserving d-path-width [1025]: A graph width parameter that has the splitting power of branching programs of bounded repetition and CNFs of bounded width.
  • Semantic path-width [454]: Linear layout parameter of Semantic tree-width.
  • Proper-path-width [1132, 1129, 1133, 1131, 1130, 884, 883]: Path-width that relates to mixed search game (A pursuit-evasion game on graphs where a searcher combines edge and node searches to capture a hidden intruder). The following relationship is known regarding proper-path-width.
    • For any graph $G$, $pw(G)\leq ppw(G)\leq pw(G)+1$ [1131, 1130].
    • For any simple graph $G$ and an integer $k$, $ppw(G)\leq k$ if and only if $G$ is a partial $k$-path [1131, 1130].
    • For any simple graph $G$, Mixed Search Number = Proper-path-width [1131, 1130].
    • edge search $\text{edge search number}(G)-1\leq\text{Proper-path-width}(G)\leq\text{edge search number}(G)$ [1131, 1130].
    • node search $\text{node search number}(G)-1\leq\text{Proper-path-width}(G)\leq\text{node search number}(G)$ [1131, 1130].
  • Connected path-width [323]: Linear layout parameter of Connected tree-width. Connected pathwidth of any graph $G$ is at most $12\times pw(G)+$ 1 [324]. Connected path-width equals to connected path vertex separation number[19].
  • Persistence path-width [377]: Persistence path-width refers to a graph path decomposition with width $k$, where each vertex of the graph is contained in at most $l$ nodes of the path.
  • Linked path decompositions [571]: Weaker variant of lean path decomposition (cf. [1145, 118]).
  • Dual path-width [1099]: The path-width of the dual of a graph (cf. [257]).
  • Radial path-width [49] : Linear layout of Radial tree-width.
  • Trellis-width [743, 462, 703, 664, 702]: Trellis-width measures the "trellis state-complexity" of a linear code, defined by the minimum dimension of intersections in a linear layout of subspaces. It is essentially equivalent to the path-width parameter.
A.2.3 Application aspect for path-width
We consider about application aspect for path-width.
  • VLSI
    VLSI (Very Large Scale Integration) refers to the process of integrating thousands to millions of transistors onto a single semiconductor chip (cf.[719, 910, 96]). It’s crucial for developing modern electronic devices, enabling complex circuits and high-performance computing. Applications for VLSI have also been studied in the context of path-width [959, 911, 909].
  • Compiler
    Applications for Compiler have also been studied in the context of path-width[161, 299].
  • Blockchain
    Applications for blockchain have also been studied in the context of path-width[746, 742].
  • Natural Language
    Applications for Natural Language have also been studied in the context of path-width[790].
  • IoT, Network
    Applications for IoT have also been studied in the context of path-width[60, 61].
A.3 Cut-width
Cut-width (also called folding number[288]), referenced in sources [290, 868, 867], measures the minimum number of edges crossing any vertical cut in a linear layout of a graph’s vertices. Related concepts include page-width [605].
Definition A.10.
Let $G=(V,E)$ be an undirected graph, where $V$ is the set of vertices and $E$ is the set of edges. The cutwidth of $G$ is the smallest integer $k$ such that there exists an ordering of the vertices $v_{1},v_{2},\dots,v_{n}$ in $V$, where for every $\ell=1,2,\dots,n-1$, the number of edges $(v_{i},v_{j})\in E$ with $i\leq\ell$ and $j>\ell$ is at most $k$.
A.3.1 Hierarchy for cut-width
We consider about Hierarchy for cut-width.
Theorem A.11.
The relationships between cut-width and other graph parameters are as follows:
  • If a graph $G$ has bounded cut-width, it also has bounded carving-width [264].
  • Path-width is less than or equal to cut-width [788].
  • $\text{thinness}(G)\leq\text{cutwidth}(G)+1,$[875, 116]
  • Bisection width is less than or equal to cut-width[333].
  • The cut-width of a graph can provide a lower bound on another parameter, the crossing number [360]. [^33] The crossing number of a graph is the minimum number of edge crossings in any drawing of the graph in the plane(cf.[1078, 785]).
  • In subcubic graphs (graphs with a maximum degree of three), the cut-width equals the path-width plus one [869].
[…]
Proof A.12.
Please refer to each reference.
A.3.2 Related concepts for cut-width
Related parameter is Weighted cut-width.
  • Weighted cut-width
    Cut-width which considers the weight of the graph[217, 216].
  • Circuit cut-width
    Circuit cut-width is a graph-width of circuits that characterizes the complexity of solving ATPG (Automatic Test Pattern Generation), demonstrating that circuits with smaller cut-widths can be solved more efficiently [999, 285].
  • Cyclic Cutwidth
    Cyclic Cutwidth is the minimal value of the maximum number of edges crossing any cut when the vertices are arranged on a circle[1024].
A.3.3 Application aspect for cut-width
We consider about application aspect for cut-width.
  • VLSI
    Applications for VLSI have been studied in the context of cut-width [123].
A.4 MIM-width
MIM-width (maximum induced matching width) [690, 691] measures the largest induced matching in any bipartite graph induced by the cuts in the decomposition.
Definition A.13.
[690, 691] A branch decomposition is a pair $(T,L)$, where $T$ is a subcubic tree and $L$ is a bijection from $V(G)$ to the set of leaves of $T$. For each edge $e$ of $T$, let $T^{e}_{1}$ and $T^{e}_{2}$ be the two connected components of $T-e$. Define $(A^{e}_{1},A^{e}_{2})$ as the vertex bipartition of $G$ such that, for each $i\in\{1,2\}$, the set $A^{e}_{i}$ contains all vertices in $G$ that are mapped by $L$ to leaves within $T^{e}_{i}$.
For a vertex set $A\subseteq V(G)$, let $\text{mim}(A)$ denote the maximum size of an induced matching in $G[A,V(G)\setminus A]$. The mim-width of $(T,L)$, denoted by $\text{mimw}(T,L)$, is defined as $\max_{e\in E(T)}\text{mim}(A^{e}_{1})$. The minimum mim-width over all branch decompositions of $G$ is referred to as the mim-width of $G$. If $|V(G)|\leq 1$, then $G$ does not admit a branch decomposition, and the mim-width of $G$ is defined to be 0.
A.4.1 Hierarchy for MIM-width
Theorem A.14.
The relationships between MIM-width and other graph parameters are as follows:
  • If a graph $G$ has bounded MIM-width, then it also has bounded Sim-width [1058].
  • If a graph $G$ has bounded Independent set or Dominating Number, then it also has bounded MIM-width [1058].
  • If a graph $G$ has bounded Boolean-width, then it also has bounded MIM-width [690, 691].
  • If a graph $G$ has bounded Clique-width, then it also has bounded MIM-width [689].
Proof A.15.
Please refer to each reference.
A.4.2 Related concepts for MIM-width
Related concepts are following:
  • One-sided maximum induced matching-width [126]: Graph width parameters which solves maximum independent set problem. Any graph with tree-independence number $k$ has o-mim-width at most $k$ [126].
  • SIM width [728, 463]: Graph width parameters which is a more useful parameter of MIM-width. Any graph with minor-matching hypertree-width $k$ has sim-width at most $k$ [126].
A.5 Linear MIM-width
Linear MIM-width [690, 691] is the linear restriction of MIM-width.
A.5.1 Hierarchy for linear-MIM-width
We consider about Hierarchy for linear-mim-width.
Theorem A.16.
The relationships between linear-MIM-width and other graph parameters are as follows:
  • If a graph $G$ has bounded linear-MIM-width, then it also has bounded MIM-width (trivially).
  • If a graph $G$ has bounded linear-MIM-width, then it also has bounded o-mim-width and sim-width[116].
  • If a graph $G$ has bounded tree-independence-number, then it also has bounded linear-MIM-width[116].
  • If a graph $G$ has bounded path-independence-number, then it also has bounded linear-MIM-width[116].
  • If a graph $G$ has bounded simultaneous interval number, then it also has bounded linear-MIM-width[116].
  • If a graph $G$ has bounded thinness, then it also has bounded linear-MIM-width[116].
Proof A.17.
Please refer to each reference.
A.5.2 Related concepts for Linear-MIM-width
Related concept is following:
  • Linear SIM width [728]: Linear layout parameter of SIM width. If a graph $G$ has bounded linear-SIM-width, then it also has bounded SIM-width (trivially).
A.6 Boolean-width
Boolean-width [230] measures the width of a graph decomposition based on the number of different unions of neighborhoods across cuts.
Definition A.18.
[230] A decomposition tree of a graph $G$ is a pair $(T,\delta)$, where $T$ is a tree with internal nodes of degree three, and $\delta$ is a bijection between the leaves of $T$ and the vertices of $G$. Removing an edge from $T$ yields two subtrees, and correspondingly, a cut $\{A,\bar{A}\}$ of $G$, where $A\subseteq V(G)$ corresponds to the leaves of one subtree.
Let $f:2^{V(G)}\rightarrow\mathbb{R}$ be a symmetric cut function, meaning $f(A)=f(\bar{A})$ for all $A\subseteq V(G)$. The f-width of $(T,\delta)$ is defined as the maximum value of $f(A)$ over all cuts $\{A,\bar{A}\}$ corresponding to the removal of an edge from $T$.
For rooted trees, subdivide an edge of $T$ to introduce a new root $r$, yielding a binary tree $T_{r}$. The subtree of $T_{r}$ rooted at a node $u$ corresponds to a subset $A_{u}^{r}\subseteq V(G)$, denoted simply by $A_{u}$ when the root is clear. For an edge $\{u,v\}$ with $u$ as a child of $v$ in $T_{r}$, the associated cut is denoted $\{A_{u},\bar{A}_{u}\}$.
A divide-and-conquer approach on $(T,\delta)$, following the edges of $T_{r}$ bottom-up, solves the problem recursively by combining solutions from the cuts given by edges from parent nodes to their children. In the context of independent sets, if two sets $X\subseteq A$ and $X^{\prime}\subseteq A$ have identical neighborhood unions across the cut (i.e., $N(X)\cap\bar{A}=N(X^{\prime})\cap\bar{A}$), they can be treated equivalently, suggesting the minimization of distinct neighborhood unions across cuts.
Definition A.19 (Boolean-Width).
[230] Given a graph $G$ and a subset $A\subseteq V(G)$, define the set of neighborhood unions across the cut $\{A,\bar{A}\}$ as:
$U(A)=\{Y\subseteq\bar{A}:\exists X\subseteq A\text{ such that }Y=N(X)\cap\bar{A}\}.$
The bool-dim function for $G$ is defined by:
[…]
The boolean-width of a decomposition tree $(T,\delta)$, denoted by $\text{boolw}(T,\delta)$, and the boolean-width of a graph $G$, denoted by $\text{boolw}(G)$, are defined using the bool-dim function as $f$.
The relationships between boolean-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Tree-width, then it also has bounded Boolean-width [690, 691].
[…]
  • If a graph $G$ has bounded Boolean-width, then it also has bounded MIM-width [690, 691].
A.7 Linear-Boolean-width
Linear Boolean-width [230] is the linear restriction of Boolean-width. If a graph $G$ has bounded Linear Boolean-width, then it also has bounded Boolean-width (trivially).
For any graph $G$, it holds that $\text{lboolw}(G)\leq\text{pw}(G)+1$ (Path-width) [215].
A.8 Band-width
Band-width [284, 1165] measures the minimum width of a band or interval in an optimal linear arrangement of a graph’s vertices. It is a width parameter introduced from VLSI design and networking. Related concepts include split-band-width [463].
Definition A.20.
[284, 1165] Let $f:V(G)\to\mathbb{Z}$ be a layout of a graph $G=(V,E)$. The bandwidth of this layout $f$ is defined as $\max_{\{u,v\}\in E(G)}|f(u)-f(v)|$. The bandwidth of the graph $G$ is the minimum bandwidth over all possible layouts of $G$, and this graph parameter is denoted by $\text{BW}(G)$.
A.8.1 Hierarchy for Band-width
We consider about hierarchy for Band-width.
Theorem A.21.
The relationships between band-width(bw) and other graph parameters are as follows:
  • $\text{ppw}(G)\leq\text{bw}(G)\leq 2\times\text{ppw}(G)$ (ppw: path-partition-width)
  • Max-leaf number is a minimal upper bound for band-width [1154].
  • $\text{bw}(G)\leq 2\times\text{pdw}(G)$ (pdw: path-distance-width)
  • If a graph $G$ has bounded Band-width, then it also has bounded bisection-width [1154].
  • If a graph $G$ has bounded Band-width, then it also has bounded Tree-width.
  • If a graph $G$ has bounded Band-width, then it also has bounded path-width [9].
  • If a graph $G$ has bounded Band-width, then it also has bounded proper thinness[191].
  • If a graph $G$ has bounded band-width, then it also has bounded maximum degree, $c$-closure [782]  [^34] A graph $G$ is $c$-closed if, for all pairs of nonadjacent vertices $u$ and $v$, the size of the intersection of their neighborhoods $|N(u)\cap N(v)|$ is less than $c$. The $c$-closure of $G$ is the smallest integer $c$ for which $G$ is $c$-closed.[783, 476, 727], acyclic chromatic number, and $h$-index [425] [1154]  [^35] The $h$-index of a graph is the maximum integer $h$ such that the graph has at least $h$ vertices, each with a degree of at least $h$(cf.[426, 148])..
  • Treespan is less than or equal to band-width [469].
  • For each $\Delta\geq 2$, every graph $G$ on $n$ vertices with maximum degree $\Delta(G)\leq\Delta$ has bandwidth $\text{bw}(G)\leq\frac{6n}{\log_{\Delta}\left(\dfrac{n}{\text{separator}(G)}\right)}$[199].
  • The slope number $\text{sn}(G)$ satisfies: $\text{sn}(G)\leq\frac{1}{2}\text{bw}(G)\cdot(\text{bw}(G)+1)+1.$[391]  [^36] The slope number of a graph is the minimum number of distinct edge slopes required in a straight-line drawing of the graph[1182, 700].
Proof A.22.
Please refer to each reference.
A.8.2 Related concepts for band-width
Related parameter is Weighted band-width.
  • Edge band-width: Edge bandwidth measures the smallest maximum difference between labels of incident edges when numbering all edges of a graph; it reflects how closely edges can be labeled without large gaps[989, 835, 25].
  • Cyclic Bandwidth: Cyclic Bandwidth is the smallest possible maximum distance between adjacent vertices when arranging the vertices around a circle to minimize that maximum distance[1024].
A.9 Boundary-width
Vertex-boundary-width [965] measures the minimum boundary of vertices in a discrete isoperimetric problem. Also edge-boundary-width [965] measures the minimum boundary of edges in a discrete isoperimetric problem.
A.10 Carving-width
Carving-width [1073, 1142, 847] measures a graph’s complexity by focusing on edge cuts, similar to tree-width, but with an emphasis on edges rather than vertices.
Definition A.23.
[1073, 1142, 847] A tree $T$ is called a carving of a graph $G$, and the pair $(T,w)$ is referred to as a carving decomposition of $G$, where $w(e)$ is a weight function defined on the edges $e\in E(T)$. The width of a carving decomposition $(T,w)$ is the maximum weight $w(e)$ over all edges $e$ in $T$. The carving-width of $G$, denoted by $\text{cw}(G)$, is the minimum width among all possible carving decompositions of $G$. If $|V(G)|=1$, we define $\text{cw}(G)=0$.
A.10.1 Hierarchy for carving-width
We consider about Hierarchy for carving-width.
Theorem A.24.
The relationships between carving-width and other graph parameters are as follows:
  • If a graph $G$ has bounded cut-width, it also has bounded carving-width [264].
  • Let $G$ be a graph with maximum degree $d$. Then, the following inequality holds:
    $\frac{2}{3}\left(\text{tw}(G)+1\right)\leq\text{cng}(G)\leq d\left(\text{tw}(G)+1\right),$
    where $\text{tw}(G)$ is the tree-width of $G$ and $\text{cng}(G)$ is the carving-width of $G$.[144]
A.10.2 Related concept for Carving-width
Related concept is following:
  • Dual Carving-width [848]: The carving-width of the dual of a graph.
A.10.3 Application for carving-width
Carving-width is closely related to the concept of congestion in network theory. For example, consider a real-world scenario where a network with low performance is used—this often leads to frequent delays and congestion. The level of congestion experienced in such cases is analogous to the carving-width. Extensive research has been conducted to minimize delays and congestion, making this measure particularly relevant in network optimization studies [1073]. This graph width parameter is a minimal upper bound for maximum degree and tree-width [1058, 119].
A.11 Branch-width
Branch-width [565, 966, 1038, 177] measures the smallest width of a branch decomposition of a graph. The width is the minimum number of edges connecting subgraph pairs. This graph width parameter is known to be extendable to matroids and connectivity systems. Examples of obstructions are Tangle [565, 661], Profile [344, 345, 421, 341, 352], $k$-block [409, 430, 778], Ultrafilter [529], Maximal ideal [1231], loose tangle [966], loose tangle kit [966], Quasi-Ultrafilter [512], Weak-Ultrafilter [497], tangle kit[685], Bramble[858, 70], (k,m)-obstacle [473]. Related concepts include sphere-cut width [151] and Rooted Branch-width[704].
A.11.1 Hierarchy for branch-width
We consider about Hierarchy for branch-width.
Theorem A.25.
The relationships between branch-width and other graph parameters are as follows:
  • For all matroids $M$, $\text{bw}(M)\leq\text{bd}(M)$ (Branch-depth) [961].
  • $\text{bw}(M)\leq\text{cdd}(M)\leq\text{cd}(M)$ and $\text{bw}(M)\leq\text{cdd}(M)\leq\text{dd}(M)$ [1058] (cdd:contraction-deletion-depth, cd:contraction-depth, and dd:deletion-depth).
  • $\text{bw}(G)\leq\text{tw}(G)+1\leq\frac{3}{2}\cdot\text{bw}(G)$ for a graph $G$ with $\text{bw}(G)\geq 2$ (tree-width) [164].
[…]
  • The branch-width of a graph is a lower bound on the zero-forcing number [438].
  • $bw(G)\leq tw(G)+1\leq Z(G)+1$ ($Z(G)$: zero-forcing number)[438].
  • If a graph $G$ has bounded Linear Branch-width, then it also has bounded Branch-width (trivially).
Proof A.26.
Please refer to each reference.
A.12 Linear Branch-width
Linear Branch-width (Also called Linear-width[525] or caterpillar width[152]) is the linear restriction of Branch-width. This graph width parameter is known to be extendable to matroids and connectivity systems. Examples of obstructions are Linear Tangle [527], Single Ideal [524], Linear Loose Tangle [511], Linear-obstacle [473], Ultra Matroid [492], Ultra Antimatroid [505], Ultra Greedoid [505], Ultra Quasi-matroid [498] (cf. [747]).
A.12.1 Hierarchy for Linear-branch-width
We consider about Hierarchy for Linear-branch-width.
Theorem A.27.
The relationships between Linear Branch-width and other graph parameters are as follows:
  • $\text{pw}(G)\leq\text{lbw}(G)\leq 1+\text{pw}(G)$ (path-width) [954].
  • If a graph $G$ has bounded Linear Branch-width, then it also has bounded Branch-width (trivially).
Proof A.28.
Please refer to each reference.
A.13 Rank-width
Rank-width [682] measures the complexity of a graph based on the rank of adjacency matrices over cuts, closely linked to Matroid theory [967, 1205, 1212] and the Rank function. Examples of obstructions are $\rho$-Tangle [681].
The concept of a rank function is introduced from the world of matroids, and parameters such as Branch-width, Tree-width, and Tangles have been extended from graphs to matroids[565, 966, 529, 662, 663]. Therefore, we will briefly introduce the concept of matroids. For those interested in learning more about matroids, please refer to books, lecture notes, or surveys on the topic[968, 1208, 1029, 1213].
Definition A.29 (Matroid).
(cf.[968, 1208]) A matroid is a collection of subsets $\mathcal{F}$ of a set $E$ that satisfies the following conditions:
  1. 1.
    $E\in\mathcal{F}$,
  2. 2.
    If $F_{1},F_{2}\in\mathcal{F}$, then $F_{1}\cap F_{2}\in\mathcal{F}$,
  3. 3.
    If $F\in\mathcal{F}$ and $\{F_{1},F_{2},\ldots,F_{k}\}$ is the set of minimal members of $\mathcal{F}$ that properly contain $F$, then the sets $F_{1}\setminus F,F_{2}\setminus F,\ldots,F_{k}\setminus F$ partition $E\setminus F$.
Matroid can also be defined as follows[968, 1208].
Definition A.30 (Matroid of Rank $d+1$).
(cf.[968, 1208])
A matroid on $E$ of rank $d+1$ is a function $r:2^{E}\rightarrow\mathbb{Z}$ satisfying the following conditions:
  1. 1.
    $0\leq r(S)\leq|S|$ for any $S\subseteq E$,
  2. 2.
    If $S\subseteq U\subseteq E$, then $r(S)\leq r(U)$,
  3. 3.
    $r(S\cup U)+r(S\cap U)\leq r(S)+r(U)$ for any $S,U\subseteq E$,
  4. 4.
    $r(\{0,\ldots,n\})=d+1$.
Based on the above, we define Rank-width as follows [682] .
Definition A.31.
[682] Let $G=(V,E)$ be a graph. The cut-rank function $\rho_{G}$ is defined for any subset $X\subseteq V$ as the rank of the $|X|\times|V\setminus X|$ binary matrix $A_{X}$, where the entry at position $(i,j)$ is 1 if and only if the $i$-th vertex in $X$ is adjacent to the $j$-th vertex in $V\setminus X$. We set $\rho_{G}(X)=0$ when $X=\emptyset$ or $X=V$.
A rank-decomposition of $G$ is a pair $(T,L)$, where $T$ is a subcubic tree and $L:V\to\text{Leaves}(T)$ is a bijection. For each edge $e\in E(T)$, removing $e$ partitions the leaves of $T$ into two sets $A_{e}$ and $B_{e}$. The width of $e$ is defined as $\rho_{G}(L^{-1}(A_{e}))$. The width of the rank-decomposition $(T,L)$ is the maximum width over all edges $e\in E(T)$.
The rank-width of $G$, denoted by $\text{rw}(G)$, is the minimum width over all rank-decompositions of $G$. If $|V|<2$, we define $\text{rw}(G)=0$.
A.13.1 Hierarchy for rank-width
We consider about hierarchy for rank-width.
Theorem A.32.
The relationships between rank-width and other graph parameters are as follows:
  • $\text{rw}(G)\leq\text{tw}(G)+1\leq\text{pw}(G)+1$ (tw: tree-width and pw: path-width) [1144].
  • Neighborhood Diversity is greater than or equal to rank-width [544].
Proof A.33.
Please refer to each reference.
A.13.2 Related concept for rank-width
Related concept is following:
  • Bi-rank-width [729]: A graph width parameter that discusses possible extensions of the notion of rank-width to colored graphs.
  • Signed-rank-width [542, 558, 548, 552]: Rank-width used for model counting.
  • $\mathbb{Q}$-Rank-Width [686]: A modified width parameter that utilizes the rank function over the rational field, as opposed to the binary field traditionally used. The $\mathbb{Q}$-rank-width does not exceed its corresponding clique-width. Furthermore, the $\mathbb{Q}$-rank-width is consistently less than or equal to the rank-width.
And Linear Rank-width is the linear restriction of Rank-width. If a graph G has bounded Linear Rank-width, then it also has bounded Rank-width (trivially).
A.14 Clique-width
Clique-width [684] measures the complexity of a graph based on a composition mechanism using vertex labels. Note that a clique in a graph is a subset of vertices where every pair of distinct vertices is connected by an edge, forming a complete subgraph (cf.[1128]).
Definition A.34 (Clique-width).
Let $k$ be a positive integer. A $k$-graph $(G,\text{lab})$ consists of a graph $G$ and a labeling function $\text{lab}:V(G)\rightarrow\{1,2,\ldots,k\}$, which assigns one of $k$ labels to each vertex in $G$. In this context, all graphs are finite, undirected, and do not contain loops or parallel edges. The label of a vertex $v$ is denoted by $\text{lab}(v)$.
We define the following operations on $k$-graphs:
  1. 1.
    For $i\in\{1,\ldots,k\}$, $\cdot_{i}$ denotes an isolated vertex labeled by $i$.
  2. 2.
    For $i,j\in\{1,2,\ldots,k\}$ with $i\neq j$, the unary operator $\eta_{i,j}$ adds edges between all vertices labeled $i$ and all vertices labeled $j$. Formally, if $(G,\text{lab})$ is a $k$-graph, then $\eta_{i,j}(G,\text{lab})=(G^{\prime},\text{lab})$, where $V(G^{\prime})=V(G)$ and $E(G^{\prime})=E(G)\cup\{vw:\text{lab}(v)=i,\text{lab}(w)=j\}$.
  3. 3.
    The unary operator $\rho_{i\rightarrow j}$ relabels all vertices labeled $i$ to $j$. Formally, if $(G,\text{lab})$ is a $k$-graph, then $\rho_{i\rightarrow j}(G,\text{lab})=(G,\text{lab}^{\prime})$, where
    $\text{lab}^{\prime}(v)=\begin{cases}j&\text{if }\text{lab}(v)=i,\\ \text{lab}(v)&\text{otherwise}.\end{cases}$
  4. 4.
    The binary operation $\oplus$ denotes the disjoint union of two $k$-graphs. If $G_{1}$ and $G_{2}$ are two graphs, then $G_{1}\oplus G_{2}$ is their disjoint union.
A well-formed expression $t$ using these operations is called a $k$-expression. The value $\text{val}(t)$ of a $k$-expression $t$ is the $k$-graph produced by performing the operations in $t$. If $\text{val}(t)=(G,\text{lab})$, we say that $t$ is a $k$-expression of $G$.
The clique-width of a graph $G$, denoted by $\text{cwd}(G)$, is the minimum $k$ such that there exists a $k$-expression for $G$.
A.14.1 Hierarchy for clique-width
We consider about hierarchy for clique-width.
Theorem A.35.
The relationships between clique-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Clique-width, then it also has bounded twin-width [536].
  • If a graph $G$ has bounded linear clique-width, then it also has bounded clique-width (trivially).
[…]
  • The stretch-width of any graph is at most twice its clique-width[1255].
  • If a graph $G$ has bounded Path-width, then it also has bounded clique-width [448].
  • Fusion-width(G) $\leq$ Clique-width(G) [532].
  • If a graph $G$ has bounded edge clique cover number, then it also has bounded Clique-width (and simultaneous interval number)[116].
[…]
  • Neighborhood Diversity is greater than or equal to clique-width [544].
  • If a graph $G$ has bounded Twin-cover, then it also has bounded Rank-width/Clique-width [543].
  • For any graph $G$, For any graph $G$, $\text{sd}(G)\leq 2\text{clique-width}(G)-2$. (sd: symmetric difference). [51]
  • For any graph $G$, $\text{fun}(G)\leq 2\text{clique-width}(G)-1$. (fun: functionality). [51, 399]  [^37] Functionality of a vertex $y$ is the minimum number of vertices $k$ such that $y$ can be represented as a Boolean function of $k$ vertices[51, 399].
Proof A.36.
Please refer to each reference.
A.14.2 Related concepts for clique-width
Related concepts are following:
  • H-Clique-Width [667]: A graph width parameter which aims to bridge the gap between classical hereditary width measures and the recently introduced graph product structure theory [636].
  • Signed Clique-width [1125]: Clique-width used for model counting.
  • Multi-Clique-Width (MCW) [533]: A graph width parameter that provides a natural extension of tree-width. For every graph $G$, it holds that $\text{mcw}(G)\leq\text{tw}(G)+2$ [533] (tw:Tree-width), $\text{boolw}(G)\leq\text{mcw}(G)\leq 2^{\text{boolw}(G)}$ [533] (boolw:Boolean-width), and $\text{mcw}(G)\leq\text{cw}(G)$ [533] (cw:Clique-width).
  • Clique-Partitioned Treewidth [567]: A parameter for graphs that is similar to tree-width but additionally considers the structure (clique) of each bag.
  • courcelle2004clique: Symmetric clique-width is a graph parameter that measures complexity using equivalence classes of vertex adjacency relationships within subsets.
A.15 Linear-Clique-width
Linear Clique-width (Also called Sequential clique-width[446, 447]) is the linear restriction of Clique-width.
A.15.1 Hierarchy for Linear-Clique-width
The relationships between Linear Clique-width and other graph parameters are as follows:
  • If a graph $G$ has bounded linear clique-width, then it also has bounded clique-width (trivially).
  • If a graph $G$ has bounded shrub-depth, then it also has bounded linear clique-width [690, 691].
  • If a graph $G$ has bounded linear clique-width, then it also has bounded clique-tree-width [617].
A.16 NLC-width
NLC-width and NLCT-width [617, 1197] measure the complexity of a graph based on a composition mechanism similar to clique-width.
Definition A.37.
[617, 1197] Let $k$ be a positive integer. The class $\text{NLC}_{k}$ of labeled graphs is defined recursively as follows:
  1. 1.
    A single vertex graph labeled $a$, where $a\in[k]$, belongs to $\text{NLC}_{k}$.
  2. 2.
    Let $G=(V_{G},E_{G},\text{lab}_{G})$ and $J=(V_{J},E_{J},\text{lab}_{J})$ be two vertex-disjoint labeled graphs in $\text{NLC}_{k}$, and let $S\subseteq[k]\times[k]$ be a relation. The graph $G\times_{S}J$ is defined as $(V^{\prime},E^{\prime},\text{lab}^{\prime})$, where:
    $V^{\prime}:=V_{G}\cup V_{J},$
    $E^{\prime}:=E_{G}\cup E_{J}\cup\{\{u,v\}\mid u\in V_{G},v\in V_{J},(\text{lab}_{G}(u),\text{lab}_{J}(v))\in S\},$
    and the labeling function $\text{lab}^{\prime}$ is given by:
    $\text{lab}^{\prime}(u):=\begin{cases}\text{lab}_{G}(u)&\text{if }u\in V_{G},\\ \text{lab}_{J}(u)&\text{if }u\in V_{J}.\end{cases}$
    Then, $G\times_{S}J$ is in $\text{NLC}_{k}$.
  3. 3.
    Let $G=(V_{G},E_{G},\text{lab}_{G})$ be a labeled graph in $\text{NLC}_{k}$, and let $R:[k]\to[k]$ be a function. The graph $\circ R(G)$ is defined as $(V_{G},E_{G},\text{lab}^{\prime})$, where the labeling function $\text{lab}^{\prime}$ is given by:
    $\text{lab}^{\prime}(u):=R(\text{lab}_{G}(u))\quad\text{for all }u\in V_{G}.$
    Then, $\circ R(G)$ is in $\text{NLC}_{k}$.
The NLC-width of a labeled graph $G$ is the smallest integer $k$ such that $G\in\text{NLC}_{k}$.
A.16.1 Hierarchy for NLC-width
We consider about hierarchy for NLC-width.
Theorem A.38.
The relationships between NLC-width and other graph parameters are as follows:
[…]
Proof A.39.
Please refer to each reference.
A.17 Linear NLC-width
Linear NLC-width [617] is the linear restriction of NLC-width.
A.17.1 Hierarchy for Linear-NLC-width
We consider about Hierarchy for Linear-NLC-width.
Theorem A.40.
The relationships between Linear NLC-width and other graph parameters are as follows:
  • If a graph $G$ has bounded linear NLC-width, then it also has bounded NLC-width (trivially).
  • If a graph $G$ has bounded linear NLC-width, then it also has bounded NLCT-width [617].
[…]
Proof A.41.
Please refer to each reference.
A.18 Hypertree-width
Hypertree-width [17] measures the complexity of hypergraphs by extending tree-width concepts to cover hyperedges.
First, we introduce the concept of a hypergraph [214]. Hypergraphs have numerous real-world applications across various fields. For readers interested in more detailed information or exploring these applications, we recommend consulting lecture notes, surveys, or books on the subject (e.g., [214, 449, 753, 75, 915, 213, 124]).
Definition A.42 (Hypergraph).
[214] A hypergraph $H=(V,E)$ consists of a finite set $V$ of vertices and a finite family $E=\{e_{i}\}_{i\in I}$ of non-empty subsets of $V$, called hyperedges, where $I$ is a finite index set. The set $V$ is also denoted by $V(H)$, and the family $E$ is denoted by $E(H)$.
The order of the hypergraph $H$, denoted by $|V|$, is the cardinality of the vertex set $V$. The size of the hypergraph, denoted by $|E|$, is the cardinality of the hyperedge family $E$.
  • The empty hypergraph is defined as the hypergraph where:
    • $V=\emptyset$,
    • $E=\emptyset$.
  • A trivial hypergraph is defined as a hypergraph where:
    • $V\neq\emptyset$,
    • $E=\emptyset$.
Example A.43.
Let $H$ be a hypergraph with vertex set $V(H)=\{A,B,C,D,E\}$ and hyperedge set $E(H)=\{e_{1},e_{2},e_{3}\}$, where:
$e_{1}=\{A,D\},\quad e_{2}=\{D,E\},\quad e_{3}=\{A,B,C\}.$
Thus, $H$ is represented by the pair $H(V,E)=(\{A,B,C,D,E\},\{\{A,D\},\{D,E\},\{A,B,C\}\})$.
Next, we introduce the concept of a hypertree-width.
Definition A.44.
[17] A hypertree decomposition of a hypergraph $H=(V,E)$ is a triple $D=(T,\chi,\lambda)$, where:
  • $T$ is a tree,
  • $\chi$ is a function that assigns each tree node $t\in V(T)$ a set $\chi(t)\subseteq V$, known as a bag,
  • $\lambda$ is a function that maps each tree node $t\in V(T)$ to a set $\lambda(t)\subseteq E$, known as an edge cover of the bag $\chi(t)$.
The decomposition must satisfy the following conditions:
  1. 1.
    Edge Cover Condition (P1): For every hyperedge $e\in E(H)$, there exists a tree node $t\in V(T)$ such that $e\subseteq\chi(t)$.
  2. 2.
    Connectedness Condition (P2): For every vertex $v\in V(H)$, the set of tree nodes $t\in V(T)$ with $v\in\chi(t)$ induces a connected subtree of $T$.
  3. 3.
    Special Condition (P3): For any two nodes $t,t^{\prime}\in V(T)$ where $t^{\prime}$ is a descendant of $t$ in $T$, and for every hyperedge $e\in\lambda(t)$, it holds that $(e\setminus\chi(t))\cap\chi(t^{\prime})=\emptyset$.
The width of a hypertree decomposition $D=(T,\chi,\lambda)$ is defined as the maximum size of any edge cover $\lambda(t)$ over all tree nodes $t\in V(T)$. The hypertree width $\text{htw}(H)$ of the hypergraph $H$ is the minimum width over all possible hypertree decompositions of $H$.
The known obstructions are Hypertangles [16] and Hyperbrambles [16]. In the future, we will consider a hyperultrafilter analogous to [529] (cf.[872, 509]).
A.18.1 Related concepts for Hypertree-width
Related concepts are following:
  • Hyperbranch width [16]: A generalized graph width parameter that extends the concept of hypertree-width.
  • Hyper-T-width and Hyper-D-width [1047]: These graph width parameters were introduced as the first stable connectivity measures for hypergraphs, providing tools for designing algorithms to address various problems in hypergraph theory. They are closely related to Hyper-Directed-width.
  • Fractional hypertree-width [877]: A hypergraph measure that is analogous to tree-width and hypertree-width.
  • FAQ-width [2, 759] : FAQ-width is a width parameter used to optimize the evaluation of functional aggregate queries (FAQs) with free variables over semirings, generalizing fractional hypertree width[2, 759].
  • m-width [710] : m-width is a graph width parameter used in join processing, often smaller than fractional hypertree width, aiding in identifying subquadratically solvable joins[710].
  • Hyperpath-width and Hypertree-depth [15]: Path-width and tree-depth in hypergraph.
  • Closure tree-width and Closure hypertree-width [13] : This is a width parameter used for applications in databases and similar fields.
  • Biconnected width [480]: Biconnected width measures the maximum vertex count in biconnected components of a hypergraph’s primal graph.
  • TCLUSTER-width [321, 586]: TCLUSTER-width measures the tree-likeness of a hypergraph, defined as the maximum clique size in its chordal graph representation.
  • Hinge-width [624, 625]: Hinge-width measures a hypergraph’s complexity by generalizing acyclic hypergraphs, computable in polynomial time.
  • Hierarchical hypertree width [671]: Hierarchical hypertree width measures how far a join query is from being hierarchical, influencing the complexity of temporal join algorithms [671].
  • Nest set-width [820] : Nest set-width is a graph width parameter defined by the smallest width of a nest set elimination order, which organizes vertices into subsets ordered by inclusion[820].
  • $\beta$-Hyperorder width [587, 242]: $\beta$-Hyperorder width is a graph width parameter that generalizes hypertree width, tailored for analyzing the complexity of negative conjunctive queries [587, 242]. A related parameter is $\beta$-fractional hyperorder width.
  • # hypertree-width [273, 595] : # hypertree-width is a structural measure of query instances that determines the tractability of counting solutions in conjunctive queries and constraint satisfaction problems[273, 595].
  • Primal treewidth [82] : Primal treewidth is the treewidth of the primal graph of a CSP instance, where vertices represent variables and edges connect variables appearing together in constraints[82].
    Related parameters are Incidence tree-width, modular incidence treewidth[559], Incidence clique-width, twinclass-pathwidth[650], twinclass-treewidth[650], modular tree-depth[939], prefix pathwidth[405], (generalized) modular tree-width[802], modular pathwidth[895], and Incidence MIM-width[1126, 1214].
    Various hierarchies are also known. For example, $\beta$-hypertree width is less than or equal to incidence clique-width[559]. Incidence clique-width is less than or equal to signed incidence clique-width[457] and modular incidence treewidth[559]. Signed incidence clique-width[457] and modular incidence treewidth[980, 559] are less than or equal to incidence treewidth[1126, 1054]. Incidence treewidth is less than or equal to primal treewidth[1054].
Question A.45.
Is there any new characteristic that emerges by extending each of these width parameters to SuperHyperGraphs? (cf.[521, 509])
A.18.2 Applications for Hypertree-width
We consider about Applications for Hypertree-width.
  • Artificial intelligence
    Similar to tree-width, the concept of hypertree-width is applied in the techniques and concepts of artificial intelligence that utilize graphs[406, 962].
  • Database
    The relationship between queries and hypertree-width has been particularly studied in the context of databases[570, 1160, 582, 594, 709].
A.19 Modular-width
Modular-width [5] is defined using modular decompositions of the graph.
Definition A.46.
[5] A graph $G$ is said to have a tree-model with $m$ colors and depth $d\geq 1$ if there exists a rooted tree $T$ of height $d$ such that:
  1. 1.
    The set of leaves of $T$ is exactly $V(G)$.
  2. 2.
    Every root-to-leaf path in $T$ has length exactly $d$.
  3. 3.
    Each leaf of $T$ is assigned one of $m$ colors (this is not a graph coloring in the traditional sense).
  4. 4.
    The existence of an edge between $u,v\in V(G)$ depends only on the colors of $u$ and $v$, and the distance between $u$ and $v$ in $T$.
The class of all graphs that have a tree-model with $m$ colors and depth $d$ is denoted by $\text{TM}_{m}(d)$.
Definition A.47.
[5] A class of graphs $\mathcal{G}$ has shrub-depth $d$ if there exists some $m$ such that $\mathcal{G}\subseteq\text{TM}_{m}(d)$, and for every natural number $m$, we have $\mathcal{G}\nsubseteq\text{TM}_{m}(d-1)$.
A.19.1 Hierarchy for Modular-width
We consider about hierarchy for Modular-width.
Theorem A.48.
The relationships between modular-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Modular-width, then it also has bounded Clique-width [535].
  • If a graph $G$ has bounded vertex cover or neighborhood diversity, then it also has bounded Modular-width [535].  [^38] Neighborhood diversity is a graph parameter that measures the number of distinct vertex neighborhoods in a graph, where vertices with identical neighborhoods are grouped together(cf.[780, 545]).
  • If a graph $G$ has bounded twin-cover, then it also has bounded Modular-width [535].  [^39] The twin-cover of a graph is the smallest set of vertices such that every edge is either covered by this set or connects two vertices with identical neighborhoods[535].
Proof A.49.
Please refer to each reference.
A.20 Submodular-width
Submodular-width [583] is defined using the submodular condition property. And sharp submodular width is a graph width parameter that refines submodular width, allowing improved algorithm runtimes for evaluating queries on arbitrary semirings[758].
A.21 Amalgam-width
Amalgam-width [860] is a new matroid width parameter based on matroid amalgamation. Note that matroid amalgamation is a process of combining two matroids on overlapping ground sets into a single matroid, preserving their independent sets while maintaining matroid properties, particularly submodularity (cf.[993, 941]).
A.22 Kelly-width
Kelly-width [893, 768] is a parameter for directed graphs, analogous to tree-width. Related concepts include Kelly-path-width [765].
Definition A.50.
[893] A Kelly-decomposition of a directed graph $G$ is a triple $D:=(D,(B_{t})_{t\in V(D)},(W_{t})_{t\in V(D)})$ that satisfies the following conditions:
  • $D$ is a directed acyclic graph (DAG) and $(B_{t})_{t\in V(D)}$ is a partition of $V(G)$.
  • For all $t\in V(D)$, $W_{t}\subseteq V(G)$ is a set of vertices that guards $B_{t}^{\downarrow}:=\bigcup_{t^{\prime}\leq_{D}t}B_{t^{\prime}}$, where $\leq_{D}$ denotes the partial order induced by $D$.
  • For each node $s\in V(D)$, there exists a linear order on its children $t_{1},\ldots,t_{p}$ such that for all $1\leq i\leq p$, $W_{t_{i}}\subseteq B_{s}\cup W_{s}\cup\bigcup_{j<i}B_{t_{j}}^{\downarrow}$. Similarly, there is a linear order on the roots such that $W_{r_{i}}\subseteq\bigcup_{j<i}B_{r_{j}}^{\downarrow}$.
The width of $D$ is defined as $\max\{|B_{t}\cup W_{t}|:t\in V(D)\}$. The Kelly-width of $G$ is the minimum width over all possible Kelly-decompositions of $G$.
A.22.1 Hierarchy for Kelly-width
We consider about hierarchy for Kelly-width.
Theorem A.51.
The relationships between Kelly-width and other graph parameters are as follows:
  • A directed graph $G$ has directed elimination width $\leq k$ if and only if $G$ has Kelly-width $\leq k+1$.[893]
  • The Kelly-width is less than or equal to the Directed Path-width plus one [550].
  • If a digraph $G$ has bounded Kelly-width, then it also has bounded Directed Tree-width [129, 71].
  • If a graph has bounded Kelly-path-width, then it has also bounded Kelly-width (trivially).
Proof A.52.
Please refer to each reference.
A.23 Monoidal Width
Monoidal Width [823] measures the complexity of morphisms in monoidal categories. Note that monoidal categories are categories equipped with a binary operation (tensor product) and an identity object, allowing the composition of objects and morphisms, similar to how multiplication works in algebra (cf.[1161])  [^40] The tensor product of two vector spaces combines them into a new vector space, where each element is a pair from the original spaces, allowing multilinear maps to be expressed as linear maps(cf.[277, 1107])..
A.24 Tree-cut Width
Tree-cut width [554] is a graph width parameter that serves as an analogue to tree-width, specifically for edge cuts. This concept is accompanied by related width parameters, including 0-Tree-cut width [555].
Definition A.53.
[554] A tree-cut decomposition of a graph $G$ is a pair $(T,\mathcal{X})$, where $T$ is a tree and $\mathcal{X}=\{X_{t}\subseteq V(G):t\in V(T)\}$ is a near-partition of $V(G)$. Each set $X_{t}$ is called a bag.
For an edge $e=(u,v)$ in $T$, let $T_{u}$ and $T_{v}$ be the two connected components of $T-e$ containing $u$ and $v$, respectively. The corresponding partition of $V(G)$ is $(\bigcup_{t\in T_{u}}X_{t},\bigcup_{t\in T_{v}}X_{t})$, and the set of edges with one endpoint in each part is denoted by $\text{cut}(e)$.
A tree-cut decomposition is rooted if one node $r\in V(T)$ is designated as the root. For any $t\in V(T)\setminus\{r\}$, let $e(t)$ be the unique edge on the path from $r$ to $t$. The adhesion of $t$, denoted $\text{adh}(t)$, is defined as $|\text{cut}(e(t))|$. For the root $r$, set $\text{adh}(r)=0$.
The torso of a tree-cut decomposition $(T,\mathcal{X})$ at a node $t$ is the graph $H_{t}$ obtained as follows: if $T$ has a single node, then $H_{t}=G$. Otherwise, for each component $T_{i}$ of $T-t$, define the vertex set $Z_{i}=\bigcup_{b\in V(T_{i})}X_{b}$. The torso $H_{t}$ is formed by contracting each $Z_{i}$ into a single vertex $z_{i}$, adding an edge between $z_{i}$ and any vertex $v\in V(G)\setminus Z_{i}$ that was adjacent to a vertex in $Z_{i}$. This may create parallel edges.
The 3-center of $(G,X)$ is obtained by repeatedly suppressing vertices in $V(G)\setminus X$ with degree at most 2. For a node $t$, the 3-center of its torso $H_{t}$ with respect to $X_{t}$ is denoted $\widetilde{H}_{t}$, and the torso-size is defined as $\text{tor}(t)=|\widetilde{H}_{t}|$.
The width of a tree-cut decomposition $(T,\mathcal{X})$ is defined as $\max_{t\in V(T)}\{\text{adh}(t),\text{tor}(t)\}$. The tree-cut width of $G$, denoted $\text{tcw}(G)$, is the minimum width over all tree-cut decompositions of $G$.
A.24.1 Related Concepts for tree-cut width
Several related concepts further explore the idea of tree-cut width:
  • Slim Tree-cut Width [555]: Slim tree-cut width is a graph width parameter that meets all structural and algorithmic criteria for an edge-cut-based analogue of tree-width, while being less restrictive than edge-cut width.
    • If a graph $G$ has bounded slim tree-cut width, then it also has bounded tree-cut width [555].
  • Edge-crossing Width and $\alpha$-Edge-crossing Width [266, 825]: These parameters are defined by the number of edges that cross a bag in a tree-cut decomposition.
  • Edge-cut Width [209]: Edge-cut width is an algorithmically driven analogue of tree-width that focuses on edge cuts.
    • If a graph $G$ has bounded feedback edge set number, then it also has bounded edge-cut width [266].
  • Edge-Tangle [508]: Edge-Tangle is another related concept that further refines the understanding of edge cuts in graph decompositions.
A.25 Resolution width
Resolution width [89] measures proof complexity using the existential pebble game in finite model theory.
Related concept is Linear Resolution width. This width is linear layout of Resolution width.
A.26 Twin-width
Twin-width [1256, 1260, 1258, 1259, 1257, 187] is a graph width parameter that measures the complexity of graphs by evaluating their ability to avoid a fixed minor. Intuitively, twin-width reflects the extent to which a graph resembles a cograph—a type of graph that can be reduced to a single vertex by repeatedly merging twin vertices, which are vertices sharing the same neighbors. This width includes classes such as planar graphs and graphs with bounded tree-width or clique-width.
Definition A.54.
(cf.[1256, 1260, 1258, 1259, 1257, 187] ) Let $G=(V,E,R)$ be a trigraph, where $V$ is the vertex set, $E$ is the set of black edges, and $R$ is the set of red edges. For any two vertices $u,v\in V$, the contraction of $u$ and $v$ in $G$, denoted $G/u,v=(V^{\prime},E^{\prime},R^{\prime})$, is defined as follows:
  • The vertex set is $V^{\prime}=(V\setminus\{u,v\})\cup\{w\}$, where $w$ is the new vertex representing the contracted pair $u,v$.
  • The edges of $G/u,v$ are defined such that:
    $wx\in E^{\prime}\quad\text{if and only if}\quad ux\in E\ \text{and}\ vx\in E,$
    $wx\notin E^{\prime}\cup R^{\prime}\quad\text{if and only if}\quad ux\notin E\cup R\ \text{and}\ vx\notin E\cup R,$
    $wx\in R^{\prime}\quad\text{otherwise}.$
In other words, contracting $u$ and $v$ results in red edges for any vertex $x$ that is not consistently adjacent to both $u$ and $v$. The trigraph $G/u,v$ is called a d-contraction if both $G$ and $G/u,v$ are $d$-trigraphs.
A trigraph $G$ on $n$ vertices is said to be d-collapsible if there exists a sequence of $d$-contractions that reduces $G$ to a single vertex. Formally, this means there is a sequence of trigraphs $G=G_{n},G_{n-1},\dots,G_{2},G_{1}$, where each $G_{i-1}$ is a contraction of $G_{i}$, and $G_{1}$ is a single vertex graph.
The minimum $d$ for which $G$ is $d$-collapsible is called the twin-width of $G$, denoted by $\text{tww}(G)$.
A.26.1 Hierarchy for Twin-width
We consider about hierarchy for Twin-width.
Theorem A.55.
The relationships between Twin-width and other graph parameters are as follows:
  • Every graph class of bounded Genus has bounded Twin-width [1154]. [^41] The genus of a graph is the minimum number of ”holes” in a surface required to embed the graph without edge crossings. It quantifies the graph’s topological complexity (cf.[1229, 1095]).
  • Every graph of bounded rank-width, also has bounded twin-width[189].
  • Every graph of bounded stack number or bounded queue number also has bounded twin-width[188]
  • Every graph class of bounded Distance to Planar, Sparse-Twin-width, or Feedback Edge Set has bounded Twin-width [1154]  [^42] The Feedback Edge Set of a graph is the smallest set of edges whose removal makes the graph acyclic, eliminating all cycles (cf.[22, 627])..
  • If a graph $G$ has bounded Clique-width, then it also has bounded Twin-width [536].
  • If a graph $G$ has bounded Twin-width, then it also has bounded monadically dependent [536].
Proof A.56.
Please refer to each reference.
A.26.2 Related concepts for Twin-width
Related concepts for Twin-width are the following:
  • Sparse Twin-width [537, 187]: A graph width parameter that generalizes results for classes with bounded linear clique-width and clique-width. If a graph $G$ has bounded Sparse Twin-width, then it also has bounded nowhere dense [536].
  • Signed Twin-width [549]: A graph width parameter specifically applied to CNF formulas, demonstrating that BWMC (Bounded Width Monotone Circuits) is fixed-parameter tractable.
A.27 Decomposition width
Decomposition width [808] is related to matroid theory and second-order logic.
A.28 Minor-matching hypertree width
Minor-matching hypertree width [1239] measures the complexity of a tree decomposition for graphs and hypergraphs, ensuring polynomially-sized independent sets in each decomposition bag.
A.29 Tree-clique width
Tree-clique width [80] extends the algorithmic benefits of tree-width to more structured and dense graphs.
A.30 Tree-distance-width
Tree distance width (TDW) [1233] measures the width from a distance perspective, restricted by tree-width. Related concepts are Rooted Tree-distance-width (RTDW) [1233] and Connected Tree-distance-width [844, 964].
Definition A.57.
[1233] A tree distance decomposition of a graph $G=(V,E)$ is a triple ($(\{X_{i}\}_{i\in I},T=(I,F),r)$ where:
  • $\bigcup_{i\in I}X_{i}=V(G)$ and $X_{i}\cap X_{j}=\emptyset$ for all $i\neq j$.
  • For each vertex $v\in V$, if $v\in X_{i}$, then the distance from $v$ to the root set $X_{r}$ in $G$ is equal to the distance from $i$ to the root $r$ in the tree $T$.
  • For each edge $\{v,w\}\in E$, there exist $i,j\in I$ such that $v\in X_{i}$, $w\in X_{j}$, and either $i=j$ or $\{i,j\}\in F$.
The node $r$ is called the root of the tree $T$, and $X_{r}$ is called the root set of the tree distance decomposition.
The width of a tree distance decomposition ($(\{X_{i}\}_{i\in I},T,r)$ is defined as $\max_{i\in I}|X_{i}|$. The tree distance width of a graph $G$ is the minimum width over all possible tree distance decompositions of $G$, denoted by $\text{TDW}(G)$.
A rooted tree distance decomposition of $G$ is a tree distance decomposition where the root set $X_{r}$ contains exactly one vertex (i.e., $|X_{r}|=1$). The rooted tree distance width of a graph $G$ is the minimum width over all rooted tree distance decompositions of $G$, denoted by $\text{RTDW}(G)$.
A.30.1 Hierarchy for Tree-distance-width
We consider about Hierarchy for Tree-distance-width
Theorem A.58.
The relationships between Tree-distance-width and other graph parameters are as follows:
[…]
  • $\text{TDW}(G)\leq\text{RTDW}(G)$  (RTDW: Rooted Tree-distance width).
  • $\text{TDW}(G)\leq\text{PDW}(G)\leq\text{RPDW}(G)$  (PDW: Path-distance-width and RPDW: Rooted Path Distance width).
Proof A.59.
Please refer to each reference.
A.31 Path-distance-width
Path distance width (PDW) [1233, 64, 1230] measures the width from a distance perspective, constrained by bandwidth. Related concepts include Rooted Path Distance width (RPDW) [1233], Connected Path-distance-width, and c-connected Path-distance-width [844, 964]. Distance width is particularly useful in problems like Graph Isomorphism, where the goal is to determine if two graphs are identical by checking if there is a one-to-one correspondence between their vertices and edges that preserves the connections (cf.[1028, 601, 831]).
Definition A.60.
[1233] A (rooted) path distance decomposition of a graph $G=(V,E)$ is defined similarly to the (rooted) tree distance decomposition and (rooted) tree distance width, but with the requirement that the tree $T$ is a path.
For simplicity, a (rooted) path distance decomposition is denoted by $(X_{1},X_{2},\dots,X_{t})$, where $X_{1}$ is the root set of the decomposition.
The corresponding graph parameters are the path distance width (PDW) and rooted path distance width (RPDW).
A.31.1 Hierarchy for Path-distance-width
We consider about Hierarchy for Path-distance-width.
Theorem A.61.
The relationships between Path-distance-width and other graph parameters are as follows:
  • $\text{bandwidth}(G)\leq 2\times\text{Path-distance-width}(G)-1$
  • If a graph $G$ has bounded Path distance width, then it also has bounded Tree distance width (trivially).
  • If a graph $G$ has bounded Rooted Path distance width, then it also has bounded Rooted Tree distance width (trivially).
Proof A.62.
Please refer to each reference.
A.32 Tree-partition-width
Tree-partition-width [1215] is the minimum width of a tree-partition of a graph. This width is also called 0-quasi-tree-partitions-width [839].
Definition A.63.
[1215] Let $G$ be a graph. A graph $H$ is called a partition of $G$ if the following conditions hold:
  1. 1.
    Each vertex of $H$ is a subset of the vertices of $G$, referred to as a bag.
  2. 2.
    Every vertex of $G$ belongs to exactly one bag in $H$.
  3. 3.
    Two bags $A$ and $B$ in $H$ are adjacent if and only if there is an edge in $G$ with one endpoint in $A$ and the other endpoint in $B$.
The width of a partition is defined as the maximum number of vertices in any bag.
Informally, the graph $H$ is obtained from a proper partition of $V(G)$ by identifying the vertices in each bag, deleting loops, and replacing parallel edges with a single edge.
If a forest $T$ is a partition of a graph $G$, then $T$ is called a tree-partition of $G$. The tree-partition width of $G$, denoted by $\text{tpw}(G)$, is the minimum width of a tree-partition of $G$.
A.32.1 Hierarchy for Tree-partition-width
We consider about Hierarchy for Tree-partition-width .
Theorem A.64.
The relationships between Tree-partition-width and other graph parameters are as follows:
[…]
  • 2×Tree partition $12\times\text{Tree partition width}(G)\geq\text{Tree-width}(G)+$ 1 [1215].
  • $tn(G)\leq 3\times\text{tpw}(G)$ (tn: track-number) [387].
  • $qn(G)\leq 3\times\text{tpw}(G)-1$ (qn: queue-number) [387].
Proof A.65.
Please refer to each reference.
Related concepts are Directed Tree partition width [696], which considers Tree partition width in Digraphs, and star-partition-width [1216]
A.33 Path-partition-width
Path-partition-width is the minimum width of a path-partition of a graph. This width is also called 0-quasi-path-partitions-width [839]. Related concepts are Directed Path partition width [696], which considers Path partition width in Digraphs.
A.33.1 Hierarchy for Path-partition-width
We consider about Hierarchy for Path-partition-width.
Theorem A.66.
  • For any graph $G$: $\frac{1}{2}(\text{bw}(G)+1)\leq\text{ppw}(G)\leq\text{bw}(G).$ ($bw$:Band-width)[391]
  • If a graph $G$ has bounded Path partition width, then it also has bounded Tree partition width (trivially).
Proof A.67.
Please refer to each reference.
A.34 CV-width
CV-width [969] is specific to CNFs and dominates the tree-width of the CNF incidence graph. CNFs, or Conjunctive Normal Forms, are Boolean formulas expressed as a conjunction (AND) of clauses, where each clause is a disjunction (OR). And linear CV width is the linear restriction of CV width.
A.35 Dominating-set-width
Dominating-set-width [899] measures the number of different dominating sets on a subgraph.  [^43] A dominating set in a graph is a subset of vertices such that every vertex in the graph is either in the dominating set or adjacent to a vertex in it [1194].
A.36 Point width
Point width [243] for hypergraphs provides a condition for the tractability of Max-CSPs, generalizing bounded MIM-width and $\beta$-acyclicity.  [^44] In graph algorithms, tractability refers to whether a problem can be solved efficiently, typically in polynomial time. Tractable problems have algorithms that run in reasonable time, while intractable problems are computationally difficult. And Max-CSPs (Maximum Constraint Satisfaction Problems) involve finding an assignment to variables that maximizes the number of satisfied constraints in a given set. These problems are NP-hard and widely studied(cf.[1042]).
A.37 Neighbourhood width
Neighbourhood width [611, 609] considers all neighborhoods of the vertices. Neighbourhood width is less than path-width plus one[611, 609].
Definition A.68.
[611, 609] A linear layout for a graph $G=(V,E)$ is a bijection $\varphi:V\rightarrow\{1,\dots,|V|\}$. We denote the set of all linear layouts for $G$ by $\Phi(G)$. For a given $\varphi\in\Phi(G)$ and $1\leq i\leq|V|$, we define the left and right sets as follows:
$L(i,\varphi,G)=\{u\in V\mid\varphi(u)\leq i\}\quad\text{and}\quad R(i,\varphi,G)=\{u\in V\mid\varphi(u)>i\}.$
Let $G=(V,E)$ be a graph, and let $U,W\subseteq V$ be two disjoint vertex sets. The restricted neighborhood of a vertex $u$ in $W$ is denoted by $N_{W}(u)=\{v\in W\mid\{u,v\}\in E\}$. The set of all neighborhoods of the vertices in $U$ with respect to $W$ is denoted by $N(U,W)=\{N_{W}(u)\mid u\in U\}$.
This allows us to define the neighborhood-width of a graph $G$, denoted by $\text{nw}(G)$, as follows:
$\text{nw}(G)=\min_{\varphi\in\Phi(G)}\max_{1\leq i\leq|V|-1}|N(L(i,\varphi,G),R(i,\varphi,G))|.$
A.38 Fusion-width
Fusion-width [532] generalizes tree-width and encompasses graphs of bounded tree-width and clique-width.
Definition A.69.
[532] A k-fusion-tree expression is an expression constructed from the following elements:
  • $i^{(m)}$, which creates a graph consisting of $m$ isolated vertices labeled $i$, where $i$ is an element of the set $\{1,\dots,k\}$.
  • $\eta_{i,j}$, a unary operation that creates an edge between every vertex labeled $i$ and every vertex labeled $j$ for $i\neq j$.
  • $\rho_{i\rightarrow j}$, a unary operation that changes all labels $i$ to $j$.
  • $\theta_{i}$, a unary operation that merges all vertices labeled $i$ into a single vertex. The new vertex is labeled $i$ and is adjacent to every vertex not labeled $i$ that was adjacent to any vertex labeled $i$ before the operation.
  • $\oplus$, a binary operation that forms the disjoint union of two graphs.
Finally, the generated graph is obtained by deleting all labels.
Definition A.70.
[532] The fusion-width $\text{fw}(G)$ of a graph $G$ is the smallest $k$ such that there exists a $k$-fusion-tree expression that generates $G$.
A.38.1 Hierarchy for Fusion-width
We consider about Hierarchy for Fusion-width.
Theorem A.71.
The relationships between Fusion-width and other graph parameters are as follows:
  • Fusion-width(G) $\leq$ Tree-width(G) + 2 [532],
  • Fusion-width(G) $\leq$ Clique-width(G) [532].
A.39 Directed NLC width
Directed NLC width [618] is NLC width on directed graphs.
Definition A.72 (Directed NLC-width).
[618] Let $k$ be a positive integer. And the new operation $\otimes(\overrightarrow{S},\overleftarrow{S})$ utilizes two relations $\overrightarrow{S}$ and $\overleftarrow{S}$ for label pairs, where $\overrightarrow{S}$ defines the arcs directed from left to right and $\overleftarrow{S}$ defines the arcs directed from right to left. The class $\text{dNLC}_{k}$ of labeled digraphs is defined recursively as follows:
  1. 1.
    The single vertex digraph $\bullet_{a}$ for some $a\in[k]$ is in $\text{dNLC}_{k}$.
  2. 2.
    Let $G=(V_{G},E_{G},\text{lab}_{G})\in\text{dNLC}_{k}$ and $J=(V_{J},E_{J},\text{lab}_{J})\in\text{dNLC}_{k}$ be two vertex-disjoint labeled digraphs, and let $\overrightarrow{S},\overleftarrow{S}\subseteq[k]^{2}$ be two relations. Then the operation $G\otimes(\overrightarrow{S},\overleftarrow{S})J:=(V^{\prime},E^{\prime},\text{lab}^{\prime})$ is defined by:
    $V^{\prime}:=V_{G}\cup V_{J},$
    $E^{\prime}:=$ $\,E_{G}\cup E_{J}\cup\{(u,v)\mid u\in V_{G},v\in V_{J},(\text{lab}_{G}(u),\text{lab}_{J}(v))\in\overrightarrow{S}\}$
    $\,\cup\{(v,u)\mid u\in V_{G},v\in V_{J},(\text{lab}_{G}(u),\text{lab}_{J}(v))\in\overleftarrow{S}\},$
    $\text{lab}^{\prime}(u):=\begin{cases}\text{lab}_{G}(u)&\text{if }u\in V_{G},\\ \text{lab}_{J}(u)&\text{if }u\in V_{J},\end{cases}$
    for every $u\in V^{\prime}$, which results in a digraph in $\text{dNLC}_{k}$.
  3. 3.
    Let $G=(V_{G},E_{G},\text{lab}_{G})\in\text{dNLC}_{k}$ and $R:[k]\rightarrow[k]$ be a function. Then the operation $\circ_{R}(G):=(V_{G},E_{G},\text{lab}^{\prime})$ is defined by:
    $\text{lab}^{\prime}(u):=R(\text{lab}_{G}(u))$
    for every $u\in V_{G}$, resulting in a digraph in $\text{dNLC}_{k}$.
[…]
A.40 Directed Tree-width
Directed Tree-width [711, 12] is Tree-width on directed graphs. Examples of obstructions are directed tangle [572] and directed ultrafilter [523].
The definition of Directed Tree-width is provided below. Note that various papers have proposed and examined alternative definitions of Directed Tree-width.
Definition A.73 (Directed Tree-width, [711, 12] ).
A (arboreal) tree-decomposition of a digraph $G=(V_{G},E_{G})$ is a triple $(T,X,W)$, where:
  • $T=(V_{T},E_{T})$ is an out-tree,
  • $X=\{X_{e}\mid e\in E_{T}\}$ and $W=\{W_{r}\mid r\in V_{T}\}$ are sets of subsets of $V_{G}$,
such that the following conditions hold:
  1. (dtw-1)
    $W=\{W_{r}\mid r\in V_{T}\}$ is a partition of $V_{G}$ into non-empty subsets.
  2. (dtw-2)
    For every $(u,v)\in E_{T}$, the set $\bigcup_{r\in V_{T},v\leq r}W_{r}$ is $X_{(u,v)}$-normal.
The width of an arboreal tree-decomposition $(T,X,W)$ is defined as:
$\max_{r\in V_{T}}\left\{|W_{r}\cup\bigcup_{e\sim r}X_{e}|\right\}-1,$
where $e\sim r$ means that $r$ is one of the endpoints of the arc $e$.
[…]
A similar concept to tree-width is DAG-width, which we define below [130]. If a graph has entanglement $k$, its DAG-width is at most $k+1$[130].
Definition A.74.
[130] Let $G=(V,E)$ be a graph. A set $W\subseteq V$ guards a set $V^{\prime}\subseteq V$ if, whenever there is an edge $(u,v)\in E$ such that $u\in V^{\prime}$ and $v\notin V^{\prime}$, then $v\in W$.
Definition A.75 (DAG-decomposition).
[130] Let $G=(V,E)$ be a directed graph. A DAG-decomposition is a tuple $D=(D,\{X_{d}\}_{d\in V(D)})$ such that:
  • (D1)
    $D$ is a directed acyclic graph (DAG).
  • (D2)
    $\bigcup_{d\in V(D)}X_{d}=V$.
  • (D3)
    For all $d\preceq d^{\prime}\preceq d^{\prime\prime}$ in $D$, we have $X_{d}\cap X_{d^{\prime\prime}}\subseteq X_{d^{\prime}}$.
  • (D4)
    For a root $d$, $X_{d}$ is guarded by $\emptyset$.
  • (D5)
    For all $(d,d^{\prime})\in E(D)$, $X_{d}\cap X_{d^{\prime}}$ guards $X_{d^{\prime}}\setminus X_{d}$, where $X_{d^{\prime}}:=\bigcup_{d^{\prime}\preceq d^{\prime\prime}}X_{d^{\prime\prime}}$.
The width of a DAG-decomposition $D$ is defined as $\max\{|X_{d}|:d\in V(D)\}$. The DAG-width of a graph is the minimal width over all possible DAG-decompositions of $G$.
A.40.1 Hierarchy for Directed Tree-width
We consider about Hierarchy for Directed Tree-width .
Theorem A.76.
The relationships between Directed Tree-width and other graph parameters are as follows:
  • If a digraph $G$ has bounded DAG-width, then it also has bounded Directed Tree-width [129]  [^45] DAG-width measures how close a directed graph is to being acyclic. It is defined as the minimal width of a DAG-decomposition, which is similar in concept to tree-decomposition in undirected graphs[129, 957, 130, 647]. DAG-width is also used in Parity game. Related parameters of parity game are alternation-depth, nesting depth [756, 755]..
  • If a digraph $G$ has bounded Kelly-width, then it also has bounded Directed Tree-width [129, 71].
  • For every digraph $D$, it holds that $\text{dtw}(D)\geq\text{cycle-degeneracy}(D)$. [^46] Cycle-degeneracy (or c-degeneracy) of a directed graph is the minimum number of vertices that must be removed to eliminate all directed cycles containing any given vertex[953].
  • For every digraph $G$, $1/3(dtw(G)+2)\leq\textit{entanglement}(G)$[712].
Proof A.77.
Please refer to each reference.
A.41 Directed path-width
Directed Path-width (d-pw) [109, 615] is Path-width on directed graphs.
Definition A.78 (Directed Path-width).
[109, 615] Let $G=(V,E)$ be a digraph. A directed path-decomposition of $G$ is a sequence $(X_{1},\dots,X_{r})$ of subsets of $V$, called bags, that satisfies the following conditions:
  1. 1.
    $X_{1}\cup\dots\cup X_{r}=V$,
  2. 2.
    For each edge $(u,v)\in E$, there exists a pair $i\leq j$ such that $u\in X_{i}$ and $v\in X_{j}$,
  3. 3.
    For all indices $1\leq i<j<\ell\leq r$, it holds that $X_{i}\cap X_{\ell}\subseteq X_{j}$.
The width of a directed path-decomposition $X=(X_{1},\dots,X_{r})$ is defined as:
$\max_{1\leq i\leq r}|X_{i}|-1.$
[…]
A.41.1 Hierarchy for Directed Path-width
We consider about Hierarchy for Directed Path-width .
Theorem A.79.
[486] Let $D$ be a directed graph. The Directed Path-width $\text{dpw}(D)$ has the following relationships with various other graph parameters:
  1. 1.
    Directed Vertex Separation Number: The Directed Path-width is equal to the Directed Vertex Separation Number [161].  [^47] The Directed Vertex Separation Number of a digraph measures the smallest number of vertices needed to separate the remaining vertices into disjoint sets, ensuring no directed path crosses between these sets[161, 171].
    $\text{dpw}(D)=\text{Directed Vertex Separation Number}(D)$
  2. 2.
    Directed feedback vertex set number: If $G$ has bounded Directed feedback vertex set number, then $G$ has bounded Directed Path-width [950].
  3. 3.
    Kelly-width: The Kelly-width is less than or equal to the Directed Path-width plus one [550].
[…]
  4. 4.
    Entanglement: The entanglement of $D$ is less than or equal to the Directed Path-width [1009] . [^48] Entanglement of a graph measures the complexity of its cycles, defined by the minimum number of detectives required to catch a thief in a graph-based game[247, 132].
    $\text{entanglement}(D)\leq\text{dpw}(D)$
  5. 5.
    Cycle Rank: The Directed Path-width is less than or equal to the Cycle Rank of $D$ [604].  [^49] Cycle rank is a measure of the complexity of a connected graph, defined as the number of independent cycles in the graph.
    $\text{dpw}(D)\leq\text{Cycle Rank}(D)$
  6. 6.
    Kelly Path-width: The Kelly Path-width of $D$ is equal to its Directed Path-width [766].
[…]
  7. 7.
    DAG-path-width: The Directed Path-width is less than or equal to the DAG-path-width [742].
    $\text{dpw}(D)\leq\text{DAG-path-width}(D)$
  8. 8.
    Diblockage: A directed graph has Directed Path-width $\geq k-1$ if and only if it has a diblockage of order $\geq k$ [429].  [^50] Diblockage is directed version of blockage.
    $\text{dpw}(D)\geq k-1\iff\text{diblockage order}(D)\geq k$
  9. 9.
    Directed Path-Distance-Width: The Directed Path-width is bounded above by twice the Directed Path-Distance-Width minus one [486].
    $\text{dpw}(D)\leq 2\times\text{DPDW}(D)-1$
  10. 10.
    Directed Proper-Path-Width: The Directed Proper-Path-Width is bounded by the Directed Path-width and is at most one greater [486].
    $\text{dpw}(D)\leq\text{dppw}(D)\leq\text{dpw}(D)+1$
  11. 11.
    Directed Tree-width: The Directed Path-width is greater than or equal to the Directed Tree-width [613].
    $\text{dtw}(D)\leq\text{dpw}(D)$
  12. 12.
    Directed Cut-width: The Directed Path-width is less than or equal to the Directed Cut-width [613].
    $\text{dpw}(D)\leq\text{dcw}(D)$
  13. 13.
    Linear Width Parameters: For any directed graph $G$, the Directed Path-width is related to various linear width parameters as follows [613]:
    $\text{dpw}(G)\leq\min(\Delta^{-}(G),\Delta^{+}(G))\cdot\text{d-nw}(G),$
    $\text{dpw}(G)\leq\min(\Delta^{-}(G),\Delta^{+}(G))\cdot\text{d-lnlcw}(G),$
    $\text{dpw}(G)\leq\min(\Delta^{-}(G),\Delta^{+}(G))\cdot\text{d-lcw}(G),$
[…]
Proof A.80.
Please refer to each reference.
A.42 Directed Branch-width
Directed Branch-width [232] is Branch-width on directed graphs.
A.43 Directed Cut-width
[…]
Definition A.81.
[287] The directed cut-width of a digraph $G=(V,E)$ is defined as:
[…]
where $\Phi(G)$ denotes the set of all linear layouts of $G$, $L(i,\varphi,G)$ represents the set of vertices mapped to positions $\leq i$ by the linear layout $\varphi$, and $R(i,\varphi,G)$ represents the set of vertices mapped to positions $>i$.
A.44 Directed Clique-width
Directed Clique-width [303, 618] is clique width on directed graphs. And Directed Linear-Clique-width (d-lcw) [618] is linear-concept of clique width on directed graphs. We introduce about the definitions of directed clique-width and directed linear-clique-width.
Definition A.82.
[618] Directed Clique-width. Let $k$ be a positive integer. The class $\text{dCW}_{k}$ of labeled digraphs is defined recursively as follows:
  1. 1.
    A single vertex digraph $\bullet_{a}$ for some $a\in[k]$ is in $\text{dCW}_{k}$.
  2. 2.
    Let $G=(V_{G},E_{G},\text{lab}_{G})\in\text{dCW}_{k}$ and $J=(V_{J},E_{J},\text{lab}_{J})\in\text{dCW}_{k}$ be two vertex-disjoint labeled digraphs. Then the disjoint union $G\oplus J:=(V^{\prime},E^{\prime},\text{lab}^{\prime})$ defined by:
    $V^{\prime}:=V_{G}\cup V_{J},$
    $E^{\prime}:=E_{G}\cup E_{J},$
    $\text{lab}^{\prime}(u):=\begin{cases}\text{lab}_{G}(u)&\text{if }u\in V_{G},\\ \text{lab}_{J}(u)&\text{if }u\in V_{J},\end{cases}$
    for every $u\in V^{\prime}$, is in $\text{dCW}_{k}$.
  3. 3.
    Let $a,b\in[k]$ be two distinct integers, and let $G=(V_{G},E_{G},\text{lab}_{G})\in\text{dCW}_{k}$ be a labeled digraph.
    1. (a)
      The operation $\rho_{a\rightarrow b}(G):=(V_{G},E_{G},\text{lab}^{\prime})$ defined by:
      $\text{lab}^{\prime}(u):=\begin{cases}\text{lab}_{G}(u)&\text{if }\text{lab}_{G}(u)\neq a,\\ b&\text{if }\text{lab}_{G}(u)=a,\end{cases}$
      for every $u\in V_{G}$, results in a digraph in $\text{dCW}_{k}$.
    2. (b)
      The operation $\alpha_{a,b}(G):=(V_{G},E^{\prime},\text{lab}_{G})$ defined by:
      $E^{\prime}:=E_{G}\cup\{(u,v)\mid u,v\in V_{G},u\neq v,\text{lab}(u)=a,\text{lab}(v)=b\},$
      results in a digraph in $\text{dCW}_{k}$.
[…]
Definition A.83 (Directed Linear Clique-width).
[…]
  1. 1.
    Creation of a new vertex with label $a$ (denoted by $\bullet_{a}$).
  2. 2.
    Disjoint union of a labeled digraph $G$ and a single vertex labeled $a$ (denoted by $G\oplus\bullet_{a}$).
  3. 3.
    Inserting an arc from every vertex with label $a$ to every vertex with label $b$ (where $a\neq b$, denoted by $\alpha_{a,b}$).
  4. 4.
    Changing label $a$ to label $b$ (denoted by $\rho_{a\rightarrow b}$).
A.45 Directed Linear-NLC-width
Directed Linear NLC-width (d-lnlcw) [618, 612] is Linear NLC width on directed graphs.
A.45.1 Hierarchy for Linear-NLC-width
We consider about Hierarchy for Linear-NLC-width.
Theorem A.84.
The relationships between Directed Linear NLC-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Directed Linear NLC-width, then it also has bounded Directed NLC-width (trivially).
  • For every digraph $G$, we have d-lnlcw($G$) $\leq$ d-lcw($G$) $\leq$ d-lnlcw($G$) + 1 (Directed Linear-Clique-width) [612].
A.46 Directed Neighbourhood width
Directed Neighbourhood width (d-nw) [612] is Neighbourhood width on directed graphs.
Definition A.85.
[612] Let $G=(V,E)$ be a digraph, and let $U,W\subseteq V$ be two disjoint vertex sets. The set of all out-neighbors of $u$ into the set $W$ is defined by $N^{+}_{W}(u)=\{v\in W\mid(u,v)\in E\}$, and the set of all in-neighbors of $u$ from the set $W$ is defined by $N^{-}_{W}(u)=\{v\in W\mid(v,u)\in E\}$. The directed neighborhood of a vertex $u$ into the set $W$ is given by $N_{W}(u)=(N^{+}_{W}(u),N^{-}_{W}(u))$. The set of all directed neighborhoods of the vertices in the set $U$ into the set $W$ is denoted by $N(U,W)=\{N_{W}(u)\mid u\in U\}$.
For a layout $\varphi\in\Phi(G)$, we define
[…]
The directed neighborhood-width of a digraph $G$ is given by
[…]
A.46.1 Hierarchy for Neighbourhood width
We consider about Hierarchy for Neighbourhood width.
Theorem A.86.
For every digraph $G$, we have:
  • d-nw($G$) $\leq$ d-lnlcw($G$) $\leq$ d-nw($G$) + 1 (Directed Linear NLC-width) [612],
  • d-nw($G$) $\leq$ d-lcw($G$) $\leq$ d-nw($G$) + 1 (Directed Linear-Clique-width) [612].
Proof A.87.
Please refer to each reference.
A.47 Directed Rank-width
Directed Rank-width [730] is Rank-width on directed graphs.
A.48 Directed Linear-Rank-width
Directed Linear Rank-width (d-lrw) [612] is Linear rank width on directed graphs.
Definition A.88.
[612] Let $G=(V,E)$ be a digraph, and let $V_{1},V_{2}\subseteq V$ be a disjoint partition of the vertex set of $G$. Define the adjacency matrix $M^{V_{1}}_{V_{2}}=(m_{ij})$ over the four-element field $\mathbb{GF}(4)$ for the partition $V_{1}\cup V_{2}$, where
$m_{ij}=\begin{cases}0&\text{if }(v_{i},v_{j})\notin E\text{ and }(v_{j},v_{i})\notin E,\\ a&\text{if }(v_{i},v_{j})\in E\text{ and }(v_{j},v_{i})\notin E,\\ a^{2}&\text{if }(v_{i},v_{j})\notin E\text{ and }(v_{j},v_{i})\in E,\\ 1&\text{if }(v_{i},v_{j})\in E\text{ and }(v_{j},v_{i})\in E.\end{cases}$
In $\mathbb{GF}(4)$, the field elements are $\{0,1,a,a^{2}\}$, with the properties $01+a+a^{2}=$ 0 and $a^{3}=1$.
[…]
A.48.1 Hierarchy for Directed Linear-Rank-width
We consider about Hierarchy for Directed Linear-Rank-width.
Theorem A.89.
The relationships between Directed Linear Rank-width and other graph parameters are as follows:
  • If a graph $G$ has bounded Directed Linear Rank-width, then it also has bounded Directed Rank-width (trivially).
  • For every digraph $G$, we have d-lrw($G$) $\leq$ d-nw($G$) [612] (d-nw: Directed Neighbourhood width).
Proof A.90.
Please refer to each reference.
A.49 Elimination width
Elimination width os a graph width parameter to address issues with acyclic digraphs [450].
A.50 Flip-width
The flip-width [185] parameters are defined using variants of the Cops and Robber game. Related concepts include radius-one flip-width [1153].
A.50.1 Hierarchy for Flip-width
We consider about Hierarchy for Flip-width.
Theorem A.91.
The relationships between Flip-width (or radius-one flip-width) and other graph parameters are as follows:
  • Every class of bounded twin-width has bounded flip-width [1153].
  • If a graph $G$ has bounded radius-one flip-width, then it also has bounded symmetric difference, functionality, and VC-dimension [1153].
Proof A.92.
Please refer to each reference.
A.51 Stretch-width
Stretch-width [1255] lies strictly between clique-width and twin-width.
A.51.1 Hierarchy for Stretch-width
We consider about Hierarchy for Stretch-width.
Theorem A.93.
The relationships between Stretch-width and other graph parameters are as follows:
  • The stretch-width of any graph is at most twice its clique-width[1255].
  • There is a constant $c$ such that for every graph $G$, $\text{tw}(G)\leq c\Delta(G)^{4}\text{stw}(G)^{2}\log|V(G)|$. (Tree-width and maximal degree)
Proof A.94.
Please refer to each reference.
A.52 Cops-width
The Cop-width [185, 659, 3] parameters are defined using variants of the Cops and Robber game.
A.52.1 Hierarchy for cops-width
We consider about Hierarchy for cops-width.
Theorem A.95.
The relationships between (radius-)cop-width and other graph parameters are as follows:
[…]
Proof A.96.
Please refer to each reference.
A.52.2 Related concepts for cops-width
Related concepts are the following:
  • Marshal width and Monotone marshal width [10]: Graph width parameters related to a winning strategy in the Cops and Robber game.
  • Game-width [1071, 423]: A graph width parameter of the Cops and Robber game. The game width of a graph equals its tree-width plus 1 [1071].
A.53 Perfect matching width
Perfect matching width [574] is a width parameter for matching covered graphs based on a branch decomposition. 515151 A perfect matching in a graph is a set of edges in which every vertex is paired with exactly one other vertex. A matching covered graph is a connected graph where every edge is part of at least one perfect matching, meaning all edges are ”admissible.” These concepts are central to matching theory, a topic extensively studied in graph theory research [1127, 648, 715, 249, 1248].
Definition A.97 (Matching-Porosity).
[646] Let $G$ be a matching covered graph and $X\subseteq V(G)$. We define the matching-porosity of $\partial(X)$ as follows:
$\text{mp}(\partial(X)):=\max_{M\in\mathcal{M}(G)}|M\cap\partial(X)|.$
A perfect matching $M\in\mathcal{M}(G)$ is maximal with respect to a cut $\partial(X)$ if there is no perfect matching $M^{\prime}\in\mathcal{M}(G)$ such that $\partial(X)\cap M\subset\partial(X)\cap M^{\prime}$.
A perfect matching $M\in\mathcal{M}(G)$ maximizes a cut $\partial(X)$ if $\text{mp}(\partial(X))=|M\cap\partial(X)|$.
Definition A.98 (Perfect Matching Width).
[646] Let $G$ be a matching covered graph. A perfect matching decomposition of $G$ is a tuple $(T,\delta)$, where $T$ is a cubic tree and $\delta:L(T)\to V(G)$ is a bijection. The width of $(T,\delta)$ is given by $\max_{e\in E(T)}\text{mp}(e)$ and the perfect matching width of $G$ is then defined as
$\text{pmw}(G):=\min_{\begin{subarray}{c}(T,\delta)\\ \text{perfect matching}\\ \text{decomposition of }G\end{subarray}}\max_{e\in E(T)}\text{mp}(\partial(e)).$
A.54 Bisection-width
The Bisection problem seeks to partition the vertices of a graph into two equally sized sets while minimizing the cut size. The Bisection Width is a width parameter closely related to this Bisection problem [135, $1$, 137]. If a graph $G$ has bounded Band-width, then it also has bounded bisection-width [1154].
Definition A.99.
[137]. A bisection $\xi$ of an undirected graph $G=(V,E)$, where $|V|=n$, is a partition $V=V_{0}\cup V_{1}$ such that $|V_{0}|,|V_{1}|\in\left\{\left\lfloor\frac{n}{2}\right\rfloor,\left\lceil\frac{n}{2}\right\lceil\right\}$. For simplicity, we assume that $n$ is even throughout this paper. The number of edges between $V_{0}$ and $V_{1}$ for a given bisection $\xi$ is called the cut size of $\xi$. The minimum cut size among all possible bisections of $G$ is called the bisection width of $G$ and is denoted by $\text{bw}(G)$. Formally, this is defined as
$\text{bw}(G):=\min\{|\{\{v,w\}\in E\mid\xi(v)\neq\xi(w)\}|\mid\xi\text{ is a bisection of }G\}.$
A.55 Maximum matching width
Maximum matching width is a graph width parameter used to consider fast algorithms for the dominating set problem [704].
Definition A.100.
A rooted branch decomposition is a branch decomposition $(T,\delta)$ where an edge of $T$ is subdivided to create a root $r$. For an internal vertex $v\in V(T)$, $\delta(v)$ is the union of $\delta(l)$ for all leaves $l$ that have $v$ as an ancestor.
Given a symmetric function $f:2^{X}\to\mathbb{R}$, the $f$-value of an edge $e$ in $(T,\delta)$ is $f(A)=f(B)$, where $A$ and $B$ are the parts of the partition induced by $e$. The $f$-width of $(T,\delta)$ is the maximum $f$-value over all edges of $T$, denoted by $f(T,\delta)$. The $f$-width of $X$ is the minimum $f$-width over all branch decompositions of $X$. If $|X|\leq 1$, the $f$-width is defined as $f(\emptyset)$.
For a graph $G$ and a subset $S\subseteq E(G)$, the branchwidth $\text{bw}(G)$ is the $f$-width of $E(G)$ for $f(S)$, defined as the number of vertices incident with edges in both $S$ and $E(G)\setminus S$.
The maximum matching-width of a graph $G$, denoted by $\text{mmw}(G)$, is defined as the $f$-width of $V(G)$ for $f=\text{mm}$, where $\text{mm}(S)$ is the size of the maximum matching in $G[S,V(G)\setminus S]$.
A.55.1 Hierarchy for Maximum matching width
We consider about Hierarchy for Maximum matching width.
Theorem A.101.
The relationships between Maximum matching width and other graph parameters are as follows:
  • If a graph $G$ has bounded Maximum matching width, then it also has bounded tree-depth [1058].
  • If a graph $G$ has bounded Maximum matching width, then it also has bounded Maximum induced matching-width [1058].
  • For any graph $G$, we have $\text{mmw}(G)\leq\text{branch-width}(G)\leq\text{tree-width}(G)+1\leq 3\cdot\text{mmw}(G).$
Proof A.102.
Please refer to each reference.
A.56 Linear Maximum Matching width
Linear Maximum Matching width is the linear restriction of Maximum Matching width.
A.56.1 Hierarchy for Linear-Maximum matching width
We consider about Hierarchy for Linear-Maximum matching width.
Theorem A.103.
The relationships between Linear MIM-width and other graph parameters are as follows:
  • $\text{lmmw}(G)\leq\text{pw}(G)\leq 2\times\text{lmmw}(G)$. (Path width) [954].
  • If a graph $G$ has bounded Linear MIM-width, then it also has bounded MIM-width (trivially).
Proof A.104.
Please refer to each reference.
A.57 Clustering-width
Clustering-width is graph width parameter of the smallest number of variables whose deletion results in a variable-disjoint union of hitting formulas[952].
A.58 Query-width
Query-width is a graph width parameter that generalizes both acyclicity and tree-width, measuring the complexity of hypergraphs through restricted hypertree-width[178, 461, 585, 270].
A.59 Universal width
Universal width [749, 637] is a width parameter for an alternating finite automaton (AFA). Note that an alternating finite automaton (AFA) is a computational model where states can make simultaneous existential or universal transitions, allowing parallel computation paths.
A.59.1 Related concepts for Universal width
Related concepts are the following:
  • Maximal universal width [749]: Maximal Graph width parameter of universal width.
  • Combined width [752]: Width parameter for an alternating finite automaton (AFA).
  • Maximal combined width [752]: Maximal Graph width parameter of combined width.
  • Split-width [23, 309]: Width parameter for an alternating finite automaton (AFA). Also Split-width is a graph parameter used to measure the complexity of the behavior graphs of computational models, aiding in decidability results for systems like multi-pushdown systems [308, 307, 24].
A.60 Degree-width
Degree-width [315] is graph width parameter for tournaments that we can be seen as a measure of how far is the tournament from being acyclic.
A.61 Median-width
Median-width [1118, 1119] is a graph width parameter used for median graphs (A median graph is an undirected graph where any three vertices have a unique median vertex that lies on the shortest paths between each pair [781, 772].) as the underlying graph of the decomposition. For any graph $G$, Median-width(G) $\leq$ Tree-width(G) + 1 [1119].
The definition of a median graph is provided below for reference.
Definition A.105.
A median graph is an undirected graph $G=(V,E)$ such that, for any three vertices $a,b,c\in V$, there exists a unique vertex $m(a,b,c)\in V$ (called the median) that lies on the shortest paths between each pair of the three vertices. That is, $m(a,b,c)$ belongs to the shortest paths between $a$ and $b$, $b$ and $c$, and $a$ and $c$.
Formally, a vertex $m(a,b,c)$ satisfies:
$d_{G}(a,m(a,b,c))+d_{G}(m(a,b,c),b)=d_{G}(a,b),$
$d_{G}(b,m(a,b,c))+d_{G}(m(a,b,c),c)=d_{G}(b,c),$
$d_{G}(a,m(a,b,c))+d_{G}(m(a,b,c),c)=d_{G}(a,c),$
where $d_{G}(x,y)$ denotes the shortest path distance between vertices $x$ and $y$ in $G$.
A.61.1 Related concepts for Median-width
Lattice-width [1119] refers to a parameter that constrains the median graph of a decomposition to be isometrically embeddable into the Cartesian product of $i$ paths.
A.62 Directed Modular-width
Directed Modular-width [1120] is Modular-width on directed graphs.
A.63 Cycle-width
Cycle-width [646, 1203] measures graph complexity, focusing on cycle structures and highlighting differences between directed and undirected tree-width by emphasizing local graph properties.
Question A.106.
Is there a corresponding width parameter for Walk-width and Circuit-width? If not, can it be defined?
A.63.1 Related concepts for cycle-width
Related concepts is following.
  • radial cycle-width [49]: Radial cycle-width measures the smallest radial width in a cycle-decomposition of a graph, where the radial width is the largest radius of the cycle’s parts.
Question A.107.
Is it possible to define Directed Cycle-Width by considering directed cycles?
Question A.108.
Can Cycle-width be extended to Fuzzy Graphs or similar structures?
A.63.2 Hierarchy for Cycle-width
We consider about Hierarchy for Cycle-width.
Theorem A.109.
The relationships between cycle-width and other graph parameters are as follows:
  • The cycle-width of a graph $G$ is at most $k-1$ if and only if its arc thickness is at most $k$ [1203].
  • For any graph $G$, $\frac{1}{2}\cdot\text{pathwidth}(G)\leq\text{cyclewidth}(G)\leq\text{pathwidth}(G)$[1203].
  • Let $T$ be a tree. Then $\text{cyclewidth}(T)=\text{pathwidth}(T)$[1203].
Proof A.110.
Please refer to each reference.
A.64 Arc-width
Arc-width [110, 83] is the maximum number of arcs.
Definition A.111.
[110, 83] The arc-representation of a graph $G$ is a mapping $\phi$ from the vertex set $V(G)$ to the set of arcs on a base circle, such that adjacent vertices of $G$ are mapped to intersecting arcs. The width (in a representation) of a point $P$ on the base circle is the number of representing arcs containing $P$. The width of $\phi$ is the maximum width of the points on the base circle. The arc-width of a graph $G$ is the minimal possible width of such arc-representations, denoted as $\text{aw}(G)$.
A.64.1 Related concepts for arc-width
Vortex-width [1035, 1146] measures how closely a graph can be "nearly drawn" on a surface with respect to a given cyclic ordering of vertices. Vortex-width is equal to Arc-width.
Question A.112.
Is it possible to define Directed Arc-Width by considering directed arcs?
Question A.113.
Can Arc-width be extended to Fuzzy Graphs or similar structures?
A.65 Match width and braid width
Match width and braid width [235] are parameters used to measure structural complexity in XML documents; they help determine XPath satisfiability based on document depth.
A.66 Clique cover width
The clique cover width of $G$, denoted by $ccw(G)$, is the minimum value of the bandwidth of all graphs that are obtained by contracting the cliques in a clique cover of G into a single vertex [1076, 1077, 1075].
A.67 Questionable-width
Questionable-width [856, 857] relates to graph width measurements concerning questionable-width. A “question” identifies the initial divergence between two sequences of elements from orders or sets, indexed ordinally. In a questionable representation with finite width of an order $O$, comparisons are resolved by examining the “question” that evaluates elements within the finite order $O$ [856, 857].
A.68 Plane-width
The plane-width of a graph is defined as the minimum diameter of the image of the graph’s vertex set [723, 723].
A.69 Scan-width
Scan-width[670, 669] is a width parameter for directed acyclic graphs (DAGs) that measures their tree-likeness by considering the direction of arcs(cf.[1069]).
A.70 V-width
V-width is a complexity measure in graph structures that is smaller than, and potentially much smaller than, the maximum clique size and the maximum number of parents per node in a DAG[283].
A.71 Cross-width
Cross-width is a graph parameter used in kernelization, particularly in analyzing the complexity of reducing problems, often related to polynomial equivalence relations[695].
A.72 Guidance-width
Guidance-width measures the minimal number of colors needed to color a guidance system in a tree, ensuring that conflicting guides are distinctly colored [179].
A.73 Star-width
Star-width measures the complexity of a graph’s decomposition into "stars" (graphs with a shared central vertex) where the width is determined by the size of the interfaces connecting the stars[158, 1172, 1171]. Related parameters are radial star-width[49], costar-width[158], and atomic star-width[158]. If a graph has bounded starwidth, it has bounded tree-depth[1171].
Definition A.114.
[1172, 1171] A star decomposition of a graph $G=(V,E)$ is a star with a core $X_{0}$ and leaves $X_{1},\dots,X_{m}$, where each $X_{i}$ is associated with a subset of $V$, satisfying the following properties:
  1. 1.
    The union of all sets $X_{i}$ equals $V$.
  2. 2.
    For every edge $(v,w)\in E$, there exists a subset $X_{i}$ that contains both $v$ and $w$.
  3. 3.
    For $i\neq j$, if $X_{i}$ and $X_{j}$ both contain a vertex $v$, then $X_{0}$ contains $v$ as well.
The width of a star decomposition is defined as $\max_{i}|X_{i}|-1$. The starwidth $sw(G)$ of a graph $G$ is the minimum width among all possible star decompositions of $G$.
Question A.115.
Is it possible to define Directed Star-Width?
Question A.116.
Can Star-width be extended to Fuzzy Graphs or similar structures?
Appendix B Various Length Parameters
In this section, we consider about various length parameters.
B.1 Tree-length and Branch-length
We begin by discussing Tree-length and other well-known parameters. Tree-length measures the maximum distance between any two vertices within a single bag of a tree-decomposition, minimized over all possible decompositions[301, 368, 369]. The formal definition is provided below.
Definition B.1.
[301, 368, 369] The length of a tree-decomposition $T(G)=(\{X_{i}\mid i\in I\},T=(I,F))$ of a graph $G$ is defined as:
$\lambda(T)=\max_{i\in I}\max_{u,v\in X_{i}}d_{G}(u,v),$
where $d_{G}(u,v)$ denotes the distance between vertices $u$ and $v$ in the graph $G$. The tree-length of a graph $G$, denoted by $\text{tl}(G)$, is the minimum length over all possible tree-decompositions of $G$. Chordal graphs have tree-length 1.
A parameter closely related to tree-length is tree-breadth. Tree-breadth measures how tightly the vertices in each bag of a tree-decomposition can be clustered around a central vertex within a specific radius, minimized over all decompositions[828, 827]. The formal definition is provided below.
Definition B.2.
[828, 827] The breadth of a tree-decomposition $T(G)=(\{X_{i}\mid i\in I\},T=(I,F))$ of a graph $G$ is the minimum integer $r$ such that for every $i\in I$, there exists a vertex $v_{i}\in V(G)$ with $X_{i}\subseteq D_{r}(v_{i},G)$, where $D_{r}(v_{i},G)=\{u\in V(G)\mid d_{G}(u,v_{i})\leq r\}$ is the disk of radius $r$ centered at $v_{i}$ in $G$. The tree-breadth of $G$, denoted by $\text{tb}(G)$, is the minimum breadth over all possible tree-decompositions of $G$.
It is known that for any graph $G$:
$1\leq\text{tb}(G)\leq\text{tl}(G)\leq 2\cdot\text{tb}(G).$
A similar concept known as Branch-length is equivalent to Tree-length [1164]. Additionally, linear-length is equal to path-length [1164].
Definition B.3.
[1164] Let $B=(T,\mu)$ be a branch decomposition of a graph $G$. The branch-length of $B$, denoted by $\text{bl}(B)$, is defined as
$\text{bl}(B)=\max_{e\in E(T)}\text{diam}_{G}(\text{mid}_{B}(e)).$
The branch-length of $G$, denoted by $\text{bl}(G)$, is defined as
$\text{bl}(G)=\min_{B}\text{bl}(B),$
where the minimum is taken over all branch decompositions of $G$.
And similar concepts are tree-stretch and tree-distortion. Also, please provide the statement to be written as a theorem.
Theorem B.4.
[378] Let $G=(V,E)$ be an undirected, unweighted graph with $n$ vertices. The following inequalities hold for various graph parameters:
  1. 1.
    The tree-breadth $\text{tb}(G)$ is bounded by the tree-length $\text{tl}(G)$ as follows:
    $\text{tb}(G)\leq\text{tl}(G)\leq 2\cdot\text{tb}(G).$
    Similarly, the cluster-radius $R_{s}(G)$ and cluster-diameter $\Delta_{s}(G)$ are related as:
    $R_{s}(G)\leq\Delta_{s}(G)\leq 2\cdot R_{s}(G).$
  2. 2.
    The hyperbolicity $\text{hb}(G)$ of the graph is related to the tree-length and cluster-diameter by:
    $\text{hb}(G)\leq\text{tl}(G)\leq O(\text{hb}(G)\cdot\log n),$
    and
    $\text{hb}(G)\leq\Delta_{s}(G)\leq O(\text{hb}(G)\cdot\log n).$
    [280, 281]
  3. 3.
    The tree-stretch $\text{ts}(G)$ and tree-distortion $\text{td}(G)$ are related to the cluster-diameter as follows:
    $\text{ts}(G)\geq\text{td}(G)\geq\frac{1}{3}\Delta_{s}(G),$
    and
    $\text{td}(G)\leq 2\cdot\Delta_{s}(G)+2,\quad\text{for every }s\in V.$
    [282]
  4. 4.
    The cluster-radius $R_{s}(G)$ is bounded by the tree-distortion as follows:
    $R_{s}(G)\leq\max\{3\cdot\text{td}(G)-1,2\cdot\text{td}(G)+1\},\quad\text{for every }s\in V.$
    [282]
  5. 5.
    The tree-length $\text{tl}(G)$, cluster-diameter $\Delta_{s}(G)$, and cluster-radius $R_{s}(G)$ are related by:
    $\text{tl}(G)-1\leq\Delta_{s}(G)\leq 3\cdot\text{tl}(G),$
    and
    $R_{s}(G)\leq 2\cdot\text{tl}(G),\quad\text{for every }s\in V.$
    [370, 369]
  6. 6.
    The tree-breadth $\text{tb}(G)$ is related to the cluster-radius as follows:
    $\text{tb}(G)-1\leq R_{s}(G)\leq 3\cdot\text{tb}(G).$
    [379]
  7. 7.
    The tree-length $\text{tl}(G)$, tree-distortion $\text{td}(G)$, and tree-stretch $\text{ts}(G)$ satisfy:
    $\text{tl}(G)\leq\text{td}(G)\leq\text{ts}(G),$
    and
    $\text{tb}(G)\leq\frac{\text{ts}(G)}{2}.$
    [379]
  8. 8.
    The tree-stretch is bounded by the tree-breadth and tree-distortion as:
    $\text{ts}(G)\leq 2\cdot\text{tb}(G)\cdot\log_{2}n,$
    and
    $\text{ts}(G)\leq 2\cdot\text{td}(G)\cdot\log_{2}n.$
    [379]
  9. 9.
[…]
  10. 10.
[…]
  11. 11.
    $\text{Slimness}\leq\left\lfloor\dfrac{3}{2}\,\text{treelength}(G)\right\rfloor$ [906].
  12. 12.
    $\text{Thinness}\leq 3\,\text{treelength}(G)$ [906].
  13. 13.
    Branch-length is equivalent to Tree-length[1164].
B.2 Future tasks for length parameters
We will consider the following length parameters. These are merely suggestions, and we plan to explore them further in the future. Mathematical verification of these parameters is forthcoming.
  • Directed tree length: The length of a directed tree decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The directed tree length of a graph $G$ is the minimum length over all directed tree decompositions of $G$.
  • Directed Path length: The length of a directed path decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The directed path length of a graph $G$ is the minimum length over all directed path decompositions of $G$.
  • Rank length: The length of a rank decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The rank length of a graph $G$ is the minimum length over all rank decompositions of $G$.
  • Linear Rank length: The length of a linear rank decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The linear rank length of a graph $G$ is the minimum length over all linear rank decompositions of $G$.
  • Boolean length: The length of a boolean decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The boolean length of a graph $G$ is the minimum length over all boolean decompositions of $G$.
  • Linear Boolean length: The length of a linear boolean decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The linear boolean length of a graph $G$ is the minimum length over all linear boolean decompositions of $G$.
  • Arc length: The length of an arc decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The arc length of a graph $G$ is the minimum length over all arc decompositions of $G$.
  • Star length: The length of a star decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The star length of a graph $G$ is the minimum length over all star decompositions of $G$.
  • Hypertree length: The length of a hypertree decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The hypertree length of a graph $G$ is the minimum length over all hypertree decompositions of $G$.
  • Carving length: The length of a carving decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The carving length of a graph $G$ is the minimum length over all carving decompositions of $G$.
  • Cut length: The length of a cut decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The cut length of a graph $G$ is the minimum length over all cut decompositions of $G$.
  • Tree-cut length: The length of a tree-cut decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The tree-cut length of a graph $G$ is the minimum length over all tree-cut decompositions of $G$.
  • Tree distance length: The length of a tree distance decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The tree distance length of a graph $G$ is the minimum length over all tree distance decompositions of $G$.
  • Path distance length: The length of a path distance decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The path distance length of a graph $G$ is the minimum length over all path distance decompositions of $G$.
  • Clique length: The length of a clique decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The clique length of a graph $G$ is the minimum length over all clique decompositions of $G$.
  • Linear Clique length: The length of a linear clique decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The linear clique length of a graph $G$ is the minimum length over all linear clique decompositions of $G$.
  • NLC length: The length of a NLC decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The NLC length of a graph $G$ is the minimum length over all NLC decompositions of $G$.
  • Linear NLC length: The length of a linear NLC decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The linear NLC length of a graph $G$ is the minimum length over all linear NLC decompositions of $G$.
  • Modular length: The length of a modular decomposition is the largest diameter of the subgraphs induced by the vertex subsets used in the decomposition. The modular length of a graph $G$ is the minimum length over all modular decompositions of $G$.
Appendix C Comparing Various Graph Parameters (Over 70 Parameters)
We examine the relationships among various graph parameters. Please refer to the supplemental file “Supplemental Figure: Comparing Graph Width Parameters” for more details. Understanding these relationships is highly beneficial for developing algorithms[522].
Note C.1.
Note that unless otherwise specified, we consider any graph when comparing graph parameters. We say that a parameter $p$ upper bounds a parameter $q$ if there exists a non-decreasing function $f$ such that $f(p(G))\geq q(G)$ for all graphs $G$. Conversely, if $p$ does not upper bound $q$, then $q$ is considered unbounded with respect to $p$. Additionally, if value $a$ upper bounds value $b$ and $b$ upper bounds value $c$, then value $a$ upper bounds $c$.
Note C.2.
Comparing graph width parameters mathematically means evaluating how different measures of graph complexity, such as tree-width, path-width, and clique-width, relate to each other. These comparisons help in determining which parameter is more restrictive or more general, and thus, influence the choice of techniques and tools in algorithm design.
Appendix D Various Graph Depth Parameters
Many graph depth parameters are known. Examining the relationships among these depth parameters, such as inequalities, upper bounds, and lower bounds, is a well-known research topic.
D.1 Tree-depth
Tree-depth is a graph depth parameter introduced under several names as a measure of the sparsity of a graph [401, 945, 599, 666, 258].
Definition D.1.
The tree-depth of a graph $G$, denoted by $\text{td}(G)$, is defined based on the concept of a rooted forest. A rooted forest is a collection of trees where each tree has a designated root and all edges are directed away from the root. The closure of a rooted forest is the undirected graph obtained by connecting two vertices if there is a directed path between them in the forest. The height of a rooted forest is the length of the longest directed path from any root. The tree-depth of $G$ is the minimum height of a rooted forest whose closure contains $G$ as a subgraph.
D.1.1 Hierarchy for Tree Depth
We coinsider about hierarchy for Tree Depth. Related concept is a Block Treedepth[575].
Theorem D.2.
The relationships between Tree-depth and other graph parameters are as follows:
  • Cluster vertex deletion number is greater than or equal to Tree-depth [215].
  • Bridge-depth (a graph width parameter which measures the kernelization complexity of structural parameterizations of the Vertex Cover problem) is less than or equal to Tree-depth [202].
  • If a graph $G$ has bounded Tree-depth, then it also has bounded Path-width [1058].
  • If a graph $G$ has bounded Tree-depth, then it also has bounded Tree-width [1058].
  • If a graph $G$ has bounded vertex integrity, then it also has bounded Tree-depth [638].
  • If a graph $G$ has bounded vertex cover, then it also has bounded Tree-depth [638].
[…]
  • For any graph $G$, Tree-depth(G) = The minimum size of a centered coloring for $G$ minus 1 [940].
  • If tree-depth(G) is $k$, then for every shelter $S$ of $G$, $th(S)\leq k+1$(thickness of a shelter), stationary cop number of $G$ is $k+1$, and lifo cop number of $G$ is $k+1$[575].
  • The branch-depth of a connected graph is less than or equal to its tree-depth [328].
  • Let $G$ be a graph, $k$ be its branch-depth, and $t$ be its tree-depth. Then $k-1\leq t\leq\max(2k^{2}-k+1,2)$ [328].
  • If $G$ is of tree-depth $\leq k$, then $G$ is of shrub-depth $\leq k$ [546].
Proof D.3.
Please refer to each reference.
D.2 Branch-depth
Branch-depth is a graph depth parameter generalizing tree-depth of graphs [328]. Note that the radius of a tree is the minimum integer $r$ such that there exists a node in the tree with a distance of at most $r$ from every other node.
Definition D.4.
[328] A decomposition of a connectivity function on a set $V$ is a pair $(T,\phi)$, where $T$ is a tree with at least one internal node, and $\phi$ is a bijection from $V$ to the set of leaves of $T$. The radius of a decomposition $(T,\phi)$ is defined as the radius of the tree $T$. For an internal node $v$ of $T$, the components of the graph induced by $v$ result in a partition $\{V_{i}\}$ of $V$ by removing $v$. The width of $v$ is defined as $\max_{i}|V_{i}|$. The width of the decomposition $(T,\phi)$ is the maximum width among all internal nodes of $T$. We say that a decomposition $(T,\phi)$ is an $(r,d)$-decomposition of $V$ if its width is at most $d$ and its radius is at most $r$. The branch-depth of $V$ is the minimum integer $r$ such that there exists an $(r,d)$-decomposition of $V$.
D.2.1 Hierarchy for branch-depth
We coinsider about hierarchy for branch-depth.
Theorem D.5.
The relationships between branch-depth and other graph parameters are as follows:
  • The branch-width is less than or equal to its branch-depth [328].
  • The branch-depth of a connected graph is less than or equal to its tree-depth [328].
  • Let $G$ be a graph, $k$ be its branch-depth, and $t$ be its tree-depth. Then $k-1\leq t\leq\max(2k^{2}-k+1,2)$ [328].
  • For all matroids $M$, the branch-depth of $M$ is less than or equal to contraction-deletion depth of $M$. And
[…]
    [328].
[…]
Proof D.6.
Please refer to each reference.
D.2.2 Related Parameters for Branch-depth
Deletion-depth [328], Contraction-depth [328], Contraction-deletion Depth [328], Contraction*-depth [734], and Contraction-deletion Depth [734] are graph depth parameters analogous to operations on matroids.
D.3 Rank-depth
Rank depth [328] is a graph depth parameter defined by the Rank depth-decomposition. Note that Rank depth-decomposition of a graph $G$ is a branch-depth-decomposition of the cut-rank function [716, 955, 761, 717]. A class of simple graphs has bounded rank-depth if and only if it has bounded shrub-depth [328].
Definition D.7.
[328] The cut-rank function of a simple graph $G=(V,E)$ is a function $\rho_{G}$ defined on the subsets of $V$. For a subset $X\subseteq V$, $\rho_{G}(X)$ is the rank of the $X\times(V\setminus X)$ matrix $A_{X}=(a_{ij})_{i\in X,j\in V\setminus X}$ over the binary field. Here, $a_{ij}=1$ if and only if vertices $i$ and $j$ are adjacent, and $a_{ij}=0$ otherwise. This cut-rank function serves as an instance of a connectivity function on the vertex set of a simple graph, as discussed in a paper by Oum and Seymour [30]. We define the rank-depth of a simple graph $G$, denoted by $rd(G)$, as the branch-depth of $\rho_{G}$.
D.4 Shrub-depth
Shrub-depth is a graph depth parameter that captures the height of dense graphs [716, 547]. A related parameter is SC-depth [716, 547].
D.4.1 Hierarchy for Shrub-depth
We consider about Hierarchy for Shrub-depth.
Theorem D.8.
The relationships between Shrub-depth and other graph parameters are as follows:
  • If $G$ is of tree-depth $\leq k$, then $G$ is of shrub-depth $\leq k$ [546].
  • If $G$ is of bounded shrub-depth, then $G$ is of bounded clique-width [546].
  • If $G$ is of bounded twin-cover, then $G$ is of bounded Shrub-depth [535].
Proof D.9.
Please refer to each reference.
D.5 Directed Tree-depth
Directed Tree-depth is tree-depth on directed graphs [804].
Related parameter is DAG-depth. DAG-depth is a structural depth measure of directed graphs, which naturally extends the tree-depth of ordinary graphs [136, 551].
If $\text{If }G\text{ is a symmetric digraph, then }\text{ddp}(G)=\text{cr}(G)+1$. Definition of DAG-depth is following.
Definition D.10 (DAG-depth).
[136, 551] Let $D$ be a digraph, and let $R_{1},\ldots,R_{p}$ be the reachable fragments of $D$. The DAG-depth $\text{dd}_{p}(D)$ is defined inductively as follows:
$\text{dd}_{p}(D)=\begin{cases}1&\text{if }|V(D)|=1,\\ 1+\min_{v\in V(D)}(\text{dd}_{p}(D-v))&\text{if }p=1\text{ and }|V(D)|>1,\\ \max_{1\leq i\leq p}\text{dd}_{p}(R_{i})&\text{otherwise}.\end{cases}$
Appendix E Proposal for Width-Parameters for Handling Uncertainty
In the real world, Tree-width and related parameters are often adapted and applied in more suitable forms. Concepts such as Fuzzy Graphs[859, 150, 1092, 996], SuperHyperGraphs[1101, 1106, 635, 1103, 634], Fuzzy Directed Graphs[111], Fuzzy HyperGraphs[1180], Rough Graphs[272, 908, 1057, 931, 881, 140, 833], and Neutrosophic Graphs[504, 520, 679, 726, 223, 42, 221, 331, 1104, 225] are used to address uncertainties and probabilistic aspects. We are conducting studies and investigations into Width Parameters for these concepts, and an introduction to these is provided.
E.1 SuperHyperTree-width
A Superhypergraph is a generalization of the concept of a hypergraph. The definition of SuperHypergraph is as follows[1103, 1101, 635, 1106, 634]. Similarly, Tree-width has been extended to SuperHypertree-width, as detailed in the literature [509, 521]. Please refer to these sources as necessary.
Definition E.1.
[1102] A SuperHyperGraph (SHG) is an ordered pair $\text{SHG}=(G\subseteq P(V),E\subseteq P(V))$, where:
  1. 1.
    $V=\{V_{1},V_{2},\dots,V_{m}\}$ is a finite set of $m\geq 0$ vertices, or an infinite set.
  2. 2.
    $P(V)$ is the power set of $V$ (all subsets of $V$). Therefore, an SHG-vertex may be a single (classical) vertex, a super-vertex (a subset of many vertices) that represents a group (organization), or even an indeterminate-vertex (unclear, unknown vertex); $\emptyset$ represents the null-vertex (a vertex that has no element).
  3. 3.
    $E=\{E_{1},E_{2},\dots,E_{m}\}$, for $m\geq 1$, is a family of subsets of $V$, and each $E_{j}$ is an SHG-edge, $E_{i}\in P(V)$. An SHG-edge may be a (classical) edge, a super-edge (an edge between super-vertices) that represents connections between two groups (organizations), a hyper-super-edge that represents connections between three or more groups (organizations), a multi-edge, or even an indeterminate-edge (unclear, unknown edge); $\emptyset$ represents the null-edge (an edge that means there is no connection between the given vertices).
Example E.2.
We define a finite set of vertices:
$V=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$
In a SuperHypergraph, vertices can be single vertices, super-vertices (subsets of vertices), or even the empty set. Here are some examples:
  • Single Vertices:
    $V_{1}$ =$=\{v_{1}\}$
    $V_{2}$ =$=\{v_{2}\}$
    $V_{3}$ =$=\{v_{3}\}$
    $V_{4}$ =$=\{v_{4}\}$
    $V_{5}$ =$=\{v_{5}\}$
  • Super-Vertices (Subset Vertices):
    $\text{SV}_{1,2}$ =$=\{v_{1},v_{2}\}$
    $\text{SV}_{3,4,5}$ =$=\{v_{3},v_{4},v_{5}\}$
  • Null Vertex:
    $\emptyset_{V}=\emptyset$
Edges in a SuperHypergraph can be single edges, super-edges, hyper-edges, hyper-super-edges, or the null edge. Below are definitions of each type:
  • Single Edges:
    • $E_{1,2}=\{V_{1},V_{2}\}$: An edge connecting $v_{1}$ and $v_{2}$
    • $E_{4,5}=\{V_{4},V_{5}\}$: An edge connecting $v_{4}$ and $v_{5}$
  • Hyper-Edges:
    $\text{HE}_{1,3,5}=\{V_{1},V_{3},V_{5}\}$
    An edge connecting $v_{1},v_{3},$ and $v_{5}$
  • Super-Edges (Subset Edges):
    $\text{SE}_{(1,2),(3)}=\{\text{SV}_{1,2},V_{3}\}$
    An edge connecting super-vertex $\text{SV}_{1,2}$ and vertex $v_{3}$
  • Hyper-Super-Edges (Hyper Subset Edges):
    $\text{HSE}_{(1,2),(3,4,5)}=\{\text{SV}_{1,2},\text{SV}_{3,4,5}\}$
    An edge connecting super-vertices $\text{SV}_{1,2}$ and $\text{SV}_{3,4,5}$
  • Null Edge:
    $\emptyset_{E}=\emptyset$
    Represents the absence of a connection
E.2 Neutrosophictree-width
In recent years, Neutrosophic Graphs[222, 220, 223, 42, 221, 331, 1104, 224, 810, 1244] have been actively studied within Neutrosophic Set Theory[58, 57]. Neutrosophic[1188] refers to a mathematical framework that generalizes classical and fuzzy logic[774, 628, 894, 1044], simultaneously handling degrees of truth, indeterminacy, and falsity within an interval. These graphs, as generalizations of Fuzzy Graphs[1043, 915], have garnered attention for their potential applications similar to those of Fuzzy Graphs.
First, we introduced the concept of the Neutrosophic Set[1100, 57]. The Neutrosophic Set is an extension of set theory that incorporates the principles of Neutrosophic logic.
Definition E.3 (Neutrosophic Set).
(cf.[1100, 57]) Let $X$ be a space of points and let $x\in X$. A neutrosophic set $S$ in $X$ is characterized by three membership functions: a truth membership function $T_{S}$, an indeterminacy membership function $I_{S}$, and a falsity membership function $F_{S}$. For each point $x\in X$, $T_{S}(x),I_{S}(x),$ and $F_{S}(x)$ are real standard or non-standard subsets of the interval $]0^{-},1^{+}[$, where:
$T_{S},I_{S},F_{S}:X\to[0^{-},1^{+}].$
The neutrosophic set $S$ can be represented as:
$S=\{(x,T_{S}(x),I_{S}(x),F_{S}(x))\mid x\in X\}.$
There are no restrictions on the sum of $T_{S}(x),I_{S}(x),$ and $F_{S}(x)$, so:
$0\leq T_{S}(x)+I_{S}(x)+F_{S}(x)\leq 3^{+}.$
The definition of a Neutrosophic graph is as follows [1101, 1105, 1104, 633]. A Neutrosophic graph is an extension of the Neutrosophic set concept applied to graphs. Similarly, Tree-width has been extended to Neutrosophictree-width, as detailed in the literature [521]. Please refer to these sources as necessary.
Definition E.4.
(cf.[1104]) A neutrosophic graph $NTG=(V,E,\sigma=(\sigma_{1},\sigma_{2},\sigma_{3}),\mu=(\mu_{1},\mu_{2},\mu_{3}))$ is defined as a graph where $\sigma_{i}:V\rightarrow[0,1]$, $\mu_{i}:E\rightarrow[0,1]$, and for every $v_{i}v_{j}\in E$, the following condition holds: $\mu(v_{i}v_{j})\leq\sigma(v_{i})\land\sigma(v_{j})$.
  1. 1.
    $\sigma$ is called the neutrosophic vertex set.
  2. 2.
    $\mu$ is called the neutrosophic edge set.
  3. 3.
    $|V|$ is called the order of $NTG$, and it is denoted by $O(NTG)$.
  4. 4.
    $\sum_{v\in V}\sigma(v)$ is called the neutrosophic order of $NTG$, and it is denoted by $On(NTG)$.
  5. 5.
    $|E|$ is called the size of $NTG$, and it is denoted by $S(NTG)$.
  6. 6.
    $\sum_{e\in E}\mu(e)$ is called the neutrosophic size of $NTG$, and it is denoted by $Sn(NTG)$.
Definition E.5.
  1. (i)
    A sequence of vertices $P:x_{0},x_{1},\cdots,x_{n}$ is called a path where $x_{i}x_{i+1}\in E$, $i=0,1,\cdots,n-1$.
  2. (ii)
    The strength of the path $P:x_{0},x_{1},\cdots,x_{n}$ is $\bigwedge_{i=0,\cdots,n-1}\mu(x_{i}x_{i+1})$.
  3. (iii)
    The connectedness between vertices $x_{0}$ and $x_{n}$ is defined as:
    $\mu_{\infty}(x,y)=\bigwedge_{P:x_{0},x_{1},\cdots,x_{n}}\bigwedge_{i=0,\cdots,n-1}\mu(x_{i}x_{i+1}).$
Example E.6.
(cf.[224]) Consider a neutrosophic graph $NTG=(V,E,\sigma=(\sigma_{1},\sigma_{2},\sigma_{3}),\mu=(\mu_{1},\mu_{2},\mu_{3}))$ with four vertices $V=\{v_{1},v_{2},v_{3},v_{4}\}$, as shown in the diagram.
The neutrosophic membership degrees of the vertices are as follows:
$\sigma(v_{1})=(0.5,0.1,0.4),\quad\sigma(v_{2})=(0.6,0.3,0.2),\\ \sigma(v_{3})=(0.2,0.3,0.4),\quad\sigma(v_{4})=(0.4,0.2,0.5)$
The neutrosophic membership degrees of the edges are as follows:
$\mu(v_{1}v_{2})=(0.2,0.3,0.4),\quad\mu(v_{2}v_{3})=(0.3,0.3,0.4),\\ \mu(v_{3}v_{4})=(0.2,0.3,0.4),\quad\mu(v_{4}v_{1})=(0.1,0.2,0.5)$
In this case, the neutrosophic graph $NTG$ has the following properties:
  • Vertices $v_{1},v_{2},v_{3},v_{4}$ are connected by edges with varying neutrosophic membership degrees.
  • The neutrosophic relations ensure that for every edge $v_{i}v_{j}\in E$, $\mu(v_{i}v_{j})\leq\sigma(v_{i})\land\sigma(v_{j})$, where $\land$ denotes the minimum operation.
E.3 Rough Tree-width from an Information System
A rough graph represents relationships between objects, utilizing rough set theory to handle uncertainty, with edges determined by the rough membership values of vertices[272, 908, 1245, 1057, 931, 881, 140, 833, 275]. The definition of rough graph in information system [327] is as follows. Similarly, Tree-width has been extended to Roughtree-width, as detailed in the literature [516]. Please refer to these sources as necessary.
Definition E.7.
[327] Assume $\mathbb{M}=(U,F)$ is an information system, and let $\emptyset=G\subseteq U$. The rough membership function for the set $\omega_{G}$ is defined as:
$\omega_{G}^{F}=\frac{|[f]_{G}\cap G|}{|[f]_{F}|}$
for some $f\in U$.
Definition E.8.
[327] Let $\mathcal{U}=\{V,E,\omega\}$ be a triple consisting of a non-empty set $V=\{v_{1},v_{2},\dots,v_{n}\}=\mathcal{U}$, where $\mathcal{U}$ is a universe, $E=\{e_{1},e_{2},\dots,e_{n}\}$ is a set of unordered pairs of distinct elements of $V$, and $\omega$ is a function $\omega:V\rightarrow[0,1]$. A Rough graph is defined as:
$(v_{i},v_{j})=\begin{cases}\text{edge}&\text{if }\max\left(\omega_{V}^{G}(v_{i}),\omega_{V}^{G}(v_{j})\right)>0,\\ \text{no edge}&\text{if }\max\left(\omega_{V}^{G}(v_{i}),\omega_{V}^{G}(v_{j})\right)=0.\end{cases}$
E.4 Bipolar fuzzytree width
A bipolar fuzzy graph represents relationships where each edge and vertex has two fuzzy membership values: positive (indicating satisfaction of a property) and negative (indicating partial counter-property satisfaction). It helps model systems with both positive and negative traits simultaneously[1234, 46, 1195, 1023, 41, 26, 1020, 42, 30, 904, 905]. The definitions of bipolar fuzzy graph and bipolar fuzzy tree are presented as follows.
Definition E.9.
[26] Let $V$ be a non-empty set. A bipolar fuzzy graph is a triple $G=(V,A,B)$, where:
  • $A=(\mu_{A}^{P},\mu_{A}^{N})$ is a bipolar fuzzy set on $V$, where $\mu_{A}^{P}:V\rightarrow[0,1]$ and $\mu_{A}^{N}:V\rightarrow[-1,0]$ represent the positive and negative membership degrees of the vertices in $V$.
  • $B=(\mu_{B}^{P},\mu_{B}^{N})$ is a bipolar fuzzy relation on $E\subseteq V\times V$, where $\mu_{B}^{P}:E\rightarrow[0,1]$ and $\mu_{B}^{N}:E\rightarrow[-1,0]$ represent the positive and negative membership degrees of the edges in $E$.
For all edges $xy\in E$, the following conditions hold:
$\mu_{B}^{P}(xy)\leq\min(\mu_{A}^{P}(x),\mu_{A}^{P}(y))\quad\text{and}\quad\mu_{B}^{N}(xy)\geq\max(\mu_{A}^{N}(x),\mu_{A}^{N}(y)),$
where $V$ is the set of vertices and $E$ is the set of edges. Here, $\mu_{A}^{P}(x)$ represents the positive membership degree of vertex $x$, and $\mu_{A}^{N}(x)$ represents the negative membership degree of vertex $x$, while $\mu_{B}^{P}(xy)$ and $\mu_{B}^{N}(xy)$ represent the positive and negative membership degrees of the edge $xy$, respectively.
Example E.10.
(cf.[1052]) Let $V=\{v_{1},v_{2},v_{3}\}$ be a set of vertices, where the bipolar fuzzy membership degrees for each vertex are defined as follows:
  • For vertex $v_{1}$, the positive membership degree is $\mu_{A}^{P}(v_{1})=0.6$ and the negative membership degree is $\mu_{A}^{N}(v_{1})=-0.2$.
  • For vertex $v_{2}$, the positive membership degree is $\mu_{A}^{P}(v_{2})=0.2$ and the negative membership degree is $\mu_{A}^{N}(v_{2})=-0.5$.
  • For vertex $v_{3}$, the positive membership degree is $\mu_{A}^{P}(v_{3})=0.1$ and the negative membership degree is $\mu_{A}^{N}(v_{3})=0$.
The edges between these vertices have the following bipolar fuzzy membership degrees:
  • For the edge between $v_{1}$ and $v_{2}$, the positive membership degree is $\mu_{B}^{P}(v_{1}v_{2})=0$ and the negative membership degree is $\mu_{B}^{N}(v_{1}v_{2})=-0.2$.
  • For the edge between $v_{2}$ and $v_{3}$, the positive membership degree is $\mu_{B}^{P}(v_{2}v_{3})=0.1$ and the negative membership degree is $\mu_{B}^{N}(v_{2}v_{3})=0$.
  • For the edge between $v_{3}$ and $v_{1}$, the positive membership degree is $\mu_{B}^{P}(v_{3}v_{1})=0.3$ and the negative membership degree is $\mu_{B}^{N}(v_{3}v_{1})=0$.
This example satisfies the conditions for a Bipolar Fuzzy Graph, as the edge membership degrees adhere to the restrictions based on the vertex membership degrees:
$\mu_{B}^{P}(v_{i}v_{j})\leq\min(\mu_{A}^{P}(v_{i}),\mu_{A}^{P}(v_{j}))\quad\text{and}\quad\mu_{B}^{N}(v_{i}v_{j})\geq\max(\mu_{A}^{N}(v_{i}),\mu_{A}^{N}(v_{j})).$
Definition E.11.
Let $G=(V,A,B)$ be a connected bipolar fuzzy graph, where $A=(\mu_{A}^{P},\mu_{A}^{N})$ is a bipolar fuzzy set on $V$ and $B=(\mu_{B}^{P},\mu_{B}^{N})$ is a bipolar fuzzy relation on $E\subseteq V\times V$. The graph $G$ is called a bipolar fuzzy tree if there exists a bipolar fuzzy spanning subgraph $H=(V,A,C)$, where $C=(\mu_{C}^{P},\mu_{C}^{N})$ is a bipolar fuzzy relation on $E_{H}\subseteq E$, such that:
  1. 1.
    $H$ is a tree (i.e., a connected, acyclic graph).
  2. 2.
    For all edges $(x,y)\in E\setminus E_{H}$ (i.e., edges not in the spanning tree $H$):
    $\mu_{B}^{P}(x,y)<\mu_{C}^{P}(x,y)\quad\text{and}\quad\mu_{B}^{N}(x,y)>\mu_{C}^{N}(x,y),$
    where $\mu_{C}^{P}(x,y)$ and $\mu_{C}^{N}(x,y)$ denote the maximum positive and minimum negative membership degrees for edges in $H$, respectively.
A bipolar fuzzy graph $G$ is connected if any two vertices in $V$ are joined by a path in $G$, where the path consists of edges $xy\in E$ such that $\mu_{B}^{P}(x,y)>0$ or $\mu_{B}^{N}(x,y)<0$.
Here is the definition, although it is still in the conceptual stage.
Definition E.12.
Let $G=(V,E)$ be a graph where $V$ is the set of vertices and $E$ is the set of edges. A positive tree-decomposition of $G$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a tree with nodes $I$ and edges $F$.
  • $\{B_{t}\}_{t\in I}$ is a collection of subsets of $V$ (called bags) associated with the nodes of $T$, such that:
    1. 1.
      For each vertex $v\in V$, the set $\{t\in I:v\in B_{t}\}$ is connected in the tree $T$.
    2. 2.
      For each edge $(u,v)\in E$, there exists at least one node $t\in I$ such that both $u$ and $v$ belong to the bag $B_{t}$.
The width of a positive tree-decomposition is defined as:
$\text{width}=\max_{t\in I}\left(|B_{t}|-1\right),$
where $|B_{t}|$ is the size of the bag $B_{t}$.
The Positive Tree-width of the graph $G$ is the minimum width among all possible positive tree-decompositions of $G$.
Definition E.13.
Let $G=(V,E)$ be a graph where $V$ is the set of vertices and $E$ is the set of edges. A negative tree-decomposition of $G$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a tree with nodes $I$ and edges $F$.
  • $\{B_{t}\}_{t\in I}$ is a collection of subsets of $V$ (called bags) associated with the nodes of $T$, such that:
    1. 1.
      For each vertex $v\in V$, the set $\{t\in I:v\in B_{t}\}$ is connected in the tree $T$.
    2. 2.
      For each edge $(u,v)\in E$, there exists at least one node $t\in I$ such that both $u$ and $v$ belong to the bag $B_{t}$.
The width of a negative tree-decomposition is defined as:
$\text{width}=\max_{t\in I}\left(|B_{t}|-1\right),$
where $|B_{t}|$ is the size of the bag $B_{t}$.
The Negative Tree-width of the graph $G$ is the minimum width among all possible negative tree-decompositions of $G$.
Definition E.14.
Let $G=(V,E)$ be a graph where $V$ is the set of vertices and $E$ is the set of edges. A positive path-decomposition of $G$ is a pair $(P,\{B_{p}\}_{p\in I})$, where:
  • $P=(I,F)$ is a path with nodes $I$ and edges $F$.
  • $\{B_{p}\}_{p\in I}$ is a collection of subsets of $V$ (called bags) associated with the nodes of $P$, such that:
    1. 1.
      For each vertex $v\in V$, the set $\{p\in I:v\in B_{p}\}$ is connected in the path $P$.
    2. 2.
      For each edge $(u,v)\in E$, there exists at least one node $p\in I$ such that both $u$ and $v$ belong to the bag $B_{p}$.
The width of a positive path-decomposition is defined as:
$\text{width}=\max_{p\in I}\left(|B_{p}|-1\right),$
where $|B_{p}|$ is the size of the bag $B_{p}$.
The Positive Path-width of the graph $G$ is the minimum width among all possible positive path-decompositions of $G$.
Definition E.15.
Let $G=(V,E)$ be a graph where $V$ is the set of vertices and $E$ is the set of edges. A negative path-decomposition of $G$ is a pair $(P,\{B_{p}\}_{p\in I})$, where:
  • $P=(I,F)$ is a path with nodes $I$ and edges $F$.
  • $\{B_{p}\}_{p\in I}$ is a collection of subsets of $V$ (called bags) associated with the nodes of $P$, such that:
    1. 1.
      For each vertex $v\in V$, the set $\{p\in I:v\in B_{p}\}$ is connected in the path $P$.
    2. 2.
      For each edge $(u,v)\in E$, there exists at least one node $p\in I$ such that both $u$ and $v$ belong to the bag $B_{p}$.
The width of a negative path-decomposition is defined as:
$\text{width}=\max_{p\in I}\left(|B_{p}|-1\right),$
where $|B_{p}|$ is the size of the bag $B_{p}$.
The Negative Path-width of the graph $G$ is the minimum width among all possible negative path-decompositions of $G$.
E.5 Bipolar fuzzy hypertree-width
A Bipolar Fuzzy Hypergraph extends traditional hypergraphs by incorporating two membership functions—positive and negative—to represent both positive and negative relationships within a vertex set. Various studies have been conducted on this topic[928, 37, 38, 33, 1079, 849, 1051]. The definition of a Bipolar Fuzzy Hypergraph is provided below.
Definition E.16.
[1051] Let $X$ be a finite set of vertices, and let $\xi$ be a finite family of bipolar fuzzy subsets on $X$. A bipolar fuzzy hypergraph is a pair $H=(X,\xi)$, where:
1. The set $X$ is the crisp vertex set.
2. Each hyperedge $B\in\xi$ is a bipolar fuzzy subset of $X$, represented by two membership functions:
  • $\mu_{B}^{+}:X\to[0,1]$, the positive membership function,
  • $\mu_{B}^{-}:X\to[-1,0]$, the negative membership function.
These functions satisfy:
$0\leq\mu_{B}^{+}(x)+|\mu_{B}^{-}(x)|\leq 1$
for all $x\in X$ and $B\in\xi$.
3. The support of a bipolar fuzzy hyperedge $B$ is defined as:
$\text{supp}(B)=\{x\in X\ |\ \mu_{B}^{+}(x)\neq 0\text{ or }\mu_{B}^{-}(x)\neq 0\}$
The bipolar fuzzy hypergraph satisfies:
$X=\bigcup_{B\in\xi}\text{supp}(B)$
4. The height of the bipolar fuzzy hypergraph $H$, denoted $h(H)$, is defined as:
$h(H)=\max_{B\in\xi}\max_{x\in X}\mu_{B}^{+}(x)$
The depth of the bipolar fuzzy hypergraph $H$, denoted $d(H)$, is defined as:
$d(H)=\min_{B\in\xi}\min_{x\in X}\mu_{B}^{-}(x)$
5. The incidence matrix of the bipolar fuzzy hypergraph is given by $[a_{ij}]$, where $a_{ij}=(\mu_{B}^{+}(x_{i}),\mu_{B}^{-}(x_{i}))$ for each vertex $x_{i}$ and hyperedge $B_{j}$.
Definition E.17.
A Bipolar Fuzzy Hypertree is a bipolar fuzzy hypergraph $H=(X,\xi)$, where:
  1. 1.
    $X$ is a finite set of vertices.
  2. 2.
    $\xi=\{B_{1},B_{2},\dots,B_{m}\}$ is a family of bipolar fuzzy hyperedges, each $B_{j}\in\xi$ represented by two membership functions:
    • $\mu_{B_{j}}^{+}:X\to[0,1]$, the positive membership function,
    • $\mu_{B_{j}}^{-}:X\to[-1,0]$, the negative membership function.
    These functions satisfy:
    $0\leq\mu_{B_{j}}^{+}(x)+|\mu_{B_{j}}^{-}(x)|\leq 1\quad\text{for all }x\in X\text{ and }B_{j}\in\xi.$
  3. 3.
    The support of a bipolar fuzzy hyperedge $B_{j}$ is defined as:
    $\text{supp}(B_{j})=\{x\in X\ |\ \mu_{B_{j}}^{+}(x)\neq 0\text{ or }\mu_{B_{j}}^{-}(x)\neq 0\}.$
Although still in the conceptual stage, the definition is provided below.
Definition E.18.
A Bipolar Fuzzy Hypertree Decomposition of a bipolar fuzzy hypergraph $H=(X,\xi)$ is a triple $(T,\{B_{t}\},\{C_{t}\})$, where:
  • $T$ is a tree with vertex set $V(T)$ and edge set $E(T)$.
  • $B_{t}\subseteq X$ is a bipolar fuzzy subset associated with each node $t\in V(T)$, with positive and negative membership functions $\mu_{B}^{+}(x),\mu_{B}^{-}(x)$, satisfying:
    $0\leq\mu_{B}^{+}(x)+|\mu_{B}^{-}(x)|\leq 1$
  • $C_{t}\subseteq\xi$ is a bipolar fuzzy subset of hyperedges associated with each node $t\in V(T)$.
The decomposition must satisfy:
  1. 1.
    Vertex Coverage: For each vertex $x\in X$, the set $\{t\in V(T)\ |\ x\in B_{t}\}$ forms a connected subtree of $T$.
  2. 2.
    Edge Coverage: For each bipolar fuzzy hyperedge $B_{j}\in\xi$, there exists a node $t\in V(T)$ such that $\text{supp}(B_{j})\subseteq B_{t}$.
The width of a bipolar fuzzy hypertree decomposition is:
$\text{width}(T,\{B_{t}\},\{C_{t}\})=\max_{t\in V(T)}\left(\sup_{x\in B_{t}}\left(\mu_{B}^{+}(x)+|\mu_{B}^{-}(x)|\right)-1\right).$
The Bipolar Fuzzy Hypertree-Width of $H$ is the minimum width over all bipolar fuzzy hypertree decompositions:
[…]
Definition E.19.
In a Bipolar Fuzzy Hypertree $H=(X,\xi)$, we define two types of widths: positive-width and negative-width, which capture the contributions from the positive and negative membership functions, respectively.
  1. 1.
    Positive-Width: The positive-width of a bipolar fuzzy hypertree decomposition $(T,\{B_{t}\},\{C_{t}\})$ is defined as:
[…]
    where $\mu_{B}^{+}(x)$ is the positive membership degree of vertex $x$ in the fuzzy bag $B_{t}$.
    The Positive-Fuzzy-Hypertree-Width of $H$ is the minimum positive-width over all possible bipolar fuzzy hypertree decompositions:
[…]
  2. 2.
    Negative-Width: The negative-width of a bipolar fuzzy hypertree decomposition $(T,\{B_{t}\},\{C_{t}\})$ is defined as:
[…]
    where $\mu_{B}^{-}(x)$ is the negative membership degree of vertex $x$ in the fuzzy bag $B_{t}$.
    The Negative-Fuzzy-Hypertree-Width of $H$ is the minimum negative-width over all possible bipolar fuzzy hypertree decompositions:
[…]
E.6 Single valued neutrosophic tree-width
Single-Valued Neutrosophic Graphs are graphs where each vertex and edge has three membership values: truth, indeterminacy, and falsity, allowing for modeling uncertainty and imprecision in relationships between vertices [43, 871, 1134, 932, 329, 842, 890, 314, 45, 226, 732, 27, 933].
Definition E.20.
Let $V$ be a set of vertices. A single valued neutrosophic graph (SVN-graph) is defined as a pair $G=(A,B)$, where:
1. For each vertex $v\in V$, we have three functions:
  • Truth-membership function $T:V\to[0,1]$
  • Indeterminacy-membership function $I:V\to[0,1]$
  • Falsity-membership function $F:V\to[0,1]$
These satisfy the condition:
$0\leq T(v)+I(v)+F(v)\leq 3\quad\text{for all }v\in V$
2. For each edge $e=(v_{i},v_{j})\in E$, we define:
  • Truth-membership function $T_{E}:E\to[0,1]$
  • Indeterminacy-membership function $I_{E}:E\to[0,1]$
  • Falsity-membership function $F_{E}:E\to[0,1]$
These satisfy the following conditions:
$T_{E}(v_{i},v_{j})\leq\min(T(v_{i}),T(v_{j}))$
$I_{E}(v_{i},v_{j})\geq\max(I(v_{i}),I(v_{j}))$
$F_{E}(v_{i},v_{j})\geq\max(F(v_{i}),F(v_{j}))$
Additionally, we have:
$0\leq T_{E}(v_{i},v_{j})+I_{E}(v_{i},v_{j})+F_{E}(v_{i},v_{j})\leq 3$
3. The edge set $E$ forms a symmetric single valued neutrosophic relation on $V$.
Example E.21.
Let the vertex set be $V=\{v_{1},v_{2},v_{3},v_{4}\}$. The truth-membership, indeterminacy-membership, and falsity-membership functions for the vertices are given as follows:
$T(v_{1})=0.3,\quad I(v_{1})=0.4,\quad F(v_{1})=0.6$
$T(v_{2})=0.3,\quad I(v_{2})=0.2,\quad F(v_{2})=0.5$
$T(v_{3})=0.1,\quad I(v_{3})=0.6,\quad F(v_{3})=0.6$
$T(v_{4})=0.4,\quad I(v_{4})=0.4,\quad F(v_{4})=0.3$
The edge set $E=\{(v_{1},v_{2}),(v_{2},v_{3}),(v_{3},v_{4}),(v_{4},v_{1})\}$ has the following corrected truth, indeterminacy, and falsity membership values for the edges:
$T_{E}(v_{1},v_{2})=0.3,\quad I_{E}(v_{1},v_{2})=0.5,\quad F_{E}(v_{1},v_{2})=0.6$
$T_{E}(v_{2},v_{3})=0.1,\quad I_{E}(v_{2},v_{3})=0.6,\quad F_{E}(v_{2},v_{3})=0.6$
$T_{E}(v_{3},v_{4})=0.1,\quad I_{E}(v_{3},v_{4})=0.7,\quad F_{E}(v_{3},v_{4})=0.6$
$T_{E}(v_{4},v_{1})=0.3,\quad I_{E}(v_{4},v_{1})=0.4,\quad F_{E}(v_{4},v_{1})=0.6$
  1. 1.
    Truth-membership condition:
    $T_{E}(v_{i},v_{j})\leq\min(T(v_{i}),T(v_{j}))$
    For example, for edge $(v_{1},v_{2})$:
    $T_{E}(v_{1},v_{2})=0.3\leq\min(0.3,0.3)=0.3$
    Similarly, the conditions hold for all edges.
  2. 2.
    Indeterminacy-membership condition:
    $I_{E}(v_{i},v_{j})\geq\max(I(v_{i}),I(v_{j}))$
    For example, for edge $(v_{4},v_{1})$:
    $I_{E}(v_{4},v_{1})=0.4\geq\max(0.4,0.4)=0.4$
    Similarly, the conditions hold for all edges.
  3. 3.
    Falsity-membership condition:
    $F_{E}(v_{i},v_{j})\geq\max(F(v_{i}),F(v_{j}))$
    For example, for edge $(v_{4},v_{1})$:
    $F_{E}(v_{4},v_{1})=0.6\geq\max(0.6,0.3)=0.6$
    Similarly, the conditions hold for all edges.
Thus, after corrections, the graph satisfies the conditions of a Single Valued Neutrosophic Graph.
Although still in the conceptual stage, the definition is provided below.
Definition E.22.
A Single-Valued Neutrosophic Tree (SVN-Tree) is a Single-Valued Neutrosophic Graph $G=(A,B)$ with vertex set $V$ and edge set $E$, where:
  1. 1.
    For each vertex $v\in V$, there are three membership functions:
    • Truth-membership function $T(v)\in[0,1]$,
    • Indeterminacy-membership function $I(v)\in[0,1]$,
    • Falsity-membership function $F(v)\in[0,1]$,
    satisfying:
    $0\leq T(v)+I(v)+F(v)\leq 3$
  2. 2.
    For each edge $e=(v_{i},v_{j})\in E$, there are three membership functions:
    • Truth-membership function $T_{E}(e)\in[0,1]$,
    • Indeterminacy-membership function $I_{E}(e)\in[0,1]$,
    • Falsity-membership function $F_{E}(e)\in[0,1]$,
    satisfying:
    $T_{E}(e)\leq\min(T(v_{i}),T(v_{j})),$
    $I_{E}(e)\geq\max(I(v_{i}),I(v_{j})),$
    $F_{E}(e)\geq\max(F(v_{i}),F(v_{j})),$
    and:
    $0\leq T_{E}(e)+I_{E}(e)+F_{E}(e)\leq 3$
  3. 3.
    The SVN-tree is acyclic and connected, meaning the underlying crisp graph of the SVN-graph is a tree.
Definition E.23.
The Single-Valued Neutrosophic Tree-Width (SVN-Tree-Width) of an SVN-graph $G$ is defined by a tree-decomposition $(T,\{W_{t}\})$, where:
  • $T$ is a tree, and $\{W_{t}\}$ is a collection of subsets $W_{t}\subseteq V$ satisfying the tree-decomposition conditions.
  • The width of the decomposition is:
    $\text{width}(T,\{W_{t}\})=\max_{t\in V(T)}\left(|W_{t}|-1\right).$
  • The SVN-Tree-Width is the minimum width over all tree-decompositions:
[…]
E.7 Single-Valued Neutrosophic Hypertree-width
A Single-Valued Neutrosophic Hypergraph (SVN-HG) represents uncertainty through three membership functions—truth, indeterminacy, and falsity—assigned to each vertex and hyperedge. Extensive research has been conducted on this topic[27, 39, 329, 43, 34, 33, 45, 40, 632, 870]. The definition of a Single-Valued Neutrosophic Hypergraph is provided below.
Definition E.24.
Let $H=\{v_{1},v_{2},\dots,v_{n}\}$ be a finite set of vertices, and let $P^{*}(H)$ denote the power set of $H$ excluding the empty set. A Single-Valued Neutrosophic Hypergraph (SVN-HG) is defined as follows:
1. For each vertex $v_{j}\in H$, three membership functions are defined:
  • $\alpha_{E}(v_{j}):H\to[0,1]$, the truth-membership function,
  • $\beta_{E}(v_{j}):H\to[0,1]$, the indeterminacy-membership function,
  • $\gamma_{E}(v_{j}):H\to[0,1]$, the falsity-membership function.
These functions satisfy:
$0\leq\alpha_{E}(v_{j})+\beta_{E}(v_{j})+\gamma_{E}(v_{j})\leq 3$
for all $v_{j}\in H$.
2. A single-valued neutrosophic hyperedge is a set $E_{i}=\{(v_{j},\alpha_{E}(v_{j}),\beta_{E}(v_{j}),\gamma_{E}(v_{j}))\}\subseteq P^{*}(H)$, where each $v_{j}\in E_{i}$ has associated neutrosophic membership values. The family of hyperedges is denoted $\{E_{i}\}_{i=1}^{m}$, where $m$ is the number of hyperedges, and each $E_{i}$ is a non-trivial neutrosophic subset of $H$. The support of $E_{i}$ is defined as:
$\text{supp}(E_{i})=\{v_{j}\in H\ |\ \alpha_{E}(v_{j})\neq 0\ \text{or}\ \beta_{E}(v_{j})\neq 0\ \text{or}\ \gamma_{E}(v_{j})\neq 0\}$
The hypergraph satisfies:
$H=\bigcup_{i=1}^{m}\text{supp}(E_{i})$
3. For any $0\leq\epsilon_{1},\epsilon_{2},\epsilon_{3}\leq 1$, the $(\epsilon_{1},\epsilon_{2},\epsilon_{3})$-level subset of a single-valued neutrosophic set $A$ is defined as:
$A(\epsilon_{1},\epsilon_{2},\epsilon_{3})=\{x\in X\ |\ \alpha_{A}(x)\geq\epsilon_{1},\beta_{A}(x)\geq\epsilon_{2},\gamma_{A}(x)\leq\epsilon_{3}\}$
Although still in the conceptual stage, the definition is provided below.
Definition E.25.
A Single-Valued Neutrosophic Hypertree (SVN-Hypertree) is a neutrosophic hypergraph $H=(V,E)$, where:
  1. 1.
    For each vertex $v_{j}\in V$, we define three membership functions:
    • Truth-membership function $\alpha_{H}(v_{j})\in[0,1]$,
    • Indeterminacy-membership function $\beta_{H}(v_{j})\in[0,1]$,
    • Falsity-membership function $\gamma_{H}(v_{j})\in[0,1]$,
    satisfying:
    $0\leq\alpha_{H}(v_{j})+\beta_{H}(v_{j})+\gamma_{H}(v_{j})\leq 3\quad\text{for all }v_{j}\in V.$
  2. 2.
    For each hyperedge $E_{i}\subseteq V$, we define a set of neutrosophic membership functions $\alpha_{E}(v_{j}),\beta_{E}(v_{j}),\gamma_{E}(v_{j})$ for each vertex $v_{j}\in E_{i}$. The support of $E_{i}$ is defined as:
    $\text{supp}(E_{i})=\{v_{j}\in V\ |\ \alpha_{E}(v_{j})\neq 0\ \text{or}\ \beta_{E}(v_{j})\neq 0\ \text{or}\ \gamma_{E}(v_{j})\neq 0\}.$
  3. 3.
    The hypertree is connected and acyclic, with all vertices and hyperedges satisfying the neutrosophic membership conditions.
Definition E.26.
The Single-Valued Neutrosophic Hypertree-Width (SVN-Hypertree-width) of a single-valued neutrosophic hypertree is defined by a hypertree decomposition $(T,\{B_{t}\},\{C_{t}\})$, where:
  • $T$ is a tree with vertex bags $B_{t}\subseteq V$ and edge bags $C_{t}\subseteq E$.
  • The width of the decomposition is:
    $\text{width}(T,\{B_{t}\},\{C_{t}\})=\max_{t\in V(T)}\left(\sup_{v\in B_{t}}\left(\alpha_{B}(v)+\beta_{B}(v)+\gamma_{B}(v)\right)-1\right).$
  • The Single-Valued Neutrosophic Hypertree-Width is the minimum width over all possible decompositions:
[…]
E.8 Vague tree-width
Vague graphs extend traditional graphs by incorporating true and false membership functions, enabling representation of uncertainty and imprecision in relationships between vertices[1019, 1135, 1053, 193, 195, 31, 194, 192, 1017, 1022]. We provide the definition of a vague graph below.
Definition E.27.
Let $G^{*}=(V,E)$ be a crisp graph, where $V$ is the set of vertices and $E\subseteq V\times V$ is the set of edges. A vague graph $G=(A,B)$ is defined as follows:
1. For each vertex $v\in V$, two membership functions are defined:
  • $t_{A}(v):V\to[0,1]$, the true membership function,
  • $f_{A}(v):V\to[0,1]$, the false membership function.
These functions satisfy:
$0\leq t_{A}(v)+f_{A}(v)\leq 1$
for all $v\in V$. The interval $[t_{A}(v),1-f_{A}(v)]$ represents the degree of membership of $v$ in the vague set $A$.
2. For each edge $e=(v_{i},v_{j})\in E$, two membership functions are defined:
  • $t_{B}(e):E\to[0,1]$, the true membership function for the edge,
  • $f_{B}(e):E\to[0,1]$, the false membership function for the edge.
These functions satisfy:
$t_{B}(v_{i},v_{j})\leq\min(t_{A}(v_{i}),t_{A}(v_{j})),$
$f_{B}(v_{i},v_{j})\geq\max(f_{A}(v_{i}),f_{A}(v_{j}))$
for all $e=(v_{i},v_{j})\in E$.
3. The order of the vague graph $G$ is defined as:
$O(G)=\left(\sum_{v\in V}t_{A}(v),\sum_{v\in V}f_{A}(v)\right)$
4. The size of the vague graph $G$ is defined as:
$S(G)=\left(\sum_{(v_{i},v_{j})\in E}t_{B}(v_{i},v_{j}),\sum_{(v_{i},v_{j})\in E}f_{B}(v_{i},v_{j})\right)$
5. The open degree of a vertex $u\in V$ is defined as:
$d(u)=\left(d_{t}(u),d_{f}(u)\right)$
where
$d_{t}(u)=\sum_{u\neq v,v\in V}t_{B}(u,v),\quad d_{f}(u)=\sum_{u\neq v,v\in V}f_{B}(u,v)$
6. The vague graph $G$ is called $n$-regular if all vertices have the same open degree $n$.
Example E.28.
Consider the vague graph $G_{1}$ with three vertices $v_{1},v_{2},v_{3}$ and three edges as shown in the figure below.
$t_{A}(v_{1})=0.2,\quad f_{A}(v_{1})=0.3$
$t_{A}(v_{2})=0.2,\quad f_{A}(v_{2})=0.2$
$t_{A}(v_{3})=0.2,\quad f_{A}(v_{3})=0.3$
The edges have the following true and false membership values:
$t_{B}(v_{1},v_{2})=0.1,\quad f_{B}(v_{1},v_{2})=0.3$
$t_{B}(v_{2},v_{3})=0.1,\quad f_{B}(v_{2},v_{3})=0.6$
$t_{B}(v_{3},v_{1})=0.1,\quad f_{B}(v_{3},v_{1})=0.4$
Although still in the conceptual stage, the definition is provided below.
Definition E.29.
A Vague Tree is a vague graph $G=(A,B)$, where:
  1. 1.
    For each vertex $v\in V$, there are two membership functions:
    • True membership function $t_{A}(v)\in[0,1]$,
    • False membership function $f_{A}(v)\in[0,1]$,
    satisfying:
    $0\leq t_{A}(v)+f_{A}(v)\leq 1$
  2. 2.
    For each edge $e=(v_{i},v_{j})\in E$, there are two membership functions:
    • True membership function $t_{B}(e)\in[0,1]$,
    • False membership function $f_{B}(e)\in[0,1]$,
    satisfying:
    $t_{B}(e)\leq\min(t_{A}(v_{i}),t_{A}(v_{j})),$
    $f_{B}(e)\geq\max(f_{A}(v_{i}),f_{A}(v_{j})).$
  3. 3.
    The vague tree $G=(A,B)$ is connected and acyclic, like a crisp tree.
Definition E.30.
The Vague Tree-Width of a vague tree $G=(A,B)$ is defined by a tree-decomposition $(T,\{W_{t}\})$, where:
  • $T$ is a crisp tree, and $\{W_{t}\}$ is a collection of vague subsets $W_{t}\subseteq V$.
  • The width of the decomposition is:
    $\text{width}(T,\{W_{t}\})=\max_{t\in V(T)}\left(\sup_{v\in W_{t}}\left(t_{A}(v)+f_{A}(v)\right)-1\right).$
  • The Vague Tree-Width is the minimum width over all tree-decompositions:
[…]
E.9 Vague hypertree-width
Vague hypergraphs extend traditional hypergraphs by incorporating truth and falsity membership functions, enabling the representation of uncertain, imprecise, or vague data in complex systems[1018, 36, 1018]. Similar to vague graphs, vague hypergraphs have been the subject of extensive research.
Definition E.31.
[36] Let $V$ be a finite set of vertices, and let $E=\{E_{1},E_{2},\dots,E_{m}\}$ be a finite family of non-trivial vague subsets of $V$, where each vague subset is defined by true and false membership functions. A vague hypergraph is a pair $H=(V,E)$, where:
1. The set $V$ is the crisp vertex set of the hypergraph.
2. Each hyperedge $E_{j}\in E$ is a vague subset of $V$, represented by a pair $(t_{E_{j}},f_{E_{j}})$, where:
  • $t_{E_{j}}:V\to[0,1]$, the true membership function,
  • $f_{E_{j}}:V\to[0,1]$, the false membership function,
These functions satisfy:
$0\leq t_{E_{j}}(v)+f_{E_{j}}(v)\leq 1$
for all $v\in V$ and $E_{j}\in E$.
3. The support of a vague hyperedge $E_{j}$ is defined as:
$\text{supp}(E_{j})=\{v\in V\ |\ t_{E_{j}}(v)\neq 0\text{ or }f_{E_{j}}(v)\neq 0\}$
The vague hypergraph satisfies:
$V=\bigcup_{j=1}^{m}\text{supp}(E_{j})$
4. A vague hypergraph $H=(V,E)$ is called elementary if each vague hyperedge is single-valued on its support.
5. A vague hypergraph $H=(V,E)$ is called simple if for any two hyperedges $A=(t_{A},f_{A})$ and $B=(t_{B},f_{B})$ in $E$, $t_{A}\leq t_{B}$ and $f_{A}\geq f_{B}$ imply that $t_{A}=t_{B}$ and $f_{A}=f_{B}$.
6. A vague hypergraph $H=(V,E)$ is called strongly support simple if for any two vague hyperedges $A=(t_{A},f_{A})$ and $B=(t_{B},f_{B})$, $\text{supp}(A)=\text{supp}(B)$ implies that $A=B$.
Although still in the conceptual stage, the definition is provided below.
Definition E.32.
A Vague Hypertree is a vague hypergraph $H=(V,E)$, where:
  1. 1.
    $V$ is a finite set of vertices.
  2. 2.
    $E=\{E_{1},E_{2},\dots,E_{m}\}$ is a finite family of vague hyperedges, each represented by true and false membership functions $t_{E_{j}}(v),f_{E_{j}}(v)$ for all $v\in V$, satisfying:
    $0\leq t_{E_{j}}(v)+f_{E_{j}}(v)\leq 1\quad\text{for all }v\in V\text{ and }E_{j}\in E.$
  3. 3.
    The support of a vague hyperedge $E_{j}$ is defined as:
    $\text{supp}(E_{j})=\{v\in V\ |\ t_{E_{j}}(v)\neq 0\text{ or }f_{E_{j}}(v)\neq 0\}.$
The vague hypertree must be connected and acyclic.
Definition E.33.
A Vague Hypertree Decomposition of a vague hypergraph $H=(V,E)$ is a triple $(T,\{B_{t}\},\{C_{t}\})$, where:
  • $T$ is a tree with vertex set $V(T)$ and edge set $E(T)$.
  • $B_{t}\subseteq V$ is a vague subset of vertices associated with each node $t\in V(T)$, with true and false membership functions $t_{B}(v),f_{B}(v)$, satisfying:
    $0\leq t_{B}(v)+f_{B}(v)\leq 1$
  • $C_{t}\subseteq E$ is a vague subset of hyperedges associated with each node $t\in V(T)$.
The decomposition must satisfy:
  1. 1.
    Vertex Coverage: For each vertex $v\in V$, the set $\{t\in V(T)\ |\ v\in B_{t}\}$ forms a connected subtree of $T$.
  2. 2.
    Edge Coverage: For each vague hyperedge $E_{j}\in E$, there exists a node $t\in V(T)$ such that $\text{supp}(E_{j})\subseteq B_{t}$.
The width of a vague hypertree decomposition is:
$\text{width}(T,\{B_{t}\},\{C_{t}\})=\max_{t\in V(T)}\left(\sup_{v\in B_{t}}\left(t_{B}(v)+f_{B}(v)\right)-1\right).$
The Vague Hypertree-Width of $H$ is the minimum width over all vague hypertree decompositions:
[…]
E.10 Intuitionistic Fuzzy Tree-width
An intuitionistic fuzzy graph (IFG) extends a fuzzy graph by assigning both membership and non-membership degrees to vertices and edges, allowing for uncertainty and hesitation in defining graph properties[28, 1021, 809, 1031, 998, 88, 977, 122, 738]. The definitions of intuitionistic fuzzy graph and intuitionistic fuzzy tree are presented as follows.
Definition E.34.
Let $G=(V,E)$ be a graph where $V$ is the set of vertices and $E\subseteq V\times V$ is the set of edges. An intuitionistic fuzzy graph (IFG) is defined by the pair $G=(V,E)$, along with two functions $\mu_{1}:V\to[0,1]$ and $\nu_{1}:V\to[0,1]$ representing the degree of membership and non-membership of the vertices, and two functions $\mu_{2}:E\to[0,1]$ and $\nu_{2}:E\to[0,1]$ representing the degree of membership and non-membership of the edges.
The following conditions hold for all vertices $v\in V$ and edges $(v_{i},v_{j})\in E$:
  1. 1.
    For each vertex $v\in V$, $0\leq\mu_{1}(v)+\nu_{1}(v)\leq 1$.
  2. 2.
    For each edge $(v_{i},v_{j})\in E$:
    • $\mu_{2}(v_{i},v_{j})\leq\min(\mu_{1}(v_{i}),\mu_{1}(v_{j}))$,
    • $\nu_{2}(v_{i},v_{j})\geq\max(\nu_{1}(v_{i}),\nu_{1}(v_{j}))$,
    • $0\leq\mu_{2}(v_{i},v_{j})+\nu_{2}(v_{i},v_{j})\leq 1$.
An intuitionistic fuzzy subgraph (IFSG) of $G=(V,E)$ is a subgraph $H=(V^{\prime},E^{\prime})$, where $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$, such that:
  • For all $v\in V^{\prime}$, $\mu_{1}(v)$ and $\nu_{1}(v)$ are the same as in $G$,
  • For all $(v_{i},v_{j})\in E^{\prime}$, $\mu_{2}(v_{i},v_{j})$ and $\nu_{2}(v_{i},v_{j})$ are the same as in $G$.
An intuitionistic spanning fuzzy subgraph (ISFS) is an IFSG where $V^{\prime}=V$.
An intuitionistic fuzzy graph $G=(V,E)$ is said to be:
  • A $\mu$-tree if the underlying crisp graph $G^{*}_{\mu}=(V^{*}_{\mu},E^{*}_{\mu})$ formed by considering only the membership degrees $\mu$ is a tree (i.e., a connected, acyclic graph).
  • A $\nu$-tree if the underlying crisp graph $G^{*}_{\nu}=(V^{*}_{\nu},E^{*}_{\nu})$ formed by considering only the non-membership degrees $\nu$ is a tree.
An intuitionistic fuzzy graph $G=(V,E)$ is called an intuitionistic fuzzy tree if it satisfies the following properties:
  • $G$ is both a $\mu$-tree and a $\nu$-tree.
  • The underlying graphs $G^{*}_{\mu}$ and $G^{*}_{\nu}$ are equal, i.e., $G^{*}_{\mu}=G^{*}_{\nu}$, meaning that the structure remains consistent when viewed from both the membership and non-membership perspectives.
Here is the definition, although it is still in the conceptual stage.
Definition E.35.
Let $G=(V,E,\mu_{1},\nu_{1},\mu_{2},\nu_{2})$ be an intuitionistic fuzzy graph. An intuitionistic fuzzy tree-decomposition of $G$ is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a tree with nodes $I$ and edges $F$,
  • $\{B_{t}\}_{t\in I}$ is a collection of subsets of $V$ (called bags) associated with the nodes of $T$, such that:
    1. 1.
      For each vertex $v\in V$, the set $\{t\in I:v\in B_{t}\}$ is connected in the tree $T$.
    2. 2.
      For each edge $(v_{i},v_{j})\in E$, there exists at least one node $t\in I$ such that both $v_{i}$ and $v_{j}$ belong to the same bag $B_{t}$, and the membership and non-membership degrees satisfy:
      $\mu_{2}(v_{i},v_{j})\leq\min\{\mu_{1}(v_{i}),\mu_{1}(v_{j})\},\quad\nu_{2}(v_{i},v_{j})\geq\max\{\nu_{1}(v_{i}),\nu_{1}(v_{j})\}.$
The width of an intuitionistic fuzzy tree-decomposition is defined as:
$\text{width}=\max_{t\in I}\left(\sup_{v\in B_{t}}(\mu_{1}(v)+\nu_{1}(v))-1\right),$
where $\mu_{1}(v)$ and $\nu_{1}(v)$ are the membership and non-membership degrees of vertex $v$ in bag $B_{t}$.
The intuitionistic fuzzy tree-width of the graph $G$ is the minimum width among all possible intuitionistic fuzzy tree-decompositions of $G$.
Definition E.36.
Let $G=(V,E,\mu_{1},\nu_{1},\mu_{2},\nu_{2})$ be an intuitionistic fuzzy graph. A path-decomposition is a special case of a tree-decomposition where the underlying tree $T$ is a path. Specifically, it is a pair $(P,\{B_{p}\}_{p\in I})$, where:
  • $P=(I,F)$ is a path with nodes $I$ and edges $F$,
  • $\{B_{p}\}_{p\in I}$ is a collection of bags associated with the nodes of $P$, such that:
    1. 1.
      For each vertex $v\in V$, the set $\{p\in I:v\in B_{p}\}$ is connected in the path $P$,
    2. 2.
      For each edge $(v_{i},v_{j})\in E$, there exists at least one node $p\in I$ such that both $v_{i}$ and $v_{j}$ belong to the same bag $B_{p}$, and the membership and non-membership degrees satisfy:
      $\mu_{2}(v_{i},v_{j})\leq\min\{\mu_{1}(v_{i}),\mu_{1}(v_{j})\},\quad\nu_{2}(v_{i},v_{j})\geq\max\{\nu_{1}(v_{i}),\nu_{1}(v_{j})\}.$
The width of an intuitionistic fuzzy path-decomposition is defined as:
$\text{width}=\max_{p\in I}\left(\sup_{v\in B_{p}}(\mu_{1}(v)+\nu_{1}(v))-1\right).$
The intuitionistic fuzzy path-width of the graph $G$ is the minimum width among all possible intuitionistic fuzzy path-decompositions of $G$.
E.11 Dombi FuzzyTree-width
A Dombi fuzzy graph uses Dombi’s t-norm to model fuzzy relationships between vertices, representing edge memberships in terms of degrees, making it suitable for complex networks and decision-making processes[85, 693, 633, 44, 234, 678, 32, 271, 29]. The definitions are presented as follows.
Definition E.37 (Dombi Fuzzy Graph).
[85] Let $V$ be a finite set of vertices. A Dombi fuzzy graph $G=(\eta,\zeta)$ consists of:
  • $\eta:V\rightarrow[0,1]$, a function assigning a degree of membership to each vertex $v\in V$,
  • $\zeta:V\times V\rightarrow[0,1]$, a symmetric fuzzy relation assigning a degree of membership to each edge $(x,y)\in V\times V$.
For each pair of vertices $x,y\in V$, the degree of edge membership $\zeta(x,y)$ must satisfy the following condition based on Dombi’s t-norm:
$\zeta(x,y)\leq\frac{\eta(x)\eta(y)}{\eta(x)+\eta(y)-\eta(x)\eta(y)},$
where $\eta(x)$ and $\eta(y)$ are the degrees of membership of vertices $x$ and $y$, respectively.
Example E.38.
(cf.[85]) Consider a Dombi fuzzy graph $G=(\eta,\zeta)$ over the set of vertices $V=\{v_{1},v_{2},v_{3},v_{4}\}$ with the following membership degrees for the vertices:
$\eta(v_{1})=0.4,\quad\eta(v_{2})=0.5,\quad\eta(v_{3})=0.9,\quad\eta(v_{4})=0.7.$
The degree of membership for the edges is given as:
$\zeta(v_{1},v_{2})=0.1,\quad\zeta(v_{2},v_{3})=0.4,\quad\zeta(v_{1},v_{3})=0.3,\quad\zeta(v_{2},v_{4})=0.2.$
We now check whether these values satisfy the Dombi t-norm condition for edge memberships, which requires that for each pair $(x,y)\in V\times V$, the following inequality holds:
$\zeta(x,y)\leq\frac{\eta(x)\eta(y)}{\eta(x)+\eta(y)-\eta(x)\eta(y)}.$
For $\zeta(v_{1},v_{2})=0.1$, we have:
$\frac{\eta(v_{1})\eta(v_{2})}{\eta(v_{1})+\eta(v_{2})-\eta(v_{1})\eta(v_{2})}=\frac{0.4\times 0.5}{0.4+0.5-0.4\times 0.5}=\frac{0.2}{0.7}\approx 0.286.$
Since $0.1\leq 0.286$, this condition is satisfied.
For $\zeta(v_{2},v_{3})=0.4$, we have:
$\frac{\eta(v_{2})\eta(v_{3})}{\eta(v_{2})+\eta(v_{3})-\eta(v_{2})\eta(v_{3})}=\frac{0.5\times 0.9}{0.5+0.9-0.5\times 0.9}=\frac{0.45}{1.4-0.45}=\frac{0.45}{0.95}\approx 0.474.$
Since $0.4\leq 0.474$, this condition is also satisfied.
For $\zeta(v_{1},v_{3})=0.3$, we have:
$\frac{\eta(v_{1})\eta(v_{3})}{\eta(v_{1})+\eta(v_{3})-\eta(v_{1})\eta(v_{3})}=\frac{0.4\times 0.9}{0.4+0.9-0.4\times 0.9}=\frac{0.36}{1.3-0.36}=\frac{0.36}{0.94}\approx 0.383.$
Since $0.3\leq 0.383$, this condition is satisfied.
For $\zeta(v_{2},v_{4})=0.2$, we have:
$\frac{\eta(v_{2})\eta(v_{4})}{\eta(v_{2})+\eta(v_{4})-\eta(v_{2})\eta(v_{4})}=\frac{0.5\times 0.7}{0.5+0.7-0.5\times 0.7}=\frac{0.35}{1.2-0.35}=\frac{0.35}{0.85}\approx 0.412.$
Since $0.2\leq 0.412$, this condition is satisfied.
Thus, all the edge memberships in the given Dombi fuzzy graph satisfy the Dombi t-norm condition.
Definition E.39 (Dombi Fuzzy Tree).
A Dombi fuzzy tree is a Dombi fuzzy graph $G=(\eta,\zeta)$ that satisfies the following properties:
  • $G$ is connected: for any pair of vertices $x,y\in V$, there exists a path of edges with positive membership degrees between $x$ and $y$.
  • $G$ is acyclic: there is no sequence of edges $(e_{1},e_{2},\dots,e_{n})$ such that the first and last vertices coincide, and all edges have positive membership degrees.
  • The graph $G$ forms a tree-like structure in its crisp version, where edges are defined using a threshold on the membership degrees.
The Dombi fuzzy tree-width extends classical tree-width to Dombi fuzzy graphs, incorporating both the vertex and edge membership degrees in the decomposition. Here is the definition, although it is still in the conceptual stage.
Definition E.40 (Dombi Fuzzy Tree-Decomposition).
Let $G=(\eta,\zeta)$ be a Dombi fuzzy graph. A Dombi fuzzy tree-decomposition is a pair $(T,\{B_{t}\}_{t\in I})$, where:
  • $T=(I,F)$ is a tree with nodes $I$ and edges $F$,
  • $\{B_{t}\}_{t\in I}$ is a collection of fuzzy subsets of $V$ (called bags), such that:
    1. 1.
      For each vertex $v\in V$, the set $\{t\in I:v\in B_{t}\}$ is connected in $T$,
    2. 2.
      For each edge $(x,y)\in V\times V$, there exists at least one node $t\in I$ such that both $x$ and $y$ are in $B_{t}$, and the fuzzy membership degree $\zeta(x,y)$ satisfies:
      $\zeta(x,y)\leq\frac{\eta(x)\eta(y)}{\eta(x)+\eta(y)-\eta(x)\eta(y)}.$
Definition E.41 (Dombi Fuzzy Tree-Width).
The Dombi fuzzy tree-width of a Dombi fuzzy graph $G=(\eta,\zeta)$ is defined as:
$\text{width}(T)=\max_{t\in I}\left(\sup_{v\in B_{t}}\eta(v)-1\right),$
where $\eta(v)$ represents the membership degree of vertex $v$ in the fuzzy bag $B_{t}$. The Dombi fuzzy tree-width of $G$ is the minimum width over all possible Dombi fuzzy tree-decompositions.
The Dombi fuzzy path-width extends the concept of path-width to Dombi fuzzy graphs by incorporating fuzzy memberships into the decomposition where the underlying structure is a path rather than a tree. Here is the definition, although it is still in the conceptual stage.
Definition E.42 (Dombi Fuzzy Path-Decomposition).
A Dombi fuzzy path-decomposition of a Dombi fuzzy graph $G=(\eta,\zeta)$ is a special case of the tree-decomposition, where the underlying structure $T$ is a path. Specifically, it is a pair $(P,\{B_{p}\}_{p\in P})$, where:
  • $P=(I,F)$ is a path with nodes $I$ and edges $F$,
  • $\{B_{p}\}_{p\in I}$ is a collection of fuzzy subsets of $V$ (called bags), such that:
    1. 1.
      For each vertex $v\in V$, the set $\{p\in I:v\in B_{p}\}$ is connected in $P$,
    2. 2.
      For each edge $(x,y)\in V\times V$, there exists some node $p\in I$ such that both $x$ and $y$ belong to $B_{p}$, and the membership degree $\zeta(x,y)$ satisfies:
      $\zeta(x,y)\leq\frac{\eta(x)\eta(y)}{\eta(x)+\eta(y)-\eta(x)\eta(y)}.$
Definition E.43 (Dombi Fuzzy Path-Width).
The Dombi fuzzy path-width of a Dombi fuzzy graph $G=(\eta,\zeta)$ is defined as:
$\text{width}(P)=\max_{p\in I}\left(\sup_{v\in B_{p}}\eta(v)-1\right),$
where $\eta(v)$ is the membership degree of vertex $v$ in $B_{p}$. The Dombi fuzzy path-width of $G$ is the minimum width over all possible Dombi fuzzy path-decompositions.
E.12 Picture Fuzzy Tree-width
A Picture Fuzzy Graph (PFG) is a generalization of fuzzy graphs that incorporates three membership degrees: positive membership, neutral membership, and negative membership(cf.[1086, 1045, 760, 1093, 66, 65, 1094, 1087, 1254, 1222]). It is defined as follows:
Definition E.44.
Let $G^{*}=(V,E)$ be a classical (crisp) graph where $V$ is the set of vertices and $E\subseteq V\times V$ is the set of edges. A Picture Fuzzy Graph is a pair $G=(A,B)$, where:
  • $A=(\mu_{A},\eta_{A},\nu_{A})$ is a picture fuzzy set on the vertex set $V$, with the following properties:
    • $\mu_{A}:V\to[0,1]$ represents the positive membership degree of a vertex.
    • $\eta_{A}:V\to[0,1]$ represents the neutral membership degree of a vertex.
    • $\nu_{A}:V\to[0,1]$ represents the negative membership degree of a vertex.
    • These functions satisfy the condition:
      $0\leq\mu_{A}(v)+\eta_{A}(v)+\nu_{A}(v)\leq 1\quad\forall v\in V.$
  • $B=(\mu_{B},\eta_{B},\nu_{B})$ is a picture fuzzy relation on the edge set $E$, where:
    • $\mu_{B}:E\to[0,1]$ represents the positive membership degree of an edge.
    • $\eta_{B}:E\to[0,1]$ represents the neutral membership degree of an edge.
    • $\nu_{B}:E\to[0,1]$ represents the negative membership degree of an edge.
    • These functions satisfy the following conditions for all $(u,v)\in E$:
      $\mu_{B}(u,v)\leq\min(\mu_{A}(u),\mu_{A}(v)),$
      $\eta_{B}(u,v)\leq\min(\eta_{A}(u),\eta_{A}(v)),$
      $\nu_{B}(u,v)\geq\max(\nu_{A}(u),\nu_{A}(v)),$
      ensuring consistency between the edge and vertex membership degrees.
Additionally, the refusal membership degree for each vertex $v\in V$ and edge $(u,v)\in E$ is given by:
$\pi_{A}(v)=1-(\mu_{A}(v)+\eta_{A}(v)+\nu_{A}(v)),$
$\pi_{B}(u,v)=1-(\mu_{B}(u,v)+\eta_{B}(u,v)+\nu_{B}(u,v)).$
Example E.45.
(cf.[1254]) Consider a Picture Fuzzy Graph $G=(A,B)$ defined over the vertex set $V=\{v_{1},v_{2},v_{3}\}$ and edge set $E=\{(v_{1},v_{2}),(v_{2},v_{3}),(v_{1},v_{3})\}$. The picture fuzzy membership degrees for each vertex are given as follows:
  • $v_{1}:(\mu_{A}(v_{1}),\eta_{A}(v_{1}),\nu_{A}(v_{1}))=(0.3,0.5,0.2)$
  • $v_{2}:(\mu_{A}(v_{2}),\eta_{A}(v_{2}),\nu_{A}(v_{2}))=(0.5,0.2,0.3)$
  • $v_{3}:(\mu_{A}(v_{3}),\eta_{A}(v_{3}),\nu_{A}(v_{3}))=(0.4,0.2,0.3)$
For the edges, the picture fuzzy membership degrees are as follows:
  • $(v_{1},v_{2}):(\mu_{B}(v_{1},v_{2}),\eta_{B}(v_{1},v_{2}),\nu_{B}(v_{1},v_{2}))=(0.3,0.1,0.25)$
  • $(v_{2},v_{3}):(\mu_{B}(v_{2},v_{3}),\eta_{B}(v_{2},v_{3}),\nu_{B}(v_{2},v_{3}))=(0.3,0.1,0.1)$
  • $(v_{1},v_{3}):(\mu_{B}(v_{1},v_{3}),\eta_{B}(v_{1},v_{3}),\nu_{B}(v_{1},v_{3}))=(0.3,0.1,0.1)$
The picture fuzzy graph satisfies the following conditions:
$\mu_{B}(v_{1},v_{2})\leq\min\bigl{(}\mu_{A}(v_{1}),\mu_{A}(v_{2})\bigr{)},\\ \eta_{B}(v_{1},v_{2})\leq\min\bigl{(}\eta_{A}(v_{1}),\eta_{A}(v_{2})\bigr{)},\\ \nu_{B}(v_{1},v_{2})\geq\max\bigl{(}\nu_{A}(v_{1}),\nu_{A}(v_{2})\bigr{)},$
and similarly for the other edges, ensuring consistency between the vertex and edge membership degrees.
Thus, the Picture Fuzzy Graph $G$ represents a triangular structure with the given membership degrees for both vertices and edges.
Next, the definition of a Picture Fuzzy Tree is introduced below. A Picture Fuzzy Tree is a graph in which each vertex and edge is assigned three fuzzy membership values: positive, neutral, and negative. The graph must adhere to the properties of a tree (i.e., being connected and acyclic) within the framework of picture fuzzy sets.
Definition E.46.
Let $G=(A,B)$ be a Picture Fuzzy Graph where:
  • $A=(\mu_{A},\eta_{A},\nu_{A})$ is the picture fuzzy set on the vertex set $V$, where:
    • $\mu_{A}:V\to[0,1]$ is the positive membership degree,
    • $\eta_{A}:V\to[0,1]$ is the neutral membership degree,
    • $\nu_{A}:V\to[0,1]$ is the negative membership degree.
  • $B=(\mu_{B},\eta_{B},\nu_{B})$ is the picture fuzzy set on the edge set $E\subseteq V\times V$, where:
    • $\mu_{B}:E\to[0,1]$ is the positive membership degree for edges,
    • $\eta_{B}:E\to[0,1]$ is the neutral membership degree for edges,
    • $\nu_{B}:E\to[0,1]$ is the negative membership degree for edges.
A Picture Fuzzy Tree satisfies:
  1. 1.
    Acyclicity: The graph contains no cycles. Formally, there is no cycle where the total membership (positive, neutral, or negative) of any path forms a closed loop.
  2. 2.
    Connectedness: Any two vertices $u,v\in V$ are connected by a path in $G$ such that:
    $\mu_{B}(u,v)>0\quad\text{or}\quad\eta_{B}(u,v)>0\quad\text{or}\quad\nu_{B}(u,v)>0.$
  3. 3.
    Membership Conditions: For any edge $(u,v)\in E$:
    $\mu_{B}(u,v)\leq\min(\mu_{A}(u),\mu_{A}(v)),\quad\eta_{B}(u,v)\leq\min(\eta_{A}(u),\eta_{A}(v)),\quad\nu_{B}(u,v)\geq\max(\nu_{A}(u),\nu_{A}(v)).$
The Picture Fuzzy Tree-Width generalizes the classical tree-width by incorporating picture fuzzy membership values. It measures the "tree-likeness" of a picture fuzzy graph by associating the graph with a tree structure where the bags contain fuzzy sets of vertices. Although still in the conceptual stage, the definition is provided below.
Definition E.47.
A Picture Fuzzy Tree-Decomposition of a graph $G=(A,B)$ is defined as a pair $(T,\{B_{t}\}_{t\in T})$, where:
  • $T=(I,F)$ is a tree,
  • $\{B_{t}\}_{t\in T}$ is a collection of picture fuzzy subsets of $V$ (called bags), where each bag contains positive, neutral, and negative membership degrees:
    $B_{t}=(\mu_{B_{t}},\eta_{B_{t}},\nu_{B_{t}}).$
The decomposition satisfies:
  1. 1.
    Connectivity: For each vertex $v\in V$, the set of nodes $\{t\in I:v\in B_{t}\}$ forms a connected subtree of $T$.
  2. 2.
    Edge Coverage: For each edge $(u,v)\in E$, there exists a node $t\in T$ such that both $u$ and $v$ belong to $B_{t}$, and the fuzzy membership degrees satisfy:
    $\mu_{B}(u,v)\leq\min(\mu_{B_{t}}(u),\mu_{B_{t}}(v)),\quad\eta_{B}(u,v)\leq\min(\eta_{B_{t}}(u),\eta_{B_{t}}(v)),\quad\nu_{B}(u,v)\geq\max(\nu_{B_{t}}(u),\nu_{B_{t}}(v)).$
The width of a picture fuzzy tree-decomposition is:
$\text{width}=\max_{t\in T}\left(\sup_{v\in B_{t}}(\mu_{B_{t}}(v)+\eta_{B_{t}}(v)+\nu_{B_{t}}(v))-1\right).$
The Picture Fuzzy Tree-Width of $G$ is the minimum width among all possible tree-decompositions of $G$.
The Picture Fuzzy Path-Width is a special case of the Picture Fuzzy Tree-Width, where the underlying tree in the decomposition is a path.
Definition E.48.
A Picture Fuzzy Path-Decomposition of a graph $G=(A,B)$ is a pair $(P,\{B_{p}\}_{p\in P})$, where:
  • $P=(I,F)$ is a path,
  • $\{B_{p}\}_{p\in P}$ is a collection of picture fuzzy subsets of $V$ (called bags), where each bag $B_{p}=(\mu_{B_{p}},\eta_{B_{p}},\nu_{B_{p}})$.
The decomposition satisfies:
  1. 1.
    Connectivity: For each vertex $v\in V$, the set $\{p\in I:v\in B_{p}\}$ forms a connected sub-path of $P$.
  2. 2.
    Edge Coverage: For each edge $(u,v)\in E$, there exists a node $p\in I$ such that both $u$ and $v$ belong to $B_{p}$, and the fuzzy membership conditions are satisfied.
The width of a picture fuzzy path-decomposition is:
$\text{width}=\max_{p\in I}\left(\sup_{v\in B_{p}}(\mu_{B_{p}}(v)+\eta_{B_{p}}(v)+\nu_{B_{p}}(v))-1\right).$
The Picture Fuzzy Path-Width of $G$ is the minimum width among all possible path-decompositions of $G$.
Appendix F Graph Games related to Width Parameters
As frequently discussed throughout this paper, width parameters have been examined in various games. The interpretation of these parameters through game theory is highly beneficial for enhancing the intuitive understanding of width and its obstructions. Below, we provide examples of common games where such interpretations are frequently explored.
F.1 Cops and Robbers (Pursuit-evasion games)
Cops and Robbers is a pursuit-evasion game on a graph, where cops aim to capture a robber. Numerous variations, including visibility and strategy constraints, have been studied[529, 792, 286, 1190, 466, 892, 4, 714, 19, 891, 424, 1136, 433, 534, 326, 396, 62]. An example of the definition is presented below.
Definition F.1 (Helicopter Cops and Robber Game (Jump-Searching Version)).
[1072] Let $X\subseteq V(G)$. An X-flap is the vertex set of a connected component of $G-X$. Two subsets $X,Y\subseteq V(G)$ touch if $N(X)\cap Y\neq\emptyset$. A position is a pair $(X,R)$, where $X\subseteq V(G)$ and $R$ is an $X$-flap. Here, $X$ represents the set of vertices currently occupied by the cops, and $R$ indicates the component of $G-X$ where the robber is located—since the robber can move arbitrarily fast, only the component containing her matters.
At the start of the game, the cops choose an initial subset $X_{0}$, and the robber chooses an $X_{0}$-flap $R_{0}$. If there are $k$ cops in the game, then $|X_{0}|\leq k$.
At the beginning of round $i$, the game is in position $(X_{i-1},R_{i-1})$. The cops then choose a new set $X_{i}\subseteq V(G)$ such that $|X_{i}|\leq k$, with no other restrictions, and announce their choice. The robber, knowing $X_{i}$, selects an $X_{i}$-flap $R_{i}$ that touches $R_{i-1}$. If the robber cannot choose such an $R_{i}$, the cops win. Otherwise, if the robber always has a valid move, she wins.
F.2 Cops and invisible robbers
A "Cops and Invisible Robber" game is a pursuit-evasion game where cops try to capture a robber who is invisible to them, with strategic movement on a graph[325, 754, 1032, 294]. Various studies explore its relation to graph width parameters, such as pathwidth.
Definition F.2 (Zero-Visibility Cops and Robber Game).
[325] The zero-visibility cops and robber game is a pursuit-evasion game played on a simple connected graph $G$ between two players: the cop player and the robber player. The cop player controls a fixed number $k$ of cops, and the robber player controls a single robber. The goal of the cop player is to capture the robber by moving one of the cops to the same vertex as the robber, while the robber tries to evade capture indefinitely.
Rules:
  • Setup: The game is played on a simple connected graph $G=(V(G),E(G))$. The cop player places $k$ cops on vertices of $G$, and the robber is placed on a vertex, unknown to the cops.
  • Gameplay: The players alternate turns, with the cops moving first. The cop player may move one or more cops to adjacent vertices, and the robber may move to an adjacent vertex. The cops win if they land on the same vertex as the robber. The robber wins if this never occurs.
  • Visibility: The cops have zero visibility of the robber’s location and movements, while the robber has full visibility of the cops’ moves.
  • Strategies: A strategy for the cops is defined as a set of walks $\mathcal{O}=\{l_{i}\}_{i=1}^{k}$, representing the movement of each cop. A strategy is successful if it guarantees the capture of the robber regardless of the robber’s movements.
  • Winning Condition: The cops win if a cop occupies the same vertex as the robber. The robber wins if it can avoid capture indefinitely.
  • Zero-Visibility Cop Number: The zero-visibility cop number of $G$, denoted $c_{0}(G)$, is the minimum number of cops required to guarantee capture of the robber.
F.3 Parity games
A parity game is a two-player game on a directed graph where players (Even and Odd) move a token, aiming to achieve a winning condition based on the minimum recurring vertex priority (even or odd). Extensive research has been conducted, particularly on Directed Width and related topics[130, 1117, 1007, 439, 956, 440, 541, 756, 1006, 958].
Definition F.3 (Parity Game).
[130] A parity game is a tuple $P=(V,V_{0},E,\Omega)$, where:
  • $V$ is a finite set of vertices,
  • $V_{0}\subseteq V$ is the set of vertices controlled by player Even (the remaining vertices $V\setminus V_{0}$ are controlled by player Odd),
  • $E\subseteq V\times V$ is a set of directed edges,
  • $\Omega:V\to\omega$ is a priority function that assigns a non-negative integer to each vertex.
Two players, Even and Odd, take turns moving a token along the edges of the graph. A play of the game is an infinite sequence $\pi=(v_{0},v_{1},v_{2},\dots)$, where $(v_{i},v_{i+1})\in E$ for all $i\geq 0$.
Winning condition: A play $\pi$ is winning for Even if
$\liminf_{i\to\infty}\Omega(v_{i})\text{ is even}.$
Otherwise, it is winning for Odd if
$\liminf_{i\to\infty}\Omega(v_{i})\text{ is odd}.$
A strategy for Even is a function $f:V^{<\omega}\to V$, and a strategy is memoryless if it depends only on the current vertex. Parity games are determined, meaning one of the players always has a winning strategy, and memoryless strategies always exist.
F.4 Spanning-tree games
A spanning tree is a subgraph that includes all the vertices of the original graph with the minimum number of edges, forming no cycles(cf.[590, 988]). The Spanning-Tree Game is a two-player game where players alternately select edges to form a spanning tree in a weighted graph, aiming to maximize or minimize the total weight. Spanning-tree games have been widely studied[649, 934] and are also related to graph width parameters[244]. Tree-width measures a graph’s decomposition complexity, which influences strategies for efficient spanning tree construction in the Spanning-Tree Game.
Definition F.4 (Spanning-Tree Game).
[244] The Spanning-Tree Game is played on a weighted graph $G=(V,E,w)$, where $V$ is the set of vertices, $E$ is the set of edges, and $w:E\to\mathbb{R}^{+}$ is a weight function that assigns a non-negative real number to each edge.
Game Setup:
  • A configuration is a forest $F\subseteq E$ in the graph $G$.
  • The set of legal moves from a configuration $F$ is given by $M(F)=\{e\in E\setminus F:(V,F\cup\{e\})\text{ has no cycles}\}$, i.e., adding $e$ to $F$ does not form a cycle.
Players: There are two players, Max and Min:
  • Max aims to maximize the total weight of the resulting spanning tree.
  • Min aims to minimize the total weight of the spanning tree.
  • The players alternate turns, with Max moving first.
Strategy: A strategy for a player is a function $\pi:\mathcal{F}_{G}\to E$ that selects a legal move from each configuration $F\in\mathcal{F}_{G}$. The outcome of the game, $T(\pi_{\text{max}},\pi_{\text{min}})$, is the spanning tree produced when both players follow their respective strategies.
Winning Condition: The total weight of the resulting spanning tree is given by:
$w(\pi_{\text{max}},\pi_{\text{min}})=\sum_{e\in T(\pi_{\text{max}},\pi_{\text{min}})}w(e).$
F.5 Mixed search games
Mixed Search Games involve clearing a graph’s edges by placing and moving searchers, with the goal of minimizing recontamination while preventing adversarial recapture of previously cleared edges[147, 133, 651, 468, 1131]. Extensive research has been conducted on this topic. Related games include Edge Search Game and Node Search Game[576].
Definition F.5.
The Mixed Search Game is played on a graph $G=(V,E)$, where the graph is treated as a system of tunnels. Initially, all edges in the graph are contaminated with gas. The objective is to clear the graph of gas using a sequence of operations that involve placing, removing, and moving searchers. The game is defined by the following rules:
  • Clearing an Edge: An edge $e\in E$ is considered cleared if either:
    1. 1.
      Searchers are placed at both endpoints of the edge simultaneously, or
    2. 2.
      A searcher slides along the edge, from one endpoint to the other.
  • Recontamination: A cleared edge is recontaminated if there exists a path from an uncleared edge to the cleared edge without any searchers occupying any vertex or edge along the path.
  • Search Operations: The search progresses through a sequence of allowed operations:
    1. (a)
      Placing a new searcher on a vertex.
    2. (b)
      Removing a searcher from a vertex.
    3. (c)
      Sliding a searcher from a vertex along an adjacent edge to another vertex and placing the searcher at the destination vertex.
    4. (d)
      Sliding a searcher from a vertex along an edge but leaving it on the vertex (this operation clears the edge during the movement).
    5. (e)
      Sliding a new searcher along an edge and placing the searcher at its endpoint.
    6. (f)
      Sliding a new searcher along an edge without it stopping at any vertex.
Objective: The objective of the game is to clear all the edges of the graph while minimizing the number of searchers used and ensuring that no cleared edge gets recontaminated.